A Bibliography of Publications in ACM SIGOPS Operating Systems Review

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

07 August 2025 Version 1.55

Title word cross-reference

3 [ASR^{+}17, GPY^{+}17, KDS^{+}06, MAS^{+}06]. < [Zho16]. > [Zho16]. ^{T M} [FPG89]. ≈ [KLK17]. δ [ZLX01b]. μ [Dru92, HHLS97, LBB^{+}91].

* [TYKZ07]. */ [TYKZ07]. *icomment [TYKZ07].

-core [ZLX01b]. -fully [FPG89]. -kernel [Dru92, LBB^{+}91]. -kernel-based [HHLS97].

/*icomment [TYKZ07]. /evolution [Pat02a].

0 [Woo85]. 0-932376-52-5. [Woo85].

1 [SF80, SDV^{+}87]. 10 [BBMT72, SHSB75]. 11 [MA79, PK75, Ros78, ZDP83]. 11/45 [HB80]. 13th [Laz92a, Laz92b]. ’16 [Tsa16]. 17th [LE00]. 191 [VHAA25]. 1981 [Wai83a]. 1982 [Had83, OSV82, Sta83, Wai83b]. 1983 [Fêa83, Had84, Had85, OST83]. 1984} [San86]. 1986 [OSV86]. 1988 [Lev88]. 1994 [Sat95]. 1st [BY08, CM13].

2 [Als72, WCW^{+}04]. 2.0 [KL07]. 2000 [Pas92]. 2007 [Isa08]. 2009 [Bel10]. 2012 [MvR13]. 2013 [SG14, VZ14]. 2015 [DK15]. 21364 [MSB^{+}02]. 2L [SDH^{+}97].

3 [HBG^{+}06]. 370 [San81]. 3G [JHC^{+}11]. 3rd [PP09, Sub11, WTC09].

4 [PCD91]. 4.0 [Vog99, Vog00]. 432 [PKW81]. 4381 [GPR87]. 45 [HB80]. 4th [BCC^{+}94].

5. [Woo85]. 5th [BR10, LS09, OFB16, RAVC12].

60 [BB75]. 6000 [HO91]. 64 [ZRMH00]. 64-bit [CBHLL92, THK95]. 6th [BM17].

736 [AD07]. 7th [Gue87].

80 [BSUH87]. 802.11n [AHB15]. 8800 [Cla87]. 8th [CJRV15].

9 [Gan08, PPT^{+}93]. 9th [CvR14].

AAMP [Bec90]. Abadi [Nes90]. Abraham [Bla95]. Abstract [BAD^{+}11, Fab73, HKU79, Bor98, CV93, CW92, CG85, CMMS77, CSS^{+}91, ECS73, FR85, Her87, Hol82, KLS85, LLS91, LGJS91, Mcd77, RK77, Ten96, Van96]. Abstraction [Bal24, DDK^{+}16, LSAS77, SWL77, Bak95, CF89, FH85, Mad81a, Mad81b, RCL01]. abstractions [KYB^{+}07, WFHJ07]. abstracts [Laz92b]. Accelerating [BSSM08, CFR98, KXWB17]. Acceleration [FZY^{+}23, GPY^{+}17, NS16]. Accelerator [OHW17, TPO06, YZG^{+}11]. Accelerators [CYMT16, CYG^{+}17]. accent [FR85, RR81]. Access [Wol02, BHLM94, BSL08, BD91, EM06, GS90, Gór78, Gsc94, JdLT^{+}95, KS09, KKC02, LTQZ06, LPH^{+}07, MC91, MS91b, MES95, Nee79, PSC^{+}07, SHV01, SFL^{+}94, WS06, WS92]. access/execute [BD91]. accessibility [Lie94a]. accessing [ACM02, OB10, ZXMJ04]. accountability [Gup05, HKD07, HBP06]. accountable [Hae10]. Accounting [LSV^{+}19]. Accuracy [CKN^{+}19, VYW^{+}02, YS94]. Accurate [CPT08, LB06, LZH^{+}22, PRD10, VGX17, EEKS06, GAT13]. Achieving [KMK16, Les04, KGGK09, WTKW08]. ACM [Bac99, BK08, Bre08, DNT10, Laz92a, Laz92b, Sub11, Tan97, WTC09, Bab91, MM92, MM93]. ACM/ONR [MM92, MM93]. Acquiring [SETB08]. acquisition [LK08]. Acrobat [Wai97a]. Acrobat/PDF [Wai97a]. across [ATMZ01, GS95]. Action [CJM15]. actions [AM85, BRR^{+}00, Küh98, Lom77, Nat80, OLS85, RZ97, VA96]. activation [CHCmWH00]. activations [ABLL91]. Active [AUS98, GJXJ03a, Ten96, Wet99, Wet00, ACD^{+}14, EENV02, GUB^{+}08, LW01, SADAD02, TNA12]. ActivePointers [SBS18]. Activities [ZFP^{+}21, JRR97]. activity [Ens75, GNB^{+}09, MB08]. Ad [EM06, BBD^{+}02, BBBAN04, MFHH02, ÖGA06]. ad-hoc [MFHH02]. Ada [Hil92, Taf82]. Adaptable [Lux95, LLK96, MM81]. adaptation [BAMM77, FS99, FS00, NSN^{+}97, SNKP95, TS06, dLWZ00a, dLWZ00b, MCdL06]. adaptations [SHA02]. adapter [VFH98]. Adapting [FGBA96, WCS08]. Adaptive [LMV12, Led97, MRH^{+}16, PSZ^{+}07, RF17, SKI08, WQA^{+}24, dSFdAM13, CPM10, CALM97, ESB^{+}06, HC04, KBK02, LP01, MR07, MRC^{+}97, MPC08, PG96, PLH98, RA06, WAB^{+}89, XXMC05, YRC05]. adaptivity [LB08]. Adding [AR07, Nee79]. additive [LC04b]. Address [CBHLL92, CB17, EMZ^{+}16, Lie95b, SBS18, ACM02, BMvdV93, CIL93, CBD^{+}98, Est02, Goo87, KSS^{+}96, Lie94b, LNBZ08, MS00, Ros94, SS95, THK95, ZZP04]. address-based [KSS^{+}96]. addressed [IKWS92, Lie95c]. addressing [CKD94, CCH^{+}87, MB80]. administration [Del80, GDRT13]. administrative [HK00]. admission [NXQ05]. advanced [ST01, Cri91]. Advances [AC23]. advantages [WSH94]. adversaries [CMSK07]. advertised [PSB06]. affects [RR72]. affinity [VZ91]. AFS [SS97]. after [KBB^{+}06]. Against [AYQ^{+}16, BS15, BK12, Mit00, PB08, TNL^{+}07]. Age [YBQ^{+}25, LC04b]. Agent [YG25, CWL05, GXJJ03, KXD00, LCKFA24, SH00]. Agents [YBQ^{+}25, Jon93, KG99]. aggregate [Ste97, VFMM08, WK05]. aggregates [Str12]. AGgregation [MFHH02]. aggressive [SKKM02]. Agile [NSN^{+}97]. agreement [Che04, LKKY03a, LKKY03b, WYC03b, WY04, YWC04, YS02, MSF85]. Ahead [KKB^{+}16, CR12, DRTT24, MA10, SJSM96]. AI [BKGT25, BH21, JXQ^{+}22, LCKFA24, YBQ^{+}25]. aid [AEG^{+}91]. aided [HLL^{+}02]. Aims [Ano75]. air [EKF^{+}14, Wal73]. AjaxScope [KL07]. Akamai [Bel10, NSS10, NSW10, RCSW10]. alchemy [Pra86]. algebraic [GHM77]. Algol [BB75]. Algol-60 [BB75]. Algorithm [DDM^{+}18, BMD94, BL00, CH81, Cha96, CCB^{+}06, Fog74, Fon72, Gre72, Gup01, Hof90, KSS73, LS75, Lei89, LB81, MPC08, Mil92, Nai93, RH97, Riz97, Sad75, SJGY94, SD86, WJ98, Woo90, XHJB99]. Algorithm/Architecture [DDM^{+}18]. algorithmic [DH10]. Algorithms [SHW^{+}15, AUS98, BBBAN04, BM90, CPM10, DGH^{+}88, DMD13, ELG95, FFM07, GLL04, Kai75, KY02, KTP^{+}96, LRW91, LA94, MSB^{+}02, Ray91, RL96, SS94]. Alibaba [Che17]. alignment [CG94]. alike [Hol82]. all-software [JKW95]. allergies [QTSZ05]. allocating [WC02]. Allocation [KXWB17, WM16, BL00, BEH91, CFL73, ELG95, LFZE00, Mah98, SSS01, Ste83, UHMB94]. allocator [BMBW00]. allocators [ROLV06]. Alpha [MSB^{+}02]. Alphard [SWL77]. AlphaServer [GSSV00]. alternative [BMW02b, GLG93, MSC^{+}06, SPF^{+}07]. alternatives [BVR^{+}00, HM93]. Amazon [DHJ^{+}07]. Amber [CAL^{+}89]. Amdahl [SBH^{+}10].

Amdahl-balanced [SBH^{+}10]. AMNESIAC [AK17]. Amnesic [AK17]. Amoeba [TM81, ZSK97]. among [Bre83, HZ09, SJ95]. AMPI [ZHK06]. Amsterdam [Lit87]. Analyses [WHZ^{+}17]. Analysis [ASR^{+}17, BS15, CCM96, CKN^{+}19, Duc89, FXZ^{+}17, KL98, Küh99, LS75, LML00, MCN^{+}17, NHH^{+}17, Nut94a, Pot77, PBM22, RL96, Rob96, dSM16, WP91, BBFH07, BBC^{+}06, Bod11, BBM09, BMER14, CHV04, CKDK91, DS06, DH10, DBH^{+}06, DKC^{+}02, GMM98, GFPcF08, Had85, HGB^{+}80, LST^{+}06, LFWL10, MT96, NXQ05, OST83, ODH^{+}85, REL00, Sad75, SK96, SGD^{+}02, Sny77, SS98, SAF07, TLD^{+}11, TP72, TACT08, WGL^{+}08, WAC^{+}81, ZWWL01, ZL04b]. Analytics [Che17, HDGP21, RD12]. analyzers [RR04]. Analyzing [FD10, NS16, ENCH96, ME08]. Andrew [dV96]. Ann [Wai83a]. Annotated [YM93, FHL95, Gan92, New79, WKT^{+}13, ZK88]. annotation [QPP02]. annotations [Wei98]. Annual [BY08]. Anomaly [HT15, CG06, MZI08, MC91]. anti [CXMX05, RSW08]. anti-phishing [RSW08]. anti-spam [CXMX05]. antichains [CL01]. Anticipatory [ID01]. Antiquity [WECK07]. ANVIL [AYQ^{+}16]. any [CSBA17b, Den74b]. AOP [MFGSP12]. AP1000 [HDH^{+}94]. aperiodic [LLK96]. APls [WYW^{+}25]. apparently [Rou84]. append [CMSK07]. append-only [CMSK07]. Appends [DRTT24]. Appliances [RHMR15, BSM^{+}12]. Application [AW17, ALM^{+}18, BMP^{+}04, CDY^{+}17, HC92, JSDG08, KEG^{+}97, SNKP95, Sha95, SRS22, TZZ^{+}18, ATSV06, BvS00, Bec90, BRR^{+}00, BCE^{+}95, Bod11, CLM^{+}07, CB95, DZ95, EKO95a, GS78, HL92, KLS^{+}10, LLS^{+}08, MPHD06, NSN^{+}97, PCH^{+}14, Pra87, SFW99, USR02, YWC04, ZJS^{+}11]. Application-aware [SNKP95, NSN^{+}97]. Application-Controlled [HC92]. Application-level [BMP^{+}04, JSDG08, EKO95a, PCH^{+}14, ZJS^{+}11]. Application-specific [CDY^{+}17, BCE^{+}95]. Application-Transparent [ALM^{+}18, AW17]. Applications [BHD19, DJS^{+}17, FZY^{+}23, Had93, HJrCH16, ID25, MAHK16, NTC^{+}21, Sub11, UJE^{+}22, Val94, Wai83b, ACT94, BDMS98, BFSG94, BMBW00, Bla85, BGS04, CCZ07a, Cos13, FURM00, FS99, FS00, GAK^{+}02, GS89, Hop90, JBDP08, KSP09, KS85, KL07, LCJV^{+}11, LHPL87, LGN07, LP01, MT02, MCM07, MDK96, NL95, NL97, NXQ05, NSS10, dOL12, RA06, RBLP07, Sat95, SPHH06, SAG06, SSR^{+}10a, Sta83, Tai13, Tri82, Tri02, VE08, WBB02, WYA^{+}07, YS98, ZWZ01, Tan97]. applicative [FW77]. applying [BDDMR11, MT02]. Approach [DDOL16, HS16, VJ19, XD17, Bac81, Bec90, Bos06, CGL^{+}08, Che85, CXMX05, CGS96a, Edi13, ECH^{+}01, FS95, GHP^{+}08, JT90, JW01, Kah72, Lor86, MSA^{+}00, Moo92, MPC^{+}02, NB91, NBW87, OCLN14, OMCB07, PSC^{+}07, RB93, Rei85, Rob08, Rou84, SGT96, SW00, Svo81a, TPH12, TNA12, War76, WDH89, Wei98, Won93, ZLX99, ZL04a, Zim94, dJKH93]. approaches [KXD00, NRS13, ÖGA06, SH87]. appropriate [AYK08]. Approximate [GSCM16, JSCM17, PAM^{+}16, SLFP16]. Approximations [VGX17]. APSys’15 [HKPvR16]. Apt [RWS^{+}15]. arbitrary [GMM98]. Arbitration [SKJ^{+}17, MSB^{+}02]. Arbor [Wai83a]. ARC [Wis05]. Archipelago [LNBZ08]. Architectural [BF87, BMA00, CB95, DLLN18, KKK^{+}17, TML^{+}00, DBMZ08, HO91, HDH^{+}94, ÏMC^{+}06, KSS^{+}96, RBH^{+}95, Rou84, TNNI87]. Architecture [ACAAT16, CJM15, DDM^{+}18, FXZ^{+}17, GSSV00, KK84, LJdL^{+}16, MCN^{+}17, MRH^{+}16, PC75, ALBL91, Ant90, AB75b, BJ81, BMTW91, BKN05, BC91a, BSF^{+}91, BMR^{+}09, BJM^{+}96, CM87, CNO^{+}87, CLDA07, DS09, DB96, EKF^{+}14, EKO95a, Est02, EEKS06, Fle81, GNA^{+}98, GBBL85, GB93, Har85, HFWZ87, Her78, HSW^{+}00, Jon80, Kie87, KBC^{+}00, LLA^{+}81, LCTK01, LE96, LWMX05, MPPZ87, MSP^{+}06, MP81, Mil77, OPSS93, PJDL06, PK75, REL00, RBLP07, Ros06, San81, SLN00, SLN99, SS72, SV06, SHC73, SZII11, WCB01, Wil80, WH94, WS91b, XFO08, YZJ02, YW06, YLE02, YTM^{+}91, ZRMH00, LaR92]. Architecture-Adaptive [MRH^{+}16]. architecture-dependent [LE96]. Architectures [CSBA17a, CB17, Dru92, GF15, KKK^{+}17, BJM^{+}91, BH75, CGKM11, CCB^{+}06, DSBK04, EBP16, GTK^{+}02, KGS06, MCR^{+}09, PBH^{+}07, PG73, RTY^{+}87, RKBH11, Rip03, SS95, TF04, WBDF97, Wir87, CBHLL92]. Archival [BLC^{+}16, Sal91]. archiving [RCC01]. area [AEE^{+}94, BvS00, DKK^{+}01, Fab98, GS95, HM91, MPC08, SKKM02, Sha00, WECK07]. ARGOS [DH73, PSB06]. argument [GKS11]. Argus [LCJS87]. Ari [Dos88]. Arithmetic [EPG^{+}20]. ARM [DLLN18]. array [TNA12]. array-oriented [TNA12]. Arrays [ACS15, DBP^{+}04, LK91, SFV^{+}04, ZCT^{+}05]. Art [WP91, EFL07]. Artifacts [Eid15, SRH^{+}06]. Artificial [KEF^{+}19, YWKYS15]. ARTS [TM89]. Artur [Bar14]. asbestos [EKV^{+}05]. ASCI [BBH^{+}00]. Asia [HKPvR16]. ASIC [KZVT17, XDM^{+}18]. Aspect [BS15]. Aspect-Oriented [BS15]. Aspects [FW77, Had83, LST^{+}06, McN77, McN82, McN88]. ASPLOS [Tsa16]. Assembly [HS16]. assertion [PW93, Per92]. assess [HD12]. Assessing [AFB95, GTSS11]. Assessment [NCBB14, VPH^{+}15]. assist [BKT87, KKM^{+}06]. assisted [LHL04, RSW08, SN94]. associated [HM90]. associative [Ger72]. associativity [GMM98]. Assurance [AHC^{+}16, RAF07]. asymmetric [Bla83, KF09, RRBN09]. AsyncClock [HNK^{+}17]. asynchronization [WM80]. Asynchronous [Col73, HNK^{+}17, VTGH17, WYC03a, ATK92, BDM97, BN78a, BJL^{+}06, DDYM99, FV06, GN80, JXHQ02]. Asynchronously [LL16]. asynchrony [ES10, Ste97]. ATM [CKMV99, PS98, PS99a, PS99c, PS99b, PS99e, Pat02a, Pat02b]. Atomic [IKK16, MA91, Nat80, SKB^{+}17, BRR^{+}00, HV92, JLR^{+}05, LB91, Lom77, OLS85, RZ97]. atomicity [Her86, LTQZ06]. attached [Van96]. Attacking [Zay87]. Attacks [AYQ^{+}16, MMT16, DH95, HL05, KTC03, LSH00, PSB06, SW00, YS02, dVdVI98]. Attaining [GKS11]. attempt [Fou74]. Attested [CMSK07]. auction [Tai13]. auction-based [Tai13]. Audio [BCC^{+}94, JH93, RV91, Her92b, Jef92]. auditable [PS09]. augmenting [PG06, dORF12]. August [Lev88]. authenticated [Che04, LKKY03a, LKKY03b, SMBA10, YS02]. Authentication [AC97, CV93, CGM97, LABW91, Lie93a, LSH03b, NS87, WABL93, ABC^{+}98, BO99, BAN89, Car94, CCK04a, CL04c, CC04, CCK04b, CC05, CH07, DH96, Gan08, Gif81, HYS03, HLL04, HL05, KC95, KLY03, KTC03, KCL03, Ku04, KC05, KCC05, LHY02, LLH02, LKY04, LW04, LHL04, LFW04, LC04a, MC96, Mit00, ME08, OR87, PS98, Sco04, SY96, Syv93, WL94, YW04, YRY04, YbJf04, LSH03a]. authentications [KSL92]. Author [Den07]. Authorisation [CL01, LM97]. authorization [VA96, YbJf04]. Authorizing [WYA^{+}07]. Auto [KSP09]. Auto-tuning [KSP09]. AutoBash [SAF07]. Autograph [PKM^{+}09]. Automated [Arn10, KP97, LWPG17, TAH^{+}22, VPH^{+}15, YLW^{+}06, VM07]. Automatic [AK17, ACD^{+}14, APGG00, BAMM77, BA06, CG00, FVDS20, MDK96, RS91, ZBN07, CLM^{+}07, FM02, GBZP10, HB06, HCZ97, Isa07, JM95, PSB06, ZWG^{+}97, ZHK06]. Automatically [LLL^{+}17, SPHC02, LPH^{+}07, PKM^{+}09, RR04]. automating [PLHM08]. Automation [Cri91, WKL07]. Autonet [RS91]. autonomic [SWC08]. Autonomous [GS95, BM99, Sal78b]. Autopilot [Isa07]. AutoRAID [WGSS95]. Availability [BO91, AGM93, Bro00a, Bro00b, Cri94, yL91, SBL99, SBL00, WS91a, YD02, Yu00a, Yu00b, YV01, ZSS08]. available [ABC^{+}02, DHJ^{+}07, Kil00, NLO95]. average [SLCG89]. AVIO [LTQZ06]. avoidance [Lev05, Pea89]. Avoiding [BLRC94, Fon72]. Award [Mog08, vR14]. Aware [BLI17, HABZ17, ID25, JV21, KSCK17, LSL^{+}17, LCCZ17, PFK^{+}22, BDMS98, CEV00, CCHV11, DB11, DB97, EDZ07, FNRC^{+}07, FS99, FS00, GS13, HEKSP11, KAI^{+}13, LFZE00, LSKK08, MVKA06, MB08, NSN^{+}97, PAB^{+}98, RF17, SNKP95, TAS07, TLL03, WBB02, MCdL06]. Awareness [CYMT16]. AxGames [PAM^{+}16].

B2 [Loe85]. back [Mat06, Mit96]. backfilling [LCJV^{+}11]. background [VKD02]. backing [Del80]. backoff [Gup01]. backup [COS^{+}08, CMN02, Rus77]. bad [CCZ^{+}07b, HYM10, TYKZ07]. balanced [GSM08, LH04, PHYO96, SBH^{+}10]. balancer [JXY95]. Balancing [MB06, AEP^{+}97, BMD94, EDZ07, HBD95, LL98, PL95, ZWL09, ZSK97]. ballot [Lee99]. BAN [XZZ97]. BAN-like [XZZ97]. band [PBYH^{+}08]. bandwidth [BSR06b, DP93, Fab98, GNA^{+}98, HS96, LWY^{+}04, LBJ03, MCM01, SGK^{+}04, SS07, SF91]. bandwidth-minimizing [SS07]. Banking [Tai13]. banks [SCL96]. Bankshot [BXS14]. Bantam [Val94]. BAR [AAC^{+}05]. Bare [RB24]. Bare-metal [RB24]. barrier [Joh91, VBLM07]. Baruchi [Bar14]. base [ACC^{+}08, HPM93, SDE85, WH08, RCL01]. Based [AYQ^{+}16, BLC^{+}16, NCBB14, PG16, SKJ^{+}17, SLD15, WM16, ADG^{+}07, AEP^{+}97, AMA^{+}11, BJ81, Bab90, BL00, BK12, BMR^{+}09, BS95b, BP91, CKD94, Cec00, Cha96, CL04c, CKmWH16, CGL^{+}08, CGKM11, Che84, CG06, CLC05, CXMX05, CNV^{+}06, CGJ^{+}07, DB99, DB00a, EFL07, Fab98, Fab73, FAH^{+}06, FFM07, FGC^{+}97, Gai72, GSGN00, Gon89, GBCH00, GXJJ03, GLL04, Gup05, HLL^{+}02, HHLS97, HCZ98, HDL^{+}02, HF08, HFC^{+}06, HLFZ97, HL96, HSPC01, JY98, JKH^{+}00, KCD^{+}81, Kam13, KSL92, KLY03, KEP07, KSS^{+}96, KCL03, Ku04, KCC05, KKC02, LMG^{+}07, Lan89, LH04, LSKK08, LJX97b, LBB^{+}91, LLS^{+}08, LBJ03, LT11, LCH^{+}81, LF13, LM96, LG04, LZ03, LJW^{+}06, MMN08, MW75, MXXC05, MPC08, MA06, MC91, MD81, Mit00, MT85, MB80, Nai93, OCLN14, ÖGA06, PAB^{+}98, PSG06, RS08, RG02, RCC01, RB93, RRP06, RBLP07, RMS98]. based [Rou84, SBL99, SBL00, SHV01, SMS11, SGNG00, SG10a, Sco04, STYC02, SCG01, SPF^{+}07, Son05, SH00, SKPG01, Sto07, ST01, Taf82, Tai13, TSF90, VA96, WC02, WLAG93, WG08, WCW^{+}04, WHZ^{+}17, Wet99, Wet00, Wis05, XXMC05, YW04, YW06, YD02, ZZ03, ZHK06, ZS06, ZIL96, ZXMJ04, dLWZ00a, dLWZ00b]. Basic [Nut94a, dVdVI98]. batch [AVZR11]. batching [dSFdAM13]. Bate [Woo85]. Battery [HT15]. Baymax [CYMT16]. Bayou [TTP^{+}95]. BBFS [HR92]. be [Gur07, MMTW10]. beast [FGR^{+}07]. Beating [OD89]. beauty [FGR^{+}07]. Bechtolsheim [Lig94, Lig95]. Bedford [Had83, Woo85]. been [HH88]. before [Spr85]. behave [CGKM11]. Behavior [KTG^{+}17, NS16, BSL08, ECH^{+}01, GSA10, KL07, Mog06, PLH98, REL00, RR72, SZ98, SPHC02, XFO08, YZZZ06]. behaviour [BR09, HKL^{+}06]. behavioural [MMN08]. Behind [JW24, vR14]. Bell [BM06]. Ben [Dos88]. Ben-Ari [Dos88]. Benchmark [BCR^{+}14, CKN^{+}19, GKV07]. benchmarking [FLM^{+}08, RB24]. benchmarks [Bro00a, Bro00b, CKDK91, VE08]. benefit [Fes07]. benefits [KPR^{+}08]. best [SPBP06]. Better [CCM08, DL15, JS08, KF09]. Between [KHG^{+}17, PVB17, Cos13, GHW07, GC96, GC05, GSW^{+}17, Klo80, LWY^{+}04, RK83, SZII11, SMM^{+}09, ZCSM02]. Beyond [BBB^{+}17, Löh77]. BFXM [LJX97b]. Bibliography [Smi78, CNL89, Esk96, Gan92, Lie93a, New79, WKT^{+}13, YM93, Zöb83, ZK88]. Big [Che17, Wai95b, Lee99]. bilinear [BDDMR11]. billions [SMBA10]. Binary [MCN^{+}17, Kot88, Le98]. binding [Lin81, PB96]. bit [BM06, CBHLL92, Lon93, THK95]. bit-encoding [BM06]. bit-mapped [Lon93]. bits [BJ81]. BitVault [ZLL^{+}07]. Black [CSBA17a, KJ08]. Black-box [CSBA17a, KJ08]. Blade [WL15]. blades [SBH^{+}10]. Block [ID25, DY01, FFBG08, GJXJ03a, GUB^{+}08, LH04, SJGY94, SJSM96, WF07, WSH94, YVM13]. block-device [FFBG08]. blocked [LRW91]. blocking [CCZ^{+}07b, FD10, GC96, RB93]. blocks [IBY^{+}07, Lie95c]. BLR [BDDMR11]. BLR-D [BDDMR11]. Blue [AUW08]. board [BCRS10, SMS11]. Boki [JW24]. Bolt [DK17]. Book [Bec75, Bla95, Had84, Had93, Heu97, Kad95b, Kad95a, Lig94, Lig95, Lit87, Nut94a, San86, Val94, WP91, Wai86, Wai94, Wai95b, Wai97a, Wai97c, Wai97b, dV96, Bla95]. BookKeeper [JKR13]. Books [Val94]. Boosting [YVM13]. bootstrapping [HB06]. BorderPatrol [KJ08]. Borrowed [DC99, DC00]. Borrowed-virtual-time [DC99, DC00]. BOS} [RP07]. both [Har82, SPR00]. bother [vR94]. Bothnia [CCW^{+}11]. bottleneck [OD89, Zay87]. bottom [BC01]. bottom-up [BC01]. Bouncer [CCZ^{+}07b]. Boundaries [BLJ^{+}17, HK00]. boundary [Pet93]. bounded [BL89]. bounding [SGK^{+}04]. Bounds [GG73, YW05]. box [ADAD01, CSBA17a, KJ08]. Bozyigit [RS02]. Brain [Wil16]. brainiacs [Gur07]. branch [CG94, CCM96, CPT08, KT91b, SJSM96, SEP98, YS94]. Brazil [LGMF14]. Brazilian [BM17, OFB16, dSBP11, BOB15, FBL^{+}12, LGMF14, dOS08]. break [PW98]. Breaking [BLJ^{+}17, GMT16]. brick [LG04]. brick-based [LG04]. Bridging [Cos13, GSW^{+}17, PG06, PVB17, RKBH11]. brief [Nut94b, Sch73b, Van96]. bring [ZUW^{+}09]. Bringing [PWT^{+}19, PPS^{+}18]. broadcast [KTH89, LB91, MA91, Oes91, PP83]. broadcasts [AN02, EGE02]. brokers [LWY^{+}04]. Brother [Wai95b, Lee99]. Browser [PVB17]. browsers [WFHJ07]. Browsix [PVB17]. Bruce [Val94]. BRUWIN [MM81]. BSD [ODH^{+}85]. buddies [WTLS^{+}09]. buddy [BL89]. Budget [WM16, Tai13]. budget-driven [Tai13]. Buffer [SEF^{+}16, BRW89, GCM^{+}94, GKL95, JADAD06, Pea89, Ros89, TL96]. buffering [BS96, DD12]. bug [CCM08, LPSZ08]. Bugs [ECH^{+}01, HABZ17, LLLG16, LLL^{+}17, MCXS16, CG06, LPH^{+}07, QTSZ05, TYKZ07, WGL^{+}08]. Build [BNE16, QPP02]. Building [BJKT15, DDOL16, HSI^{+}01, LKvR^{+}99, LKvR^{+}00, LZC^{+}17, SKPG01, TSP17, Wai95b, AMS^{+}07, AUW08, BWV^{+}12, IBY^{+}07, KSL90, KAR^{+}06, LOM^{+}09, LP01, MPHD06, SFV^{+}04, SG10b, WH08, Wil93, MW92, Wei92]. bullet [KSDC14]. Bulletin [BCRS10]. Burns [HV92]. Burroughs [WAC^{+}81]. Burrows [Nes90]. Burst [SEF^{+}16]. bus [ZZP04, OPSS93]. business [DKW^{+}09, YWC04]. BVT [DC99, DC00]. Bytecode [OKN02]. Byzantine [AEMGG^{+}05, BACF08, DY10, HGR07, KAD^{+}07, LS86, MSF85, Ric88, VBLM07, WQA^{+}24, dSFdAM13].

C [Had85, CDG^{+}17, DBMZ08, DMB87]. Cache [BFZ^{+}25, CIP^{+}23, CKJA98, GF15, HC92, KTG^{+}17, LLN^{+}17, MBS16, AEH^{+}08, CKA91, CNC^{+}96, GC89, GJXJ03b, Gup05, HP95, JADAD06, KBK02, LRW91, LC04b, MDO94, MC91, MB91, MS94, OCLN14, PHYO96, PLH98, PEA^{+}96, SS94, SM89, VZ91, WZWZ10, WSH94, YW06, ZYG00]. cache-based [MC91]. Cache-Coherent [GF15, WSH94]. Cache-conscious [CKJA98]. cache-consistency [SM89]. caches [BLRC94, BRW89, GMM98, Goo87, IKWS92, KBK02, KGGK09, SKI08]. Caching [NWO87, Sat00, BXS14, CALM97, CG91, CD95b, yKPR02, KTP^{+}96, MA06, PGG^{+}95, RL96, RD01, SH96, SGN85, SDP^{+}00, Son05, WVS^{+}99, WVS^{+}00, ZZ03]. Caernarvon [TKP^{+}08]. call [BALL89, Cas91, Coo86, CKR08, FS08a, MK91, San81, Spi94]. calls [ATK92, BN83, Had77, JP78, Lis77, Par78, TA90, VMBM12, Wet78]. Cambridge [Dio80, GN80, Her78, NW77, WN80]. Can [BC06, Fle07, MPP^{+}08a]. Canada [San86]. cannot [GS89]. Cap [ZH16, BN78a, Coo78, Del80, NW77, NB77, Nee77]. capabilities [HH88, HB80, Rus88]. Capability [JKS^{+}15, Jon80, CKD94, Fab73, Gon89, Her78, Lan89, MB80, Nee79, SSF99, SSF00, Wil80]. capability-based [CKD94, Gon89, Lan89, MB80]. Capable [Ott18]. capsule [Wet99, Wet00]. capsule-based [Wet99, Wet00]. Capturing [CZG^{+}05, PLH98]. cards [KLY03, PV95, CL04c, CCK04b, HL05, Ku04, KC05, LHY02, Sco04, YW04]. care [HBB13]. Cary [Gra14]. Case [DIS19, KSCK17, LXYZ19, Ser21, SBS18, AKGR10, BJK^{+}06, Bor98, BCDN87, CII^{+}10, DH10, Fab73, Fes07, GUB^{+}08, Hae10, HJT^{+}93, Joh91, Lio78, Mat04, MW08, ONH^{+}96, OD89, OAE^{+}09, PK75, SPHH06, SHSB75, WA09, ZWWL01]. Cashmere [SDH^{+}97]. Cashmere-2L [SDH^{+}97]. CASPAR [GMT16]. Cassandra [LM10]. Catastrophe [Pra87]. CATOCS [Shr94, vR94]. Causal [RMSB01, vR93, AN02, Bir94, CGS96a, SB91]. causal-consistency [CGS96a]. causal-phase} [AN02]. Causality [HNK^{+}17, KKS^{+}16, SAF07]. causally [CS93, Coo94, Toi92]. Causes [dSM16]. cautionary [Coo94]. CC [VDGR96]. CC-NUMA [VDGR96]. CDN [WPP02]. Cedar [Hag87]. cell [RJK^{+}14]. Cells [GSCM16]. Cellular [GTHR99, GTHR00]. Center [UJE^{+}22, Zha23, Isa07, RRT^{+}08]. centers [AVZR11, CAT^{+}01, Cos13, GSM08, RJK^{+}14, WTB10, WKT^{+}13, WTLS^{+}09]. central [Bas72, SDV^{+}87]. centralized [PG06]. centric [ZYG00]. CERN [PBM22]. certain [Kno74, Kno75]. certificates [DS90]. certified [CJ05]. CFS [DKK^{+}01]. CGI [Wag98]. chairman [Lam75]. Chairs [BK08, DK15]. Challenge [San86]. Challenges [AC23, HZ09, JSS^{+}15, SG10b, VAK^{+}11, Wit16, AAA^{+}23, CR12, DW07b, Est02, FM98, MA10, WTB10]. Change [DHK^{+}15, KSDC14]. changing [JBDP08, Wil94]. Channel [HMK20, Loe89]. channels [Loe85]. CHAOS [GS89]. Character [BN78b]. Characteristics [SHW^{+}15, BM99, HO91, LPSZ08, MDO94]. Characterization [LPM17, PBM22]. Characterizing [BvS00, SPHC02, SJ05]. Charge [HT15]. Charm [ZHK06]. CHART [BMR^{+}09]. charter [Sop84]. Cheap [CL87, CMN02, Lie96]. check [CCEH00, KKN00]. checkers [HV08]. Checking [BR09, BKL^{+}16, BNE16, HABZ17, MCXS16, FL77, KL02, MPC^{+}02, NL96, PW93, SH87, Per92]. Checkpoint [RS02, BBHL08, BW01, RPM97, YD02, ZWZ05, ZHK06]. checkpoint-based [YD02, ZHK06]. checkpoint/restart [BBHL08]. checkpointing [BMP^{+}04, Kan83, TSLBYF08, VGBT14, WJ98, ZWZ01]. checks [NPCF08]. CHERI [CDG^{+}17]. Cheriton [Bir94]. Chip [ACAAT16, BSL08, HWO98, KDS^{+}06, KBK02, ONH^{+}96, SKI08, vdWMH11]. chips [MAS^{+}06]. choice [RR72]. Choices [CJR87]. chosen [LGSN89]. chronics [KJH^{+}11]. Chunked [AKP^{+}25]. cinematic [CSJZ08]. ciphers [DY01]. circuit [LCH^{+}81, PV95]. CIRCUMFLEX [WBR^{+}12]. CISC [BC91a]. CISCs [BCDN87]. CLANGER [Ros95]. Clarifying [KS99]. class [CJR87, Loe85, MSLM91, Mou96, MRS09, Smo95, SS17, ZELV02]. classic [Hof90]. Classification [MBS16, AWW08, Küh98, Lev05, MMN08]. classifying [Zöb83, ZK88]. cleaning [Rob96]. click [MKJK00, MKJK99]. client [BKN05, DPW^{+}09, EBS01, FGBA96, KL07, SLN00, SLN99, CSBA17c]. client-initiated [EBS01]. client-side [KL07]. ClientVisor [CJS^{+}09]. Cliffs [Sta83, Wai83b]. clock [LS86, Ric88, UHMB94, WJMC04]. clocks [DM90, Gon92, Ray92]. clone [Tan87]. closely [KLS85]. closely-coupled [KLS85]. ClosestNode.com [WS06]. closing [GKO^{+}00]. closure [Neu89]. Cloud [Bas12, BHD19, CM14, DK16, DK17, GZH^{+}19, GLD^{+}22, GSW^{+}17, KHG^{+}17, LZH^{+}22, LLL^{+}17, YBQ^{+}25, YJX^{+}16, BTMS10, BKP^{+}12, BCC^{+}13, BK12, CM13, DZP^{+}11, Edi13, Hae10, HYM10, MRS09, dOL12, OB10, RRCC10, SG10a, SK13, VESM10, WL09, vdWMH11, Che17, RRCC10]. cloud-based [BK12]. Cloud-Hosted [BHD19]. Cloud-TM [RRCC10]. Cloud9 [CZB^{+}09]. Cloudifying [SW10]. Clouds [KMK16, KZVT17, XDM^{+}18, BJK^{+}06, KMK10, LMV12, PPO14, SSR^{+}10a, Tai13]. CloudSeer [YJX^{+}16]. Cloudsim [OBSR16]. CLU [LSAS77]. Cluster [FGC^{+}97, ELG95, FMP^{+}95, GW04, GBCH00, JKH^{+}00, PAB^{+}98, SBL99, SBL00, STYC02]. Cluster-based [FGC^{+}97, GBCH00, JKH^{+}00, PAB^{+}98, SBL99, SBL00, STYC02]. Clustered [DJS^{+}17, AEH^{+}08, ENCH96, SDH^{+}97]. Clustering [LSL^{+}17, CZG^{+}05, TAS07]. Clusters [GSW^{+}17, HJrCH16, Cec00, GTHR99, GTHR00, HCJ07, JXY95, LK10, LZJ03, MSF85, PL95, RRBN09, TDM12, YD02]. CMC [MPC^{+}02]. CMP [TAS07, CWS06, GPV04, MCR^{+}09, SATG^{+}07]. CMPs [SQP08]. CNNs [RHR^{+}17]. Co [AVN^{+}16, DDM^{+}18, Had84, Had85, KSCK17, LXYZ19, San86]. Co-Design [AVN^{+}16, DDM^{+}18, KSCK17, LXYZ19]. coalescing [BL89]. coarse [Dub00, GTA06]. coarse-grained [Dub00, GTA06]. Coasting [GB01]. COATCheck [LSMB16]. Coda [KS91a]. Code [BD91, BNE16, EPG^{+}20, MRH^{+}16, PB09, WHZ^{+}17, CCEH00, EP94, ECH^{+}01, GA98, Jon93, MPP^{+}08a, MRA87, MPC^{+}02, NAR08, SFW99, SLS^{+}05, SLQP07, SW10, SJ95, Tan87, TACT08, VE08, Jon92]. Code-partitioning [PB09]. codes [JKL^{+}13]. Coding [ID25]. CoGENT [AHC^{+}16]. Coherence [CIP^{+}23, HCBS04, OHW17, YVCB18, BKP^{+}96, CKA91, HCW^{+}04, HP95, MS94, SHT97, SS94]. Coherency [Goo87, PK96]. Coherent [GF15, CF89, FP89, SDH^{+}97, WSH94]. collaboration [HDL^{+}02]. Collaborative [KHG^{+}17, VJ19]. collaborators [SS97]. collateral [PLM06, PLHM08]. Collection [LW01, Bar79, CHV04, ONG93]. collector [BN78a, GN80, JHT^{+}07, SN94, WK08]. collectors [GTSS11, KPS09, SMTZ09]. colocation [WTLS^{+}09]. Colony [CMK^{+}06]. coloring [BAM^{+}96, GP05]. Coloured [Nut94a]. column [Fle07]. COMANDOS [TCH^{+}91]. Combinatorial [SLTB^{+}06]. combined [CGS96a]. combined-consistency [CGS96a]. Combining [CG91, Cri94, JHK^{+}16, ACM02, Str12]. commands [Ste73]. Commensal [SF12]. comment [Küh99, Lip75]. Comments [Hem89, JW01, Kot88, LL04, MC96, Ng99, RS02, Tro00, Hsi89, TYKZ07, TT00]. commercial [GWSY08, JBDP08, MDO94, Oes01, PAB^{+}95]. Commercially [EENV02]. commit [Hag87, ML85, MSF85, PG96, VBLM07]. committee [Isa08, Sop84]. Commodity [SHP^{+}16, VMM20, ZLJ16, BDR97, CGL^{+}08, CLDA07, GAK^{+}02, MR07, NPCF08, RPNT08, SFV^{+}04, SLQP07, WZWZ10, XLDB09].

Common [RB24, CM06, GW04, WDH89]. communicating [Hab72, Mou96, PL95]. Communication [ACAAT16, Boc75, DB75, MDR^{+}00, OA08, WL15, ADG^{+}07, BS95a, BHM77, BVR^{+}00, BKP^{+}96, Bir94, Bla83, Cer75, CC05, Che75b, Che84, CS93, Coo94, CCLP81, DBRD91, FAH^{+}06, FR85, FH85, GW04, GTK^{+}02, GC05, Had83, HYS03, KT91a, KSS^{+}96, LHWY83, LKvR^{+}99, LKvR^{+}00, MK91, McN77, McN82, McN88, MW75, Neg00, Oes01, OCF00, Opd75, PFGD02, PP83, RR81, Rus88, Sco96, Sor73, WZZ93, WFHJ07, YTR^{+}87, ZCSM02, ZZ03, Fin92]. communication-exposed [GTK^{+}02]. communications [AEH75, Car94, CNL89, LWY^{+}04, LC04a, Owe84, WV02, vdWMH11]. Community [CJM15, WdSA^{+}08]. Community-Supported [CJM15]. CoMon [PP06]. compact [KDS^{+}06]. compacting [ONG93]. compaction [WK08]. Company [Wai86]. Comparative [OSV86, PSK08, DS92, GS90, MSB^{+}02, OSV82, TP72]. Comparing [Her86, PBH^{+}07, BC91a]. Comparison [LCTK01, ZH16, AA06, KTP^{+}96, MMTW10, MD81, MMB96, TSF90, TF04]. Comparisons [AHB15]. compatibility [Gue87]. Competitive [LSP07, KLMO91]. competitors [SS97]. compilation [CCEH00, WS87, Won93]. compile [DCZ96]. compile-time [DCZ96]. compile-time/run-time [DCZ96]. Compiler [BAM^{+}96, CMT94, CH98, LM96, MP85, RSEW04, ZCSM02, CNO^{+}87, CHCmWH00, CBC^{+}08, CSS^{+}91, GTK^{+}02, HDH^{+}94, KY02, MDK96, SS94, ZRMH00]. Compiler-based [LM96]. Compiler-controlled [CH98, CSS^{+}91]. Compiler-directed [BAM^{+}96, CHCmWH00]. compiler-inserted [MDK96]. Compilers [HS16, HZ09, KSP09]. Compiling [BSUH87]. Complete [Gar07, KAR^{+}06, KGGK09, Ull73]. Complex [ACS15, Mog06]. complexity [DV87, FS08b, Sal00, SPHH06, ZK88]. component [GSM08, LP01, LF13, WV02, dLWZ00a, dLWZ00b]. component-based [LF13, dLWZ00a, dLWZ00b]. Components [RF17, EEKS06, Fes07, FRL00, LKvR^{+}99, LKvR^{+}00, MFGSP12, SFV^{+}04, YW05]. Compositional [MCN^{+}17, RBLP07, Bor98]. compound [VMBM12]. Comprehensive [LZH^{+}22, LWPG17, Esk96, LB08, LPSZ08]. Comprehensively [KJS^{+}06]. Compressed [JSCM17]. Compression [BFZ^{+}25, PFK^{+}22, CG91, CCM96, Dou93, Riz97, WSW05]. compromise [PCP00]. Compromises [EPG^{+}20]. Computation [BFZ^{+}25, CWS06, LHWY83, LLN^{+}17, BVCG04, CHCmWH00, HN81, JL75, Kie87, LC04a, MCC^{+}06, Por10]. Computational [BB75, FZL16, Cho77]. Computations [VGX17, BAI93, BK12, FR94, NSKS11]. Compute [GSW^{+}17, CDV^{+}94, EJD13, VDGR96]. Computer [AK17, CJM15, ELR15, Lam00, LGMF14, Lit87, MW09, Mog09, PBM08, RWS^{+}15, San86, Voe98, WP91, Wai83b, Wai97c, AFB95, AUW08, AB75b, AC97, Bas72, CS77, CEC^{+}95, Coo78, CJM^{+}75, DH73, Ell73, Gai72, GS78, GSGN00, Her78, Hol72, HH08, KCD^{+}81, KS95, KSS73, LBP^{+}07, Lam83, LB81, Mad81a, Mad81b, MP75, MV86, NW77, NHM83, NXQ05, Nut74, Opd75, Pop75, Ros06, Rou84, SGNG00, Spe81, Sta83, Svo81a, Svo81b, Tri82, Tri02, vdWMH11]. Computers [CYMT16, CYG^{+}17, BBH96, Fab73, GB93, Han83, JS08, KP97, LHPL87, Rei85, SCP^{+}02, SGGB99, SGGB00, SJ95, Tan79, Wai86]. Computing [BOB15, BR10, BM17, CM14, FBL^{+}12, OFB16, PAM^{+}16, RLD^{+}17, SJS^{+}23, TZZ^{+}18, VMM20, Wil16, XDM^{+}18, AUW08, Bab90, BKN05, BKP^{+}12, BLNS81, BMK06, Cec00, CM13, DHRS91, DB11, ESB^{+}06, EEKS06, Gan92, Gar07, GNB^{+}09, HK99, HdRC95, HCZ98, HEK^{+}07, Hog88, HC95, HL96, JXT93, JOW^{+}02, Lac00, Lev07, LS94, Mah98, MUKX06, Nic87, OLLY02, OVS^{+}06, OSSN02, PSZ^{+}07, Pra86, SNKP95, Sat95, SB10a, SS83b, ST93, SBH^{+}10, TBM^{+}06, VESM10, VAK^{+}11, WH08, Wel95, WL09, Yan92, YD96, dLWZ00a, dLWZ00b, vEBBV95, CM14, Duc92, YGG^{+}03]. concentrated [XX00]. Concept [BCR^{+}14, AN02, Gai72, Lux95, Smo95, WM80, YTM^{+}91]. Concepts [Nut94a, CG91]. conceptual [RBLP07]. Concierge [RA07]. Concurrency [LLLG16, LLL^{+}17, Her87, KHL^{+}07, Lam85, LPS10, LPH^{+}07, LPSZ08, MT85, Wei85]. Concurrent [CSBA17a, CC21, Her92a, ONG93, CSBA17b, Dos88, Fon72, Hol82, KPS09, Kru82, KGGK09, LSP07, Löh77, Rom95, TMW10, WK08, Hei78, SB78]. conditioned [WCB01]. conditions [Dun91, YRC05]. Conducting [AHB15]. Conference [And09, BY08, Fêa83, OSV82, OST83, OSV86, San86, Voe98, Had84, LH04, Ter14]. Conferences [Mog09]. confidentiality [ZZNM01]. Configurable [PKB^{+}16, LAAW00, Maf94, WSW05]. Configuration [TLD^{+}11, GGL^{+}09, SAF07, TDM12, ZBN07]. Configurations [RB24, KMC02]. Confined [VTGH17]. confinement [Lip75]. confirmation [MXXC05]. conflict [BLRC94, GCM^{+}94]. conflicts [TTP^{+}95]. Confused [HH88]. congestion [LBJ03]. connect [NHM83]. connected [DB97, TTP^{+}95, WYC03a]. connection [yKR06, MMN08]. Connections [LCL^{+}16, CGJ^{+}07, SG05]. connectivity [MES95]. conscious [CKJA98]. Consensus [Bal24, HSMC15, WQA^{+}24, ACC^{+}09, FV06, WYC03a, WCYJ05, YW05]. consequences [LK91]. considerations [MW75, YN12, ZRMH00]. considered [And09, Hof07, MPLH06]. Considering [CPM10]. consistence [GJXJ03b]. Consistency [BKL^{+}16, BGS04, DB85, Ell77, LPS10, CGS96a, GGH91, GC89, Gup05, HCW^{+}04, HSPC01, LJX97a, LLY05, Mos93, PRAH96, PCH^{+}14, RMSB01, Ros89, SHT97, SM89, Yu00a, Yu00b, ZIL96]. Consistent [DJS^{+}17, BCRS10, CGS^{+}96b, PST^{+}97, WJ98]. console [BEW75, BEW76]. Consolidated [HJrCH16]. Constrained [KEF^{+}19, RA07]. constraints [AEH75, AS10, JRR97, NCL12]. construct [KS82, SS83b]. construction [HV92, JM95, Lie95a]. consumer [Hil92, HYM10, RB75, Rus77]. consumption [HHS05, KS95, MB06, SCM05]. Container [SPF^{+}07, EKF^{+}14, SG10a, And95]. Container-based [SPF^{+}07]. Containerization [HSL17]. Containerized [HSL17]. containment [CRD^{+}95, CCC^{+}05, VMC^{+}05]. Content [MS91b, MCdL06, BL03, CEV00, CJG02, LJW^{+}06, OB10, SGD^{+}02, Sat00, SCG01, Son05, SAG06]. content-based [LJW^{+}06]. Content-dependent [MS91b]. content-directed [CJG02]. contention [DD80, MCS91]. Context [BFZ^{+}25, KGS06, KKC02, Bla91, DB75, MB91, SG05]. Context-specific [KGS06]. contexts [TE94]. Continual [SRA^{+}04]. continuations [DBRD91]. Continuous [ABD^{+}97, LJdL^{+}16, GA91, HSS^{+}06, TSLBYF08]. contract [WK05]. Contrasting [MDO94]. contribution [CCAP06]. Control [AS10, BR10, GCJ17, LSV^{+}19, Mil77, SLFP16, Arn10, ADAD01, BHLM94, CW92, DB85, DH73, EVvdW89, EKF^{+}14, EM06, Gór78, Gue88, GDRT13, GA08, Her87, HKU79, How72, KHL^{+}07, KYB^{+}07, KKC02, LPS10, LJS^{+}02, Lie95c, LBJ03, LSA^{+}00a, LSA^{+}00b, MCD^{+}08, Mal10, MS91b, MT85, ML97, NB91, Opd75, PSZ^{+}07, PSC^{+}07, ROJS09, Sal73, SHV01, SFL^{+}94, Ste73, TLL94, TG89, Wal73, WKL07, Wei85, Wol02, WJMC04, ZUW^{+}09]. controllable [FS95]. Controlled [HC92, NAR08, Cha73, CH98, CSS^{+}91, Sto07, Wol02]. controller [NXQ05]. Controllers [AMH^{+}16, LER^{+}17]. Controlling [Wag98]. controversy [vR93]. conventional [Nee79]. converged [DPW^{+}09]. Conversational [Rom97]. convertible [CL04a]. Converting [BJ81]. Convolutional [RLD^{+}17]. convoy [BGMP79]. Cool [ACM02]. Cool-Mem [ACM02]. cooling [EKF^{+}14, WTB10]. cooperating [Woo73]. cooperation [RK83, WZZ93]. Cooperative [EKF^{+}14, WBB02, AAC^{+}05, BDT00, DKK^{+}01, DBH^{+}06, JXHQ02, SH96, Son05, WVS^{+}99, WVS^{+}00, XXM04, ZZ03]. coordinated [BRR^{+}00, NS07, RRT^{+}08, RZ97]. Coordinating [GNB^{+}09, GDRT13, MAHK16]. coordination [BACF08, FG91, KLS^{+}10, Pet76]. coprocessor [GPR87]. Copy [MMT16, HDG09, TML^{+}00]. Copying [Sal00]. CoRAL [VTGH17]. CORBA [YD96, YS98]. Core [MGT^{+}17, SHP^{+}16, CAW08, DD12, DMD13, FD10, GKS11, KF09, LCWM08, MDK96, RRBN09, ZLX01b]. cores [CWS06, GTSS11, NBB09, SMM^{+}09, VZ14]. CORFU [MBD^{+}12, DJS^{+}17]. Corporation [Had83, Woo85]. Correct [Hem89, TT00, Hem88, Kea88, SP00, Shi00, ZLX01b]. Correction [JHK^{+}16]. correctly [CGKM11, SAL20]. correctness [Ell77, Kah72]. correlation [YS94]. correlations [LPH^{+}07]. Corroboration [Fei15]. cosmos [AJG07]. Cost [AMH^{+}16, Coo78, Fes07, GNA^{+}98, KSK09, KBC94, PS09, SCP^{+}06, SW10]. cost-effective [GNA^{+}98, KSK09]. Costs [WL15, CG94, YV01]. Cosy [BBH96]. COTS [CJS^{+}09]. COTSon [AFF^{+}09]. council [Sop84]. Count [MCXS16]. Counter [KTG^{+}17, EEKS06, Ger77, SZD^{+}08]. Counterexamples [HV92]. counters [EBP16, Wei98]. coupled [KLS85, LWQ09, Pea89, PR83]. course [Atw84, BC01, CS00, HV08, Met82]. courses [AEG^{+}91, DS80, Han83, HP93, Tan87]. coverage [RRP06, TLD^{+}11]. covert [Loe85, Loe89]. Cowell [Heu97]. CPI [EEKS06]. CPME [RD87]. CPU [FS96, GGV96, JRR97, Kam13, NXQ05, YZG^{+}11, FRL00]. CPU- [NXQ05].

CPU-accelerator [YZG^{+}11]. CPUs [WDA^{+}08]. Crash [BKL^{+}16, FV06]. Crash-Consistency [BKL^{+}16]. crash-tolerant [FV06]. crashes [CNC^{+}96]. crawler [BPP12]. CRAY [Rei85]. Crazy [Tsa16]. Creating [ELR15, TZZ^{+}18, MHD^{+}07]. credential [CM06]. CRISP [DMB87]. Critical [ZSG^{+}17, ZE16, KS82, Nai93, PGZ08, Spi74, YS98]. criticism [Bir94]. critique [Nes90, WB07]. CRL [JKW95]. cross [DP93, EER12, PFGD02]. cross-domain [DP93, PFGD02]. cross-vendor [EER12]. Crossing [OHW17]. Crowd [DIS19, LOM^{+}09]. Crowdsourcing [PAM^{+}16]. Cryptanalysis [KC05, LKY04, LLH04, Sco04]. Cryptographic [EPG^{+}20, Gif81, Ng99, PS99d, Wai95b, JW01, SW00, XZZ97, XZZ98, ZLX99, ZWWL01, ZL04a]. cryptography [BMA00, JY98]. CS [AD07]. CSP [BR09, ZL86]. CSR [SHP^{+}16]. CTA [LSL^{+}17]. CTRON [Moo92]. CUBIC [HRX08]. cuckoo [SF12]. current [JT90, Van96]. curve [ZPS^{+}04]. CuSP [HDGP21]. custom [CJR87, MSC^{+}06]. Customer [PPO14, Bro75]. Customer-oriented [PPO14]. customers [PM03]. Customizable [HDGP21]. customization [ZSS08]. customizations [SFW99]. Cybersecurity [BH21, Ott18]. cycles [ABD^{+}97, WL82].

D [Wai94, ASR^{+}17, BDDMR11, GPY^{+}17, KDS^{+}06, LG04, MAS^{+}06]. D-SPTF [LG04]. DACIA [LP01]. Dagstuhl [SK13]. DAOS [LLSK24]. Data [CSBA17a, CKmWH16, Che17, HWO98, HLL^{+}02, Her92a, ID25, MBS16, NTC^{+}21, Owe84, PR15, PFK^{+}22, TZZ^{+}18, UJE^{+}22, WYD^{+}21, Wei85, Woo85, You92, YWKYS15, ZLJ16, ZJL17, Zha23, AVZR11, Als72, AAMV09, BFHW75, BC08, Buc77, CSBA17b, CKJA98, CMT94, CCM96, Col73, CJG02, Cos13, DVS12, DZ95, DBH^{+}06, Gan77, GKD91, GTA06, GBCH00, GSM08, Had83, HN08, Her87, HSS^{+}06, HHS05, Isa07, IBY^{+}07, KBB^{+}06, KB84, KGB88, yKPR02, KSLA08, KPR^{+}08, LM96, LJW^{+}06, Mad81a, Mad81b, McN77, McN82, McN88, MRS09, MMAS08, PGZ08, PHYO96, PK96, Pop75, RRT^{+}08, RKV11, RB93, Rei85, RJK^{+}14, RR72, RMS98, SCL96, Sal93, SBN^{+}97, SP00, Shi00, SF91, SDE85, SETB08, Svi83, Svo81b, SBH^{+}10, TSLBYF08, TPO06, TLL94, TGR^{+}21, TL96, Tug83, Tur80, VT01, VL87, VM07, VDGR96, WTB10]. data [WKT^{+}13, Wed88, WS91a, WSW05, WSH94, WTLS^{+}09, YVM13, YRC05, ZYG00, ZLL^{+}07, ZJS^{+}11, WS92]. Data-Aware [PFK^{+}22]. Data-dependent [Wei85]. data-flow [Rei85]. Data-Intensive [ID25, NTC^{+}21]. data-memory [SCL96]. Data-parallel [CKmWH16, IBY^{+}07]. data-race [PK96]. Database [SAG06, BJW87, CK86, DGH^{+}88, Ell77, EDP06, GWSY08, GKL95, LHWY83, PR83, RGAB98, Tra82]. Databases [LS09, EDZ07, yL91]. Datacenter [Bia17, LLLG16, XDM^{+}18, BCP^{+}08, SG10b]. Datacenters [BLJ^{+}17, KGGS18, CII^{+}10]. dataflow [AGP77, MSP^{+}06]. DataMesh [WCE^{+}92, Wil93]. Datapath [TSP17]. dataref [DL15]. DataSeries [AAMV09]. Dave [Had93]. DAWNBench [CKN^{+}19]. day [PSB06, PB08]. days [AD07]. DCatch [LLL^{+}17]. DCFS [XXM04]. DCG [EP94]. DCNN [RLD^{+}17]. de-facto [Rus08]. dead [BS02]. dead-instruction [BS02]. deadline [Mil90]. deadlines [SLCG89]. Deadlock [New79, Pea89, Ell73, Hol72, Lei89, Lev03a, Lev03b, Lev05, Zöb83]. Deadlocks [Dim98, Fon72, ZK88]. Dealing [SESS96]. debate [Bak95, Wai95b]. debug [FD10, KL02]. Debugger [CHLS16]. Debugging [GLD^{+}22, CL87, MM92, MM93]. December [Sat95]. decentralized [Che85, Cra83, KLS08, LM10, LG04, ML97, RF98, Sal78b]. deceptive [ID01]. Decidability [Mou96]. Deciding [SFH^{+}99, SFH^{+}00]. declarative [LCH^{+}05, Mao09, TPH12]. declarativity [Mao09]. declare [ACC^{+}09]. decoding [AS10]. Decomposing [JKS^{+}15]. decomposition [CL04b, MB93, SFS13]. decompression [VHAA25]. Deconstructing [DBP^{+}04]. Decoupling [LZC^{+}17, Por10, HCBS04, KGGK09, Tai13]. dedicated [TBM^{+}06]. Deduplication [SJS^{+}23, YN12]. Deep [GZH^{+}19, HABZ17, JCY^{+}19, LLSK24, MKL^{+}19, PWT^{+}19, RLD^{+}17]. Deeper [JV21]. Default [MGT^{+}17]. Default-On [MGT^{+}17]. defect [PJDL06, SCP^{+}06]. defending [TNL^{+}07]. Defensive [QPP02]. deferred [BHB^{+}08]. define [Den74a, Fou74, Zel74]. Defined [Ott18, BBM^{+}81, Gsc94, RB24]. Defining [Lev03a, Lev03b, SWL77]. definitions [FHL95]. degree [ZSS08]. DejaView [LBP^{+}07]. delay [KBK02, LCJV^{+}11, PS99a]. delay-tolerant [LCJV^{+}11]. delays [BL89, KS99]. delegated [CL04b]. delegation [CLC05, HK00]. delete [Gar07]. deleted [BC08]. delivery [SGD^{+}02]. Delta [PCD91]. Delta-4 [PCD91]. demand [CPM10, CSJZ08, FGBA96, HFC^{+}06, LGN07, PB08, Pot77, VM07, WMH72, YGG^{+}03, YZZZ06]. demands [BB75, CM75]. demons [Gur07]. DEMOS [BHM77, Pow77, PM83]. DEMOS/MP [PM83]. Dena [GSA10]. Denali [WSG02]. deniable [CCK04a]. Dennis [vR14]. Density [GSCM16, GPV04]. Dependability [CM13, CM14, BKP^{+}12, HAF^{+}07, SK13, VW08]. Dependable [CvR14, DBR09, Gan92, MC11, MDB01]. Dependence [RMS98, Bas72]. dependencies [FMK^{+}07, NPC06]. dependency [GLL04, LFWL10]. dependent [LE96, MS91b, PG03a, Wei85]. depending [Gon92]. deployment [CKK^{+}07]. DepSpace [BACF08]. Deputy [HH88]. deriving [RR04]. describing [SRH^{+}06]. description [WP87]. deserve [KMA^{+}14]. Design [AEE^{+}94, ACS15, AVN^{+}16, Bor92, BR10, DHK^{+}15, DMB87, DDM^{+}18, GF15, HF08, KGGS18, KRS97, KY02, LCL^{+}16, Maf94, Rus81, SGGB99, SGGB00, Val94, vR92, AMPS73a, AMPS73b, AMPS74, AWW08, AWSBL99, AWSBL00, AEH75, ALBL91, Atw84, BMvdV93, BVR^{+}00, BITW07, Dij05, FLR77, FP89, GRB^{+}08, GSSV00, GKS11, Had85, HKL^{+}06, HdRC95, IKWS92, ÏMC^{+}06, JOW^{+}02, KSCK17, KGGK09, Lam83, LH04, LRS^{+}08, LXYZ19, Lie93b, Lis72, MSAD91, MM81, MP91, NMS^{+}00, NBW87, NL97, OSV82, OST83, OSV86, OSSN02, PHOA89, PR06, PHL^{+}77, Pra87, RO91, SHN^{+}85, SNV10, SCS77, SHSB75, SMI80, ST01, TWL05, Toi92, WL02, WL94, WXX08, YN12, YAK93, ZRMH00, ZYG00]. designed [EVvdW89]. designer [HH89]. designers [FM98]. Designing [BPP12, RV91, SGNG00, GLL04, RBLP07]. designs [Moh78]. desk [HM91]. desktop [CJS^{+}09, FURM00]. Desperanto [MLB83]. detectable [PW98]. Detecting [JKDC05, LLL^{+}17, CWdO^{+}06, LHL04, LTQZ06, LPH^{+}07, WG08, ZXMJ04]. Detection [Bre83, LZH^{+}22, ZLJ16, ZJL17, AMA^{+}11, BM06, BS02, CG06, Dim98, FES09, FLM^{+}08, HLL^{+}02, HC04, Lei89, MZI08, MC91, New79, PBYH^{+}08, PK96, SGK^{+}04, YRC05]. detector [SBN^{+}97]. detectors [SS07]. Determination [PAM^{+}16, CC77]. Determining [CDY^{+}17, Won93]. determinism [Ste97]. Deterministic [LLLG16, PM03]. Developer [LJdL^{+}16, Pen09]. developers [SS17]. Developing [Had93, PP09, SP00, Shi00, SXZ^{+}88, OT95]. Development [LWQ09, Wai86, BvS00, DBR09, Her10, Lau81, Sal74, TBM^{+}06, WLP75]. deviant [ECH^{+}01]. Device [Hol88, SLLP^{+}10, ACD^{+}14, BBC^{+}06, CCG95, FFBG08, HF08, Hei78, KPG93, KHL^{+}07, KL02, MZWZ02, PLM06, PLHM08, Rya98, Rya99, SRH^{+}06, TF04, WS91a]. Devices [XD17, BTK11, DPW^{+}09, DZP^{+}11, KS09, MCdL06, Neg00, PSMB16, RA07, Rus08, Sch95, WDA^{+}08, XLDB09]. Devirtualizable [LSS04]. DFTS [WLZ03]. DGates [ASR^{+}17]. DGDBM [Fra95]. DGSA [FM98]. DHEKE [LSH01]. diagnosing [TLH^{+}07]. diagnosis [BDDMR11, PPO14, Wal73, WY04, YLW^{+}06]. dictionary [WB86]. Did [DK17]. didactical [AEG^{+}91]. difference [Fle07]. differencing [BPP12]. different [GLC99, LZJ03]. Differential [KBPM10]. Differentially [LSV^{+}19]. Differentiated [CEV00, MA11, GC08]. difficult [Nee72]. Digest [Sat99, Sat95]. Digital [BCC^{+}94, Had83, Woo85, BSR^{+}06a, CS08, CJ05, HCK08, MKY08, RV91, SCL96, Sal78a, Her92b, Jef92]. Dijkstra [Kos73]. dilemmas [ES10]. dimension [CPM10]. dimensional [BSSM08]. dining [Ran82]. Direct [RKV11, BLRC94, GLL04, HFWZ87, KS09, MSP98]. direct-access [KS09]. direct-mapped [BLRC94]. directed [BAM^{+}96, CHCmWH00, CJG02, Lei89, LLD^{+}04, MP85, Nai93, RP07]. directions [Fiu06, HSW^{+}00, PV95]. director [Fle07, KMK10]. directories [CKA91, Pon97, SD86]. directory [LEH86, SMBA10]. Dirigent [ZE16]. Disaggregated [CIP^{+}23, FZY^{+}23, Zha23]. Disaggregation [AC23, AAA^{+}23]. disambiguation [GCM^{+}94]. disaster [SESS96]. disasters [KBB^{+}06]. DISCCO [CM13]. discipline [Wir77]. disciplines [Ful73]. disco [GTHR00, BDR97, GTHR99]. Disconnected [KS91a, KS91b]. discovery [HLL^{+}02, KJH^{+}11, dGdB10]. Discrete [WKL07, GDRT13]. Disk [WHZ^{+}17, BC10, CS08, Ful73, GJXJ03a, HXL01, HHS05, ID01, Jin99, KS95, LK91, PKB^{+}08, SFV^{+}04, TP72, WLRZ03, ZCT^{+}05, dJKH93]. Disk-based [WHZ^{+}17]. diskless [CZ83]. disks [AUS98, BITW07, Gur07, HJ10, LT96, LLD^{+}04, hTMAC^{+}08]. Dispersing [VE08]. display [BKN05, SK96]. Dissertation [vR14]. distillation [FGBA96]. Distributed [BBBAN04, BFD97, Cec00, Cha90, CC21, CJRV15, FKZ17, Hac85, HJrCH16, HDGP21, JBW^{+}87, KvRvST92, LLLG16, LLL^{+}17, MAHK16, McD00, MV86, Mul87, Nai93, Nai96, PR15, Pow89, Rei92, RAVC12, SDD^{+}85, VTGH17, Wai83a, WS92, WTC09, WN80, Yan92, YG25, dV96, Aba93, AMS^{+}07, APGG00, AMMR92, AEP^{+}97, Bab91, BS95a, BDM97, Bac81, Bac91, BAI93, BO91, BHK^{+}91, BBH^{+}00, BFSG94, BMD94, BJ87, BLNS81, Bir91, Bla85, BDF^{+}08, Bor92, Bos06, Bou94, BL00, BGS04, Bre83, BHJ^{+}93, CW92, CS77, Cas91, CALM97, Cha96, CC97, CZ83, Che85, CK86, Coo85, CGS^{+}96b, CB95, Cri94, DHRS91, Dou93, DFS00, DCZ96, ENCH96, Esk96, ELG95, Fle81, FP89, Fra95, FdAM14, FR94, Gan92, GBZP10, GB90, GC89, GBCH00, GXJJ03, GA08, Gup01, GLL04, HKD07]. distributed [HK99, HPM93, HdRC95, HCZ97, HZCC97, HCZ98, HM90, HL92, HKM^{+}87, HL96, HSPC01, IBY^{+}07, JZZW02, JKW95, JH93, JLZx90, KKS89, KS85, KvS07, KSL90, KSLA08, KLS85, Kut84, Lac00, LLS91, LABW91, LB91, LT96, LCTK01, LAAW00, LMM93, LJX97b, LGN07, Lie93a, LHWY83, LP01, Lit88, LFWL10, LT11, LCH^{+}81, LB81, MK91, Mah98, MLB83, MO85, Mcd77, MM92, MM93, MS00, ML85, MSF85, MDB01, MM91, MT85, NSKS11, NB91, Nes82, Neu89, NB00, NCF05, OPSS93, Oli90, Ous81, PG96, PRD10, PWC^{+}81, Pra86, Pu93, Ray91, Ray92, RL96, RRCC10, SFV^{+}04, Sal93, SHN^{+}85, SJL^{+}87, SG97, SFL^{+}94, SBN83, SDP^{+}00, Sei90, STM^{+}07, SSS01, SF80, SMRD06, SGGB99, SGGB00, Son05, SS83b, SJ05, SSR^{+}10a, Str93, SMI80, SXZ^{+}88, Svo81b, TSF90, TM81, TCH^{+}91, TLL94, TML97]. distributed [TM89, Tur87, WC02, WPE^{+}83, WZZ93, WECK07, WAB^{+}89, WLZ03, WCL^{+}04, WPLP85, WLS^{+}02, Wil93, WBC^{+}83, WS06, XXM04, YM93, YD96, YbJf04, ZLL^{+}07, ZDP83, ZXMJ04, Zim94, dGdB10, vRvST88, vEBBV95, GB01, HTW01]. Distribution [CIL93, AEP^{+}97, Bas72, BC06, CKK^{+}07, LH04, LG04, PAB^{+}98, PS98, RZ97, SY96, Syv93, THB06, Wed88]. distributions [HBD95]. distrusting [BDT00]. Dive [JV21]. Diversity [SG14, MdS09, Pen09, Rom95]. division [CFR98, MPPZ87]. Diwali [JR05]. DJS200 [ZLX^{+}80]. DJS200/XT1 [ZLX^{+}80]. DMA [MMT16]. DNA [BLC^{+}16, Win08]. DNA-Based [BLC^{+}16]. Do [AZEE18, HSMC15, RPNT08, ACC^{+}09, BJ81]. Do-It-Yourself [AZEE18]. Docker [Boe15]. Doctoral [vR14]. document [CL04b, Oli90]. Documenting [PLHM08]. Does [BC10, CR12, KC94, SW10, ZUW^{+}09]. domain [CS77, DBR09, DP93, Jan75, PFGD02, San81, SHC73, WJMC04]. domain-specific [DBR09]. Domains [JKS^{+}15]. dominated [KBK02]. done [HUL06]. DoS-resistant [QPP02]. double [KT91b]. double-width [KT91b]. down [CS00]. downloading [XXMC05]. dragon [AM87, WAB^{+}89]. DRAM [KSCK17, OAE^{+}09]. DREAM [YG25]. Driven [JHK^{+}16, Har82, KTP^{+}96, NSKS11, ODH^{+}85, SZD^{+}08, SQP08, Tai13, TAH^{+}22, UNMS94]. driver [CCW^{+}11, FW72, TF04]. drivers [ACD^{+}14, BBC^{+}06, Hei78, KHL^{+}07, PLM06, PLHM08, Rya98, Rya99, SRH^{+}06]. Drives [RB24]. Drop [HT15]. drum [Ful73, Gre72, Sch73a]. Dryad [IBY^{+}07]. DSM [JTG^{+}00, LJX97a, SHT97]. DSMs [BKP^{+}96]. Dual [KKS^{+}16, CCW^{+}11, Mou96, SCL96, YW05]. dual-personality [CCW^{+}11]. Duality [FS08b, LN79, YTR^{+}87]. DudeTM [LZC^{+}17]. dues [Lev90].

DUNIX [Lit88]. duplicate [Ell77]. Durability [JKR13, EDP06]. Durable [LZC^{+}17]. DVFS [Kam13]. Dynamic [BS02, CKmWH16, DS73, GCM^{+}94, Jan75, MR07, MBS16, PB96, PPM17, WCS09, ZSG^{+}17, ZWL09, ZPS^{+}04, BJK^{+}06, BWV^{+}12, BL00, CC77, Cha96, CC97, CHCmWH00, EP94, Fes07, FS08a, FGBA96, FdAM14, GS13, HBD95, HHS05, JMK^{+}08, JXY95, LMV12, LPS10, MCD^{+}08, PL01, PLH98, PS96, PS01, RN93, SBN^{+}97, SAG06, Ste83, SLZD04, TDM12, TS06, Won93, XXMC05, ZXMJ04]. dynamically [BLRC94, OMCB07]. dynamics [ACT94]. dynamite [IvdLH^{+}00]. Dynamo [DHJ^{+}07]. DySel [CKmWH16].

Early [GMS77, JOW^{+}02, Led97, WPC12]. Easy [Gai78, CMN02, LFH^{+}09]. Economic [Sib76]. economies [HCK08]. economy [TLL03]. eCos [LST^{+}06]. ECOSystem [ZELV02]. Eden [Bla85, LLA^{+}81]. Edge [HDGP21, KHG^{+}17, CCB^{+}06, DSBK04]. Edited [Had85]. edition [Gue87, Had83]. Editor [Wai83a, Hof07]. eDonkey [HKL^{+}06]. Educational [Had83, Woo85, AMO^{+}12, NB00]. Effect [Mas77, DV87, HSPC01, Lov77, MB91]. Effective [KKN00, SLD15, ABLL91, BFS89, CH81, GNA^{+}98, KSK09, Sto07]. effectiveness [TE94, WPP02]. Effects [BS96, IKWS92, LJS^{+}02]. Efficiency [AT10, BLI17, BSR^{+}15, Bia17, LB08, SJS^{+}23, WM16, ACM02, BRW89, BJL^{+}06, LK10, Opd75, YVM13]. Efficient [AKP^{+}25, BM91, BKGT25, BEL^{+}00, CB17, DK16, FL77, FES09, GPY^{+}17, Kan83, LJdL^{+}16, LSH01, ML85, OR87, PPM17, PKB^{+}08, SH96, SZII11, WLAG93, WL82, WSW05, WLZJ17, WB86, YG25, ZLJ16, AAMV09, AC06, AD99, AD00, BJW87, CC04, CC05, cCVP99, CVP00, DY10, Edi13, EP94, GC89, GN96, HS88, HSI^{+}01, JRR97, JOW^{+}02, JXG^{+}02, KTH89, KTB12, KC95, KLS08, KDS^{+}06, LB06, LLH04, Lie94a, MC91, MRA87, NAR08, OS80, PSG06, PL01, PCH^{+}14, RD12, SG97, SP00, Shi00, SQP08, TDM12, TL94, VL87, VAK^{+}11, VGBT14, WC02, WK05, YGG^{+}03, YW06, YRC05]. Efficiently [ÏMC^{+}06, KDL^{+}16, KJS^{+}06, Spe81, ZZP04]. Eighteen [MFBWW20]. Eighth [Bac99]. EINSTEIN [FW72]. elastic [TPH12]. electrical [RJK^{+}14]. electronic [AC97, LWQ09]. Elephant [SFH^{+}99, SFH^{+}00]. elimination [BS02, KKN00]. Elsevier [Lit87, San86]. email [CXMX05]. embedded [CJR87, LBvH06, LF13, MA06, PS01, RR04, TKP^{+}08, WPC12]. Embedding [HK99]. embracing [Les04]. emerald [JLHB87]. EMERALDS [ZPS99, ZPS00]. Emergent [Mog06]. emerging [Est02, GWSY08]. EMPIRIC [BFZ^{+}25]. Empirical [CJM15, Des10, KLMO91, CYC^{+}01, MCM07, RF98, Rob98, SS98]. employing [CWS06]. Empowerment [Bla95]. emulation [HCG^{+}06, HFC^{+}06, Kam13, LAAW00, NMS^{+}00, PB08]. emulator [PSB06, VYW^{+}02]. Emulators [OB86, LFH^{+}09]. Enable [XD17, KDS^{+}06]. enabled [DW07a, DW07b]. enabler [DPW^{+}09]. Enabling [ALM^{+}18, GLD^{+}22, KDP02, KMK10, MCGL17, SATG^{+}07, WLZJ17, DKC^{+}02, FIM^{+}11, LSS04]. encoding [BM06]. Encrypted [JSCM17, LSH00, STW95]. encryption [CS08, Gai78, LK01]. End [JBDP08, VMM20, Zha23, BMK06, CCC^{+}05, ESB^{+}06, GNB^{+}09, GKS11, RN83, SS17, TBM^{+}06, TNL^{+}07, WSW05]. End-to-End [VMM20, Zha23, JBDP08, CCC^{+}05, GKS11, TNL^{+}07]. end-users [SS17]. Energy [ASR^{+}17, BSR^{+}15, CCHV11, CDY^{+}17, CHLS16, FS99, FS00, GBG^{+}10, JOW^{+}02, LJdL^{+}16, OBSR16, TDM12, AVZR11, ACM02, CAT^{+}01, CII^{+}10, Edi13, HD12, HEKSP11, HHS05, KDS^{+}06, KHL^{+}07, KAI^{+}13, LK10, LLD^{+}04, NCL12, NRS13, dOL12, RP07, SHA02, VW08, WBB02, YW06, YVM13, ZELV02]. Energy-aware [CCHV11, FS99, FS00, HEKSP11, KAI^{+}13, WBB02]. Energy-efficient [JOW^{+}02, Edi13]. Energy-harvesting [CHLS16]. Energy-interference-free [CHLS16]. enforcement [Buc77, JL75, KLS08]. Enforcing [AYK08, AC06, ZE16, FS08a, SLS^{+}05]. Engineered [ACS15]. Engineering [LGMF14, PWT^{+}19, Sch75, NN75, Ano75, BOB15, BM17, FBL^{+}12, OFB16]. Englewood [Sta83, Wai83b]. Englewood-Cliffs [Sta83, Wai83b]. enhance [SG05]. enhanced [RS08]. Enhancement [CJ05, LSH03a, LSH03b, YW04]. enhancements [HPG00]. Enhancing [ATMZ01, ATSS09, OL02, DY01]. enough [CCH^{+}87, PBR^{+}08, Pio89]. enterprise [FES09, JS08, KSDC14, NS07, SFV^{+}04]. entirely [OAE^{+}09]. entries [Nai93]. entry [Gai78]. Environment [VJ19, ABC^{+}02, BAMM77, BL75, Bro75, CJS^{+}09, CWL05, CCLP81, CLDA07, FW72, HK99, HCZ98, HC95, IvdLH^{+}00, JFV^{+}96, Jan75, JH93, JADAD06, KS92, LCTK01, MPF^{+}06, Nic87, PG96, PR83, RMSB01, RD87, SATG^{+}07, ST93, Taf82, Van06, VFMM08, WLS^{+}02, WBC^{+}83, YWC04, Yan92]. Environments [KEF^{+}19, BWV^{+}12, BDK^{+}08, DFS00, Hog88, KF09, LSA^{+}00a, LSA^{+}00b, Mcd77, OSSN02, PSZ^{+}07, YbJf04]. eNVy [WZ94]. EOS [PBM22]. EPEP [EVvdW89]. Epidemic [DGH^{+}88, ÖGA06]. Epidemic-based [ÖGA06]. Episode [You92]. EPOS [WWGF08]. ePOST [MPHD06]. equilibrium [CCAP06]. Equipment [Had83, Woo85]. equivalence [Lak85]. Eraser [SBN^{+}97]. Erasure [ID25]. EROS [SSF99, SSF00]. Errata [Ano78]. erroneous [Lev05]. Error [RB75, Boc75, LHL04, SLCG89, SGK^{+}04]. Errors [BS15, CYC^{+}01, ECH^{+}01, LRS^{+}08, RK11]. Essential [Heu97]. Esterel [LBvH06]. Estimation [OBSR16]. ESX [Wal02]. Ethernet [Gup01]. Etherphone [TS87a]. EU [BKP^{+}12]. EU-funded [BKP^{+}12]. Euclid [Hol82, PHL^{+}77]. Eudaemon [PB08]. Europe [Ens75]. European [Bac99, Mul87, Sha95, Tan97, Bab91]. EuroTM [CR12]. evaluate [EWCS96, Kai75]. Evaluating [BVR^{+}00, BJL^{+}06, JXG21, GSA10, MCC^{+}06, PHOA89]. Evaluation [CJM15, GF15, GLC99, KSS^{+}96, LLSK24, RLB08, SEF^{+}16, VMBM12, AUS98, Des10, FdAM14, Gan77, GGH91, GLG93, HLR98, KPL99, KY02, Nee77, NL97, PRAH96, PS99a, PSK08, Ros78, RN00a, RN00b, ST01, TNNI87, VW08, Zea97, ZHK06, ZIL96]. evaluator [SP00, Shi00]. Event [HNK^{+}17, EKV^{+}05, YLW^{+}06]. eventcounts [RK77]. events [KJ08, PRD10, Svi83, Tug83]. eventually [BCRS10]. Eviction [NTHAB22]. Evil [HCJ07]. evolution [AGSS10, Bro00a, Bro00b, Kor06, PLM06, Pat02a, Pow89]. evolution-some [Pow89]. Evolutionary [MM91]. evolutions [PLHM08]. evolvable [AIKS00]. Evolving [SADAD02, SZN87]. examination [HN08]. example [GC05, Hof90, Smo95, Woo73]. Exception [Mac77, MSR77, TL94]. Exchange [SHSB75, DS80, LL04, LW04, LSH00, SS00, STW95, WSW05]. exclusion [BBBAN04, Bou94, Cha96, CC97, HS88, Har82, Hof90, Nai96, Ray91, Woo90]. exclusiveness [Lie94b]. execute [BD91]. executing [ACT94]. Execution [JCY^{+}19, KKS^{+}16, AYK08, BDK^{+}08, CCG95, CR75, CG00, CLDA07, DSBK04, ELG95, HFWZ87, HEKSP11, KY02, KCLZ98, Le98, MPP^{+}08b, MPP^{+}08a, MCC^{+}06, NBB09, NCF05, PS96, PG03b, RG02, RF98, SLS^{+}05, SLZD04, SQP08, TLC85, VESM10, WKL07]. executive [HP93, Sop84]. exercise [BLNS81, LE96]. exercisers [Pay77]. Existing [CCS^{+}16]. exokernel [KEG^{+}97, EKO95a, Les04]. expanded [Lor86]. Experience [Coo94, Oes01, SW91, SBN83, BC91b, Bla85, ETKF07, GMS77, LBB^{+}91, WP87, ZSK97]. Experiences [AMMR92, AMO^{+}12, GHP^{+}08, MPHD06, NV06, CF89, JOW^{+}02, KJH^{+}11, KSL90]. Experiencing [AEG^{+}91]. experiment [Che84, EVvdW89, Led97, Ric88]. experiment-control [EVvdW89]. Experimental [ACS15, Eid15, RR72, Gan77, GPR87, Hop90, Lov77, SHC73, WH99, WCW^{+}04, WLS^{+}02]. experimentation [LFH^{+}09]. Experiments [AHB15, ELR15, SM89]. experts [Owe84]. explicit [BMR^{+}09, MP96]. explicit-rate [BMR^{+}09]. explicitly [MT02]. exploitation [PSG06]. Exploiting [BSL08, BJ87, EM89, EAS^{+}17, GHW07, GTA06, HBD95, KKB^{+}16, MES95, SCL96, Ste97, SKZ07, AYK08, FC87, HEKSP11, KKM^{+}06, LLS91, SFW99, WV02, WECK07, WTLS^{+}09]. exploits [PB08]. Exploration [JV21]. Exploratory [dSM16]. Exploring [BFZ^{+}25, CL95, CGJ^{+}07, LPM17, TZZ^{+}18, WCL17, BMvdV93, ÏMC^{+}06]. exposed [GTK^{+}02, TACT08]. expressing [Pay77]. ext3 [AR07]. Extended [CM14, Fab73, Gue88, KTB12, MT17, Bor98, CV93, CG85, CMMS77, CM13, ECS73, FC87, KLS85, LLS91, LGJS91, Mcd77, RK77, Van96]. Extending [BF08, KPG93, Var97, MSA^{+}00, Spr85]. Extensibility [BSP^{+}95, EM06]. Extensible [Als72, BHLM94, KN93, TSP17, WBDF97, BCE^{+}95, CL95, OPSS93, PB96]. extension [CCW^{+}11, CBC^{+}08, Jan81, STW95, WS91b]. extension-oriented [CBC^{+}08]. extensions [cCVP99, CVP00, GUB^{+}08, GHM77, NL96, SESS96]. External [HC92]. externally [Wol02]. extracting [PKM^{+}09]. Extreme [XDM^{+}18]. EZIOTracer [NTC^{+}21].

F [Woo85]. FAB [SFV^{+}04]. face [JBDP08]. facilities [Coo78, Nut94b, TLC85]. facility [DP93, GP95, LAAW00, VL87, SHSB75]. facto [Rus08]. Factored [WA09, BDDMR11]. Fahrrad [PKB^{+}08]. fail [BHB^{+}08]. fail-in-place [BHB^{+}08]. failed [Jin99]. fails [DH96]. Failure [IKK16, SKB^{+}17, dSM16, SS07, YW05]. Failure-Atomic [IKK16, SKB^{+}17]. failures [PBYH^{+}08, QTSZ05, TLH^{+}07]. Fair [AHB15, BMR^{+}09, CAW08, Dun91]. fair-share [CAW08]. Fairness [HS91, SFS13, SKJ^{+}17, WM16, WTKW08]. false [HSPC01]. Fame [Mog08]. family [ABC^{+}98]. fan [EKF^{+}14]. fan-less [EKF^{+}14]. fans [VZ14]. farming [Hal00a]. Farsite [ABC^{+}02, BDH07]. Fast [ACAAT16, BDF^{+}15, CC21, HABZ17, SRTH15, SL98, SB91, SMM^{+}09, VGX17, BXS14, BPP12, BMA00, CKD94, Che84, FAH^{+}06, GLL04, Riz97, SSF99, SSF00, TNL^{+}07, dBB08, Heu97, Ste83]. FastAD [SMBA10]. Faster [MMT16]. fastest [vRvST88]. Fault [AEMGG^{+}05, Bab90, Cri91, KT91a, LER^{+}17, PCD91, Rom93, Sal91, WQA^{+}24, AAC^{+}05, Bab91, BJM^{+}91, BRR^{+}00, BACF08, Bir85, BC91b, Bir91, BBG83, BS95b, CC97, CRD^{+}95, DHRS91, GG91, GC89, HGR07, JT90, Kan83, KS91b, KAD^{+}07, MS91a, NB91, PL95, PNT06, RRP06, RCL01, Sad75, SNV10, SPR00, TCH^{+}91, WLAG93, WY04, WLZ03, XXM04, ZL86, ZHK06]. fault-intolerant [ZL86]. Fault-scalable [AEMGG^{+}05]. Fault-Tolerance [Cri91, PCD91, Sal91, Bir85, KS91b, WLZ03, XXM04]. Fault-Tolerant [WQA^{+}24, Bab90, AEMGG^{+}05, BACF08, CC97, DHRS91, GC89, HGR07, MS91a, SNV10]. faults [LB91, Nee72, SLM11, VBLM07, WCS08]. faulty [Bou94, YW05]. FAWN [VAK^{+}11]. FBR [RD97]. Fbufs [DP93]. feasibility [GBG^{+}10]. feature [Had85, LJW^{+}06, OST83]. feature-rich [LJW^{+}06]. features [AEE^{+}94, Als72, AM77, Fos87, HO91]. FeBID [BR10]. federated [ABC^{+}02, EER12, SK13]. Feedback [BR10, LSA^{+}00a, LSA^{+}00b, SQP08]. Feedback-driven [SQP08]. Felix [FO81]. Ferret [LJW^{+}06]. Festival [JR05]. fetch [FG91, OKN02]. fetch-and-increment [FG91]. few [FR94]. Ficus [GHP92]. fidelity [VMC^{+}05]. field [HDL^{+}02]. field-based [HDL^{+}02]. Fieldbus [RN93]. Fifth [EMS09, Pet76]. fighting [WGL^{+}08]. File [AHC^{+}16, BKL^{+}16, Dio80, FO81, GN80, GHP92, KMA^{+}14, LK01, NTHAB22, SRTH15, THB06, Vog99, Vog00, ADN^{+}95, AC06, AR07, BO91, BHK^{+}91, BC06, BPP12, BDT00, Bor98, Bor92, BS89, CPdM^{+}96, CNC^{+}96, COS^{+}08, DZP^{+}11, DB97, EM89, EBP16, EER12, Fab98, FFBG08, FC87, FES09, FW77, FMK^{+}07, GC08, GS90, GJSO91, GC89, GNB^{+}09, Gsc94, GPK^{+}07, Hac85, Hag87, HKL^{+}06, HdRC95, HO93, HSK97, HKM^{+}87, JHT^{+}07, JXG^{+}02, KN93, KS91a, KAS^{+}06, KN96, LJX97b, LGG^{+}91, LGJS91, Maf94, Mah94, MRC^{+}97, MKKW99, MKKW00, MD81, MT85, MP91, MES95, MCM01, MMGC02, Nee79, NWO87, NCF05, ODH^{+}85, OD89, Pio89, PKM^{+}09, Pow77, PBA^{+}05, RV91, RL96, RK83, RN83, Rob96, RAF07, RO91, SKKM02, SFH^{+}99, SFH^{+}00, Sat81, SHN^{+}85, SGN85, Smo95, SG05]. file [SK97, Ste97, SMI80, SSR^{+}10b, TML97, TGR^{+}21, Vag10, WH08, WS91a, WF07, XFO08, XHJB99, ZN00, ZG07, dJKH93]. FileNet [EM89]. Files [VJ19, BBM^{+}81, Bre83, DB85, EJD13, FC87, Fra95, Sch95, SBB86]. fileserver [BFD97]. Filesystem [HR92, CG91, You92]. filesystems [AEH^{+}08]. Filet [DBR09]. Filet-o-fish [DBR09]. filing [BN78a, HP95, NB77, PKW81, RN83]. filter [MRA87]. filtering [EDZ07]. Financial [And87, Ano86, Kah85, Lev88, ZFP^{+}21]. find [BTK11]. Finding [HABZ17, MCXS16, SLM11, SW00]. Fine [CSS^{+}91, EGE02, ETKF07, JXQ^{+}22, JLHB87, LC93, SFL^{+}94, EK08, Lie95b, Lie96, SGT96]. Fine-grain [CSS^{+}91, SFL^{+}94, SGT96]. Fine-grained [EGE02, JXQ^{+}22, JLHB87, EK08, Lie96]. fingerprint [CL04c]. fingerprint-based [CL04c]. fingerprinting [PSB06, SGK^{+}04]. fingerprints [KLY03, Sco04]. finishing [ECS73]. finishing-time [ECS73]. finite [Mou96, Pea89, SLTB^{+}06]. Fireflies [JAvR06]. Firefly [SB89, TS87b]. firewall [LJY04]. firmware [KL02]. First [Fêa83, MSLM91, SRS22, Had84, MRS09, SS17, ZELV02, vR14, BK08, Bre08, Ter14]. First-class [MSLM91, SS17]. fish [DBR09]. FIST [ZN00]. Fits [UJE^{+}22, Ste83]. five [Svi83]. five-step [Svi83]. fixed [LLK96]. fixed-priority [LLK96]. Fixing [MY98]. FLash [DK15, KLK17, BITW07, Des10, LMG^{+}07, LSKK08, MBD^{+}12, CCEH00, GKO^{+}00, HGDG94, HKO^{+}94]. flash-based [LMG^{+}07]. flash-speed [MBD^{+}12]. Flawed [Dru92]. FLEP [WLZJ17]. FlexDCP [MCR^{+}09]. Flexibility [KN96, HKO^{+}94, KEG^{+}97, Les04]. Flexible [PST^{+}97, WLZJ17, AAMV09, BEL^{+}00, HPM93, HvE02, KC05, LHY02, LWMX05, MK91, McD00, Rip03, SCM05]. FlexNIC [KPS^{+}16a]. Flicker [MPP^{+}08b]. FLIP [KvRvST92]. Flipstone [BHB^{+}08]. floating [LKB91]. floating-point [LKB91]. Flow [FXZ^{+}17, YSCC16, BDF^{+}08, EK08, GA08, KYB^{+}07, LJS^{+}02, ML97, MMAS08, RB93, Rei85, SRA^{+}04, SLZD04, VBHN10]. flowchart [Fra80]. Flows [GCJ17, RKV11]. Flux [FBB^{+}97]. fly [CWS06, Jin99, Kep91, SZD^{+}08]. folding [Sch73a]. forced [KS99]. forensic [MZI08]. forensically [ME08]. Forensics [HN08, CS08, HH08, MKY08, MMN08, PBM08, SMRD06, XFO08]. forensis} [PBM08]. Foreword [Eid15]. forget [SFH^{+}99, SFH^{+}00]. Formal [BBFH07, BH75, MCN^{+}17, PG73, WJMC04, BGHL87, DMD13, JW01, LF13, LZ03, TFC99, ZLX99, ZL04a]. formalised [Pay77]. formalism [Lei89]. Formalisms [Cer75]. formalization [BAD^{+}11, HZCC97]. format [AAMV09]. formats [ZFW10]. formulas [FR94]. Forth [HFWZ87]. forward [Mat06, GB01]. fos [WA09]. foundation [BYV08]. Foundations [HMS17]. fountain [WDA^{+}08]. four [JLR^{+}05]. fourth [Bab91, SN13]. FPGA [SMS11]. fragmentation [RS86]. Framework [BMF^{+}16, CC21, LZH^{+}22, TAH^{+}22, WYD^{+}21, AIKS00, Bor92, EJD13, FFM07, GW04, GGL^{+}09, ID01, KKM^{+}06, KEP07, LP01, MCR^{+}09, ROJS09, RD12, SS00, Smo95, TBM^{+}06, WXX08]. framing [Hal00b]. Frangipani [TML97]. Frank [Bla95]. fraud [PW98]. fraud-detectable [PW98]. Free [GMT16, LC97, CHLS16, HHS05, KPS09, Lon93, RG02, YWKYS15, MP92a, MP92b]. freezing [WJ98]. Freon [HCG^{+}06]. frequency [CPM10, GS13, KTB12, Kam13, MCD^{+}08, Sad75, WJMC04, ZXMJ04].

Frequent [ZYG00]. fresh [BJK^{+}06, BSR^{+}06a, EKF^{+}14]. fresh-air [EKF^{+}14]. freshness [KMSV10, LC04b]. Friendly [LJdL^{+}16, WYW^{+}25, HRX08, HL05]. front [RN83]. FS [HdRC95]. FS2 [HHS05]. FTP [Gsc94]. Fuel [VZ14, RJK^{+}14]. Full [HSL17, AFF^{+}09, CS08, MMAS08]. fully [FPG89, WYC03a]. Function [HSL17, APGG00, CHY05, Eri14, FS95, LW04]. functional [Bos06, Fra80, MHD^{+}07, RKBH11, Sat81, WZWS08]. functionalities [CJS^{+}09]. functionality [CD95b]. functions [DK75, HSK97, LLH02, LKY04, Mil77, YRY04]. funded [BKP^{+}12]. funding [GNB^{+}09]. fungible [Lev03b]. funnel [LMV12]. Further [Hsi89, TT00]. futex [BF08]. future [Bas12, BCC^{+}13, Fiu06, Fle83, JT90, KG99, Lam00, Mit96, Svi83, Tug83]. fuzzy [FLM^{+}08].

G [Had85]. Galactica [LaR92]. Game [FZL16]. games [CCAP06]. Gap [GSW^{+}17, PVB17, BYVF08, Cos13, PG06]. gaps [Gwi94]. Garbage [Bar79, BN78a, CHV04, GN80, GTSS11, JHT^{+}07, KPS09, ONG93, SN94, SMTZ09]. GB [VHAA25]. GB/s [VHAA25]. Geiger [JADAD06]. Gene [AUW08]. Gene/P [AUW08]. General [CCS^{+}16, Hem89, TT00, BAMM77, DC99, DC00, ECH^{+}01, FIM^{+}11, FS95, GCTR08, Hem88, Hsi89, Kea88, LSS04, MQW95, TPO06, WH99]. general-purpose [DC99, DC00, GCTR08, TPO06]. Generalized [FMK^{+}07, CC77, KS82]. Generating [PKB^{+}16, HZCC97]. Generation [AYQ^{+}16, BH21, AKS73, BA06, BD91, BW95, CG00, EP94, HCZ97, PG73, PSB06, War76]. generational [WK08]. generators [SWL77]. Generic [ARS89, Hil81, AUW08, FFM07]. genetic [ELG95]. Geo [BDF^{+}15]. Geo-replicated [BDF^{+}15]. geocast [WS06]. GeoGraph [WYD^{+}21]. geographic [EHD07]. Geology [DDOL16]. Geometric [WYD^{+}21]. GET [HDH^{+}94]. Git [SLD15]. GLADE [RD12]. Glitz [EER12]. Global [Had93, San86, AUW08, FMP^{+}95, JFV^{+}96, KSS^{+}96, KBC^{+}00, OA08, SHA02, Tur87, YM93]. global-scale [AUW08, KBC^{+}00]. globally [Oes91]. globally-ordered [Oes91]. GM [BEW75, BEW76]. GNU [WB07]. go [KC94, MPP^{+}08a]. goal [WL09]. goal-oriented [WL09]. goals [AMPS73a, AMPS73b, AMPS74]. Going [Bak95]. gone [ABD^{+}97]. good [CM06, HYM10]. Goodput [RHR^{+}17]. Google [CSBA17c]. gossip [ADG^{+}07, Bir07, CGJ^{+}07, EFL07, FFM07, PB09, RBLP07, VBHN10]. gossip-based [ADG^{+}07, CGJ^{+}07, EFL07, FFM07, RBLP07]. Gossiping [FGR^{+}07, KvS07, WMI^{+}07, BBFH07, GHW07]. government [GNB^{+}09, Klo80]. GPU [DS09, FZY^{+}23, VHAA25]. GPUs [LSL^{+}17, LCCZ17, PPM17, SBS18, TPO06, WLZJ17]. Grabowski [Wai97b]. graduate [Met82]. grain [CSS^{+}91, SGT96, SFL^{+}94]. grained [Dub00, EK08, EGE02, ETKF07, GTA06, JXQ^{+}22, JLHB87, Lie96, LC93]. Grant [Bis81]. granularity [Lie95b, MS94]. Grapevine [BLNS81, SBN83]. Graph [BH21, CC21, HDGP21, JCY^{+}19, JXG21, Ser21, VTGH17, WHZ^{+}17, WYD^{+}21, ZFP^{+}21, Lei89, RB93]. graphic [WYC03b]. graphics [CCW^{+}11, Gor87, LHPL87]. Graphs [VGX17, KKFB11, Nai93, SK96]. GraphZero [MRH^{+}21]. Graspan [WHZ^{+}17]. grass [MMTW10]. gray [ADAD01, Gra14]. gray-box [ADAD01]. green [AVZR11]. greener [JS08, MMTW10]. GreenFS [JS08]. Gregory [Wai94]. grid [DW07a, TLL03, BJKT15, DW07b, KLS08, YGG^{+}03]. Grid-wide [KLS08]. GRIFFIN [GCJ17]. Group [BDM97, LJX97a, Rei92, BS95a, CL04a, CNL89, Hag87, HL92, KT91a, LLH04, Oes01, Rom97, SF12]. groups [PL01, PL95]. growth [Bro00a, Bro00b, SBN83, Svi83]. growth/evolution [Bro00a, Bro00b]. GRPC [WZZ93]. GS320 [GSSV00]. GSST [VHAA25]. guarantee [RS00, WlL01, WL02]. guaranteed [PKB^{+}08]. Guarantees [VMM20, BC06, GP95, LSA^{+}00a, LSA^{+}00b, PK96]. Guard [OHW17]. Guarded [LE96]. Guarding [GCJ17]. guessing [DH95, YS02]. guests [DY10]. Guide [Wai97c, Bru86]. Guided [UJE^{+}22]. GWiQ [KLS08]. GWiQ-P [KLS08]. Gypsy [AGB^{+}77].

H [Had85]. H. [Küh99]. H.264 [AS10]. HAC [CALM97]. Hadoop [LK10, Por10]. Haifa [BY08]. Haiti [LWQ09]. Hall [Mog08, Sta83, Wai83b]. Hamlyn [BJM^{+}96]. Handbook [Wai86, NN75]. handed [Ran82]. Handhelds [Sub11]. Handling [JH93, Nee72, HC95, Mac77, MSR77, RLB08, SMTZ09, TL94]. handoff [yKR06]. handoffs [JHC^{+}11]. Hang [WGL^{+}08]. happened [Her07]. hard [LTCA89, LRS^{+}08, RK11, YS98]. Hardbound [DBMZ08]. Hardening [BS15]. Hardware [AVN^{+}16, CKD94, CHLS16, CHCmWH00, FXZ^{+}17, KSCK17, LSMB16, MSP98, PKB^{+}16, SZD^{+}08, TL94, TML^{+}17, Wir87, YVCB18, ZH16, ZLJ16, AA06, ATSS09, BC91a, CL87, CWS06, CHV04, CSS^{+}91, GP05, Har82, JBDP08, KKN00, KKM^{+}06, MQW95, MPP^{+}08a, MFGSP12, MB80, NMS^{+}00, NPCF08, RHP^{+}07, SHA02, SN94, SS72, SH87, TE94, TACT08, Wil80]. hardware-assist [KKM^{+}06]. hardware-assisted [SN94]. hardware-driven [Har82]. Hardware-OS [LSMB16]. Hardware-Software [CHLS16, KSCK17, MSP98]. hardware-supported [MPP^{+}08a]. Hardware/Operating [AVN^{+}16]. Hardware/Operating-System [AVN^{+}16]. harmful [And09, Hof07]. Harmony [PPS^{+}18]. Harnessing [BSR^{+}15, RRCC10]. Harold [Wai83a]. Harp [LGG^{+}91]. HARTOS [KKS89]. harvesting [CHLS16]. Hash [DHK^{+}15, KCL03, Ku04, KCC05, LLH02, LKY04, LW04, TMW10, YRY04]. hash-based [KCL03, Ku04, KCC05]. Hashed [VL87]. HASS [SFB^{+}09]. Hawk [Har88, HH89]. HCCM [GJXJ03b]. HCloud [DK16]. HDDs [Str12]. heap [CG06, KJS^{+}06, LLS^{+}08, ONG93, SZ98]. heap-based [CG06]. HeapMD [CG06]. Heat [GPV04]. Heat-and-run [GPV04]. Heidelberg [WH94]. help [CGKM11, Kot88]. Helper [WCW^{+}04]. helpful [MPLH06]. helping [BTK11, ZCT^{+}05]. Helsinki [MY98]. heterogeneity [GHP^{+}08, RKBH11, Tur87, WCS09]. Heterogeneous [AVN^{+}16, BLJ^{+}17, BSR^{+}15, KGGS18, LPM17, LJdL^{+}16, LL16, TZZ^{+}18, VMM20, VSST16, AEE^{+}94, AJG07, BF87, DW07a, Gir82, GKS11, LCWM08, Pra86, SZN87, SFB^{+}09, SZII11, SJ95, SWC08, SXZ^{+}88, YZG^{+}11]. Heterogeneous-ISA [BLJ^{+}17, BSR^{+}15, VSST16]. HeteroOS [KGGS18]. HetNOS [BFSG94]. hFS [ZG07]. Hibernator [ZCT^{+}05]. hidden [CWdO^{+}06]. HIDE [ZZP04]. Hiding [BKP^{+}96]. Hierarchical [Bis81, DSGP05, RS00, CJR87, Dub00, EB78, Ger72, GGV96, VL87, Var72, WGSS95, YW06]. Hierarchy [KTG^{+}17, BHLM94, MSP98, Smo95]. High [AGM93, AHC^{+}16, DM90, EPG^{+}20, GSCM16, JKH^{+}00, KPS^{+}16a, KPS^{+}16b, MRH^{+}21, SF91, Val94, AEE^{+}94, ACG86, BM91, BVR^{+}00, BSR06b, BITW07, BMR^{+}09, BMK06, CPW07, Cri94, DD12, DP93, EDP06, ESB^{+}06, Fab98, FJLC98, GNA^{+}98, GNB^{+}09, GJXJ03a, HRX08, HdRC95, HXL01, JKW95, yL91, LKvR^{+}99, LKvR^{+}00, MB93, MW75, MUKX06, MP91, NSS10, OAE^{+}09, PG96, PWC^{+}81, PN00, RRP06, RAF07, SB10a, SPF^{+}07, SQP08, TBM^{+}06, UHMB94, WLZ03, YZJ02, YW06, YD02, ZSS08]. High-Assurance [AHC^{+}16]. High-bandwidth [SF91, BSR06b, DP93, GNA^{+}98]. high-coverage [RRP06]. High-Density [GSCM16]. high-end [ESB^{+}06, TBM^{+}06]. High-Level [EPG^{+}20, BM91]. High-Performance [KPS^{+}16b, MRH^{+}21, JKH^{+}00, BITW07, BMR^{+}09, CPW07, EDP06, JKW95, LKvR^{+}99, LKvR^{+}00, MB93, MUKX06, NSS10, OAE^{+}09, PN00, SPF^{+}07, SQP08]. High-Speed [Val94, BVR^{+}00, HRX08]. high-throughput [DD12]. highlights [AD07]. Highly [HBG^{+}06, Her92a, RLD^{+}17, BBH96, DHJ^{+}07, KGGK09, LAAW00, NLO95, SBL99, SBL00, WL09, ZLL^{+}07]. Highly-Scalable [RLD^{+}17]. hings [Zho16]. Hints [Lam83, CG00, SH96]. HIP [DTR01]. HipG [KKFB11]. HIPStR [VSST16]. History [SKJ^{+}17, CZG^{+}05, Fiu06]. History-Based [SKJ^{+}17]. Hive [CRD^{+}95]. Hoard [BMBW00]. hoarding [KP97]. hoc [BBD^{+}02, BBBAN04, EM06, MFHH02, ÖGA06]. Hoffman [Wai95b]. HOIST [RR04]. Holistic [MAHK16, VFMM08, NBW87]. Holland [Had84, Had85]. Holliday [Wai86]. home [ZIL96]. home-based [ZIL96]. homogeneous [MP75, Pra86]. Honeyfarm [VMC^{+}05]. honeypots [PSB06]. Host [OHW17, OCLN14, TDM12]. Host-Accelerator [OHW17]. Hosted [BHD19, DS09]. hosting [CAT^{+}01, USR02]. hosts [DY10, MMN08, ZZNM01]. Hot [CvR14, DNT10, HN12, Sat99, SN13, DB00b]. HotDep [CvR14]. HotPower [VZ14]. HotPower’11 [DB11]. HotSWUp [SN13]. HotSWUp’09 [DNT10]. HotSWUp’11 [HN12]. house [Wil93]. house-building [Wil93]. HP [MPPZ87, MW09, WGSS95]. HPC [CMK^{+}06, HD12, HCJ07, NTHAB22, PFK^{+}22, Tai13, TDM12, TGR^{+}21]. HPC-Colony [CMK^{+}06]. HPC-jobs [TDM12]. HPCWorkloads [LLSK24]. HTM [KGGK09]. httpd [BW95]. Huge [KYP^{+}17]. human [Klo80]. Hurd [WB07]. Hwang [KTC03, KCL03]. Hybrid [DFL06, GSW^{+}17, Str12, ZH16, CALM97, CII^{+}10, DTR01, FdAM14, Svo73, ZG07]. Hydra [CJ75, LCC^{+}75, WLP75]. Hyper [MKL^{+}19]. Hyper-parameters [MKL^{+}19]. Hypercallbacks [AWT17]. hypercubes [Nai96]. Hypervisor [BS95b, KYP^{+}17, XD17, BBD^{+}10, SLQP07]. Hypervisor-based [BS95b]. hypervisors [SPF^{+}07].

I/O [And95, BJL^{+}06, BP91, BS96, CG00, CKR08, DS09, EBP16, GNB^{+}09, GPK^{+}07, HF08, HXL01, ID01, KMN^{+}16, Lak85, LSP07, MDK96, NTC^{+}21, NXQ05, OD89, PSMB16, PSK08, RB24, Rus08, Ste97, VW08, WLRZ03, WBB02, XLDB09, dBB08]. I/O-intensive [NXQ05]. I/Os [OBSR16]. IA [ZRMH00]. IA-64 [ZRMH00]. IaaS [PPO14]. IBM [GPR87, HO91, San81, SF80, WZWS08]. ICHU [SSS01]. ID [JY98, KLY03, Sco04]. ID-based [JY98, KLY03, Sco04]. ideal [Her77]. Ideas [JW24, Tsa16, TCH^{+}91]. Identifier [Nes82]. Identifying [ZSG^{+}17, CG06]. idiom [KKM^{+}06]. idle [AYK08, JXT93, Nic87, dGdB10]. idleness [ID01]. Idletime [ET05]. IDS [GXJJ03, LLY05]. IEEE [Sat99, Ano75, Cab90, Had01]. IFIP [Fêa83, Had84, OSV82, OST83, OSV86, San86, Owe84]. IFIP/Sec’83 [Fêa83, Had84]. IFIP/Sec’84 [San86]. II [UNMS94, Var97]. IKE [SS00]. Illustrated [Wai97b]. Illustrating [FV06]. ILP [PRAH96, RF98]. Image [GSCM16, OVS^{+}06]. iMAX [KCD^{+}81, PKW81]. iMAX-432 [PKW81]. Impact [BAI93, Ros06, CS08, CB93, EAS07, GBG^{+}10, HKO^{+}94, RBH^{+}95, ZSK97]. Imperative [JCY^{+}19]. Imperfect [Wel88, Hog88]. implement [CSBA17b, DBRD91, San81]. Implementation [BCR^{+}14, BR10, CBZ91, GLG93, Hem89, JHT^{+}07, LCL^{+}16, LCJS87, TT00, Zha23, AWSBL99, AWSBL00, AGB^{+}77, AMO^{+}12, BM91, BJW87, BJM^{+}96, CCLP81, CF89, FL77, Fog74, GRB^{+}08, Hem88, HB80, Hsi89, IvdLH^{+}00, KRS97, Kea88, KAS^{+}06, Kot88, yL91, LRV94, LWQ09, Lux95, MSP98, Maf94, MM91, MB80, NL97, Nut74, Oes01, OSSN02, RO91, SGGB99, SGGB00, SHC73, ST01, TWL05, VL87, YAK93, YTR^{+}87, vR92]. Implementations [AHC^{+}16, Moh78]. Implementing [BN83, FMP^{+}95, LCH^{+}05, MFGSP12, Wai95a, CFR98, GPR87, Lie94a, Lie94b, SS72, Tan79, Her92a]. Implication [Lak85]. Implications [DLLN18, MT17, HKL^{+}06, LRS^{+}08, VZ91]. importance [DB99, DB00a]. impossibility [FV06]. Improve [CYG^{+}17, EAS^{+}17, GZH^{+}19, GKL95, HAF^{+}07, RCL01]. Improved [LW04, LH04, GS13]. Improvement [Che04, CCK04a, CL04b, YW04]. improvements [CH07, Sin85]. Improving [Bia17, BRW89, GPK^{+}07, JMK^{+}08, Lie93b, MRC^{+}97, MBS16, OCLN14, PHYO96, YS94, CMT94, HHS05, LJS^{+}02, OSV86, SAF07, SPR00, SSR^{+}10b, VDGR96, ZG07, dJKH93]. in-kernel [Uhl07]. in-memory [VGBT14]. In-Network [LLN^{+}17]. in-place [SCM05]. IncBricks [LLN^{+}17]. Incident [LZH^{+}22, LCKFA24]. includes [SJ95, vEBBV95]. incoherence [HCBS04]. incomplete [LH04]. incompletely [ABC^{+}02]. inconsistencies [Bre83]. Inconsistent [MCXS16]. Incorporating [GSGN00]. Increased [CYMT16].

Increasing [yKPR02]. increment [FG91]. Incremental [ZFW10, BPP12, KPS09, PAB^{+}95, XX00]. independent [ECS73, JRR97, LFH^{+}09, MEG94, PG03a, PG03b, RTY^{+}87, SCFS98]. Indeterminacy [AGP77]. indexing [BL03, CZG^{+}05]. indirect [JMK^{+}08]. induced [PS99a]. Inference [AKP^{+}25, BFZ^{+}25, BKGT25, HNK^{+}17, KKS^{+}16]. inferring [ECH^{+}01, LBJ03, LPH^{+}07]. INFLOW [DK15]. INFLOW’15 [DK15]. influence [PM03]. Information [ADAD01, Cho77, FXZ^{+}17, Had85, KYB^{+}07, OSV82, OST83, OSV86, YSCC16, BC08, CLC05, EK08, EHD07, FM98, Gif81, GBBL85, JdLT^{+}95, KMSV10, Kil00, LW01, ML97, Sal73, Sal74, ST93, SLZD04, VBHN10, WYC03b, ZZP04, dGdB10, OPSS93, OSV82, OST83, OSV86]. Information-Flow [YSCC16]. Informed [PGG^{+}95, PGS93]. Infrastructure [JXQ^{+}22, Ott18, YBQ^{+}25, AFF^{+}09, BDS^{+}09, DKW^{+}06, FIM^{+}11, MPP^{+}08b, dOL12, Ram00, RCSW10, RJK^{+}14, ZZP04]. Infrastructures [YJX^{+}16, HSS^{+}06]. Ingens [KYP^{+}17]. inheritance [FS96]. initial [ST01]. initialization [DIN05, Jan75, War76]. initiated [BMD94, EBS01]. innovation [DVS12]. input [BP91, CCZ^{+}07b, FO72, Har88, MP89]. input/output [BP91, FO72, Har88, MP89]. inputs [SMTZ09]. inserted [MDK96]. Insider [NCBB14]. Insights [TS06, ETKF07]. Inspired [Wil16]. installation [Fos87]. INSTANCE [HPG00]. Instruction [ASR^{+}17, MSP^{+}06, BEH91, BS02, CKDK91, DV87, Kep91, LBF^{+}98, MA06, OB86, OA08, Ros78, San81, SS98, Wal91, WS91b]. Instruction-Level [ASR^{+}17, LBF^{+}98, Wal91]. instructions [KT91b, KKM^{+}06, Lie94a, OS80]. instrument [DH73, OMCB07]. Instrumentation [TAH^{+}22, Mcd77, MMB96, OMCB07]. instruments [OB86]. Integer [MPPZ87]. Integrated [BSR06b, STYC02, CCW^{+}11, CKK^{+}07, DCZ96, JT90, LK08, LK01, PV95, WLS^{+}02]. Integrating [BEH91, cCVP99, CVP00, KHL^{+}07, OCF00, OBSR16, Rei92, AMMR92, EHD07, WSH94]. Integration [HGDG94, FR85, JTG^{+}00]. integrity [KDP02, SLS^{+}05, SLQP07]. Intel [CCW^{+}11, GCJ17, Rat11, vdWMH11]. Intelligence [BHD19, Che17, DIS19, KHG^{+}17, KEF^{+}19, DKW^{+}09]. Intelligent [BP91, LZH^{+}22, JXY95, JLZx90, XDC^{+}95]. InteMon [HSS^{+}06]. Intensive [ID25, NTC^{+}21, GWSY08, NXQ05, SBH^{+}10]. intentional [AWSBL99, AWSBL00]. inter [GW04, LJX97a, WV02]. inter-component [WV02]. inter-group [LJX97a]. inter-process [GW04]. Interaction [WV02, ALBL91, AM77, SHT97, SZII11]. Interactions [DK15, OHW17, Col73, HZ09]. Interactive [JHK^{+}16, MCdL06, BGS04, DH73, EWCS96, FURM00, HJT^{+}93, SLN00, SLN99]. intercommunication [Kno74, Kno75]. Interconnect [SKJ^{+}17]. Interconnected [VMM20]. interdisciplinary [CGJ^{+}07]. interdomain [Küh98]. Interface [LSMB16, BJM^{+}96, CJR87, DTR01, FHL95, HDH^{+}94, Jon93, Kep91, yKPR02, MK91, MQW95, Moo82, MEG94, Sch73b, vEBBV95, Jon92].

Interfaces [Wit16, BSR06b, CMK^{+}06, CBD^{+}98, Gue88, Str78]. Interfacing [ACG86]. Interference [HJrCH16, CHLS16]. interim [Nee77]. Interleaved [YJX^{+}16]. interleaving [LTQZ06, LGH94]. interlock [Eas72]. Intermediate [HS16, WP87]. Intermittent [CHLS16, WCS08]. Internal [DL15, FW72, Rou84]. Internals [Woo85, GKD91, KB84, KGB88]. International [BCC^{+}94, BR10, CM13, CM14, Had93, HLR98, Her92b, LS09, San86, Voe98, Wai83a, Owe84]. INTERNET [CKMV99, Bel10, Arn10, BvS00, CSJZ08, CCC^{+}05, GBCH00, JKH^{+}00, KG99, MHD^{+}07, NSS10, OLLY02, SGD^{+}02, STYC02, VFH98, WCB01, Yu00a, Yu00b, ZBN07, dVdVI98]. Internetwork [KvRvST92]. internetworks [GS95]. Internship [HMS17]. interoperability [WDH89]. interpolation [DSGP05]. interposing [Jon93, Jon92]. Interposition [Jon93]. interpreted [Ros95]. interpreter [OKN02]. interpreters [RLV^{+}96]. Interprocedural [WHZ^{+}17]. Interprocess [Che75b, Sor73, Cer75, Che84, CCLP81, FR85, MW75, Rus88]. interprocessor [MK91]. interrupt [DTR01, HC95, RLB08]. interrupt-polling [DTR01]. Interrupts [KE95, Hat94, Hil93, Hil94]. intervals [ET05]. Interweave [SDP^{+}00]. intolerant [ZL86]. Intra [EAS^{+}17, LJX97a]. intra-group [LJX97a]. Intra-Request [EAS^{+}17]. intrinsic [HS96]. Introducing [MW08, Rob98]. Introduction [DW08, Hoh07, Rie07, Sir06, Boe15, VZ14, XDC^{+}95, ZLX^{+}80, Lam75]. introductory [HV08]. introspection [HN08]. Introspective [MAS^{+}06]. intrusion [AMA^{+}11, DKC^{+}02, GPF^{+}05, GFPcF08, HLL^{+}02, JAvR06, YbJf04]. intrusion-tolerant [JAvR06, YbJf04]. intrusions [JKDC05]. invalidation [Gup05]. invalidation-based [Gup05]. Invariant [BDF^{+}15, Buc77]. invariants [BBE^{+}11, LTQZ06]. invented [HH88]. inversions [DS92]. Investigating [Tem98, XLDB09]. investigation [Lov77, Rob98]. invisible [MZI08]. Invited [Tsa16, Lam00, Sal00]. Invited-Speakers [Tsa16]. invocation [Led97]. involuntary [PB08]. IO [PSK08]. IOMMU [MMT16]. IoT [BHD19]. IOV [XD17]. IP [BSR06b, CWL05, PN00, YLE02]. IPC [GA91, Lie93b]. IPTables [GC05]. IPwatch [LS90]. Iris [PSMB16]. IRON [PBA^{+}05]. ISA [BLJ^{+}17, BSR^{+}15, KF09, TML^{+}17, VSST16, Wit16]. ISBN [Woo85]. ISDN [NB91]. ISIS [Bir85, BC91b]. isolating [KJ08]. Isolation [LS94, JSDG08, SFS13, VGR98, WLAG93, WSG02, WRA05]. Isolation-only [LS94]. Issue [Eid15]. Issues [CM14, Lit87, SMI80, BIYC06, CL95, CM13, GA98, Gup05, MKY08, PS99c, Pat02b, TG89, VT01, YS98, YAK93]. Itanium [WCW^{+}04]. Itanium-2 [WCW^{+}04]. ITC [SHN^{+}85]. Iteration [SSK17, SWL77]. Iterative [JXG21]. iterators [Ste97]. Itrustpage [RSW08]. ITV [NLO95]. Ivy [MMGC02]. IX [FPG89].

J [Had85, Wai95b, Woo85]. Jade [WBC^{+}83]. Jain [WP91]. Jas [Bla95]. Java [Heu97, GA98, CDG^{+}17, GLC99, HvE02, Led97, NB00, NAR08, Oes01, OKN02, PG03b, WBDF97]. JavaOS [Mit96]. Jensen [Nut94a]. Jerome [Tug83]. Jini [ATMZ01]. JIT [OMCB07]. JNI [CDG^{+}17]. job [BDF^{+}08]. jobs [AVZR11, TDM12]. jobscheduling [ST00]. John [Had83, Heu97, WP91, Wai83a]. Joint [SHA02, VA96, Voe98, LM97]. journal [Spr85]. journey [Wil09]. July [OST83]. jumps [JMK^{+}08]. JUSTDO [IKK16]. justification [Sib76]. JVM [PG03a].

K-entries [Nai93]. K42 [DKW^{+}06, KAR^{+}06, WdSA^{+}08]. Kameleon [RHMR15]. Kang [Küh99]. KC95 [PCP00]. Keith [Wai94]. Kenah [Woo85]. Kernel [CKmWH16, CCS^{+}16, JKS^{+}15, LCL^{+}16, MFBWW20, NTC^{+}21, OVS^{+}06, UJE^{+}22, YN15, ACG86, ABLL91, BF08, BAD^{+}11, Bar81, BYVF08, CG85, CZ83, CD95b, DD12, Dru92, EKO95b, ETKF07, FBB^{+}97, Har82, Har88, HHLS97, HH89, JM95, Kor06, Kru82, Kut84, LBB^{+}91, Lie93b, Lie95a, LST^{+}06, MP89, MW08, Mcd77, MMB96, NL96, PRD10, RR81, SMS11, Sch75, SCS77, SESS96, SLQP07, Sil83, SHC73, SR89, TM89, Uhl07, VMBM12, WG08, WSG02, Fin92, HBB13, MP92a, MP92b, PHOA89]. Kernel-based [CKmWH16]. Kernel-level [OVS^{+}06, PRD10]. kernel/domain [SHC73]. Kernels [CCS^{+}16, ARS89, GLC99, MR07, MSC^{+}06]. Key [JW24, BMA00, Che04, DPW^{+}09, DSGP05, DS90, DHJ^{+}07, HLL04, JY98, LKKY03a, LKKY03b, LL04, LW04, LH04, LSH00, MKKW99, MKKW00, PL01, PS98, PCP00, PW98, SY96, STW95, Syv93, YS02, MC96]. key-value [DHJ^{+}07]. KeyKOS [Har85]. Keynote [Est02]. keys [CJ05, KC95, LGSN89]. KickStarter [VGX17]. Kill [KTG^{+}17]. Kishor [Sta83, Wai83b]. Kits [Küh04]. Kittyhawk [AUW08]. KLogger [ETKF07]. Know [DK17, Wed88]. knowledge [ST01]. knowledge-based [ST01]. known [Rou84, YLW^{+}06]. Krell [Val94]. Kuperee [DH96]. Kurt [Nut94a]. KV [BFZ^{+}25]. KZ2 [XDC^{+}95].

L4 [BS15, KEP07]. L4oprof [KEP07]. Labels [EKV^{+}05]. Labs [MW09]. lacking [BJ81]. LADIS [RAVC12, WTC09, MvR13]. LADIS’14 [CJRV15]. LAHNOS [AEE^{+}94, CCG95]. LAM [ZWZ05]. LAM/MPI [ZWZ05]. Lamport [Woo90]. Lance [Wai95b]. landslide [STM^{+}07]. Language [AM77, BD17, FAH^{+}06, MAHK16, WYW^{+}25, YN15, Als72, ACC^{+}09, AGB^{+}77, ACG86, BMER14, DBMZ08, DMB87, FBB^{+}97, GSA10, HFWZ87, Her77, HM93, KMC02, LRV94, Ros95, WP87, ZN00]. Languages [EMS09, EMSPS11, AH77, DBR09, Est02, GA98, JMK^{+}08, PGZ08, Wir87]. Large [CJRV15, RAVC12, WHZ^{+}17, WYW^{+}25, WTC09, WAC^{+}81, BS95a, BJK^{+}06, BLRC94, Bod11, Cec00, CMK^{+}06, EJD13, FES09, GBBL85, GBZP10, GB90, GSM08, HSS^{+}06, JLZx90, KJH^{+}11, KKFB11, KSS^{+}96, LPS10, LJX97a, LGN07, Neu89, RRBN09, ROLV06, Ros89, RD01, SATG^{+}07, Sal91, SF12, SPHC02, SSR^{+}10a, TLD^{+}11, VYW^{+}02, YZZZ06, WS92]. Large-Scale [CJRV15, RAVC12, WTC09, WHZ^{+}17, BS95a, Bod11, FES09, HSS^{+}06, KKFB11, KSS^{+}96, LGN07, RRBN09, ROLV06, Ros89, RD01, SF12, TLD^{+}11, VYW^{+}02, YZZZ06]. largely [Sal78b]. Last [DK17, LSKK08].

Latency [BFZ^{+}25, JHK^{+}16, SS07, ZE16, BKP^{+}96, DC99, DC00, EWCS96, JFV^{+}96, PSMB16, SGK^{+}04, Ste97]. Latency-Critical [ZE16]. latency-sensitive [DC99, DC00]. Later [MFBWW20]. lattices [Pon97]. Launching [RD87]. Lawrence [Woo85]. layer [GUB^{+}08, ZL04b]. layered [LBJ03, PSC^{+}07]. layers [AEE^{+}94, KC94, Lin81]. layout [GJXJ03a]. Lazy [DB96, LLS91, BL89, CGS^{+}96b, ZIL96, You92]. LazyBase [KMSV10]. LCM [LRV94]. LDX [KKS^{+}16]. leak [BM06, HC04]. leakage [ZZP04]. leaks [ZJS^{+}11]. learned [Sha00, ZH19]. Learning [CKN^{+}19, ES10, GZH^{+}19, GLD^{+}22, JCY^{+}19, LPM17, LCCZ17, LXYZ19, LLSK24, LPSZ08, MKL^{+}19, PWT^{+}19, PFK^{+}22, VJ19, Bod11, MZI08, ZFW10]. Learning-and-System [LXYZ19]. Leases [GC89]. Least [NTHAB22, Rob96]. least-utilized [Rob96]. Lee [KCL03, Küh99]. legacy [SLS^{+}05]. length [SEP98, YN12]. Less [BNE16, DB00b, EKF^{+}14, HKL^{+}06, KLS^{+}10, TH94]. lesson [WL94]. Lessons [Cas91, ROLV06, Sha00, Wet99, Wet00, WdSA^{+}08]. letter [Hof07]. Level [ASR^{+}17, EPG^{+}20, HT15, RS02, AEE^{+}94, ACG86, ABLL91, AMO^{+}12, BM91, BSM^{+}12, BBD^{+}02, BW01, BMP^{+}04, CG91, CCEH00, EB78, EKO95a, Fes07, FURM00, Hal00b, Hal00a, HSI^{+}01, HEK^{+}07, JSDG08, KAI^{+}13, LS75, LBF^{+}98, MSLM91, MT02, MQW95, MRA87, OT95, OCF00, OVS^{+}06, PCH^{+}14, PRD10, RRT^{+}08, Sch73a, Wal91, WF07, ZZ03, ZWZ05, ZJS^{+}11, vEBBV95]. levels [HZ09, dVdVI98]. leverage [CJS^{+}09]. Leveraging [GZH^{+}19, HS16, GPV04]. Li [JW01, KCL03]. libvfio [RB24]. libvfio-user [RB24]. lie [CR12]. Life [Pet93]. Lifetime [NTHAB22, DK75, GS13, HBD95, OCLN14, SZ98, SLQP07]. lifetimes [Sat81, TGR^{+}21]. Lifting [HS16]. Light [vdWMH11, MSC^{+}06]. Light-weight [vdWMH11, MSC^{+}06]. Lightweight [BALL89, CKmWH16, CGS^{+}96b, JKS^{+}15, KKK^{+}17, KKS^{+}16, MCGL17, SMK^{+}93, AMA^{+}11, CH07, MEG94, TNL^{+}07, dORF12]. like [Neu00, XZZ97]. LilyTask [TWL05]. Limitations [Kos73, Bir07, CS93, LMG^{+}07, Pu93]. limited [BC83, GG73]. limiter [Loe89]. limiting [CCK04b]. LimitLESS [CKA91]. Limits [Wal91, LB08, YV01]. Lin [KTC03]. Linda [CG85, CG93]. line [DH95, KGS06]. Linkage [Ros94]. linked [LB81, RMS98]. linking [Jan75]. Linux [Kad95b, Kad95a, AR07, BYVF08, BBHL08, DIN05, FM02, Gan08, HBB13, JKS^{+}15, JHT^{+}07, KAS^{+}06, Kor06, MW08, MFBWW20, NV06, PLM06, PLHM08, RLB08, SLM11, SG04, TF04, UJE^{+}22, VMBM12, WRA05, WTKW08, WXX08, dBB08]. LISP [SH87]. literature [Met82, Bru86]. Litmus [LWPG17]. Live [KS09, SHW^{+}15, XD17, HDG09, OB10]. Lived [LCL^{+}16]. liveness [BC08]. LLexus [LCKFA24]. LLM [AKP^{+}25, BFZ^{+}25]. LLMs [YG25]. LND [MZWZ02]. Load [AEP^{+}97, PL95, BMD94, EDZ07, HBD95, JXY95, LWS96, ZSK97]. loading [LL98]. Local [CIP^{+}23, FR94, KLK17, AEE^{+}94, Fab98, HJ10, Kan83, SHA02, Spe81, Tem98]. localities [Mas77]. Locality [LSL^{+}17, PAB^{+}98, SZD04, SSK17, WCL17, CMT94, CR72, DK75, LSKK08, LWS96, LS90, MT96, PSG06, PEA^{+}96, VDGR96, Wei98, ZYG00]. Locality-Aware [LSL^{+}17, PAB^{+}98, LSKK08]. localized [LOM^{+}09]. Locating [ACS15]. location [LB81, ST93]. Lock [GMT16, YWKYS15, KPS09, LT11, RG02, MP92a, MP92b]. lock-based [LT11, RG02]. Lock-Free [GMT16, YWKYS15, KPS09, RG02, MP92a, MP92b]. locking [Lie94b, MMTW10]. Lockless [DD12]. locks [Gil78]. LOCUS [MMP83, PWC^{+}81, WPE^{+}83]. Log [Bal24, CGKM11, FC87, JHT^{+}07, KAS^{+}06, MBD^{+}12, MRC^{+}97, OCLN14, OD89, Rob96, RO91, SS06, SK97, WECK07, WB86, ZFW10]. Log-based [CGKM11]. log-structured [JHT^{+}07, KAS^{+}06, MRC^{+}97, OCLN14, OD89, Rob96, RO91]. log-synchronization [SS06]. Logged [CD95a]. Logging [IKK16, KKB^{+}16, CGS^{+}96b, DHRS91, DKC^{+}02, ETKF07, Hag87, Spi94]. logic [ACC^{+}09, BH81, BAN89, Nes90]. logical [Ray92, dJKH93]. logics [XZZ97]. login [CCK04b]. logistic [BDDMR11]. Logs [DRTT24, JW24, YJX^{+}16, BBE^{+}11, Bod11, LFWL10]. logTM [MBM^{+}06]. Long [BFZ^{+}25, BSR^{+}06a, Eas72, RD97]. Long-Context [BFZ^{+}25]. long-term [BSR^{+}06a, Eas72]. longer [Den74b, XHB06]. Look [HMS17, BJK^{+}06, BSR^{+}06a, BKP^{+}12, Hol82, JM98, Mas87, Spi74, Syv96]. look-alike [Hol82]. lookaside [Ros89]. looking [Mat06]. Loop [CSBA17c, GKO^{+}00, MT96]. loops [GDRT13, SCFS98]. loosely [LWQ09, LB81, Pea89]. loosely-coupled [Pea89]. loss [Mit00]. Low [HC04, HGR07, Ros89, SBH^{+}10, CDY^{+}17, DM90, DB99, DB00a, EKM04, EKF^{+}14, Fes07, HSI^{+}01, LC04a, MPP^{+}08a, MCM01, PS09, PSMB16, PS01, RRP06, SGT96, SCP^{+}06]. low-bandwidth [MCM01]. low-computation [LC04a]. low-cost [PS09, SCP^{+}06]. low-importance [DB99, DB00a]. low-latency [PSMB16]. low-level [Fes07, HSI^{+}01]. Low-overhead [HC04, HGR07, RRP06]. Low-power [SBH^{+}10, EKM04, PS01]. Low-synchronization [Ros89]. LRP [DB96]. LRU [MPC08]. LSI [ZDP83]. LSI-11 [ZDP83]. Luna [HvE02]. lunch [LCJV^{+}11]. LVQ [CXMX05]. LVQ-based [CXMX05].

M [vR14]. M3 [AVN^{+}16]. Mach [Bab90, Wie92]. Machine [CKN^{+}19, KMK16, LPM17, PFK^{+}22, RTY^{+}87, VJ19, AGSS10, AMA^{+}11, BFHW75, Bod11, CSS^{+}91, DKC^{+}02, EKO95b, FLM^{+}08, Gar07, HJ10, HUL06, JADAD06, LBF^{+}98, LC02, Lie94a, MZI08, MEG94, MAK07, PK75, RN83, SGGB99, SGGB00, TSLBYF08, TNNI87, Vag10, WP87, WHZ^{+}17, BH75]. Machine-independent [RTY^{+}87]. machines [BBM09, CWdO^{+}06, FHL^{+}96, HDG09, HKU79, KSS^{+}96, LSS04, hTMAC^{+}08, Mou96, PBYH^{+}08, SNV10, SAL20, ZWL09]. Macmillan [Wai86, Tug83]. MacOS [MCM07]. Macroprogramming [AJG07]. MADCE [HDL^{+}02]. MADIDS [GXJJ03]. MAGE [Bro00b]. Magnitude [BNE16]. Mahler [WP87]. mail [SBL99, SBL00]. Main [AW17, AMH^{+}16, CCHV11, LLD^{+}04, WZ94]. mainstream [BBHL08]. maintainability [Bro00a, Bro00b]. Maintaining [MO85, HBP06]. maintenance [BHB^{+}08, DGH^{+}88, LJX97a, LSS04]. Make [Zha23, BC10, Fle07]. Making [MW91, Mit00, MP96, Mul87, YWKYS15, Zim94, CMSK07, CMN02, HCBS04, JS08, MV86]. malicious [BK12]. Mallacc [KXWB17]. malware [CWdO^{+}06]. Mambo [WZWS08]. manage [GPV04]. Manageability [SBL99, SBL00]. Manageable [EK08]. Managed [MAHK16, BJM^{+}96]. Managed-Language [MAHK16]. Management [AW17, BLI17, HJrCH16, KGGS18, MBS16, PPM17, YBQ^{+}25, YG25, ARS89, ATSV06, AH80, AMMR92, ABLL91, BDF^{+}08, BFS89, BFD97, CH81, CJS^{+}09, CKR08, Cra83, DSGP05, DBRD91, Duc89, Edi13, EKO95a, FMP^{+}95, FR85, GGL^{+}09, Gor87, GB90, GTHR99, GTHR00, Gre72, GA08, HMSC87, HCZ98, HCG^{+}06, Hol88, HM93, Isa07, JM98, KLS^{+}10, KPG93, KMSV10, KHL^{+}07, Kno74, Kno75, Kru82, KBC94, LEK91, LCKFA24, Les04, Lev07, LLD^{+}04, LGN07, LLS^{+}08, Lon93, LK01, MP85, Mas77, MKKW99, MKKW00, McD00, NS07, dOL12, PL01, RRT^{+}08, ROJS09, RP07, RTY^{+}87, RS86, Rat87, RS00, RD01, STYC02, SG10b, SWC08, Sto84, SAF07, Tra82, Tur80, Wal02, WTB10, WL82, Wel95, WWGF08, ZPS^{+}04, ZXMJ04, HC92]. manager [LHWY83, MM81, Moo82, SDE85]. manager/virtual [MM81]. Managing [BJK^{+}06, CAT^{+}01, DKW^{+}09, TS87a, TTP^{+}95, BHB^{+}08, BCP^{+}08, Bod11, NXQ05, RHP^{+}07, SKI08, ZELV02]. MANETs [FGR^{+}07]. manifesto [JLR^{+}05]. many [CCH^{+}87, GTSS11, GKS11]. many-core [GKS11]. Manycore [BMF^{+}16, KSP09]. Manycores [AVN^{+}16]. MapCruncher [EHD07]. mapped [BLRC94, Cec00, Lon93]. Mapping [LBvH06, SCFS98, WK08]. MapReduce [RRBN09, WBR^{+}12]. March [Sat99]. mark [CHV04, TSE^{+}00]. mark-sweep [CHV04]. Market [WM16, dSM16, Had93, Svi83, Tug83]. Market-Based [WM16]. marketplace [KMK10]. MARS [DRSK89]. marshalling [Fes07]. MARUTI [LTCA89]. mashup [OB10]. MashupOS [WFHJ07]. Mass [dSM16]. Mass-Market [dSM16]. Massachusetts [Had83, Woo85]. match [MV86]. match-making [MV86]. Matching [DZ95, MRH^{+}21, Sha95, XFO08]. Maté [LC02]. mathematical [CR72]. matrix [Lei89, SHV01]. matters [AT10]. Maximizing [ZH16]. May [Fêa83, OSV82, OSV86, MMTW10]. McNamara [Had83]. MCTS [Bro75, Bro76, EW76]. MDX [Sch95]. mean [ECS73]. means [CCG95]. measured [CEC^{+}95]. Measurement [Voe98, Lie96, NRS13]. Measurements [BHK^{+}91, Mon77, Gwi94, Svo73]. Measuring [LC04b, MCD^{+}08]. mechanics [Uhl07]. Mechanism [CCS^{+}16, BD91, CBD^{+}98, CJG02, FH85, GC89, GJXJ03a, GJXJ03b, HL92, JXHQ02, KPS09, LCC^{+}75, LJX97b, LK08, LBJ03, MRA87, ME08, MMP83, PP83, RPM97, VKD02, WZZ93, WLZ03, XXM04, ZWZ01, ZXMJ04]. mechanisms [GA91, Her86, HC95, JM98, Jan81, LSAS77, Loe05, RN93, Smi88, WV02]. media [CFR98, GA91, HCK08, Maf94, Son05]. Mediating [OHW17]. medical [LWQ09]. Medusa [Wai83a, Ous81]. meet [FHL^{+}96]. meeting [Sch73b]. Meets [DDK^{+}16, DIS19, LS09, CG93, CM75, SDP^{+}00]. Melange [MHD^{+}07]. Mem [ACM02]. membership [And83, BDM97, Kah85]. memif [LL16]. memoing [CFR98]. memoization [SL98]. Memories [KPS^{+}16b, BJM^{+}91, Ger72, RRCC10, Tan79]. MemorIES3 [NMS^{+}00]. Memory [AW17, AAA^{+}23, AC23, AZEE18, AMH^{+}16, CIP^{+}23, DHK^{+}15, DDK^{+}16, DDM^{+}18, GPY^{+}17, GSCM16, IKK16, KXWB17, KGGS18, LL16, LZC^{+}17, LSMB16, LWPG17, Mos93, NHH^{+}17, NP17, PG16, Rat87, SKB^{+}17, TML^{+}17, Wal02, WTLS^{+}09, ZLJ16, ARS89, AB75a, AMMR92, AL91, ACM02, BMBW00, BXS14, BFS89, BSF^{+}91, BM06, BCRS10, BMP^{+}04, CH81, Cec00, CLR94, CRD^{+}95, CB93, CCHV11, Che85, CD95a, CNV^{+}06, CMSK07, CMM^{+}06, CH98, CGS^{+}96b, CR12, CF89, DFL06, DV87, Des10, DCZ96, EDZ07, ENCH96, Esk96, FMP^{+}95, FR85, FP89, GCM^{+}94, GGH91, GTHR99, GTHR00, GKV07, HC04, HGDG94, HM93, HSPC01, IKWS92, JM98, JFV^{+}96, Jan81, JKW95, KLMO91, KT91b, KSDC14, LEK91, LMG^{+}07, LRV94, LSKK08, LJX97b, LLD^{+}04, LK08, LLS^{+}08, MSP98, MK91, MP85, Mas77, McD00, MMTW10, MBM^{+}06, NPC06]. memory [NHM83, OCF00, PRAH96, PLH98, PPO14, RK11, RGAB98, RTY^{+}87, RS86, Ros89, RHP^{+}07, SCL96, SMK^{+}93, SGT96, SG97, Sch73a, SFL^{+}94, SDP^{+}00, SJGY94, SHT97, SF91, SDH^{+}97, Sto84, TSF90, TWL05, Tem98, Tra82, TG89, VZ91, VGR98, VGBT14, WL82, WK08, WMH72, WCA02, WRA05, WZ94, XHB06, YZG^{+}11, YTR^{+}87, ZWL09, ZIL96, ZPS^{+}04, ZPS99, ZPS00, CR12, HC92]. memory-aware [EDZ07]. memory-based [LSKK08]. memory-mapped [Cec00]. MEMS [GSGN00, SGNG00]. MEMS-based [GSGN00, SGNG00]. Mercury [HCG^{+}06]. Merge [LCWM08]. MERT [BL75]. Merz [Wai97a]. Mesa [GMS77]. Mesh [SCG01, CC97]. Mesh-based [SCG01]. Message [DK15, Yan92, AWW08, BY08, BBG83, CLR94, Cha90, CJ05, Che84, FAH^{+}06, HGDG94, HLFZ97, MW75, SCM05, SHSB75]. message-based [Che84, FAH^{+}06, HLFZ97, MW75]. Message-Passing [Yan92, CLR94]. messages [dORF12]. messaging [AC97, KC94, WLRZ03]. Meta [You92, CCEH00, MW92]. Meta-Data [You92]. meta-level [CCEH00]. metacomputing [PFGD02]. Metadata [KDL^{+}16, ZG07, dORF12]. metal [RB24]. Method [MSF85, Cha73, Led97, QTSZ05, Tan79, TFC99, WC02, WG08]. methodologies [Had85, OSV82, OST83, OSV86]. Methodology [NCBB14, Her92a]. Methods [Nut94a, Val94, Dim98, MRC^{+}97, MMB96, Ste83, WJMC04]. metric [MB08, SS17, Mcd77]. Metrics [JXQ^{+}22]. Michel [vR93]. Michigan [Wai83a, HGB^{+}80]. Micro [Wai86, KAI^{+}13, Lie95a, LE96, Neg00]. micro-devices [Neg00]. micro-kernel [Lie95a]. microarchitectural [LB06, LB08]. microcomputer [Rat87, ZDP83]. microdrivers [GRB^{+}08]. Microfilms [Wai83a]. Microkernel [BS15, BCE^{+}95, CL95, KEP07, Sto07, Uhl07, ZPS99, ZPS00, dORF12]. microkernel-based [Sto07]. Microkernels [FHL^{+}96, HUL06, HEK^{+}07]. micropayments [LOM^{+}09]. microprocessor [AB75b, ACT94, DMB87, GS13, SCP^{+}06, UHMB94]. microprocessors [WJMC04]. microprogrammable [Tan79]. microsecond [AD99, AD00, DM90]. Microservices [GZH^{+}19]. Microsharding [TPH12]. Microsoft [Sch07]. middle [RA06]. middle-ware [RA06]. Middleware [DIS19, MDB01, RAVC12, WTC09, CPW07, EBS01, EAS07, GHP^{+}08, KGS06, CJRV15]. might [HH88]. migrate [LAB^{+}06]. migrating [IvdLH^{+}00, OSSN02, PL95]. Migration [CAW08, Pat02a, RS02, SHW^{+}15, Sch95, XD17, ZSK97, Bec90, BW01, CWS06, CDV^{+}94, DDYM99, HDG09, KS09, Lux95, Nut94b, PM83, RH97, SCP^{+}02, Smi88, Won93, Zay87]. MIMO [AHB15]. Minding [BYVF08]. mini [SMS11]. mini-kernel [SMS11]. minimal [CSS^{+}91, Ful73, MPP^{+}08a, ZLX01a]. minimal-total-processing-time [Ful73]. minimally [CGM97]. minimization [MPP^{+}08b]. minimize [SLCG89, TL96]. minimizing [DD80, SS07]. Mining [BBE^{+}11, LFWL10, HLL^{+}02, HSS^{+}06]. MINIX [GLG93, HBG^{+}06, Wai95a, AEG^{+}91, CG93, KPG93]. Minute [MW92]. MIPS [CKDK91, LE96]. MIRAGE [GSM08, CKK^{+}07, FP89]. mirrored [YVM13]. mis [Mog06]. misbehaved [SESS96]. miss [GMM98, ZPS^{+}04]. misses [BLRC94]. Missing [BFZ^{+}25]. mistakes [LPSZ08]. Mitigate [KSCK17]. mixed [AVZR11, Maf94, Mil92]. mixed-media [Maf94]. ML [LSV^{+}19]. MLS [RAF07]. MMS [Cas95]. Mobile [CKMV99, CH14, DIS19, Duc92, KHG^{+}17, KG99, LJdL^{+}16, SH00, Sub11, BTK11, BBD^{+}10, BBBAN04, CC05, CWL05, DZP^{+}11, FS99, FS00, GA98, GXJJ03, HYS03, JdLT^{+}95, KCLZ98, KP97, KXD00, Lac00, LC04a, LP01, LS94, MCdL06, MES95, ÖGA06, PS09, RMSB01, SNKP95, Sat95, SLLP^{+}10, Wel95, dLWZ00a, dLWZ00b, CWL05]. mobility [BAI93, DB97, JLHB87, NSN^{+}97, SJ95]. mobility-aware [DB97]. Mode [SLD15, Gue87, SS00, ZL04b]. Model [BWV^{+}12, CKmWH16, CDG^{+}17, LWPG17, NCBB14, TML^{+}17, Var97, WYW^{+}25, WN80, AUS98, CW92, CD95b, CR72, COS^{+}08, CCLP81, DS72, Fog74, HV08, HCZ97, HCZ98, HM90, HLFZ97, HL96, JLZx90, KL02, LJX97b, LJX97a, LCWM08, LZ03, MXXC05, ML85, MPC^{+}02, ML97, PFGD02, PG03b, Pra86, SSS01, SBB86, SJ05, SLLP^{+}10, VESM10, WCL^{+}04, ZLX01a, HR92]. Model-Based [NCBB14]. Model-Friendly [WYW^{+}25]. Modeling [Gil78, HCK08, JR05, SEF^{+}16, Voe98, WL15, DMD13, FFM07, ÏMC^{+}06, LB06, WZWZ10]. Modelling [PS99a, EBP16, HKU79]. Models [AB75a, BKL^{+}16, BHJ^{+}93, LCCZ17, BHB^{+}08, BGS04, FS08a, GGH91, GS90, HCJ07, Mos93, PRAH96, RF98, SHT97, WPC12]. Modern [FKZ17, LSL^{+}17, CSBA17b, Dim98]. modes [CCH^{+}87, WZWS08, YW05]. modification [Kep91]. modified [GKL95]. modify [WL82]. modular [Gór78, KMC02, MF75, MKJK99, MKJK00]. Modularity [Dru92]. module [Str78]. MOLAR [ESB^{+}06]. molecular [Win08]. Mondrian [WCA02, WRA05]. Mondrix [WRA05]. Moneo [JXQ^{+}22]. monitor [AGSS10, AMA^{+}11, Dun91, Had77, JP78, Lis77, Par78, San81, Wet78]. monitor-based [AMA^{+}11]. Monitoring [JXQ^{+}22, YJX^{+}16, ACT94, ATSS09, JADAD06, KL07, KEP07, LS90, MMB96, NG09, PP06, RCSW10, SMRD06, Svo81a]. Monitors [How82, AB82, AGP77, FLM^{+}08, HUL06, Hil92, Kee79, Svo73, Wol02]. Mont [vR93]. Moonwalk [KZVT17]. Morpheus [TZZ^{+}18]. Mosaic [ALM^{+}18]. mostly [EM89, PP06]. mostly-scalable [PP06]. MOTIF} [Wai94]. MP [PM83, Rei85]. MPI [PSK08, ZWZ05]. MPI-IO [PSK08]. MPP [CPdM^{+}96]. mTags [dORF12]. much [SW10]. Multi [CWL05, DMD13, DBH^{+}06, MGT^{+}17, YG25, AB75b, BMTW91, BWV^{+}12, BL75, CCZ07a, CHY05, CAW08, CFR98, DD12, FD10, Jan75, KF09, LBvH06, LCWM08, LWQ09, LPH^{+}07, MZWZ02, MDO94, Mil90, MP91, NSKS11, NBB09, NTHAB22, OA08, PFGD02, RRBN09, RRT^{+}08, RD87, SBN^{+}97, SFS13, SQP08, TGR^{+}21, WB07, WS91a]. Multi-Agent [YG25, CWL05]. Multi-Core [MGT^{+}17, DMD13, CAW08, DD12, FD10, KF09, LCWM08, RRBN09]. multi-cores [NBB09]. multi-device [WS91a]. multi-domain [Jan75]. multi-environment [BL75]. multi-level [RRT^{+}08]. multi-media [CFR98]. multi-microprocessor [AB75b]. multi-objective [NSKS11]. multi-processing [Mil90]. multi-protocol [PFGD02]. multi-server [WB07]. multi-service [BMTW91]. Multi-site [DBH^{+}06, LWQ09]. multi-stage [CHY05]. multi-structured [MP91]. multi-tenant [BWV^{+}12, SFS13]. multi-threaded [LBvH06, OA08, SBN^{+}97, SQP08]. multi-tier [CCZ07a, MZWZ02, NTHAB22]. multi-tiered [TGR^{+}21]. multi-user [MDO94]. multi-variable [LPH^{+}07]. multi-vendor [RD87]. multiagent [HCZ98]. multiagent-based [HCZ98]. multicast [CNL89, Das92, LBJ03, Oes01, ÖGA06, PL01, SB91, Toi92, TFC99, YLE02, HTW01, vR92]. multicastable [ATMZ01]. multicomputer [MK91]. Multicore [GMT16, GF15, Pen09, WM16, ZE16, ATSS09, CGKM11, GCTR08, HZ09, RKBH11, SFB^{+}09, WCS08, WCS09, WZWZ10, WL09]. Multicores [RHR^{+}17, CH14, NG09, WA09]. Multics [FO72, Mon77, Sal73, Sch75, SCS77]. multilanguage [BF87]. multilevel [FLR77]. Multimedia [VT01, WS92, BGS04, CB95, GB93, GGV96, HPG00, Hal00b, Hal00a, Hop90, LMM93, NL95, NL97, TL96, Zim94]. multiparty [LL04]. Multiple [ALM^{+}18, CB17, EMZ^{+}16, SJSM96, BEW75, BEW76, Che04, Fon72, GDRT13, KSL92, PR83, TE94, WJMC04]. Multiple-block [SJSM96]. multiple-key [Che04]. multiplication [CFR98, MPPZ87]. multiprocess [Fon72]. multiprocessing [AH80]. Multiprocessor [MP92a, MP92b, SZ92, Wal73, BKT87, Bec90, BGHL87, CDV^{+}94, CCLP81, CF89, Goo87, HWO98, HGDG94, HKO^{+}94, HH89, KCD^{+}81, KLMO91, KDS^{+}06, KSL90, MB06, Mil77, NMS^{+}00, ONH^{+}96, PR83, RTY^{+}87, Sco96, SJGY94, TS87b, YTR^{+}87]. Multiprocessors [LPM17, BSL08, BAM^{+}96, BDR97, CRD^{+}95, CAL^{+}89, GGH91, GTHR99, GTHR00, LGH94, LA94, Pea89, Ros89, SKI08, TAS07, TG89, VZ91, VGR98, WSH94]. multiprogrammed [TG89, VZ91]. multiprogramming [CFL73, Han72, How72, KSS73]. multireader [HV92]. multiserver [HL96]. multisignature [CL04b]. Multitasking [PPM17, Val94, HP93, Rei85]. multithreaded [BMBW00, GLC99, REL00, SP00, Shi00, ST00]. multithreading [LGH94, PSG06, WCW^{+}04]. multiuser [ROLV06]. Munin [CBZ91]. muse [YTM^{+}91]. Mushy [Wit16]. mutation [VE08]. Mutual [Har82, BBBAN04, Bou94, BO99, Cha96, CC97, HS88, Hof90, Nai96, OR87, Ray91, Woo90]. Mutually [BDT00]. Mutually-distrusting [BDT00]. MUVI [LPH^{+}07]. My [Dij05]. myths [SPBP06].

name [PPT^{+}93, SZN87, YAK93, ZL86]. names [Lau84, Pio89]. naming [AWSBL99, AWSBL00, HSI^{+}01]. NAND [Des10, LSKK08]. nanoscale [PJDL06]. NAS [JXHQ02]. Nash [CCAP06]. National [Sop84]. Native [CSBA17c, SJ95]. Navigating [SJS^{+}23]. Near [DDM^{+}18, TZZ^{+}18, SS83a]. Near-Data [TZZ^{+}18]. Near-Memory [DDM^{+}18]. near-optimality [SS83a]. Need [NP17, BBD^{+}02, FNRC^{+}07, Neu89, RPNT08, Wed88, WCS09]. need-to-know [Wed88]. Needed [Sal93]. Needham [Nes90]. Needs [Sha95, DZ95]. neighbors [BTK11]. Nemesis [Ros94]. NEPI [LFH^{+}09]. Nessett [BAN90]. nest [MT96]. Nested [Had77, SSK17, BO99, JP78, Lis77, MBM^{+}06, MMP83, Par78, Wet78]. Net [LaR92, CG85, vEBBV95]. NetDB [LS09]. Netherlands [OSV82, OSV86]. Nets [Nut94a, Kos73]. Netstation [VFH98]. Network [BNOW93, BCC^{+}94, DS90, GPY^{+}17, Hal00b, Hal00a, HSL17, Jef92, LER^{+}17, LLN^{+}17, RLD^{+}17, Ser21, XD17, AIKS00, AEE^{+}94, AEH75, ADN^{+}95, AD99, AD00, BDMS98, BFSG94, BSR06b, BJL^{+}06, CCG95, CAL^{+}89, CBD^{+}98, Che75c, CK86, CXMX05, DB75, DZP^{+}11, DTR01, DB96, EGE02, Est02, Fab98, FIM^{+}11, FGBA96, FGC^{+}97, GAT13, GS90, Gir82, GHP^{+}08, HLL^{+}02, HKL^{+}06, HO93, HM91, HN81, HKU79, HBP06, JFV^{+}96, JAvR06, yKPR02, LCTK01, LW01, LS90, LX00, MVKA06, MP75, MMN08, MD81, MRA87, MDB01, MMB96, MCM01, NWO87, NSS10, OCF00, Owe84, PAB^{+}98, Pet93, PWC^{+}81, RR81, RCSW10, SY96, SKPG01, Spe81, SPBP06, SDH^{+}97, VYW^{+}02, Van96, WYC03a, WYC03b, WCYJ05, WlL01, WL02, Wet99, Wet00, ZDP83, vEBBV95, Her92b]. network-based [HLL^{+}02, MD81]. Networked [NSW10, PP09, GB93, HSW^{+}00, LWQ09, SGGB99, SGGB00]. Networking [ELR15, LS09, Sub11, BTMS10, EENV02, KSK09, MB93, Mao09, ROJS09, SG04, Zho10]. Networks [AHB15, BR10, CKMV99, ATMZ01, ABKM01, AC97, AJG07, BJK^{+}06, BBD^{+}02, BBBAN04, BVR^{+}00, CDG^{+}02, Cec00, CCAP06, Cos13, EKM04, Gil78, HSI^{+}01, JHC^{+}11, LMG^{+}07, LCJV^{+}11, LAAW00, LC02, LW01, MFHH02, MV86, MAK07, NPB06, Opd75, ÖGA06, PS98, PS99d, PS99a, PS99c, PS99b, Pat02b, Pop75, RN93, Ten96, WLS^{+}02, ZS06]. Neural [GPY^{+}17, RLD^{+}17, Ser21, CCG95, CXMX05]. Neurosurgeon [KHG^{+}17]. Next [AYQ^{+}16, BH21, BW95, HEK^{+}07]. Next-Generation [AYQ^{+}16, BH21]. NFS [SM89]. Nice [VKD02]. nightmare [Pen09]. nights [AD07]. Niterói [LGMF14]. No [RRT^{+}08, RJK^{+}14]. node [LSS04, ZS06]. nodes [Sal78b]. Non [AMH^{+}16, BM90, CYMT16, CYG^{+}17, LLLG16, Ric88, Yan92, ATMZ01, BXS14, CCHV11, GC96, HLL04, KL98, KPL99, KBK02, KCLZ98, Küh99, Lam85, MR07, Par78, RB93, Ste97, WZ94, Yuv76]. non-blocking [GC96, RB93]. Non-Byzantine [Ric88]. non-determinism [Ste97]. Non-Deterministic [LLLG16]. non-multicastable [ATMZ01]. Non-Preemptive [CYMT16, CYG^{+}17, BM90, KL98, KPL99, Küh99]. non-problem [Par78]. non-problems [Lam85]. non-quiescent [MR07]. non-repudiation [HLL04]. non-strict [KCLZ98]. non-system [Yuv76]. Non-Uniform [Yan92, KBK02]. Non-Volatile [AMH^{+}16, BXS14, CCHV11, WZ94]. nonce [KSL92]. nonce-based [KSL92]. nonces [NS93]. Noninterference [NBK^{+}20]. Nonintrusively [JXQ^{+}22]. nonstationarity [SKZ07]. NonStop [Bar81]. Noordwijkerhout [OSV82, OSV86]. normality [WG08]. Norstar [Cas91]. North [Had84, Had85]. North-Holland [Had84, Had85]. NOSSDAV’93 [BCC^{+}94]. Note [RD97, Wel88, And81, Den78, Den79, Den80, Dos88, Hat94, Hil94, Lie94a, Lie94b, Lie95c, Lie96, Lon93, NS93, Woo90]. Notes [PHL^{+}77, Bre08]. notification [BF08]. Novel [HS16, DDYM99, GJXJ03b, GXJJ03, JXG^{+}02, LBJ03, OMCB07, WLZ03, WLRZ03, WBB02, YW06]. November [LGMF14]. NOWs [LL98]. NRE [KZVT17]. NRICS [XX00]. NSDR [PP09]. NSX [PPS^{+}18]. NT [Vog99, Vog00, PS96, WH99, YD02, ZWZ01]. nucleus [Bro76]. nuggets [Fle07]. null [KKN00]. NUMA [BFS89, BSF^{+}91, CSBA17a, CF89, LEK91, SKJ^{+}17, VDGR96]. number [Mit00]. numbers [Dal75, Tom75]. numerical [MP85]. NUMP [Yan92]. NVM [DK15]. NVM/FLash [DK15]. NVRAM [KKB^{+}16]. NVWAL [KKB^{+}16].

o [DBR09, And95, BJL^{+}06, BP91, BS96, CG00, CKR08, DS09, EBP16, GNB^{+}09, GPK^{+}07, HF08, HXL01, ID01, KMN^{+}16, Lak85, LSP07, MDK96, NTC^{+}21, OD89, PSMB16, PSK08, RB24, Rus08, Ste97, VW08, WLRZ03, WBB02, XLDB09, dBB08]. O-intensive [NXQ05]. O2S2 [RS08]. OASES [NG09]. Oasis [MVKA06]. Object [Gir82, SJ95, DFS00, GKL95, HF08, HM90, HM93, HLFZ97, JZZW02, JMK^{+}08, KCD^{+}81, Lac00, LC93, Lux95, MS91a, Nut94b, OLS85, PKW81, RS08, SDP^{+}00, Smo95, Svo81b, Taf82, TCH^{+}91, YD96, YTM^{+}91]. object-based [HF08, KCD^{+}81, RS08, Taf82]. object-oriented [GKL95, HM93, HLFZ97, JMK^{+}08, Smo95, Svo81b]. objective [NSKS11]. Objects [BW95, TZZ^{+}18, BNOW93, GPR87, Lor86, SZ98, SMBA10, VBHN10, Her92a]. Observability [VFP22]. Observations [Lau81, Pow89]. obsolete [Wai98]. OceanStore [KBC^{+}00]. Off [WM16, MSC^{+}06]. off-the-shelf [MSC^{+}06]. Offering [LZJ03]. offline [PRD10]. offload [yKR06]. offloading [VM07]. offs [MSP98]. oil [KSDC14]. old [BKP^{+}12, Syv96]. old-world sky [BKP^{+}12]. Olle [Had85]. OLTP [TPH12]. OMF [ROJS09]. omission [LB91, SLM11]. On-Chip [ACAAT16, KBK02]. on-demand [FGBA96, LGN07, PB08, VM07, YGG^{+}03]. on-line [DH95]. On-the-fly [Jin99, CWS06, Kep91, SZD^{+}08]. once [FC87]. One [UJE^{+}22, CCK04b, CHY05, GS89, HV92, LW04, LC04a, MA91]. one-time [CCK04b, LC04a]. one-way [CHY05, LW04]. Ongoing [Sal74]. Online [PK96, Svo73, WZWZ10, YG25, ZFP^{+}21, BM06, LSS04, WJMC04]. only [CMSK07, FR85, GS89, Lam00, LS94, SBL00, Sal00, SGT96, TSE^{+}00]. ONR [MM92, MM93]. ONTAP [DVS12]. Ontario [San86]. onto [LBvH06]. Opal [CBHLL92]. Open [AC23, BMF^{+}16, BYV08, CJR87, GHP^{+}08, HCZ98, Mah94, SY96, WS06, CKMV99, TSP17]. OpenBSD [DIN05]. OpenPiton [BMF^{+}16]. OPENSIG’98 [CKMV99]. operate [SAL20]. Operating [BIYC06, BCC^{+}94, BK08, Bre08, Bru86, CCS^{+}16, CJ75, DDOL16, DK15, EMS09, EMSPS11, Fle83, GF15, HBG^{+}06, JBW^{+}87, JM95, KBC94, LaR92, Laz92a, Laz92b, LE00, Lis72, Mat06, Mat07, NBW87, Sat99, SHP^{+}16, Sha95, VDGR96, Wai83a, Wai86, WAB^{+}89, Wil94, WLP75, ZH19, dSBP11, dV96, dSM16, AYK08, AMPS73a, AMPS73b, AMPS74, ARS89, AEG^{+}91, AEE^{+}94, Als72, ALBL91, Ant90, ACT94, Atw84, AMO^{+}12, ATSS09, BFSG94, BR09, BAD^{+}11, BL75, Bec90, BSP^{+}95, BCE^{+}95, Bla91, BW95, BC01, BP91, BDR97, BBH96, Cab90, CE88, Cha90, CIL93, Chá91, CB93, CEC^{+}95, CNC^{+}96, CGL^{+}08, CMMS77, CD95b, CL95, CYC^{+}01, CB95, CS00, CLDA07, DKW^{+}06, DRSK89, DH73, Del80, DS80, Dij05, Dim98, DBRD91, EKV^{+}05, EW76, EVvdW89, ESB^{+}06, EKO95a, EKO95b, FM98, Fle81, Fra80]. operating [Gai72, GPV04, GS89, Gor87, GGV96, GC96, Gue88, HPG00, HV08, Han83, HP93, Har88, HRU75, HZ09, HH89, HH08, KCD^{+}81, KKS89, KSP09, Kee79, KS85, KS92, Klo80, KSLA08, KAR^{+}06, Küh04, LN79, Lau81, Les04, LMM93, LTCA89, LJS^{+}02, LWMX05, Lio78, Lit88, LF13, LZ03, MA79, MR07, Mat04, McD00, Met82, Mil78, MP81, MS00, MPHD06, Moh78, Moo92, MP96, MM91, NIDC02, Nee72, Nes82, NB00, NV06, OMCB07, Ous81, PV95, PBR^{+}08, PS01, Pra87, PC75, PAB^{+}95, RR81, Rat87, REL00, RLB08, Rip03, Rob98, Rob08, RPM97, Ros94, RBH^{+}95, Ros06, RHP^{+}07, Sch95, SSS01, Sil83, SF80, SPF^{+}07, Spi94, SR89, SDE85, SAF07, SXZ^{+}88, SETB08, Taf82, TH94, TM81, Tan87, TS06, TLL94, TKP^{+}08, TBM^{+}06, TLL03, Tur87, Tur80, Van06]. operating [Var72, WH99, WV02, Wai98, WB07, WPE^{+}83, War76, WDA^{+}08, WPLP85, Wel95, WA09, WMH72, WABL93, XDC^{+}95, YTM^{+}91, YTR^{+}87, Yuv76, ZELV02, ZDP83, ZLX01a, ZLX01b, dOS08, vRvST88, Her92b, Jef92, Mat10, Pet76]. Operating-System [AVN^{+}16]. operation [BM99, DB97, EKF^{+}14, JR05, KS91a, KS91b, TC96]. operational [CJM^{+}75, DKW^{+}09]. operations [LGN07, MPF^{+}06, Spe81, Vog97]. Opportunistic [KMK16]. Opportunities [DW07b, JSS^{+}15, WTB10, HZ09, VAK^{+}11]. Optimal [Car94, Sch73a, Bor92, CK86, ELG95, LML00, Tem98, LSH03a, LSH03b]. optimality [SS83a]. optimism [Cri94]. Optimistic [Her87, KPR^{+}08, PAB^{+}95, JZ91, MT85]. Optimization [ASR^{+}17, KZVT17, dGdB10, FL77, GN96, JTG^{+}00, LE96, OKN02, dOL12, SFS13, TACT08, ZCSM02, ZWG^{+}97]. Optimizations [UJE^{+}22, CMT94, DS06, KMC02, LRW91, OA08]. optimize [FHL95]. optimized [PSMB16]. optimizer [WBR^{+}12]. Optimizing [Fab98, NCL12, RHR^{+}17, SCP^{+}02, VBHN10, YZG^{+}11, KGS06]. orchestrated [RSEW04]. order [DFS00, Le98, RGAB98, SL98]. Ordered [HTW01, BBE^{+}11, Bir94, CR75, CS93, Coo94, Oes91, Toi92]. Ordering [LSMB16, AN02, Das92, EDP06, TFC99]. Orders [BNE16]. ordinary [HS96]. Org [SLD15]. Org-Mode [SLD15]. organization [BC91a]. organizations [JM98]. organize [Jan81]. organizers [BY08]. Organizing [Mog09, Pon97, BC06, CM06, PJDL06, ZS06]. Oriented [BS15, Rei92, BR09, Che85, CBC^{+}08, GKL95, HM93, HLFZ97, JMK^{+}08, Mah94, Mal10, PPO14, RR81, Smo95, Svo81b, TCH^{+}91, TNA12, WL09]. Orphan [Aba93]. Orthogonal [Dru92]. OS/network [Pet93]. OSCL [Ens75, Sib76]. OSCL/OSRL [Sib76]. OSes [SLQP07]. OSKit [FBB^{+}97]. OSL [Als72]. OSL/2 [Als72]. OSP [KS92]. OSR [Küh99]. OSRL [Sib76]. other [Mat10, MMTW10]. Our [HBB13, Spi74]. Ousterhout [Wai83a]. out-of-band [PBYH^{+}08]. out-of-core [MDK96]. out-of-order [Le98, RGAB98, SL98]. out-sourcing [NBB09]. output [BP91, FO72, Har88, MP89]. overbooking [USR02]. overcome [ID01]. overhead [BKP^{+}96, HC04, HGR07, KC94, RRP06, SGT96, SS94]. Overheads [KSCK17]. overlap [PSK08]. overlapped [AN02]. Overlapping [KCLZ98]. overlay [ABKM01, BCRS10, CDG^{+}02, MVKA06, NPB06]. overlay-aware [MVKA06]. overlays [GHW07, Gor06, GHP^{+}08, JAvR06, LCH^{+}05]. Overshadow [CGL^{+}08]. Overview [Bod11, DO09, NB00, Sub11, WLP75, BBBAN04, Fiu06, SJL^{+}87, TM81, VT01]. ownership [SS94].

P [AUW08, KLS08, Kot88]. P.R.O.S.E [Van06]. P2P [BC06, MNP07, MXXC05, PG06]. P416 [BD17]. Pacific [HKPvR16]. packages [OT95]. packer [MRA87]. Packet [KPS^{+}16a, LBJ03]. packet-pair-based [LBJ03]. Page [AW17, ALM^{+}18, CB17, HC92, KYP^{+}17, LER^{+}17, MT17, BJ81, BSSM08, BAM^{+}96, CFL73, CDV^{+}94, CNV^{+}06, JM98, Kai75, LFZE00, Lie94a, LE96, Sad75, THK95, WTLS^{+}09, ZPS^{+}04]. page-based [CNV^{+}06]. Page-Cache [HC92]. page-referenced [BJ81]. paged [BBMT72, RTY^{+}87, Tan79]. Paging [SKB^{+}17, BJ81, BAMM77, cCVP99, CVP00, Fog74, Pot77, Sch73a, Smi78, WMH72]. Pair [MCXS16, LBJ03]. Pallas [HABZ17]. Panache [AEH^{+}08]. Pandora [Hop90]. panel [Bak95, Laz92a, ST93]. Pangaea [SKKM02]. panic [SSR^{+}10b]. Paper [Dou09, SRS22, Bac91, Bir91, FNRC^{+}07, Küh99, PM03]. Papers [CM14, CM13, VZ14]. Paradigm [BKGT25, AMS^{+}07, Sal93, SR89]. paradigms [BHJ^{+}93]. Parallax [hTMAC^{+}08]. Parallel [HJrCH16, JXT93, LHPL87, PKB^{+}16, VHAA25, Wei92, AEH^{+}08, BM91, BF87, BS89, BBH96, CPdM^{+}96, CKmWH16, CAL^{+}89, DMD13, EBP16, ELG95, HdRC95, IBY^{+}07, KRS97, KTP^{+}96, Klo80, KN96, KKFB11, KSS^{+}96, LRV94, LJX97b, MT02, MM92, MM93, NSKS11, Sop84, WK08, WCE^{+}92, XXMC05, vEBBV95]. Parallelism [JHK^{+}16, ABLL91, CSS^{+}91, FURM00, GTA06, KPR^{+}08, Led97, LBF^{+}98, TPO06, Wal91]. Parallelization [WZWS08]. Parallelizing [NPCF08, HDH^{+}94]. Parallizable [LKKY03a, LKKY03b]. parameters [EJD13, MKL^{+}19, RR72]. Paravirtual [KMN^{+}16]. ParFiSys [CPdM^{+}96]. parity [AR07, GC12, LK91]. part [Lau84]. Partial [DFS00, RRP06, SP00, Shi00, WYC03b]. partially [BBE^{+}11, CR75, DB97]. Partition [LLS^{+}08, BDMS98]. partition-aware [BDMS98]. Partition-based [LLS^{+}08]. partitionable [BDM97, MDB01]. partitioned [Van06]. Partitioner [HDGP21]. Partitioning [LZ03, BJL^{+}06, CK86, CLM^{+}07, CR72, DS73, KPR^{+}08, PB09, SF12, WH99, ZZNM01]. Party [Ng99, JW01, LSH00, XZZ98, ZLX99]. PASCAL [FL77, Hei78, Kru82, Löh77, SB78]. PASIS [Kil00]. passé [BC10]. Passing [Yan92, CLR94, Cha90, HGDG94]. Passive [MMN08, MMB96]. Passthrough [XD17, XLDB09]. Password [LSH03a, LSH03b, CC04, CCK04b, DH95, Gai78, KLY03, KTC03, KCL03, Ku04, KCC05, LFW04, LC04a, Sco04, Sin85, YW04, YS02]. past [ES10, JKDC05, Lam00, RD01]. Pastiche [CMN02]. Pastry [Her07]. patches [MPLH06]. Path [HABZ17, MCXS16, CCB^{+}06, DB00b, SEP98]. Pathfinding [YG25]. paths [MP96, PHYO96, Won93]. Pattern [AWS16, JV21, SCM05]. Pattern-Aware [JV21]. Pattern-Recognition [AWS16]. Patterns [PKB^{+}16, BRR^{+}00, MMN08]. Paul [Wai86]. Paxos [HMS17, MBD^{+}12]. payment [SH00]. PC [Fos87, Kad95b, SJL^{+}87]. PC-XINU [Fos87]. PDAs [Neg00]. PDF [Wai97a]. PDP [BBMT72, HB80, MA79, PK75, Ros78]. PDP-10 [BBMT72]. PDP-11 [HB80, MA79, PK75, Ros78]. PDP-11/45 [HB80]. Peak [CDY^{+}17]. peephole [BA06]. Peer [HKL^{+}06, AC06, BL03, BCRS10, CDG^{+}02, CCAP06, GC08, Gor06, HB06, MPHD06, MMGC02, RD01, COS^{+}08]. peer-to-peer [AC06, BL03, BCRS10, CDG^{+}02, CCAP06, GC08, Gor06, HB06, MPHD06, MMGC02, RD01, COS^{+}08]. peering [YLE02]. PeerReview [HKD07]. Pegasus [LMM93]. penalty [KT91b]. penetration [HGB^{+}80, WAC^{+}81]. perceived [MCD^{+}08]. PEREGRINE [JV21]. Performance [Acq16, BC91a, Chu75, CKN^{+}19, DLLN18, DDK^{+}16, EAS^{+}17, FPG89, GZH^{+}19, GLD^{+}22, GGH91, HO91, HP95, KPL99, KPS^{+}16a, KPS^{+}16b, LK91, LLD^{+}04, MRH^{+}21, MT17, NSKS11, NXQ05, OBSR16, Per92, PW93, RHR^{+}17, RP07, RGAB98, RB24, Ros78, SHW^{+}15, SJS^{+}23, SEF^{+}16, SN94, SB89, Svo81a, TNNI87, VGR98, WP91, WKT^{+}13, Wei98, ZH16, ZHK06, ZIL96, vRvST88, AWW08, And95, ATSS09, BvS00, BSP^{+}95, BITW07, BMR^{+}09, BBM09, BS96, BMER14, CPW07, CBZ91, CB93, CEC^{+}95, CZ83, Cla87, DV87, Des10, Duc89, EDP06, EWCS96, ENCH96, EEKS06, FURM00, FM02, FD10, FdAM14, FJLC98, GKL95, GLC99, GJXJ03a, GKS11, HLR98, HdRC95, HHLS97, HKO^{+}94, HKM^{+}87, HHS05, IKWS92, JBDP08, JMK^{+}08, JKW95, JKH^{+}00, KEG^{+}97, KLS^{+}10, KMSV10, KEP07, KN96, KF09, LRW91, LKB91, LB06, LKvR^{+}99, LKvR^{+}00, LSA^{+}00a]. performance [LSA^{+}00b, LT11, MB93, MCD^{+}08, MRC^{+}97, MDO94, MA10, MW75, MUKX06, Mil92, MB91, MP91, NSS10, OCLN14, OAE^{+}09, PG96, PBH^{+}07, PHYO96, PS96, PG03a, PN00, RS00, RLV^{+}96, RBH^{+}95, Sad75, SATG^{+}07, SBL99, SBL00, SLN00, SLN99, SB10a, SPF^{+}07, SKZ07, SQP08, SPR00, Svo73, TH94, Tem98, VW08, WSG02, WVS^{+}99, WVS^{+}00, WSH94, YZJ02, YW06, ZG07]. Performance-directed [RP07]. performance-monitoring-unit-based [KEP07]. performance-setting [FM02]. PERFORMANCE’98 [Voe98]. performances [Zea97]. Performing [Spe81]. periodic [BMD94]. peripherals [Van96]. permeating [ACC^{+}08]. perpetually [Kil00]. Persistent [GBCH00, IKK16, KPS^{+}16b, LZC^{+}17, NHH^{+}17, SKB^{+}17, GPR87, JZZW02, KBC^{+}00, ONG93, RD01]. personal [CN07, CEC^{+}95, Han83, LBP^{+}07, RCC01]. personality [CCW^{+}11, Neu00]. personalized [FS08a]. perspective [Fle83, HH89, JKL^{+}13, Lev07]. perspectives [KSP09, MA10]. Pervasive [KDL^{+}16]. pessimism [Cri94]. Petal [LT96]. petascale [BIYC06, OVS^{+}06]. Peterson [HV92]. Petri [Nut94a, Kos73]. PFF [Sad75]. Pharos [VFP22]. Phase [DHK^{+}15, KSDC14, MA91, SZD04, SS00, AN02]. phenomenon [BGMP79]. philosophers [Ran82]. philosophy [Bro76]. phishing [RSW08]. Phone [DIS19]. phones [PS09]. Physical [HC92]. Physically [Zha23]. PicoServer [KDS^{+}06]. PICSEL [MCD^{+}08]. Pieces [BFZ^{+}25]. PIFT [YSCC16]. pinning [SKI08]. Pioneer [SLS^{+}05]. pipe [KRS97]. pipeline [GTA06, YZJ02]. pipelined [MSB^{+}02, WS87]. pipelines [SCP^{+}06, SRA^{+}04]. Pipelining [Cla87]. PipesFS [dBB08]. PL [Mac77]. PL/I [Mac77]. place [BHB^{+}08, SCM05]. Placement [FVDS20, ID25, PR15, APGG00, CKJA98, LK91, TGR^{+}21]. Plan [CJM15, Svi83, Gan08, PPT^{+}93]. Planet [XDM^{+}18]. Planet-Scale [XDM^{+}18]. PlanetFlow [HBP06]. PlanetLab [ATSV06, Fiu06, MPF^{+}06, PP06, PR06, SPBP06]. Planning [HCZ98, Svi83, Tug83]. Platform [BCR^{+}14, LSV^{+}19, PPS^{+}18, RWS^{+}15, VFP22, AUW08, BBD^{+}10, KL07, Lac00, NSS10, PG03a, PG03b, RA07, Sha00, VBHN10, WCW^{+}04, YZG^{+}11, YD96, ZLL^{+}07]. Platforms [BSR^{+}15, VMM20, DPW^{+}09, NV06, RA06, SK13, SZII11, USR02]. platinum [CF89]. PLOS [EMS09, EMSPS11]. PLUS [Hol88]. plush [ATSV06]. PM [BS89]. pocket [BBD^{+}10]. point [LKB91, SB10b]. Pointer [SB10b, DS06, KKN00]. pointless [DKK10, SB10b]. points [GKS11, WV02]. Policies [Mog08, BSF^{+}91, CAW08, HD12, HK99, JL75, KL98, KPL99, Küh99, LJY04, Pot77, RN00a, RN00b, SAG06, TS06, TP72]. Policy [LCC^{+}75, Wai95b, GBZP10, LWMX05, MP85, NTHAB22, YVM13]. Policy/mechanism [LCC^{+}75]. polling [DTR01, RN93]. polycyclic [MSAD91]. Polynomial [Ull73, DSGP05, LFW04]. pool [LML00, Str12]. pools [Pea89]. poorly [LGSN89]. Porcupine [SBL99, SBL00]. Portability [CE88, MS91a, Moo92]. Portable [Wei92, Car94, CMMS77, Kep91, KS95, Mil78, PCH^{+}14, Rya98, Rya99, WDH89]. Position [Bir91, Bac91, FNRC^{+}07]. Positional [SRS22]. Possible [RB24, Bel75, DS92]. Post [HDG09, Wit16]. Post-copy [HDG09]. Post-ISA [Wit16]. poster [Cec00, CG00, Kil00, LSA^{+}00b, RN00b, Yu00b, dLWZ00b]. PostScript [Wai97a]. Potemkin [VMC^{+}05]. Potential [TZZ^{+}18]. Power [BLI17, CDY^{+}17, DIS19, LFZE00, WWGF08, ZH16, BSL08, CJS^{+}09, DB11, EKM04, EKF^{+}14, GPV04, KLS^{+}10, KS95, LB06, MB06, NS07, PS01, RRT^{+}08, SQP08, SBH^{+}10, WTB10, Wel95]. power-efficient [SQP08]. powered [RJK^{+}14]. Practical [BC08, FXZ^{+}17, GLD^{+}22, HFC^{+}06, KJH^{+}11, Lit87, MGT^{+}17, NIDC02, Nut94a, VMM20, ZJL17, DBR09, FRL00, Fog74, GAT13, HKD07, LSH01, SNV10, WMI^{+}07]. practice [LABW91, OSV86, Woo73, Lig94, Lig95]. practices [SPBP06]. pragmatic [BMW02b, MPC^{+}02]. praxis [Bro76]. pre [CM75, KY02]. pre-execution [KY02]. pre-specified [CM75]. Precise [CYG^{+}17, GMM98]. Precision [MCGL17, MPPZ87]. predicates [JKDC05]. predication [JMK^{+}08, RSEW04]. Predictability [GZH^{+}19]. predictable [JRR97]. Predicting [JHC^{+}11, TGR^{+}21]. Prediction [CYG^{+}17, JHK^{+}16, AVZR11, CCM96, CPT08, DB00b, KAI^{+}13, LB06, LJS^{+}02, LWS96, RRP06, SZD04, STM^{+}07, SEP98, SKZ07, YS94]. prediction-based [RRP06]. Predictive [Mil90, SS06, YSCC16, ÏMC^{+}06]. Predictor [BSMF08]. predictors [SJSM96]. Preemptable [TLC85]. Preemption [WLZJ17, ET05]. Preemptive [CYMT16, CYG^{+}17, BM90, FPG89, KL98, KPL99, Küh99, LS75]. Prefetching [Bha17, CKP91, CLS06, CG00, CHV04, CJG02, KTP^{+}96, LSP07, LM96, MDK96, PGS93, PGG^{+}95, RSEW04, RMS98]. Prefills [AKP^{+}25]. Preliminary [Che85, FW77, NN75]. Prelude [Wei92]. Prentice [Sta83, Wai83b]. Prentice-Hall [Sta83, Wai83b]. Pres [Wai83a]. presence [Ram00]. present [Bas12, BCC^{+}13, JKDC05]. presentations [Laz92a]. Preservation [BDF^{+}15, FNRC^{+}07]. Preserving [BJKT15]. Press [Had83, Woo85]. Prevention [HMK20, Dim98, Lev05, New79]. primitives [AL91, BGHL87, Har82, Kno74, Kno75, Kot88, MC11, Kos73]. principal [ZL04a]. Principle [LE00]. Principles [SHW^{+}15, CB95, GS78, PR06, SHN^{+}85, Wed88, BK08, Bre08, Laz92a, Laz92b, Pet76]. Printers [ASR^{+}17]. Priority [BC83, DS92, LLK96, Mil92]. Privacy [BJKT15, LSV^{+}19, Car94, CCM08, WK05]. Privacy-Preserving [BJKT15]. Private [LSV^{+}19, DS90, WH08]. private-key [DS90]. Privileged [MPF^{+}06]. Proactive [SLFP16]. probabilistic [CR75, DS06, XFO08]. Probability [Sta83, Tri82, Tri02, Wai83b, Ell73]. Probes [YN15]. problem [BRR^{+}00, BL00, Che85, Gai78, HS88, Hil92, Lip75, Lis77, MY98, Par78, PCP00, RD97, Sei90, SRH^{+}06, Wet78, YLW^{+}06, Zöb83, GB01, JR05]. problem-oriented [Che85]. Problems [SDE85, Aba93, BDDMR11, Bel75, FD10, HC95, KXD00, Lam85, Rou84, Sal78b, Ull73, WB86]. procedure [ATK92, BALL89, BN83, Coo86, TA90]. Procedures [Wai86, Boc75, Opd75, dGdB10]. Proceedings [OST83, San86, Fêa83, Had84, Mat10, OSV82, OSV86, Sat95, Sat99]. Process [DB00a, Eas72, FG91, Ger77, KSCK17, Lom77, PM83, RS02, Rus77, TG89, Var72, ZL86, AYK08, AM77, BR09, BW01, Che75a, GW04, GLG93, HBD95, HL92, How72, Jan75, Kno74, Kno75, Lau84, Nut94b, RH97, SK96, Smi88, ST01, Svi83, Won93, Zay87]. Process-based [DB00a]. process-oriented [BR09]. process-resource [SK96]. processes [ACG86, AKS73, DB99, DB00a, EKV^{+}05, Hab72, HAF^{+}07, ML85, MV86, PR83, PL95, Sch95, SBB86, Woo73, Yue85]. Processing [DDM^{+}18, KPS^{+}16a, TZZ^{+}18, VTGH17, WYD^{+}21, ZFP^{+}21, AD99, AD00, Bas72, BJL^{+}06, BP91, CPW07, Cas91, CFR98, Cri94, DB96, Ful73, GB93, GP95, KKFB11, MLB83, Mil90, Oli90, Sop84, Svi83, Tug83, VBLM07]. Processor [GCJ17, KTG^{+}17, Kru82, SKJ^{+}17, ZSG^{+}17, AM87, Bas72, BJL^{+}06, BB75, CLC05, Cla87, EKM04, HS91, HF08, LKB91, LBvH06, NL95, RK11, SL98, SSS01, SDV^{+}87, ST00, VZ91, WCW^{+}04, vdWMH11]. processor-based [WCW^{+}04]. Processor-Interconnect [SKJ^{+}17]. Processors [AWS16, ALM^{+}18, CDY^{+}17, MT17, ATSS09, FJLC98, GCTR08, HZ09, MA06, MSF85, PRAH96, RGAB98, RPNT08, SCL96, SF91, SKPG01, SPR00, WZWZ10, WL09]. Procrastination [PG16]. Procrastination-Based [PG16]. produce [Cri94]. producer [Hil92, RB75, Rus77]. producer-consumer [RB75, Rus77]. producer/consumer [Hil92]. product [KGS06]. product-line [KGS06]. Production [ZJL17, TLH^{+}07]. Productivity [Wit16]. Professional [Bar14, Gra14, Tet14]. Profile [UJE^{+}22]. Profile-Guided [UJE^{+}22]. profiling [ABD^{+}97, CL87, CCZ07a, DB00b, HC04, KEP07, USR02, ZWG^{+}97]. Program [BS15, Fle07, KTG^{+}17, VSST16, BSL08, DV87, DK75, ELG95, GMM98, GN96, Isa08, Lov77, Mas87, Mas77, MCC^{+}06, RR72, RD87, SV06, SMTZ09, SPHC02, SLZD04, TPO06, Won93, XFO08, ZZNM01]. programmable [EKO95b, NMS^{+}00, WDA^{+}08]. programmed [MSR77]. Programmer [Wit16, SGN85]. Programming [AWS16, BBB^{+}17, BS15, BD17, CKmWH16, EMS09, EMSPS11, EMZ^{+}16, HCW^{+}04, KMC02, LL16, Wai94, Zho16, AUS98, AH77, BC91b, BF87, Bos06, CM87, CAL^{+}89, CBC^{+}08, Den74a, Den74b, DBMZ08, DMB87, Dos88, EFL07, Fou74, Fra95, FW77, Gan77, GA98, GCTR08, HPM93, HFWZ87, Her77, HEKSP11, LCWM08, OB86, QPP02, Ric88, Ros95, Taf82, TMW10, Win08, Wir77, Wir87, Won93, Zel74]. Programs [JCY^{+}19, NP17, SLFP16, AGB^{+}77, AL91, BAMM77, BM91, BH81, BMP^{+}04, BB75, CLR94, Coo85, GTA06, Gue87, HS96, IBY^{+}07, KCLZ98, MP85, NAR08, RG02, RK83, SBN^{+}97, SP00, Shi00, SLTB^{+}06, Wag98, Yue85, ZL86, Wei92]. Progress [DB99, Laz92b, MLB83, WS92]. Progress-based [DB99]. project [AD07, AMO^{+}12, BBH^{+}00, BC91b, BDH07, HPG00, MLB83, Nee77, SMS11, SCS77, AUW08, Neu92, Pas92]. projects [KS92]. Prolog [BCDN87]. promise [Bir07]. promote [WK05]. pronged [Rob08]. Proof [Hof90, How82, AB82]. Proofs [EPG^{+}20, SAL20]. propagation [LRS^{+}08, PST^{+}97]. Properties [DS72, ZSG^{+}17, BH75, BH81, Buc77, CC77, Hol72, TFC99, XZZ97]. property [BC83]. PROPHET [WL09, CYG^{+}17]. proportionality [GBG^{+}10]. proposal [GP05, Kno74, Kno75, LK08, Rou84]. proprietary [VE08]. ProRace [ZJL17]. prospective [OB86]. prospectus [NN75]. protect [WK05]. protecting [JS08, KJS^{+}06, LJY04, PGZ08, ZJS^{+}11, ZZP04]. Protection [AYQ^{+}16, CJ75, Dru92, Gai75, Hog88, HM93, Lam74, MMT16, Oli90, Rip03, Sal73, Var97, WFHJ07, Wel88, AH77, CGL^{+}08, cCVP99, CVP00, Coo78, HRU75, HvE02, Her78, HFC^{+}06, Les04, LC93, NW77, Nes82, O’S92, Sal74, SS72, SCP^{+}06, Sny77, TSLBYF08, WCA02, WRA05, Bis81]. Protectit [KSLA08]. ProteusTM [DDK^{+}16]. Protocol [MB93, BO99, CC97, CCK04a, CC04, CC05, Che04, CWL05, CCEH00, DDYM99, Das92, GB93, GP95, KTH89, KC95, KLS08, KSL92, KTC03, LCTK01, LKKY03a, LKKY03b, LW04, LH04, LHL04, LSH03a, MY98, PG96, PCP00, PFGD02, Syv96, WYC03a, WK05, WL94, YS02, ZWWL01, ZL04b, KvRvST92, LSH03b]. Protocols [Ng99, ADG^{+}07, ABC^{+}98, BBFH07, Bir07, Boc75, CK86, CH07, DB75, HB06, JW01, LSH01, MP75, ML85, PS98, SHT97, SS94, SM89, SW00, Syv93, Toi92, XZZ97, XZZ98, ZLX99, ZL04a, ZIL96]. prototype [ZG07]. prototyping [WBC^{+}83]. Provable [HMK20, VMM20]. provably [ZLX01b]. prove [TFC99]. Provenance [MRS09]. provide [BC06, SLQP07]. provider [BWV^{+}12]. providers [BK12, SG10a]. Providing [BDS^{+}09, GC08, ST93, Nut94b, TS06]. Proving [BH81, FLR77]. Provisioning [DK16, AC06, Edi13, GSM08, PPO14, WL09]. proxy [RCC01, SFW99, Son05, WVS^{+}99, WVS^{+}00]. proxy-based [RCC01]. Prudent [PG16]. PSI [TNNI87]. Public [ELR15, CJ05]. publications [Mat10]. Publisher [Wai97b]. Publishers [Lit87]. Publishing [Had84, Had85, San86, Wai86, PP83]. Pulse [ZFP^{+}21]. pup [Fle81]. Puppeteer [dLWZ00a, dLWZ00b]. purging [BC08]. purpose [DC99, DC00, FIM^{+}11, GCTR08, TPO06, WH99]. PUT [HDH^{+}94]. PUT/GET [HDH^{+}94]. PVM [DDYM99, IvdLH^{+}00, JW96]. Pyr.mea.IT} [ACC^{+}08]. pyramid [ACC^{+}08, TNA12].

Q4 [Bel10]. QoS [BC06, CEV00, CYMT16, CYG^{+}17, GC08, GP95, Mal10, MCR^{+}09, WlL01, WL02, ZE16]. QoS-oriented [Mal10]. Quality [Gwi05, LSV^{+}19, PAM^{+}16, CEV00, KK84, McD00, NCL12, Neu00]. Quality-of-Service [Gwi05, Neu00]. quantitative [LST^{+}06, MT96]. quantum [SV06]. quarks [JLR^{+}05]. Quasi [DDYM99, ELG95]. Quasi-asynchronous [DDYM99]. quasi-optimal [ELG95]. Queries [JXG21, SMRD06]. query [GSA10]. Querying [CC21]. queue [AH80, WLRZ03]. Queueing [Wai83b, Sta83]. queues [Bas72, SCM05]. queuing [Gil78, Tri82, Tri02]. Quick [Wai97b]. QuickSilver [HMSC87, SW91]. quiescent [MR07]. quota [KLS08].

R [LHWY83]. R&D [BYVF08]. R4600 [LE96]. Race [ZLJ16, ZJL17, PK96, SBN^{+}97, XHB06, YRC05]. RaceTrack [YRC05]. Raft [HSMC15]. RAID [GC12, Jin99, KBPM10, OCLN14]. RAID-based [OCLN14]. railway [BRR^{+}00]. Raj [WP91]. Ralph [Bla95, Wai97b]. RAM [Riz97]. Rambus [MSP98]. RAMClouds [OAE^{+}09]. RAMpage [MSP98]. Random [ZS06, MCM07]. ranking [Dou09]. RAPID [AWS16]. rapidly [AIKS00, BSM^{+}12]. rate [BMR^{+}09, UHMB94]. rather [Sat00]. rating [Dou09]. ratio [ZPS^{+}04]. raw [LBF^{+}98]. RCU [MFBWW20]. RDR [CLC05]. re [DSBK04]. re-execution [DSBK04]. Reaching [WYC03b]. Reactive [LA94, MW92, MW91]. read [EM89, Joh91, MMGC02, WL82]. read-mostly [EM89]. read/modify/write [WL82]. read/write [MMGC02]. readahead [WXX08]. reader [Sei90]. readers [KL98, KPL99, Küh99]. ready [SLCG89, SSR^{+}10a]. Real [BHD19, GF15, PS01, SZG91, TL96, Zha23, AGM93, BL75, BH81, BC91b, BC01, CMMS77, DRSK89, DS92, FPG89, GP95, GS89, Gup01, HLFZ97, KKS89, LTCA89, LSA^{+}00a, LSA^{+}00b, LPSZ08, MW91, Mil92, MPC^{+}02, NMS^{+}00, NCL12, OT95, PN00, PC75, RLB08, RPM97, SN94, SZ92, Sor73, SR89, TM89, Wai95a, WAB^{+}89, WPC12, Wir77, YS98, Zea97, ZPS99, ZPS00, FPG89]. Real-Time [GF15, BHD19, PS01, SZG91, TL96, AGM93, BL75, BH81, CMMS77, DRSK89, DS92, GP95, GS89, Gup01, HLFZ97, KKS89, LTCA89, LSA^{+}00a, LSA^{+}00b, MW91, Mil92, NMS^{+}00, NCL12, OT95, PN00, RLB08, SN94, SZ92, Sor73, SR89, TM89, Wai95a, WAB^{+}89, WPC12, Wir77, Zea97, ZPS99, ZPS00]. REAL/IX [FPG89]. realities [SPBP06]. Reality [Wit16, Wet99, Wet00]. Realize [WlL01]. reallocation [Ger72]. Realtime [Gwi05]. rearranging [KT91b]. Reasoning [FVDS20]. Reassignment [WM16]. ReBudget [WM16]. receiver [DB96]. Recipient [Bar14, Gra14, Tet14]. Reclamation [PG16]. Recognition [AWS16, KKM^{+}06]. recognizers [LOM^{+}09]. recollections [Dij05]. recommendations [MPP^{+}08a]. Recommending [VJ19]. reconfigurable [RA06]. reconfiguration [RS91]. Reconstructable [RHMR15]. Reconstructing [GFPcF08, KTG^{+}17]. reconstruction [Jin99, VM07]. Record [MGT^{+}17, LWQ09]. Record/Replay [MGT^{+}17]. recorder [LBP^{+}07]. Recording [NPC06, XHB06]. Recoverability [MF75]. recoverable [SMK^{+}93]. Recovering [VM07, RK11]. Recovery [GB01, HMSC87, VTGH17, AM85, Bac91, Boc75, CW92, CJ05, COS^{+}08, DB85, Dim98, GPF^{+}05, GLL04, JSDG08, JZ91, KBB^{+}06, Lei89, Lom77, MSF85, PW98, Wal73, Wei85, ZWZ01, ZWZ05]. Recursive [SSK17, BH75, FHL^{+}96, LM96]. redesign [CHV04]. redirection [WPP02]. REDSPY [WCL17]. Reduce [JHK^{+}16, BSL08, ECS73, SS94, Ste97, WLRZ03]. Reducing [BFZ^{+}25, CG94, JFV^{+}96, KT91b, KS95, LGSN89, SPHH06]. reduction [HCJ07, XHB06]. redundancy [FES09, Rom93, YW06]. redundancy-based [YW06]. Redundant [O’S92, PSG06, RRP06]. Reed [RD97]. Referee [Pet76]. Reference [MCXS16, AKS73, EGE02, PLH98, SZ98, Wol02, Wai97b]. referenced [BJ81]. refill [JM98]. Refinement [STW95, BR09, GBZP10]. Reflection [OT95, Str93]. Reflections [Bar14, Gra14, Sil83, Tet14]. ReFlex [KLK17]. Refloated [HSMC15]. Refresh [KSCK17]. Refresh-Aware [KSCK17]. Regenerating [JKL^{+}13]. region [KS82]. Regional [YG25]. Regions [PP09]. register [BEH91, MSAD91]. registers [Che84]. regression [BDDMR11, LB06]. regular [Ant90]. regulated [XHB06]. regulation [DB99, DB00a]. Regulator [BLI17]. Reimplementing [Hag87]. Rejoinder [BAN90]. related [LPH^{+}07, Smi78, LaR92]. relation [BSF^{+}91]. relations [DFS00, FR94]. relativistic [TMW10]. Relaxing [Pu93]. release [CGS^{+}96b, SLM11, ZIL96]. Reliability [CN07, GS13, Wai83b, BSR^{+}06a, Gan77, GPK^{+}07, HL92, KBPM10, LNBZ08, MSR77, OL02, PWC^{+}81, Sta83, SSR^{+}10b, Tri82, Tri02, WK05]. Reliable [HBG^{+}06, LB91, LGN07, OLS85, ABC^{+}02, BVR^{+}00, FAH^{+}06, KTH89, LKvR^{+}99, LKvR^{+}00, MZWZ02, MW91, MPHD06, Oes91, Oes01, ÖGA06, PP83, SS83b, SDD^{+}85, Svo81b, Van06, YWC04, ZLL^{+}07]. Relocation [VSST16]. Remaining [NTHAB22]. Remarks [CL04c]. remedies [Aba93]. Remote [FLM^{+}08, KLK17, KMN^{+}16, ATK92, BALL89, CCG95, CL04c, HL05, KC05, Led97, LHY02, LLH02, Spe81, SDH^{+}97, TA90, TLC85, BN83]. remote-write [SDH^{+}97]. remotely [KL07]. Removal [SHP^{+}16]. Removing [Del80, LMG^{+}07]. rendezvous [Hil92]. RENS [AIKS00]. repair [GBZP10]. Repairing [HBG^{+}06]. Repeatability [DL15, Eid15, Fei15]. Repeatable [AHB15, ELR15, RWS^{+}15]. repeated [Mat04, Syv93]. repetition [SS98]. replacement [Kai75, MPC08, NAR08, Sad75]. Replay [MGT^{+}17, DKC^{+}02]. replays [KS99]. replica [MNP07, ZXMJ04]. Replicated [Coo85, Coo86, GHP92, BDF^{+}15, DB85, DGH^{+}88, EDZ07, Fra95, LGJS91, LWQ09, LAB^{+}06, RB93, SD86, TTP^{+}95, WB86, Yu00a, Yu00b, YV01]. replicates [Bre83]. Replication [Bir85, LGG^{+}91, ZSS08, EDP06, GS95, Her86, HHS05, LMV12, LLS91, PST^{+}97, SKKM02, SAG06, YVM13, dSFdAM13]. Reply [How82]. Report [And83, Bab91, Bac99, Bel10, CvR14, CM14, DNT10, HN12, HKPvR16, Isa08, Lev88, Mul87, SN13, Tan97, Ter14, Voe98, And87, BK08, Kah85, MvR13, MLB83, PGS93, Sch73b, SK13, Ano86]. reporting [CCM08]. REPOS [MA79]. repositories [SW10]. repository [HSK97, Svo81b]. Representation [Che75a, HS16, RN00a, RN00b, Gir82, Gór78, VT01]. Reproducibility [Fei15]. Reproducible [SLD15, Boe15]. repudiation [HLL04]. reputation [DY10]. Request [EAS^{+}17, Cha96, LG04, PAB^{+}98, PKB^{+}08, SZD^{+}08, WPP02, dSFdAM13]. requirement [TL96]. Requirements [CDY^{+}17, HS96, JT90, PG73]. Rescue [SJS^{+}23, BW95]. Research [BMF^{+}16, LaR92, Rat11, RWS^{+}15, Sal78b, SLD15, SG14, Ten17, Wai83a, BYV08, BKP^{+}12, Boe15, Bor92, CR12, DKW^{+}06, DVS12, EAS07, Est02, FBB^{+}97, GNB^{+}09, Her10, Lam00, Lev07, LLY05, MLB83, Mat04, MW09, Moh78, NSW10, PGS93, RPNT08, Sal74, Sop84, SPBP06, WCL^{+}04, ZUW^{+}09, HMS17, Sch07]. Reservation [TLL03]. reservations [JRR97]. Resilient [ABKM01, LRS^{+}08]. resistant [QPP02, TML^{+}00, YS02]. Resolution [ZK88, Bre83, DM90, HXL01, New79, Spr85, ZL86]. Resolving [Loe85, ES10]. Resource [Cra83, DK16, GB90, KEF^{+}19, PPM17, UHMB94, USR02, WH99, WM16, AC06, BL00, CJM^{+}75, CKR08, EKO95a, FS95, Fon72, GTHR99, GTHR00, GA08, LB81, RS00, RA07, ROLV06, SLM11, SK96, STYC02, VFMM08, Wal02, YGG^{+}03, ZELV02]. Resource-Constrained [KEF^{+}19, RA07]. Resource-Efficient [DK16]. resource-release [SLM11]. resources [AYK08, AS10, CAT^{+}01, Edi13, GG73, Lev03b, PSZ^{+}07, WC02, dGdB10]. Responding [BSM^{+}12]. Response [Hil94, Bir94, CM75, Den07]. responsiveness [WGL^{+}08]. restart [BBHL08]. restartable [SSR^{+}10b]. Restore [RS02, BW01]. restoring [KBB^{+}06, XHJB99]. Restricted [Buc77, HK00]. restriction [MPC08]. results [RD97, WH99, ZK88]. resynchronization [RB75]. retargetable [EP94]. retention [ZLL^{+}07]. Rethinking [DRTT24, Ott18, HL07, KBPM10]. retrieval [TL96]. retrieving [CZG^{+}05]. retrofitting [CGL^{+}08]. retrospective [BDH07, Wil09]. Reuse [JXG21, CHCmWH00]. Review [Bla95, Had83, Had85, Had93, Heu97, Kad95b, Kad95a, Lig94, Lig95, Lit87, Nut94a, OSV86, San86, Sta83, Tug83, Val94, WP91, Wai86, Wai94, Wai95b, Wai97a, Woo85, dV96, Bec75, Had84, Mat06, Mat07, NRS13, OSV82, Wai98, Mat10]. reviewing [And09]. Reviews [Wai83a, Wai83b, Wai97c, Wai97b]. ReVirt [DKC^{+}02]. revisited [Cas95, Jon80, Loe89, NS87, Wet78]. Revisiting [DHK^{+}15, HMS17, WY04, GKS11]. revocation [CV93, CM06, Var97]. RFID [CH07]. RFS [DZP^{+}11]. rich [LJW^{+}06]. Rico [Sat99]. RID [MCXS16]. Right [MSC^{+}06, Den74b, HUL06]. Right-weight [MSC^{+}06]. rights [O’S92]. Rigor [PWT^{+}19]. RIMAC [YW06]. rings [SS72]. Rio [Sat99, CNC^{+}96, LC97]. RISC [HO91, BC91a, BSUH87, Kie87]. RISCs [BCDN87, BEH91]. risk [Gon92]. risks [LGSN89]. Ritchie [vR14]. road [KBB^{+}06]. Roadmap [CJM15]. ROADS’09 [DO09]. Robert [Wai97c]. robust [ADG^{+}07, BCRS10, Mit00, SKPG01]. robustness [LEK91, MCM07, WPP02]. Role [SHV01, CLC05, Dou93, HH08, KKC02, MSR77]. Role-based [SHV01, CLC05, KKC02]. Roll [GB01]. Roll-Forward [GB01]. rollback [CW92, JZ91, ZWZ01]. Rome [Wil09]. Root [Küh04]. rootkits [WG08]. ROSY [RK11]. RouteBricks [FIM^{+}11]. router [KMC02, MKJK99, MKJK00, MSB^{+}02, PN00, SKPG01]. routers [KAI^{+}13]. routing [CDG^{+}02, NPB06, SCG01]. Rover [JdLT^{+}95]. Rowhammer [AYQ^{+}16]. RPC [Fes07, FHL95, PHOA89, SB89, SADAD02]. RT [SJL^{+}87]. RTR [XHB06]. rule [BP91, CM75]. rule-based [BP91]. Rules [How82, AB82]. run [ACT94, FL77, GPV04, NL96, TLH^{+}07, YD02]. run-time [ACT94, DCZ96, FL77, NL96, YD02]. Running [Gue87, ZL04b, BDR97, NAR08]. Running-mode [ZL04b]. Runtime [MAHK16, NG09, WM16, XX00, ESB^{+}06, GSA10, Le98, WDH89]. Russell [Bla95]. Rust [BBB^{+}17]. Rx [QTSZ05]. Ryu [KCC05].

S [Sta83, Wai83a, Wai83b, dV96, CG85, MC96, VHAA25]. S. [Küh99]. S/KEY [MC96]. S/Net [CG85]. SaberLDA [LCCZ17]. Safe [NL96, cCVP99, CVP00, CLDA07, QTSZ05, WKL07, GA98]. Safe-Tcl [GA98]. Safety [BBB^{+}17, BSP^{+}95, DBMZ08, HAF^{+}07]. Sage [LSV^{+}19]. Sampling [Ser21, BEL^{+}00, ZS06]. Samurai [PGZ08]. satellite [CC05, HYS03]. Saving [DL15, HD12, XHJB99]. saving/restoring [XHJB99]. savings [YN12]. Scalability [Acq16, KMK16, RHR^{+}17, VYW^{+}02, VMC^{+}05, GTSS11, SATG^{+}07, SJ05]. Scalable [DSBK04, Dub00, GPY^{+}17, HJ10, HNK^{+}17, LCL^{+}16, LX00, NPB06, NP17, RLD^{+}17, Ser21, TMW10, AEMGG^{+}05, AMS^{+}07, BMBW00, BCRS10, BDR97, CKA91, EDP06, FGC^{+}97, Gup05, JZZW02, LL98, NLO95, OAE^{+}09, PP06, PNT06, RD12, SBL99, SBL00, SPF^{+}07, SG10b, TML97, TNA12, Uhl07, WCB01, WA09, WS06, JAvR06]. scalar [WS87, ZCSM02]. Scale [CYMT16, CYG^{+}17, CJRV15, HKM^{+}87, RAVC12, WTC09, WSG02, XDM^{+}18, AUW08, BS95a, Bod11, FES09, GWSY08, GBBL85, Gor06, HSS^{+}06, KKFB11, KSS^{+}96, KBC^{+}00, LJX97a, LGN07, RRBN09, ROLV06, Ros89, RD01, SATG^{+}07, SF12, SPHC02, TLD^{+}11, VYW^{+}02, WHZ^{+}17, WVS^{+}99, WVS^{+}00, YZZZ06]. scale-out [GWSY08]. Scaling [PTBD16, RCSW10, GS13, KTB12, MCD^{+}08, PS01]. scanning [KPS09]. scans [WBR^{+}12]. Scenarios [BCR^{+}14]. Schedule [SCFS98, AVZR11, BFD97]. Schedule-independent [SCFS98]. Scheduler [ABLL91, DC99, DC00, GP05, GGV96, KTB12, NL95, NL97, PM03, RR72, SFB^{+}09, ST01, WTKW08, YVM13]. schedulers [GLG93, RS00]. Scheduling [CDV^{+}94, CR75, ECS73, GA91, JW96, KSCK17, LLK96, SB78, SLCG89, AB75a, Bas72, BRR^{+}00, BDF^{+}08, BC10, BEH91, BM90, CAW08, CM75, CNO^{+}87, CCB^{+}06, CKR08, DC99, DC00, Dun91, ET05, FS95, FS96, FJLC98, Ful73, GG73, HS91, Han72, ID01, JRR97, JLZx90, KL98, KPL99, KSS73, Küh99, LS75, LBF^{+}98, LSA^{+}00a, LSA^{+}00b, LX00, MSAD91, MDR^{+}00, MSP^{+}06, MB08, Mil92, NSKS11, OA08, PEA^{+}96, PKB^{+}08, Sto07, TDM12, TAS07, TP72, TL96, TG89, Ull73, VBLM07, VZ91, WBR^{+}12, XX00]. Schema [SRS22, CWL05]. Schema-First [SRS22]. Schematic [Var97]. scheme [CKA91, CL04a, CL04b, CL04c, CCK04b, CHY05, DSGP05, DD12, HYS03, HLL04, HL05, KLY03, KCL03, Ku04, KC05, KCC05, LHY02, LLH02, LL04, LKY04, LJY04, LLH04, LC04a, LM97, MC91, MC96, Sco04, WK05, YW04, YRY04, YbJf04, GB01]. schemes [VA96]. Scholarship [Bar14, Gra14, Tet14]. Science [ELR15, Lit87, RWS^{+}15, San86, Wai83b, Spi74, Sta83, Tri82, Tri02]. scope [Ano75]. Scout [MP96]. SCR [XHJB99]. Screening [ACS15]. scribe [Bre08]. Script [FH85]. Scripting [KKK^{+}17]. SCSI [VFH98]. Sealing [HAF^{+}07, Gif81]. seamlessly [HK00]. Search [VPH^{+}15, CWdO^{+}06, LJW^{+}06, SG05]. Searchlight [BTK11]. Sec’83 [Fêa83, Had84]. Sec’84 [San86]. Second [San86, Cab90, Had83, CM14, DNT10, Her92b, Mul87]. seconds [PBR^{+}08]. secrecy [Gif81]. secret [CHY05, JY98]. secret-key [JY98]. secretary [And81, Den78, Den79, Den80]. secretary-treasurer [And81, Den78, Den79, Den80]. section [Nai93]. sector [LSKK08, Lon93]. Secure [AMH^{+}16, BJKT15, CDG^{+}02, CLM^{+}07, CLDA07, LWY^{+}04, PL01, SLZD04, VPH^{+}15, WF07, BDS^{+}09, CC04, CCK04b, EKO95b, HC95, KC95, Kil00, Lac00, Lan89, LLH04, LC04a, LWMX05, Loe85, NBK^{+}20, Pop75, Rus81, SF12, TKP^{+}08, WECK07, YRY04, ZZNM01, Zim94]. Securing [BK12, CCZ^{+}07b]. Security [CH07, CDG^{+}17, Fêa83, FXZ^{+}17, GA98, HSK97, KXD00, Lan89, LSH03a, LSH03b, Lit87, Ng99, PS99c, Pat02b, Rei92, San86, SK97, ZSG^{+}17, AFB95, BTMS10, BCP^{+}08, DS90, DY01, FLR77, FM98, GS78, Gon89, Gon92, GC05, Had84, HK99, Hog88, HC95, JL75, LJY04, LNBZ08, MKKW99, MKKW00, NPCF08, PS99d, PS99a, PS99b, PS99e, Pat02a, Rou84, RN00a, RN00b, Sch75, SK13, Sil83, SPHH06, Sin85, SH00, WBDF97, XZZ98, YW04, dVdVI98, Fêa83, Had84, ZL04b]. security-sensitive [SPHH06]. SecVisor [SLQP07]. SEDA [WCB01]. sedition [Bak95]. see [SB10b]. Seeing [MZI08]. SEEP [HEKSP11]. Seer [GZH^{+}19]. Segment [MP75, Son05, Rob96]. Segment-based [Son05]. segmentation [cCVP99, CVP00]. segmented [Tan79]. Sego [KDL^{+}16]. Segregating [SZ98]. selected [VZ14]. Selecting [CM75, Tom75, Dal75]. Selection [CKmWH16, SMTZ09, Var72]. selective [ACM02, DSBK04]. Self [HBG^{+}06, dOL12, RF17, BC06, CJ05, CM06, DY10, Edi13, HSS^{+}06, NXQ05, PJDL06, SRH^{+}06, Wal73, ZS06]. Self-aware [RF17]. self-certified [CJ05]. self-describing [SRH^{+}06]. self-infrastructures [HSS^{+}06]. Self-management [dOL12]. self-managing [NXQ05]. self-organizing [BC06, CM06, PJDL06, ZS06]. self-provisioning [Edi13]. Self-Repairing [HBG^{+}06]. self-stabilizing [DY10]. SelfTalk [GSA10]. Semantic [GJSO91, HABZ17, KLS^{+}10, MPLH06, LPH^{+}07]. Semantic-Aware [HABZ17]. Semantic-less [KLS^{+}10]. Semantics [HZCC97, BSL08, BS96, LLS91, WBB02]. semaphore [AH80, HS88, WL82, Kos73]. semaphore-queue [AH80]. Semaphores [Cas95, Dun91, Hem89, TT00, DD80, Hem88, Hil92, Hsi89, Kea88, Kot88, TC96, Tro00, Xu00]. semaphoring [OS80]. seminar [SK13]. sender [BJM^{+}96]. sender-managed [BJM^{+}96]. sense [Bak95]. Sensing [DIS19, LJdL^{+}16]. sensitive [DC99, DC00, GAK^{+}02, KSLA08, SPHH06, ZJS^{+}11]. sensitivity [KKC02]. Senslide [STM^{+}07]. sensor [AJG07, BBD^{+}02, EKM04, HSI^{+}01, HSS^{+}06, LMG^{+}07, LC02, MFHH02, MAK07, Est02]. sensors [HSW^{+}00]. separates [LJX97a]. Separating [MKKW99, MKKW00, TLL94, Les04]. separation [LCC^{+}75]. September [San86]. sequence [Dal75, Tom75]. sequencers [RK77]. Sequencing [HN81]. sequential [CGS96a, IBY^{+}07, LSP07].

Sequoia [Pas92]. serial [AAMV09, Mit00]. serializable [Pu93]. Serialization [GMT16]. Series [Wai83a, SF80]. Series/1 [SF80]. Server [SWC08, CAT^{+}01, DB96, Eri14, Hal00b, Hal00a, HKL^{+}06, HCG^{+}06, yKPR02, LHL04, LZJ03, LLS^{+}08, Mal10, NMS^{+}00, PBH^{+}07, RN83, RAF07, Wal02, WB07, Dio80, FO81, GN80]. server-less [HKL^{+}06]. Serverless [ADN^{+}95, FZY^{+}23, SJS^{+}23]. Servers [SKJ^{+}17, WL15, BHLM94, BBHL08, CSBA17b, CDV^{+}94, CGM97, EKF^{+}14, HPG00, JKH^{+}00, MD81, Nee79, PAB^{+}98, Son05, VDGR96, WCE^{+}92, YZJ02]. Service [Gwi05, Hof07, AVZR11, BMTW91, BSM^{+}12, BHB^{+}08, BACF08, CZB^{+}09, DW07a, DW07b, EBS01, FC87, JZZW02, LEH86, MFHH02, MB93, McD00, MT85, NCL12, Neu00, Oes91, RCC01, RA07, Rom97, SBL99, SBL00, SZN87, WS06, vR92]. Services [Had83, JHK^{+}16, KDL^{+}16, Woo85, AEMGG^{+}05, AIKS00, AAC^{+}05, Arn10, AC97, BFHW75, BDS^{+}09, BCE^{+}95, BCC^{+}13, CMK^{+}06, DHRS91, Fle81, FGC^{+}97, GBZP10, GBCH00, Gue88, HBP06, KSLA08, LLS91, LZJ03, LAB^{+}06, MA11, MDB01, PS99b, SJL^{+}87, SF12, SAL20, STYC02, Wai95a, WCB01, Yu00a, Yu00b, YV01, ZBN07]. Serving [BKGT25]. Session [Bre08, LE00, Tsa16, Bak95, Cec00, CG00, Kil00, Lam75, Laz92a, LSA^{+}00b, PCP00, RN00b, ST93, Yu00b, dLWZ00b]. sessions [BHJ^{+}93]. set [Cha96, CKDK91, COS^{+}08, DV87, DS72, Fog74, OS80, SKI08, Ste97]. sets [MNP07, Mar97, Pot77, SS83a]. setting [FM02]. Seventh [Sat99, Tan97]. several [JM98, TSF90]. SFT [PNT06, WJ98]. Shadow [CCS^{+}16, Isa08]. share [CAW08]. Shared [Bal24, DHRS91, DK16, ELR15, JW24, KEF^{+}19, ZE16, AMMR92, BCRS10, BMP^{+}04, Cec00, CLR94, CRD^{+}95, Che85, Col73, CGS^{+}96b, DCZ96, ENCH96, Esk96, FP89, GGH91, GTHR99, GTHR00, HGDG94, HSPC01, JKW95, KLMO91, LJX97b, MBD^{+}12, NPC06, Nic87, PRAH96, RGAB98, Ros89, SGT96, SG97, SFL^{+}94, SDP^{+}00, SJGY94, SKI08, SDH^{+}97, TSF90, TWL05, TG89, USR02, VZ91, VGR98, WMH72, WBR^{+}12, WS06, YZG^{+}11, ZIL96]. shared-memory [CLR94, CRD^{+}95, GGH91, GTHR99, GTHR00, KLMO91, PRAH96, RGAB98, Ros89, TG89, VGR98]. Sharing [BFHW75, BKGT25, Eid15, AC06, BBMT72, BEW75, BEW76, CHY05, Chá91, CJM^{+}75, FW72, GC08, Gre72, HS91, HKL^{+}06, HSPC01, Mon77, PM03, RT73, Sal73, Sha00, TAS07, VGR98, Wei98, WTLS^{+}09]. sharing-aware [TAS07]. Shasta [SGT96]. shelf [MSC^{+}06]. Shen [KTC03]. shepherding [GPK^{+}07]. Shipping [And95]. Short [Han72, LCL^{+}16, CPT08, Lie94a, Lie94b, Lie95c, Lie96]. Short-Lived [LCL^{+}16]. Short-term [Han72]. shorter [WJ98]. shortest [Won93]. Should [Gur07, HBB13]. Shredder [AMH^{+}16]. Shredding [AMH^{+}16]. side [KL07, MMTW10]. Sidewinder [LJdL^{+}16]. SIGMETRICS’98 [Voe98]. SIGMETRICS’98/PERFORMANCE’98 [Voe98]. Signal [FVDS20, SCL96]. Signaling [CKMV99, BMR^{+}09]. signature [CL04a, CJ05, HLL^{+}02, LLH04, PSB06, WK05]. signatures [BMW02b, BMW02a, PKM^{+}09, Sal78a, SZD^{+}08, TACT08]. signer [WK05].

Signet [PS09]. signing [WK05]. SIGOPS [Bab91, Bac99, Mul87, Sha95, Tan97, WTC09, And83, Bir91, Kah85, Lev88, Lev90, Mog08, Sch73b, vR14]. SIGPLAN [Sch73b]. SIGPLAN/SIGOPS [Sch73b]. Silent [AMH^{+}16, LL02]. silver [KSDC14]. SIMD [PJDL06]. similar [BC91a]. similarity [LJW^{+}06]. Simon [Woo85]. Simple [BFS89, EPG^{+}20, NTHAB22, SS94, YS02, AH80, BJW87, CH81, GS90, KTC03, LKKY03a, LKKY03b, LFW04, Oes91, Ray91, War76]. simplification [FS08b]. simplified [KS85, PSC^{+}07]. simplify [GBCH00]. SIMs [PS09]. Simulated [GKO^{+}00]. Simulation [Fra80, JBW^{+}87, LLSK24, WMH72, AFF^{+}09, GKO^{+}00, LCTK01, Nut74, SL98, UNMS94, WPC12]. simulator [HD12, WZWS08]. simulators [LFH^{+}09]. simultaneous [REL00, ST00]. Sinfonia [AMS^{+}07]. Single [CBHLL92, Rus88, WHZ^{+}17, CIL93, KF09, LSS04, MS00, ONH^{+}96, OVS^{+}06, Ros94, SBB86, Str12, vdWMH11]. single-chip [ONH^{+}96, vdWMH11]. single-ISA [KF09]. Single-machine [WHZ^{+}17]. single-node [LSS04]. Single-user [Rus88]. singularity [FAH^{+}06, WYA^{+}07, HL07]. Sinking [CDG^{+}17]. SIP [AWW08]. Site [CK86, DBH^{+}06, LWQ09, TLH^{+}07]. Sixth [EMSPS11, Sha95]. size [DV87, LML00, THB06]. Sizes [ALM^{+}18, CB17, Sat81]. Skeen [Bir94]. sketching [SLTB^{+}06]. Slack [EAS^{+}17]. Slade [Wai97c]. SLAs [Bas12]. SLAYER [WAB^{+}89]. sleep [ZCT^{+}05]. slice [PSG06]. slice-based [PSG06]. SlicK [PSG06]. SLIM [SLN00, SLN99]. Slippery [SRTH15]. Slipstream [SPR00]. Slope [SRTH15]. Slotted [SKB^{+}17]. slow [BXS14]. SLRL [NTHAB22]. Small [Gor06, MC11, BJW87, JXG^{+}02, yL91, Lie95c, WH08, ZG07, ZPS99, ZPS00]. small-memory [ZPS99, ZPS00]. Small-scale [Gor06]. smallest [Mas87]. Smalltalk [BSUH87]. Smalltalk-80 [BSUH87]. smart [KLY03, WTLS^{+}09, LAB^{+}06, NL95, NL97, BJKT15, CL04c, CCK04b, HL05, Ku04, KC05, LHY02, Sco04, YW04]. SmartApps [RA06]. SmartFrog [GGL^{+}09]. Smartphone [HT15]. SMP [TAS07]. SMPs [MSA^{+}00]. SMT [GPV04, RPNT08, TAS07]. SMV [ZWWL01]. snake [KSDC14]. snoop [BSL08]. snooping [MSA^{+}00]. Snort [GC05]. social [HB06]. Socket [ZL04b]. SODA [KS85]. SODS [SF80]. SODS/OS [SF80]. Soft [AD99, AD00, BS15, LLK96, SGK^{+}04]. soft-error [SGK^{+}04]. SoftFLASH [ENCH96]. SoftSig [TACT08]. Software [Ano75, AYQ^{+}16, CKP91, CHV04, CHLS16, DB00b, DNT10, GG91, Had93, HN12, KC94, KSCK17, MKY08, MSR77, MA06, Ott18, Rom95, RHMR15, SN13, SBS18, TML^{+}17, Wai86, WCL17, ZH16, AA06, AD99, AD00, Bac81, BKP^{+}96, BMK06, CL87, CGKM11, cCVP99, CVP00, CZB^{+}09, CCZ^{+}07b, CKK^{+}07, DCZ96, FRL00, GKV07, Har82, HL07, JKW95, Kan83, KEP07, KDP02, KGS06, LRS^{+}08, Lie94a, Lin81, MSP98, MLB83, Mog06, Moo92, NN75, OL02, Pen09, QPP02, QTSZ05, Rou84, SGT96, SG97, SDP^{+}00, SKPG01, SH87, SDH^{+}97, Svo73, TLD^{+}11, TL94, TML^{+}00, TACT08, WLAG93, WBC^{+}83, Duc92]. software- [Har82]. Software-Based [AYQ^{+}16, MA06, Rou84, SKPG01, WLAG93]. software-exposed [TACT08]. software-only [SGT96]. software-profiling [KEP07]. Sol [Had85]. solitude [JSDG08]. Solution [SEF^{+}16, BAMM77, GWSY08, HS88, KL98, Küh99, LSH00, Sei90]. solutions [DS92, WB86]. Solved [Lam85]. solves [Rou84]. Solving [SRH^{+}06, GB01]. Sombrero [MS00]. Some [AEH75, EB78, GS78, GHM77, Gwi94, Hol72, MW75, TCH^{+}91, Hog88, Pow89, YS98]. Sons [WP91]. SOSP [Sub11, Bar14, Gra14, Isa08, MdS09, Tet14]. SOSP’99 [LE00]. Sound [CSBA17c]. Source [BMF^{+}16, BYV08, SFW99, SW10, Tan87]. Sources [DS92, SJ95]. sourcing [NBB09]. Space [CBHLL92, LBF^{+}98, NTC^{+}21, YN12, BMvdV93, CIL93, GN96, HHS05, Kep91, Lie94b, Lie95b, LLY05, LNBZ08, MS00, Ros94]. Space-time [LBF^{+}98]. space/time [LLY05]. SpaceJMP [EMZ^{+}16]. Spaces [EMZ^{+}16, SSK17, BMvdV93, ÏMC^{+}06, KGGK09, PPT^{+}93, THK95]. spam [CXMX05]. spanning [HK00]. SPARC [CKDK91, LKB91]. Sparsity [LCCZ17, Lie95b]. Sparsity-Aware [LCCZ17]. Spatial [BVCG04, CCB^{+}06, DBMZ08, MCC^{+}06, WCL^{+}04]. Speakers [Tsa16]. SPEC [CKDK91]. Special [Eid15]. Specialization [CCS^{+}16, XDM^{+}18, EBS01, KGS06, PAB^{+}95]. specialize [CWS06]. Specialized [BDK^{+}08, NS16]. specific [BCE^{+}95, CDY^{+}17, DBR09, JKDC05, KGS06, SP00, Shi00]. specification [AGB^{+}77, BAD^{+}11, BGHL87, Buc77]. specifications [BDM97, GHM77, NBK^{+}20]. specified [CM75]. Specifying [BKL^{+}16, WS91a, SWL77]. speculation [FJLC98, HWO98, MT02, RSEW04]. Speculative [JCY^{+}19, MT02, NCF05, ACM02, CG00, DS06, KAD^{+}07, OL02, ZCSM02]. speech [LOM^{+}09]. Speed [Val94, BVR^{+}00, COS^{+}08, Gur07, HRX08, Les04, MBD^{+}12, XXMC05]. speed-based [XXMC05]. SPEED08 [VW08]. speeding [Hal00b, Hal00a]. spend [AD07]. spent [CLR94]. Spin [SJGY94, Gil78, BSP^{+}95, BCE^{+}95]. Spin-block [SJGY94]. Spinlocks [KMK16]. spinning [KLMO91]. spite [DY10]. SPM [CV93]. spreading [CWS06]. spring [KN93, SR89]. Springer [Had93, Lig94, Nut94a, Wai94]. Springer-Verlag [Had93, Lig94, Nut94a, Wai94]. Sprint [CPW07]. Sprinting [FZL16]. Sprite [BO91, NWO87]. Spritely [SM89]. SPTF [LG04]. squeezing [WC02]. SR [XD17]. SR-IOV [XD17]. SSD [EAS^{+}17, KBPM10, OCLN14]. SSDs [DRTT24, Str12]. St. [vR93]. St.-Michel [vR93]. Stabilizing [DY10]. Stable [BJM^{+}91, NHM83]. stack [HL07, KPS09, MVKA06, PSMB16]. Stackable [Loe05, HP95, ZN00]. stacking [KDS^{+}06]. stage [CHY05]. Staged [CKK^{+}07]. stand [CR12]. standard [KYB^{+}07, Rus08]. standards [Had01, SG10a]. standards-based [SG10a]. Stanford [HGDG94, HKO^{+}94]. starting [SRH^{+}06]. State [Bel10, HT15, VSST16, GFPcF08, JT90, Mit00, Mou96, Rob96, SAL20, Spi74, Tur80, Wei98]. stateful [LAB^{+}06]. stateless [CJG02, SLN00, SLN99]. states [FR94, XHJB99, YM93]. Static [BNE16, CC77, FXZ^{+}17, MBS16, RN93, WHZ^{+}17, BBC^{+}06, RR04, YS94]. statically [ACM02]. statistical [HC04]. Statistics [Wai83b, EJD13, Sta83, Tri82, Tri02]. status [PGS93]. steady [Rob96]. steady-state [Rob96]. Stealth [CLS06]. Stefan [Kad95b, Kad95a]. step [Svi83]. Stephan [Lig94, Lig95]. steps [HN81, MM91]. STFS [JXG^{+}02]. stick [CMSK07]. still [SB10b]. STMBench7 [GKV07]. Stochastic [RLD^{+}17]. Stockholm [Fêa83]. Stone [Wai83a]. Storage [Acq16, BFZ^{+}25, BY08, BLC^{+}16, FFBG08, FKZ17, Ger72, GSCM16, GSW^{+}17, JSCM17, PBM22, RB24, RD01, VW08, ABC^{+}02, AKGR10, APGG00, BMTW91, BSR^{+}06a, BHB^{+}08, BXS14, CALM97, CN07, CR72, DKK^{+}01, DZ95, DBP^{+}04, DW07a, DW07b, DS73, FNRC^{+}07, FC87, GSGN00, GNA^{+}98, Gre72, GBG^{+}10, GA08, GSM08, Had01, Hal00b, Hal00a, HJ10, HF08, HGR07, HYM10, KSDC14, KBC^{+}00, LM10, LSKK08, LG04, MZWZ02, MA11, NTHAB22, OLS85, OAE^{+}09, PSMB16, Por10, RS08, Rob96, SGNG00, SFS13, SADAD02, Ste83, SLLP^{+}10, SCFS98, Str12, TTP^{+}95, TGR^{+}21, TNA12, VT01, WECK07, WMI^{+}07, WLZ03, WCL^{+}04, WCE^{+}92, WGSS95, WZ94, XXM04, YW06, CM87]. store [DHJ^{+}07, Del80, JZZW02]. stored [TS87a, WS92]. stores [LL02]. Storing [OB10]. story [JLR^{+}05, vR14]. straight [KS99]. strata [NPC06]. Strategies [JTG^{+}00, AS10, HD12, Mas77, PSK08]. strategy [CFL73, CM06, ELG95, HDL^{+}02, MM81]. Stream [JXG21, Bla83, DBH^{+}06, GTK^{+}02, GTA06, WS91b]. Streaming [HDGP21, VGX17, BD91, BMER14, YLE02]. streamlining [PAB^{+}95]. streams [BN78b, GCTR08, JH93]. Streamware [GCTR08]. Streets [WCYJ05]. strict [KCLZ98]. string [AKS73, VHAA25]. striped [HO93]. stripped [WCL^{+}04]. Strobel [Kad95b, Kad95a]. Strong [LSH03a, LSH03b, CC04, KTC03, Ku04]. Strong-Password [LSH03a, LSH03b, CC04, KTC03, Ku04]. structural [BM99]. structure [CSBA17b, CS77, CB93, GC96, KBK02, Lev90, Lov77, RLV^{+}96, Ste73]. Structured [Hil93, AAMV09, BS89, CDG^{+}02, Den74a, Den74b, Fou74, GHW07, Hat94, Hil94, JHT^{+}07, KAS^{+}06, LM10, MRC^{+}97, MP91, OCLN14, OD89, Rob96, Rom95, RO91, SK97, Svo81a, Zel74]. Structures [CSBA17a, Woo85, YWKYS15, EB78, GKD91, GBCH00, KB84, KGB88, LN79, LM96, RMS98, VL87]. Structuring [Fin92, MS91a, Met82, BHJ^{+}93, CL95, Cla85, Kee79, Lom77, Sal93, YTM^{+}91]. struggles [RRT^{+}08]. Student [SMS11]. students [AD07]. Studies [PS96, KLMO91, SPHH06, WMH72]. study [AH77, BCDN87, CYC^{+}01, CR72, CCAP06, DS92, DH10, DK75, DIN05, Fes07, GS90, HJT^{+}93, Kor06, KAI^{+}13, Lio78, LPSZ08, MW08, MCM07, MSB^{+}02, PSK08, PBR^{+}08, PK75, RF98, Ros78, Sat81, SG04, SMTZ09, SHSB75, WS87, ZWWL01]. Subcontract [HPM93]. Subgraph [JV21, MRH^{+}21]. subpages [JFV^{+}96]. subsidizing [LCJV^{+}11]. substrate [FBB^{+}97]. subsystem [DB96, Har88, HZCC97, WlL01, WL02]. subsystems [MR07, VT01]. successful [RD87]. Sudden [HT15]. Suez [PN00]. Suicide [CM06]. Suites [LWPG17]. summaries [LE00]. Summary [BCC^{+}94, BR10, EMSPS11, Ful73, Her92b, MM92, MM93, RAVC12, SBN83, Sha95, WTC09, BHJ^{+}93, Cab90, Lam00, SBL00, Sal00, TSE^{+}00]. Summer [DK17, HMS17]. Sun [DM90]. Supercloud [JSS^{+}15]. Supercomputer [BBH^{+}00]. supercomputers [VM07, WS87]. SuperDataNodes [Por10]. Superoptimization [CSBA17c, PTBD16]. Superoptimizer [Mas87]. superoptimizers [BA06]. superpages [NIDC02, TH94]. superpaging [Wis05]. superscalar [LKB91, SF91]. supervisory [Gai72]. Support [ALM^{+}18, BCC^{+}94, KKK^{+}17, KYP^{+}17, LER^{+}17, RF17, Tan97, Tur87, WPC12, ABLL91, AD99, AD00, AEP^{+}97, ATSS09, Bab91, BDMS98, BvS00, BBD^{+}02, Bir91, BF87, BMA00, CL87, CKD94, CHCmWH00, Coo94, CB95, CSS^{+}91, DBMZ08, DMB87, ESB^{+}06, FAH^{+}06, Fra95, GSA10, Gup01, HPG00, HWO98, HDH^{+}94, Her86, Hil81, JAvR06, KSP09, KSS^{+}96, LRV94, LMM93, LSA^{+}00a, LSA^{+}00b, MLB83, Moo92, NIDC02, OLS85, SV06, Svi83, Taf82, TH94, TCH^{+}91, TPH12, TL94, TML^{+}00, Tug83, VDGR96, WAB^{+}89, WK08, WDA^{+}08, ZWG^{+}97, Her92b, Jef92]. Supported [CJM15, MPP^{+}08a]. Supporting [Bla85, GAK^{+}02, KvRvST92, MCN^{+}17, MBM^{+}06, RRBN09, WS92, BBG83, DH73, DC99, DC00, Mad81a, Mad81b, McD00, Neu00, RK83, SGT96, SDE85]. surgery [Wal73]. Surpassing [TH94]. Surprise [SHP^{+}16]. surrounding [GA98]. Survey [AMPS73a, AMPS73b, AMPS74, Moh78, ATK92, Hac85, Nut94b, Smi88, TA90, Van96, Wel95]. survive [QTSZ05]. surviving [CNC^{+}96, SESS96]. Sustainability [SS17, Tai13]. Svigals.’ [Tug83]. SVR4.0 [Zea97]. SVR4.2 [Zea97]. swap [BJ81, HTW01]. swap-based [BJ81]. swapped [Pot77]. swapping [Bar79]. Sweden [Fêa83]. sweep [CHV04]. Sweeper [TNL^{+}07]. switch [LCH^{+}81]. Switchblade [FS08a]. switchboard [Zim94]. switches [MB91]. switching [San81, SMM^{+}09]. Symbiotic [ST00]. Symbolic [FVDS20, JCY^{+}19, HEKSP11, Kie87]. symmetric [BMA00, LH04]. symmetric-key [BMA00]. symmetrically [BMD94]. symmetrically-initiated [BMD94]. Symposia [Mog09]. Symposium [BOB15, BK08, Bre08, BM17, FBL^{+}12, Laz92a, Laz92b, LGMF14, LE00, OFB16, Pet76]. Synchronization [ACAAT16, AM85, Bel75, BGHL87, GMT16, Hab72, Hil92, Lam75, MCS91, PG16, RK77, Rya98, Rya99, Sco96, Cha73, Che75a, Eas72, EGE02, Ger77, GC96, LS86, LA94, Lom77, LT11, MNP07, MT02, PRD10, Ric88, Rom93, Ros89, SJGY94, SS06, Uhl07, Vog97, WPLP85, Woo73]. synchronized [Gon92]. Synchronizing [PR83]. synchronous [ID01]. synchrony [BDM97, BJ87]. synergy [GHW07, GC96]. Synopsis [Tsa16]. Synthesis [LWPG17, Bel75, MP89, Sny77]. syslogs [ME08]. System [AHC^{+}16, AVN^{+}16, BDMS98, BBB^{+}17, BOB15, BCC^{+}94, BLC^{+}16, BKL^{+}16, BMK06, BM17, DDOL16, FBL^{+}12, GF15, HBG^{+}06, Her92b, HSW^{+}00, Jef92, KDL^{+}16, LGMF14, LXYZ19, MAHK16, Mat10, MRH^{+}21, Neu92, OFB16, PBM22, Ram00, Rei92, WHZ^{+}17, ZWG^{+}97, dSM16, Wai83a, ARS89, AWSBL99, AWSBL00, AEE^{+}94, AKGR10, AGM93, Als72, ALBL91, And95, Ant90, AR07, AB75b, ACT94, AFF^{+}09, AEP^{+}97, ATSS09, BJ81, BFHW75, BO91, BHK^{+}91, BFSG94, BL89, BR09, BBD^{+}02, BAD^{+}11, Bas72, BL75, Bec90, BIYC06, BSP^{+}95, BCE^{+}95, Bir85, BC91b, BN78a, BPP12, Bla83, Bla91, BW95, BDF^{+}08, BBMT72, Bod11, BBG83, BS89, BMR^{+}09, BEW75, BEW76, BBH96, CE88, Car94, CPdM^{+}96, Cas91, CAL^{+}89, CIL93, Chá91, CB93, CNC^{+}96, CMMS77, CD95b, Che75c, CZG^{+}05, CJM^{+}75, CKR08, Cra83, CKK^{+}07, DKW^{+}06]. system [DRSK89, DH73, DH10, Del80, Dij05, DH96, DZP^{+}11, Dos88, DCZ96, DB97, EM89, EKV^{+}05, EBP16, EW76, EVvdW89, EWCS96, EP94, EKO95a, EKO95b, EJD13, Fab98, FFBG08, FO72, FLR77, Fes07, FS08a, FM98, Fle81, FMK^{+}07, Gai72, GSA10, GPF^{+}05, GFPcF08, GPV04, GS89, Gór78, GC96, GNB^{+}09, Gsc94, Gue88, GPK^{+}07, GKS11, Had01, Hag87, HPG00, HLL^{+}02, Han83, Har88, HdRC95, HO93, HvE02, HN08, HCZ97, HZCC97, HH89, Hop90, HKM^{+}87, How72, HXL01, JHT^{+}07, JM95, JKL^{+}13, Jon93, JLHB87, JLZx90, JXG^{+}02, KCD^{+}81, KKS89, KS85, KS92, KS91a, Klo80, KAS^{+}06, KSS73, KAR^{+}06, KLS85, KXD00, LM10, Lam83, Lan89, LRV94, LCKFA24, LN79, Lau81, LLA^{+}81, Les04, LMM93, LTCA89, LJX97b, LJS^{+}02, LRS^{+}08, LWMX05, Lio78, LGG^{+}91, LGJS91, Lit88, Loe85, LL98, LCH^{+}81, LK01, LB81]. system [LZ03, Mad81a, Mad81b, MA79, Maf94, MK91, Mah94, MR07, MO85, MKKW99, MKKW00, McD00, MW75, Mcd77, MXXC05, Mil78, MP81, MS00, Moo92, MMB96, MP96, MM91, MP91, MCM01, MMGC02, NIDC02, NW77, NB77, Neg00, NWO87, NLO95, Nes82, NB00, NBW87, NCF05, Nut74, Oli90, OVS^{+}06, OSSN02, Ous81, ODH^{+}85, PB96, PP06, PS99e, Pay77, PBR^{+}08, PKW81, PWC^{+}81, Pow77, Pow89, Pra86, Pra87, PC75, PAB^{+}95, RR81, REL00, RR04, RK83, Rip03, RT73, Rob96, Ros94, RO91, RBH^{+}95, RHP^{+}07, SKKM02, SFH^{+}99, SFH^{+}00, SHN^{+}85, SNV10, Sch73a, Sch95, SBN83, SGN85, SJGY94, SSF99, SSF00, STM^{+}07, SY96, Sil83, SF80, Smo95, SPF^{+}07, Spi94, SDE85, SMI80, SAF07, SXZ^{+}88, SETB08, Svo73, Svo81b, Taf82, TH94, TM81, TCH^{+}91, TLD^{+}11]. system [TS06, TS87a, TTP^{+}95, TLC85, TML97, TKP^{+}08, TLL03, TNL^{+}07, Tur87, Tur80, Vag10, Van06, Var72, VDGR96, VMBM12, Vog99, Vog00, WV02, WB07, WPE^{+}83, WZWS08, War76, WAB^{+}89, WLRZ03, WH08, WDA^{+}08, WPLP85, WA09, Wet99, Wet00, WWGF08, WGSS95, WMH72, WF07, WABL93, WH94, WZ94, XFO08, XDC^{+}95, YZG^{+}11, YTM^{+}91, YD02, YTR^{+}87, Yuv76, ZELV02, ZG07, ZDP83, ZLX01b, ZFW10, vRvST88, CBHLL92, CJ75, Cri91, GHP92, HO91, HGB^{+}80, JBW^{+}87, Jon92, Lis72, WN80, WAC^{+}81, WLP75]. system-level [BBD^{+}02]. System/6000 [HO91]. systematic [NAR08, Sal93]. systemic [SHSB75]. Systems [Acq16, ACS15, BDF^{+}15, BY08, BR10, BK08, Bre08, BNE16, Bru86, CIP^{+}23, CCS^{+}16, CJRV15, CHLS16, DVS12, DK16, Duc92, HKPvR16, Her10, KvRvST92, Laz92a, Laz92b, LLLG16, LE00, LLL^{+}17, MKL^{+}19, Mul87, PR15, PP09, Pet76, RF17, RAVC12, SRTH15, Sat99, Sch07, SHP^{+}16, Sha95, SG14, Sub11, Tan97, VPH^{+}15, WP91, WS92, WHZ^{+}17, WTC09, YVCB18, ZH19, ZE16, dV96, Aba93, AYK08, AMPS73a, AMPS73b, AMPS74, AMS^{+}07, AEG^{+}91, APGG00, AH80, ADN^{+}95, AC06, ADAD01, Atw84, AC97, AMO^{+}12, Bab91, BS95a, BDM97, Bac81, Bac91, BIYC06, BYV08, BMD94, BSSM08, BJ87, Bir91, Bis81, BF87, Bod11, BDT00, Bor98, Bor92, Bou94, BC01, BP91, BHJ^{+}93, BDR97, Cab90, CJR87, CW92, CS77, CALM97, CMK^{+}06, CFL73, Cha90, CN07]. systems [Cha96, CC05, Che75b, CEC^{+}95, CGL^{+}08, Che85, CL95, Cho77, CYC^{+}01, CLC05, Cla85, CM06, COS^{+}08, CGJ^{+}07, CB95, CBC^{+}08, CS00, CLDA07, DS92, DS80, DMD13, DK15, Dim98, Dou93, DBRD91, DB96, DB11, EER12, Ell73, Ell77, ESB^{+}06, ECH^{+}01, ELG95, Est02, EFL07, FRL00, Fle83, Fon72, FES09, FD10, Fra80, FV06, FdAM14, FW77, GSGN00, GC08, GBBL85, GJSO91, GBZP10, Gon89, GG91, Gor87, GB90, GGV96, Gre72, GLL04, GKS11, Hac85, Had85, HKD07, HS91, HV08, HKL^{+}06, Han72, HP93, HRU75, HHLS97, HJT^{+}93, HZ09, HCG^{+}06, HEK^{+}07, Hol72, HM90, HL92, HKU79, HH08, HYS03, KEG^{+}97, Kah72, KTB12, KSP09, KJH^{+}11, Kee79, KMA^{+}14, KvS07, KN93, Kil00, KSDC14, KSL90, KN96, Küh04, KBC94, KF09, LABW91, Lam00, LB91, LLK96, LSKK08, LAAW00, LMM93]. systems [LPS10, LJX97a, Lie93a, LCWM08, LKvR^{+}99, LKvR^{+}00, LFWL10, LF13, LG04, Mah98, Mal10, MC11, MW91, MRC^{+}97, Mat04, Mat06, Mat07, MS91a, MB06, MF75, Met82, MM81, Mog06, MW09, Moh78, MDB01, MMAS08, NCL12, NS07, Nee72, NB91, NBK^{+}20, Neu89, NTHAB22, NV06, NXQ05, Nut94b, NSW10, O’S92, OPSS93, OB86, OSV82, OST83, OSV86, OMCB07, OD89, PRAH96, PV95, PG06, PSC^{+}07, PFGD02, PS01, PBA^{+}05, Pu93, RK11, RV91, RGAB98, Rat87, Ray92, RL96, RR04, RLB08, Rob98, Rob08, RPM97, Ros95, Ros06, Rus81, RB75, Rus77, Sal78b, Sal91, SGD^{+}02, Sat95, SGNG00, SZN87, SLS^{+}05, SFB^{+}09, SHT97, SY96, SSS01, Sny77, SF91, Sor73, SDD^{+}85, SJ05, SK97, SR89, Sto07, Str93, SSR^{+}10b, Svo81a, TSF90, Tan87, THB06, TLL94, TGR^{+}21, TFC99]. systems [TBM^{+}06, Tra82, VBLM07, WH99, Wai98, WZZ93, WMI^{+}07, WAB^{+}89, WPC12, Wel95, WCS08, WA09, WLS^{+}02, WS91a, Wil93, Wil94, WS06, XHJB99, YM93, YW06, YAK93, YZZZ06, ZN00, ZIL96, ZLX01a, ZUW^{+}09, Zim94, dGdB10, dJKH93, dOS08, CvR14, EMS09, EMSPS11, GB01, LaR92, MW92, Mog09, Voe98, dSBP11]. SYSTOR [BY08].

T [Had85, Zho16]. T3E [Sco96]. Table [DHK^{+}15, JM98, RS86, THK95]. Tables [MT17, LE96, TMW10]. TACT [Yu00a, Yu00b]. tag [PHYO96, MFHH02]. Tagged [Tro00, Xu00]. Tags [SH87]. Tail [JHK^{+}16]. taint [HFC^{+}06, ZJS^{+}11]. taint-based [HFC^{+}06]. TaintEraser [ZJS^{+}11]. Tainting [DKK10, SB10b]. Take [Bis81]. Take-Grant [Bis81]. Taking [ZFP^{+}21, HEK^{+}07]. tale [Coo94]. talk [Lam00, Sal00]. Tame [AVN^{+}16]. Taming [Bal24, MKL^{+}19, SKKM02]. tamper [TML^{+}00]. Tanenbaum [dV96]. Taos [WABL93]. Tapeworm [UNMS94]. Tapir [YN15]. Tapping [WDA^{+}08]. Target [JHK^{+}16, PAM^{+}16]. Target-Driven [JHK^{+}16]. targeting [LGH94]. Tartan [MCC^{+}06]. Taser [GPF^{+}05]. Tashkent [EDP06, EDZ07]. Task [BHM77, Mah98, MB08, Bro75, GTA06, JLZx90, KSS73]. task-scheduling [KSS73]. Tasks [ZE16, CR75, DDYM99, ECS73, IvdLH^{+}00, JW96, LLK96, SLCG89, Yue85]. Taurus [MAHK16]. TaxDC [LLLG16]. Taxonomy [LLLG16, Ray91, TSF90, YAK93]. taxonomy-based [TSF90]. Taylor [Had93]. TCB [HCJ07, MPP^{+}08b, MPP^{+}08a, SPHH06]. TCC [HCW^{+}04]. Tcl [GA98]. TClouds [BCR^{+}14]. TCP [BSR06b, HRX08, yKR06, LCL^{+}16, VKD02]. TCP-friendly [HRX08]. TCP/IP [BSR06b]. teaching [AMO^{+}12, NV06, Rob08]. Technical [Cab90, Had83, McN77, McN82, McN88, MDO94]. technique [ACT94, HLFZ97, LGH94, Le98, Pay77, WC02, XXMC05]. Techniques [PS98, VGBT14, ZH16, AA06, BBFH07, Bod11, BFS89, BC08, CB95, DY01, KGS06, PS99d, PHOA89, SETB08, WS87, Wel95, ZK88]. technologies [PG06]. technology [EAS07, KDS^{+}06, MW08, Wil93]. telecommunications [GG91, KJH^{+}11]. Telemetry [SRS22]. temperature [HCG^{+}06, MB08]. temperature-aware [MB08]. Temporal [CWdO^{+}06, BH81, BBE^{+}11]. Temporally [LL02]. tenant [BWV^{+}12, SFS13]. TENEX [BBMT72]. term [BSR^{+}06a, Den74b, Eas72, Han72]. Terminal [HGB^{+}80, Hil81, MM81, Wal73]. termination [CK86, Lau84]. territory [WY04]. tertiary [VT01]. Test [LWPG17, Lie94a, Poo73]. testbed [KSK09]. Testbeds [ELR15, LFH^{+}09, ROJS09, ROLV06]. testing [CZB^{+}09, CKK^{+}07, MCM07]. TETRIS [GPY^{+}17]. Tetzlaff [Tet14]. their [AD07, BTK11, BSF^{+}91, CMSK07, Dim98, GS78]. them [CH14, JS08]. theoretical [FFM07]. theory [LABW91, MXXC05, MM91, Pra87, Woo73, ZUW^{+}09]. Thermostat [AW17]. thin [BKN05, SLN00, SLN99]. thin-client [BKN05, SLN00, SLN99]. THINC [BKN05]. third [PG73, DK15, HN12]. Thomas [Kad95b, Kad95a, Wai97a]. Thorough [BBC^{+}06]. Thoth [CMMS77]. thoughts [Che85]. Thread [FURM00, GP05, LPM17, PEA^{+}96, TAS07, DBRD91, GLC99, Lie94b, LML00, Loe05, MT02, OT95, SP00, Shi00, SJ95, Wei98]. Thread-level [FURM00, MT02]. thread-specific [SP00, Shi00]. threadbare [Bak95]. threaded [CSS^{+}91, LBvH06, OA08, SBN^{+}97, SQP08]. threading [RRP06, SQP08]. Threads [MP89, Bak95, CPT08, DC99, DC00, GP05, GN96, HJT^{+}93, KE95, MSLM91, MQW95, MEG94, OL02, PG03b, SZG91, SZ92, SCM05, SMM^{+}09, WCW^{+}04, ZCSM02]. Threat [NCBB14]. Three [LSH00, Ng99, Rob08, Sch73a, SPHH06, XZZ98, ZL04a]. Three-Party [Ng99, LSH00, XZZ98]. three-principal [ZL04a]. threshold [BMW02a]. throughout [Fab98]. Throughput [ALM^{+}18, DD12, yKPR02]. tier [CCZ07a, MZWZ02, NTHAB22]. tiered [AW17, TGR^{+}21]. Tiger [BFD97]. tightly [PR83]. tiled [MSP^{+}06]. Time [CKN^{+}19, DL15, FS95, GF15, MCGL17, AGM93, ACT94, Bas72, BL75, BH81, BHD19, BBMT72, BEW75, BEW76, BM90, CLR94, CCK04b, Chá91, CMMS77, CM75, DRSK89, DS92, Den74a, DC99, DC00, DCZ96, EGE02, ECS73, FL77, FW72, Ful73, FPG89, Gar07, GAK^{+}02, GP95, GS89, Gre72, GN96, Gup01, HS91, HLFZ97, JRR97, KKS89, KC94, LBF^{+}98, LTCA89, Lie96, LC04a, LSA^{+}00a, LSA^{+}00b, LLY05, MO85, MW91, Mil92, NMS^{+}00, NCL12, NL96, OT95, PM03, PS01, PN00, PC75, RLB08, RT73, RPM97, SN94, SZG91, SZ92, Sor73, SR89, TSLBYF08, TM89, TL96, Wai95a, WAB^{+}89, WPC12, WJ98, WMH72, Wir77, YS98, YD02, Zea97, Zel74, ZPS99, ZPS00, JBW^{+}87]. Time-function [FS95]. time-sensitive [GAK^{+}02]. time-shared [WMH72]. time-sharing [Chá91, FW72, Gre72, RT73]. Time-to-Accuracy [CKN^{+}19]. time/run [DCZ96]. timebombs [CWdO^{+}06]. timeline [Gwi94]. timeliness [RLB08]. Timely [LZH^{+}22, OR87]. timer [AD99, AD00, DM90, PBR^{+}08, VL87]. timers [AD99, AD00, Dub00]. times [CR75, CCK04b, SLCG89, YM93]. Timestamp [MSA^{+}00, YW04]. timestamp-based [YW04]. timestamps [Nat80, NS93]. Timing [HMK20, DM90, VL87]. Timing-Channel [HMK20]. tiny [LC02, SLQP07, MFHH02]. TLB [JM98, TH94]. TLB-refill [JM98]. TM [RRCC10]. token [BL00]. token-based [BL00]. Tolerance [Cri91, AAC^{+}05, Bab91, BRR^{+}00, Bir85, Bir91, BBG83, BS95b, GG91, JT90, KT91a, Kan83, KS91b, KAD^{+}07, NB91, PL95, PNT06, PCD91, RRP06, RCL01, Rom93, Sal91, SPR00, TCH^{+}91, WLZ03, XXM04, ZHK06]. tolerance-current [JT90]. Tolerant [WQA^{+}24, AEMGG^{+}05, Bab90, BJM^{+}91, BACF08, BC91b, CC97, DHRS91, FV06, GC89, HGR07, JAvR06, LCJV^{+}11, MS91a, PJDL06, SNV10, YbJf04]. Tolerating [VBLM07]. tomography [GAT13, MMAS08]. TOMP [Das92]. too [KMA^{+}14]. tool [BFSG94, FdAM14, LS90, NMS^{+}00, RSW08, SK96, Spi94]. tooling [DH10]. toolkit [EBS01, JdLT^{+}95, LJW^{+}06, QPP02, Jon92, MW92]. tools [GC05, SETB08, Wei92]. Toolset [Ott18]. top [CS00]. top-down [CS00]. Topic [LCCZ17]. Topics [CvR14, DNT10, HN12, Sat99, SN13, Smi78]. topology [PLH98]. Toronto [San86]. TOS [NB00]. total [Das92, Ful73]. totally [Bir94, CS93, Coo94, Toi92]. TPC [JHK^{+}16]. Trace [GCJ17, CNO^{+}87, DH10, EJD13, HXL01, KTP^{+}96, ODH^{+}85, Spi94]. trace-driven [KTP^{+}96, ODH^{+}85]. traces [PS96, PRD10, YLW^{+}06]. Tracing [NTC^{+}21, DD12, KJ08]. Tracking [YSCC16, JOW^{+}02, SLZD04, YRC05, ZPS^{+}04, ZJS^{+}11]. trade [MSP98]. trade-offs [MSP98]. Tradeoffs [CMM^{+}06, SJS^{+}23, AEH75, CN07, DMB87, JOW^{+}02, Yu00a, Yu00b]. Trading [WM16, LNBZ08]. tradition [dBB08]. Traffic [LCJV^{+}11, Gup01, KAI^{+}13, Wal73]. Training [BKGT25, Ser21]. transaction [CPW07, Cri94, Duc89, EDP06, MSF85, MMP83, RB93, Spr85, Sto84, SDE85, VBLM07]. Transactional [DDK^{+}16, NP17, RG02, ZLJ16, BJM^{+}91, CCZ07a, CNV^{+}06, CMM^{+}06, CR12, DFL06, GKV07, HCW^{+}04, MMTW10, MBM^{+}06, RRCC10, RHP^{+}07]. Transactions [Ano75, KPS^{+}16b, LZC^{+}17, MCGL17, WPLP85, YWKYS15, Bla91, Fra95, KGGK09, LC97, LS94, ML85, PS09, Pu93, SW91, SDD^{+}85, Spi94, SDE85, SS83b, You92]. transfer [DP93, KCLZ98, MP75, TLL94, WSH94]. transfers [VKD02]. transformation [CEV00, SV06]. Transformations [SSK17, GMM98]. transient [VM07]. transitions [EB78]. transitive [XHB06]. Translation [AZEE18, Bha17, CB17, SBS18, YVCB18, ACM02, CBD^{+}98, LSKK08, Ros89, SS95]. Translation-Triggered [Bha17]. translator [LOM^{+}09]. translators [Le98]. transmission [Cho77]. Transparency [Str93, Hof07, SLLP^{+}10]. Transparent [ALM^{+}18, Bac91, CCG95, JZ91, KS91b, RS02, ZWZ01, AW17, cCVP99, CVP00, NIDC02, PGS93, PWC^{+}81, SG97]. transparently [Jon93]. Transport [vR92, BMR^{+}09, WH94]. Trap [UNMS94, KKN00]. Trap-driven [UNMS94]. traps [HM93]. Travel [Bar14, Gra14, Tet14, TSLBYF08]. Traveling [Wil09]. treasurer [And81, Den78, Den79, Den80]. treating [QTSZ05]. tree [ML85, MP81]. trends [Fle83, LB08, RBH^{+}95]. TRIAD [Che00a, Che00b]. Triage [TLH^{+}07]. triangular [CC97]. Tribute [TSE^{+}00]. TriCheck [TML^{+}17]. Triggered [Bha17]. Trimmed [VGX17]. TRIOS [Ter14]. Triple [Ran82, GC12]. Triple-handed [Ran82]. TRIPOS [RN83]. Trisection [TML^{+}17]. Trivedi [Sta83, Wai83b]. Troubled [Lit87]. troubles [HBB13]. troubleshooting [WKT^{+}13]. True [MMT16]. Truly [WQA^{+}24]. Trust [Gup05, DY10, MXXC05, Sat00]. Trusted [DPW^{+}09, KDL^{+}16, ABC^{+}02, BCP^{+}08, CGM97, KDP02, KSLA08, MC11, WH08]. trustworthy [HEK^{+}07]. Tully [Had85]. tunable [WL09, Yu00a, Yu00b]. Tuning [MRH^{+}16, KSP09]. Tunis [Atw84, Hol82]. Tuplink [Neg00]. TURNING [Hol88]. TVDc [BCP^{+}08]. Twenty [BK08, Bre08]. Twenty-First [BK08, Bre08]. twins [HCJ07]. Two [AW17, HL05, KTC03, Lau84, LBB^{+}91, BL89, BSSM08, CG91, GS90, GLC99, HCJ07, JW01, MD81, Wed88, ZLX99, ZIL96]. two-dimensional [BSSM08]. two-level [CG91]. Two-part [Lau84]. two-party [JW01, ZLX99]. Two-tiered [AW17]. TxLinux [RHP^{+}07]. TxRace [ZLJ16]. type [CC77, Jan81, SH87]. type-extension [Jan81]. Typed [KKK^{+}17]. types [Buc77, Gan77, Her87]. TYPESET [SHSB75]. TYPESET-10 [SHSB75].

U [vEBBV95]. U-Net [vEBBV95]. U. [AD07]. U.K [OST83]. überSpark [VMM20]. ubiquitous [Ram00, ST93, YLE02]. Ubiquity [Wed88]. ugly [HYM10]. Uhl [Kad95b, Kad95a]. ULT [PG03b]. Ultra [CDY^{+}17, SCP^{+}06, CH07, EKM04]. ultra-lightweight [CH07]. Ultra-low [CDY^{+}17]. UMI [Wai83a]. UML [WPC12]. UMTS [LCJV^{+}11]. un-fully [WYC03a]. Unbounded [CNV^{+}06, DS92]. Uncertain [Zho16]. uncertified [KC95]. uncomputation [SV06]. undergraduate [Rob98, Rob08]. underlying [YWC04]. Understanding [Bla91, CS93, HT15, LJS^{+}02, LRS^{+}08, MMAS08, PLM06, RRP06, YZZZ06, Lov77]. undetachable [BMW02b]. Undetectable [DH95]. unfair [HS88]. Unicorn [OLLY02]. unified [CCLP81, VESM10, Wei98]. Uniform [Yan92, KBK02]. unify [WTB10]. Unifying [NTC^{+}21, Str78]. unions [CC77]. uniprocessor [RTY^{+}87]. unit [KEP07]. uniting [EDP06]. units [CFR98, JM98, Rat87]. universal [LEH86]. Universe [Kad95a]. University [Wai83a, GNB^{+}09]. UNIX [BRW89, CE88, FPG89, Gue87, LC93, LCH^{+}81, Mil78, ODH^{+}85, Pon97, RT73, Tan87, THB06, Vog97, Che75c, Hol82, LGJS91, Neu00, PVB17, Ros78, dBB08]. Unix-like [Neu00]. unknown [WYC03b, WCYJ05]. Unleashing [Kad95b]. unmet [FM98]. unobstructed [WC02]. unpredictable [LSA^{+}00a, LSA^{+}00b]. unreliable [BJK^{+}06, WYC03a, WCYJ05]. Unrestrictive [Hem89, TT00, Kea88]. unsafe [PGZ08]. unsecured [YWC04]. unsolved [Lam85]. unstructured [LFWL10]. Unsupervised [GLD^{+}22]. untampered [SLS^{+}05]. Untrusted [KDL^{+}16, ZZNM01, SAL20]. upcall [GP95]. upcalls [Cla85]. update [EDZ07, PST^{+}97, TTP^{+}95]. Updates [IKK16, MR07, You92]. upgrade [CKK^{+}07]. Upgrades [DNT10, HN12, SN13, HBB13]. upon [Bas72]. URICA [MCdL06]. URL [vEBBV95]. Usage [MFBWW20, PBR^{+}08, Ros78, Vog99, Vog00, MCdL06]. Usage-awaRe [MCdL06]. Use [Atw84, NHH^{+}17, ZJL17, Nut94a, ATMZ01, CH14, HCBS04, NS93, PPT^{+}93, San81, SMI80, Wol02]. User [BBM^{+}81, BW01, Jon92, MQW95, NTC^{+}21, RS02, RB24, ZZ03, ZWZ05, ACG86, ABLL91, AL91, ACT94, AMO^{+}12, BF08, Cha73, CL04c, GP05, Gsc94, HL05, Jon93, KC05, LHY02, LLH02, LKY04, LK01, MCD^{+}08, MSLM91, MDO94, MC96, MRA87, Moo82, OT95, OCF00, RSW08, Rus88, Sto07, TLH^{+}07, YRY04, YZZZ06, vEBBV95, RB24]. user-assisted [RSW08]. user-controlled [Cha73, Sto07]. User-defined [RB24, Gsc94]. User-Level [RS02, BW01, MQW95, ZZ03, ZWZ05, ABLL91, AMO^{+}12, MSLM91, MRA87, OT95, OCF00, vEBBV95]. user-perceived [MCD^{+}08]. users [SS17]. uses [MZI08, TPO06]. USIM [Moo82]. Using [BM99, BNE16, CIP^{+}23, CCEH00, COS^{+}08, DBRD91, EBP16, EWCS96, FHL95, GCJ17, GKL95, GSCM16, HV08, Han83, HJT^{+}93, Jan81, KL02, Nic87, SMRD06, SPBP06, ZWWL01, ZLJ16, AHB15, ATSV06, AJG07, ATSS09, BSM^{+}12, BR09, BC08, CL04c, CCK04b, CHY05, CGM97, CJ05, CGKM11, Che84, CG06, Cla85, CR72, Coo78, DSGP05, EGE02, ELG95, FFBG08, GCM^{+}94, GTHR99, GTHR00, GDRT13, GA08, GCTR08, GSM08, Hag87, HS88, HJ10, HC92, HC04, HN08, Hil92, HFC^{+}06, JFV^{+}96, JXT93, KT91a, KC95, KDS^{+}06, KLY03, KCLZ98, Ku04, KC05, LFH^{+}09, LHY02, LLH02, LKY04, LW04, Lei89, LFW04, Lom77, MCM07, MSF85, MFGSP12, NPC06, NV06, Oes01, PS09, PS96, PFK^{+}22, PRD10, QPP02, RP07, RLD^{+}17, RCL01, RHP^{+}07, SH96, SL98, Sco04, Smo95, SCG01, SG05, SKPG01, Svi83, TSLBYF08, TDM12, TPO06]. using [Tug83, VBLM07, WP87, WK05, WL82, WSW05, WRA05, Wol02, Won93, YW04, YRY04, YS94, ZJS^{+}11]. utility [DH73, PSZ^{+}07, RD01]. Utilization [CYMT16, CYG^{+}17, PPM17, CKDK91]. utilizations [GSM08, Rob96]. utilized [Rob96]. Utilizing [AVZR11, KKN00]. UTLB [CBD^{+}98].

V [CZ83, Kot88, TLC85]. V-system [TLC85]. VAIF [TAH^{+}22]. validation [ME08]. Value [FJLC98, JXG21, LWS96, WCL17, BMW02a, BEL^{+}00, DHJ^{+}07, ZCSM02, ZYG00]. value-centric [ZYG00]. VAMNET [Bos06]. Vanguard [Fin92]. vApp [SG10a]. variability [FGBA96]. Variable [MS94, SEP98, HV92, LPH^{+}07, WS91b, YN12]. variables [Buc77, Ger77]. Variance [TAH^{+}22]. Variance-driven [TAH^{+}22]. Variant [MRH^{+}16, HRX08]. variants [CJ05]. VAX [Woo85, Cla87, GKD91, Gue87, Gwi94, KB84, KGB88]. VAX/VMS [Woo85, GKD91, Gwi94, KB84, KGB88]. VAXclusters [KLS85]. vCloud [KMK10]. Vector [MSAD91, MNP07]. vectors [LHL04, MB08]. vendor [EER12, RD87]. Venus [Lis72]. Verifiable [SAL20, YWKYS15, AGB^{+}77]. Verification [FXZ^{+}17, TML^{+}17, ZSG^{+}17, ACD^{+}14, DMD13, JW01, KMA^{+}14, LF13, Rus81, SWL77, Sil83, WPC12, ZLX99, ZL04a]. Verified [KDL^{+}16, YN15]. Verifying [AHC^{+}16, BCC^{+}13, LSMB16, SLS^{+}05]. Verlag [Had93, Lig94, Nut94a, Wai94]. versatile [AKGR10]. version [FW77, GKD91, KGB88]. versioning [WF07]. versuchung [DL15]. versus [Bar79, Gwi05]. Vertigo [FM02]. very [CMK^{+}06, EJD13, Riz97, Sal91]. via [AKP^{+}25, Bod11, CG94, CCM96, CLM^{+}07, DS90, FGBA96, ÏMC^{+}06, IKK16, JXG21, LTQZ06, NG09, PK96, RSEW04, SLZD04, TMW10, VGX17, WCW^{+}04, WM16, YRC05, YJX^{+}16]. viable [EENV02]. Video [BCC^{+}94, AS10, BFD97, CSJZ08, JH93, RV91, YZZZ06, Her92b, Jef92]. video-on-demand [CSJZ08, YZZZ06]. Videos [JSCM17]. View [HSPC01, Acq16, BDM97]. View-based [HSPC01]. viewpoint [Küh04]. views [DS80]. Vigilant [PBYH^{+}08]. Vigilante [CCC^{+}05]. VII [dSBP11]. violations [BSM^{+}12, LTQZ06]. VIP [HdRC95]. VIP-FS [HdRC95]. viral [HCK08]. virtio [Rus08]. Virtual [AZEE18, AL91, AMA^{+}11, BBHL08, BBM09, EMZ^{+}16, KMK16, SS95, Sto84, TSLBYF08, Tra82, Vag10, VMC^{+}05, Zho10, ARS89, AGSS10, AMMR92, BFHW75, BSM^{+}12, BDS^{+}09, BKN05, BCP^{+}08, BJ87, CH81, CD95a, CWdO^{+}06, CLDA07, DPW^{+}09, DC99, DC00, DKC^{+}02, ENCH96, FR85, FHL^{+}96, FLM^{+}08, Goo87, GTHR99, GTHR00, HJ10, HdRC95, HN08, HUL06, HDG09, IKWS92, Jan81, JADAD06, KF09, LBP^{+}07, LMG^{+}07, LT96, LCTK01, LC02, LSS04, LCH^{+}81, hTMAC^{+}08, MM81, MAK07, NV06, OCF00, PBYH^{+}08, PK75, RTY^{+}87, RS86, Rus08, SCP^{+}02, SMK^{+}93, SNV10, Sch95, SGGB99, SGGB00, Taf82, Tan79, VFH98, WCW^{+}04, WK08, WH08, WMH72, XLDB09, YZG^{+}11, ZWL09, ZIL96, BH75, Neu92]. virtual-machine [DKC^{+}02, HUL06]. virtual-memory [Jan81]. virtualised [MPF^{+}06]. virtualizable [PG73]. Virtualization [DLLN18, HSL17, MA10, MUKX06, RB24, AA06, BBD^{+}10, BC10, BSMF08, CGL^{+}08, CMM^{+}06, DS09, FFBG08, FS08b, Ros06, SPF^{+}07, SWC08, VW08, WCS09]. virtualization-based [CGL^{+}08]. virtualize [TDM12]. Virtualized [MT17, YVCB18, BSSM08, CJS^{+}09, KTB12, NS07, PSZ^{+}07, PSC^{+}07, RS08, SG10b, WTLS^{+}09]. Virtualizing [BTMS10, SB10a]. virtually [IKWS92, Lie95c]. virtually-addressed [Lie95c]. VirtualPower [NS07]. Viruses [Wai97c]. VISA [VFH98]. vision [Wet99, Wet00]. Vista [LC97]. Visual [DH10, SFW99]. Visualizing [Mar97, Vog97, MMAS08]. VLIW [CNO^{+}87, WS91b]. VLSI [BKT87]. VM [SHW^{+}15, TDM12]. VMOS [Fog74]. VMS [Woo85, GKD91, Gwi94, KB84, KGB88, Wie92, RB24]. VMware [BBD^{+}10, DS09, Her10, HMS17, PPS^{+}18, Ten17, Wal02]. VMware’s [KMK10]. voice [TS87a]. Volatile [AMH^{+}16, BXS14, CCHV11, HN08, SETB08, WZ94]. Voltage [BLI17, GS13, PS01, WJMC04]. voltage/frequency [WJMC04]. volume [Nut94a]. Volumes [Lig94]. voluntary [OLLY02]. VP [JR05, TC96]. VPFS [WH08]. VRGQ [JXG21]. vs [BCDN87, Dou09, Gar07, GKO^{+}00, GA98, KMSV10, MMTW10, Mog06, PG03a, RB24, WM16]. vSAN [FKZ17]. vSwitch [TSP17]. vulnerability [AFB95, JKDC05]. vulnerability-specific [JKDC05].

W [Had85]. WACI [Tsa16]. walk [ZS06]. walks [BSSM08]. WAN [AEH^{+}08]. ware [RA06]. Warehouse [CYMT16, CYG^{+}17]. Warehouse-Scale [CYG^{+}17]. Warp [JBW^{+}87, BM90]. Wasiq [RS02]. WASS [PS99b, PS99e]. waste [CH14]. watt [KF09]. way [CHY05, LW04, LAB^{+}06, Rom95, Toi92]. weak [HS88, MES95]. weakly [PST^{+}97, TTP^{+}95]. Weaknesses [KCL03, KCC05, XZZ97]. Wearables [DDOL16]. WEB [Bla95, Wai97b, CEV00, CLM^{+}07, CLC05, Gup05, KL07, yKPR02, PBH^{+}07, RCC01, SS97, WFHJ07, WVS^{+}99, WVS^{+}00]. Web-based [CLC05]. Weight [BKGT25, MSC^{+}06, vdWMH11]. Weir [BMER14]. Weiser [TSE^{+}00]. well [BS89, Rom95, WCB01]. well-conditioned [WCB01]. well-structured [Rom95]. WG [OSV82, OST83, OSV86]. wheels [VL87]. Where [CLR94, CR12, ABD^{+}97, KC94]. which [LJX97a, Rou84]. Whirlpool [MBS16]. WHISPER [NHH^{+}17]. Whodunit [CCZ07a]. Whole [BS15, GN96, BBM09, MCC^{+}06]. Whole-Program [BS15, GN96]. Wide [BMvdV93, DKK^{+}01, BvS00, GS95, KLS08, SKKM02, Sha00, SS95, WECK07]. Wide-address [BMvdV93, SS95]. Wide-area [DKK^{+}01, BvS00, GS95, SKKM02, WECK07]. width [KT91b]. Wild [Bal24, Tsa16]. wildlife [JOW^{+}02]. Wiley [WP91]. Willard [Bla95]. William [Tet14]. Window [Gor87, MM81]. Windows [Val94, PS96, TF04, Vog99, Vog00, WH99, YD02, ZWZ01]. winter [AD07, ZCT^{+}05]. wire [KBK02]. wire-delay [KBK02]. Wireless [ACAAT16, HSI^{+}01, PS98, PS99d, PS99a, PS99b, PS99e, Pat02a, Pat02b, Duc92]. Wisconsin [AD07]. WiSync [ACAAT16]. within [Har82, Loe85, MSF85]. Without [EPG^{+}20, CCK04b, Eas72, Ku04, LW04, LPS10, MCS91, NL96, SFW99]. wonderland [AD07]. WOPSSS [Acq16]. word [CMSK07]. Work [HMS17, Laz92b, Mul87, VJ19, Sch07, Van96, WS92]. Workday [VFP22]. worker [SCM05]. Workflow [SLD15, YJX^{+}16, PKM^{+}09]. workflows [WKL07]. Working [OSV82, OST83, OSV86, SS83a, COS^{+}08, DS72, Fog74, Mar97, MPC08, Pot77]. working-set [DS72]. Workload [PBM22, WTB10, DK15]. workloads [DKW^{+}09, EM89, GWSY08, HD12, MDO94, RGAB98, SWC08, SQP08, TPH12, WBR^{+}12]. Workshop [BCC^{+}94, BR10, CvR14, CJRV15, CM13, CM14, CR12, DK15, DNT10, DB11, EMS09, EMSPS11, HN12, HKPvR16, JT90, KSL90, LS09, MvR13, Mat10, Mul87, PP09, RAVC12, Sat95, Sat99, SN13, Sha95, Sub11, Tan97, WTC09, dSBP11, Bab91, Bir91, Bod11, Cab90, DO09, HLR98, MdS09, MM92, MM93, dOS08, Acq16, Bac99, Her92b, Mog09, SG14]. Workshops [Mog09]. Workstation [Kad95b, AEP^{+}97, Cab90, Fab98, FMP^{+}95, JXY95, PL95, SGN85, TS87b]. workstation-based [AEP^{+}97, Fab98]. workstations [CZ83, DM90, JXT93, LGH94, LX00, Nic87]. World [Wit16, BBM09, EHD07, LPSZ08, Wil94, vRvST88, Lit87]. world sky [BKP^{+}12]. Worldwide [Tan97]. worms [CCC^{+}05, TNL^{+}07]. worthwhile [NBB09]. WOWCS [Mog09]. Write [DRTT24, KKB^{+}16, FC87, GWSY08, GKL95, JXHQ02, MMGC02, SDH^{+}97, WL82]. Write-Ahead [KKB^{+}16, DRTT24]. write-intensive [GWSY08]. write-once [FC87]. writer [HV92, Sei90]. writers [KL98, KPL99, Küh99]. writes [CCHV11, JXG^{+}02]. Writing [Hei78, BFSG94]. written [Hol82]. WSCLOCK [CH81]. WTM [CR12]. Wukong [CC21].

X [CLC05, MP81, Rei85, PHOA89]. x-Kernel [PHOA89]. X-MP [Rei85]. X-RDR [CLC05]. X-tree [MP81]. x86 [AGSS10, MT17, AA06]. Xiao [JW01]. XINIX [Chá91]. XINU [Fos87]. XML [SCG01]. Xoc [CBC^{+}08]. XOS [MP81]. XT1 [ZLX^{+}80]. Xu [Ng99].

year [Mat07]. Years [MFBWW20, LBB^{+}91]. Yoo [KCC05]. Yoon [KCC05]. York [Had93, Lig94, OST83, Val94, WP91, Wai94]. Yourself [AZEE18].

Z [BAD^{+}11]. Zap [OSSN02]. ZCZOS [ZDP83]. Zebra [HO93]. ZebraNet [JOW^{+}02]. Zero [AMH^{+}16, MMT16, PSB06, PB08]. Zero-Cost [AMH^{+}16]. zero-day [PSB06, PB08]. Zhang [JW01, Ng99]. Zhu [Ng99]. ZNS [DRTT24]. Zone [DRTT24]. ZS [SDV^{+}87]. ZS-1 [SDV^{+}87]. ZWAL [DRTT24]. Zyzzyva [KAD^{+}07].

References

Adams:2006:CSH

[AA06] Keith Adams and Ole Agesen. A comparison of software and hardware techniques for x86 virtualization. Operating Systems Review, 40(5):2–13, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aguilera:2023:MDW

[AAA^{+}23] Marcos K. Aguilera, Emmanuel Amaro, Nadav Amit, Erika Hunhoff, Anil Yelam, and Gerd Zellweger. Memory disaggregation: why now and what are the challenges. Operating Systems Review, 57(1):38–46, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3606557.3606563.

Aiyer:2005:BFT

[AAC^{+}05] Amitanand S. Aiyer, Lorenzo Alvisi, Allen Clement, Mike Dahlin, Jean-Philippe Martin, and Carl Porth. BAR fault tolerance for cooperative services. Operating Systems Review, 39 (5):45–58, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:2009:DEF

[AAMV09] Eric Anderson, Martin Arlitt, Charles B. Morrey III, and Alistair Veitch. DataSeries: an efficient, flexible data format for structured serial data. Operating Systems Review, 43(1):70–75, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Agrawala:1975:MMS

[AB75a] A. K. Agrawala and R. M. Bryant. Models of memory scheduling. Operating Systems Review, 9(5):217–222, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arden:1975:MMC

[AB75b] Bruce W. Arden and Alan D. Berenbaum. A multimicroprocessor computer system architecture. Operating Systems Review, 9(5):114–121, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Adams:1982:PRM

[AB82] J. M. Adams and A. P. Black. On proof rules for monitors. Operating Systems Review, 16(2):18–27, April 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See reply [How82].

Abawajy:1993:OPR

[Aba93] Djemal H. Abawajy. Orphan problems and remedies in distributed systems. Operating Systems Review, 27(1):27–32, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1998:NFA

[ABC^{+}98] Ross Anderson, Francesco Bergadano, Bruno Crispo, Jong-Hyeon Lee, Charalampos Manifavas, and Roger Needham. A new family of authentication protocols. Operating Systems Review, 32(4):9–20, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Adya:2002:FFA

[ABC^{+}02] Atul Adya, William J. Bolosky, Miguel Castro, Gerald Cermak, Ronnie Chaiken, John R. Douceur, Jon Howell, Jacob R. Lorch, Marvin Theimer, and Roger P. Wattenhofer. Farsite: federated, available, and reliable storage for an incompletely trusted environment. Operating Systems Review, 36(5S):1–14, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1997:CPW

[ABD^{+}97] Jennifer M. Anderson, Lance M. Berc, Jeffrey Dean, Sanjay Ghemawat, Monika R. Henzinger, Shun-Tak A. Leung, Richard L. Sites, Mark T. Vandevoorde, Carl A. Waldspurger, and William E. Weihl. Continuous profiling: where have all the cycles gone? Operating Systems Review, 31(5):1–14, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andersen:2001:RON

[ABKM01] David Andersen, Hari Balakrishnan, Frans Kaashoek, and Robert Morris. Resilient overlay networks. Operating Systems Review, 35(5):131–145, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1991:SAE

[ABLL91] Thomas E. Anderson, Brian N. Bershad, Edward D. Lazowska, and Henry M. Levy. Scheduler activations: effective kernel support for the user-level management of parallelism. Operating Systems Review, 25(5):95–109, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Auyong:1997:ASC

[AC97] Keok Auyong and Chye-Lin Chee. Authentication services for computer networks and electronic messaging systems. Operating Systems Review, 31(3):3–15, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Antoniadis:2006:EER

[AC06] Panayotis Antoniadis and Costas Courcoubetis. Enforcing efficient resource provisioning in peer-to-peer file sharing systems. Operating Systems Review, 40(3):67–72, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

AlMaruf:2023:MDA

[AC23] Hasan Al Maruf and Mosharaf Chowdhury. Memory disaggregation: Advances and open challenges. Operating Systems Review, 57(1):29–37, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3606557.3606562.

Abadal:2016:WAF

[ACAAT16] Sergi Abadal, Albert Cabellos-Aparicio, Eduard Alarcon, and Josep Torrellas. WiSync: an architecture for fast synchronization through on-chip wireless communication. Operating Systems Review, 50(2):3–17, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Agarwal:2008:PMI

[ACC^{+}08] Sheetal Agarwal, Dipanjan Chakraborty, Swati Challa, Nandakishore Kambhatla, Arun Kumar, Sougata Mukherjea, Amit Anil Nanavati, and Nitendra Rajput. Pyr.mea.IT : permeating IT towards the base of the pyramid. Operating Systems Review, 42 (1):108–109, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alvaro:2009:DDC

[ACC^{+}09] Peter Alvaro, Tyson Condie, Neil Conway, Joseph M. Hellerstein, and Russell Sears. I do declare: consensus in a logic language. Operating Systems Review, 43(4):25–30, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amani:2014:AVA

[ACD^{+}14] Sidney Amani, Peter Chubb, Alastair F. Donaldson, Alexander Legg, Keng Chai Ong, Leonid Ryzhyk, and Yanjin Zhu. Automatic verification of active device drivers. Operating Systems Review, 48(1):106–118, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ancona:1986:IUP

[ACG86] M. Ancona, A. Clematis, and V. Gianuzzi. Interfacing user processes and kernel in high level language. Operating Systems Review, 20(1):19–23, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ashok:2002:CMC

[ACM02] Raksit Ashok, Saurabh Chheda, and Csaba Andras Moritz. Cool-mem: combining statically speculative memory accessing with selective address translation for energy efficiency. Operating Systems Review, 36(5):133–143, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Acquaviva:2016:PSS

[Acq16] Jean-Thomas Acquaviva. Performance and scalability of storage systems, a view from the WOPSSS Workshop. Operating Systems Review, 50(3):2, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aldaco:2015:LAN

[ACS15] Abraham N. Aldaco, Charles J. Colbourn, and Violet R. Syrotiuk. Locating arrays: a new experimental design for screening complex engineered systems. Operating Systems Review, 49 (1):31–40, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Argade:1994:TMR

[ACT94] Pramod V. Argade, David K. Charles, and Craig Taylor. A technique for monitoring run-time dynamics of an operating system and a microprocessor executing user applications. Operating Systems Review, 28(5):122–131, December 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aron:1999:STE

[AD99] Mohit Aron and Peter Druschel. Soft timers: efficient microsecond software timer support for network processing. Operating Systems Review, 33(5):232–246, December 1999. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aron:2000:STE

[AD00] Mohit Aron and Peter Druschel. Soft timers: efficient microsecond software timer support for network processing. Operating Systems Review, 34(2):25–26, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arpaci-Dusseau:2007:CPH

[AD07] Remzi H. Arpaci-Dusseau. CS 736 project highlights from U. Wisconsin: how students spend their days and nights in a winter wonderland. Operating Systems Review, 41(1):54–55, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arpaci-Dusseau:2001:ICG

[ADAD01] Andrea C. Arpaci-Dusseau and Remzi H. Arpaci-Dusseau. Information and control in gray-box systems. Operating Systems Review, 35(5):43–56, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alvisi:2007:HRG

[ADG^{+}07] Lorenzo Alvisi, Jeroen Doumen, Rachid Guerraoui, Boris Koldehofe, Harry Li, Robbert van Renesse, and Gilles Tredan. How robust are gossip-based communication protocols? Operating Systems Review, 41(5):14–18, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1995:SNF

[ADN^{+}95] T. E. Anderson, M. D. Dahlin, J. M. Neefe, D. A. Patterson, D. S. Roselli, and R. Y. Wang. Serverless network file systems. Operating Systems Review, 29(5):109–126, December 1995. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aguirre:1994:DFH

[AEE^{+}94] G. Aguirre, M. Errecalde, S. Esquivel, G. Leguizamon, and R. Gallard. Design features of high level layers in LAHNOS, a local area heterogeneous network operating system. Operating Systems Review, 28(2):34–50, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aguirre:1991:EMD

[AEG^{+}91] G. Aguirre, M. Errecalde, R. Guerrero, C. Kavka, G. Leguizamon, M. Printista, and R. Gallard. Experiencing Minix as a didactical aid for operating systems courses. Operating Systems Review, 25(3):32–39, July 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Akkoyunlu:1975:SCT

[AEH75] E. A. Akkoyunlu, K. Ekanadham, and R. V. Huber. Some constraints and tradeoffs in the design of network communications. Operating Systems Review, 9(5):67–74, November 1975. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ananthanarayanan:2008:PPW

[AEH^{+}08] R. Ananthanarayanan, M. Eshel, R. Haskin, M. Naik, F. Schmuck, and R. Tewari. Panache: a parallel WAN cache for clustered filesystems. Operating Systems Review, 42(1):48– 53, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abd-El-Malek:2005:FSB

[AEMGG^{+}05] Michael Abd-El-Malek, Gregory R. Ganger, Garth R. Goodson, Michael K. Reiter, and Jay J. Wylie. Fault-scalable Byzantine fault-tolerant services. Operating Systems Review, 39(5):59–74, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arredondo:1997:LDB

[AEP^{+}97] D. Arredondo, M. Errecalde, F. Piccoli, M. Printista, R. Gallard, and s. Flores. Load distribution and balancing support in a workstation-based distributed system. Operating Systems Review, 31(2):46–59, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alves-Foss:1995:ACS

[AFB95] Jim Alves-Foss and Salvador Barbosa. Assessing computer security vulnerability. Operating Systems Review, 29(3):3–13, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Argollo:2009:CIF

[AFF^{+}09] Eduardo Argollo, Ayose Falcón, Paolo Faraboschi, Matteo Monchiero, and Daniel Ortega. COTSon: infrastructure for full system simulation. Operating Systems Review, 43(1):52– 61, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ambler:1977:GLS

[AGB^{+}77] Allen L. Ambler, Donald I. Good, James C. Browne, Wilhelm F. Burger, Richard M. Cohen, Charles G. Hoch, and Robert E. Wells. Gypsy: A language for specification and implementation of verifiable programs. Operating Systems Review, 11(2):1–10, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Almeida:1993:HAR

[AGM93] Carlos Almeida, Brad Glade, and Keith Marzullo. High availability in a real-time system. Operating Systems Review, 27(2): 82–87, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arvind:1977:IMD

[AGP77] Arvind, Kim P. Gostelow, and Wil Plouffe. Indeterminacy, monitors, and dataflow. Operating Systems Review, 11(5):159–169, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Agesen:2010:EXV

[AGSS10] Ole Agesen, Alex Garthwaite, Jeffrey Sheldon, and Pratap Subrahmanyam. The evolution of an x86 virtual machine monitor. Operating Systems Review, 44(4):3–18, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ambler:1977:SPP

[AH77] Allen L. Ambler and Charles G. Hoch. A study of protection in programming languages. Operating Systems Review, 11(2): 25–40, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amit:1980:SSQ

[AH80] Neta Amit and Micha Hofri. A simple semaphore-queue management for multiprocessing systems. Operating Systems Review, 14(3):13–15, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abedi:2015:CRE

[AHB15] Ali Abedi, Andrew Heard, and Tim Brecht. Conducting repeatable experiments and fair comparisons using 802.11n MIMO networks. Operating Systems Review, 49(1):41–50, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amani:2016:CVH

[AHC^{+}16] Sidney Amani, Alex Hixon, Zilin Chen, Christine Rizkallah, Peter Chubb, Liam O’Connor, Joel Beeren, Yutaka Nagashima, Japheth Lim, Thomas Sewell, Joseph Tuong, Gabriele Keller, Toby Murray, Gerwin Klein, and Gernot Heiser. CoGENT: Verifying high-assurance file system implementations. Operating Systems Review, 50(2):175–188, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Acharya:2000:RFR

[AIKS00] Anurag Acharya, Maximilian Ibel, Matthias Koelsch, and Michael Schmitt. RENS: a framework for rapidly evolvable network services. Operating Systems Review, 34(2):37, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Awan:2007:MHS

[AJG07] Asad Awan, Suresh Jagannathan, and Ananth Grama. Macroprogramming heterogeneous sensor networks using cosmos. Operating Systems Review, 41(3):159–172, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Akturk:2017:AAA

[AK17] Ismail Akturk and Ulya R. Karpuzcu. AMNESIAC: Amnesic automatic computer. Operating Systems Review, 51(2):811–824, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Al-Kiswany:2010:CVS

[AKGR10] Samer Al-Kiswany, Abdullah Gharaibeh, and Matei Ripeanu. The case for a versatile storage system. Operating Systems Review, 44(1):10–14, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Agrawal:2025:ELI

[AKP^{+}25] Arney Agrawal, Nitin Kedia, Ashish Panwar, Jayashree Mohan, Nipun Kwatra, Bhargav S. Gulavani, Alexey Tumanov, and Ramachandran Ramjee. Efficient LLM inference via chunked prefills. Operating Systems Review, 59(2):9–16, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arvind:1973:RSG

[AKS73] Arvind, R. Y. Kain, and E. Sadeh. On reference string generation processes. Operating Systems Review, 7(4):80–87, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Appel:1991:VMP

[AL91] Andrew W. Appel and Kai Li. Virtual memory primitives for user programs. Operating Systems Review, 25(3S):96–107, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1991:IAO

[ALBL91] Thomas E. Anderson, Henry M. Levy, Brian N. Bershad, and Edward D. Lazowska. The interaction of architecture and operating system design. Operating Systems Review, 25(3S):108– 120, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ausavarungnirun:2018:MEA

[ALM^{+}18] Rachata Ausavarungnirun, Joshua Landgraf, Vance Miller, Saugata Ghose, Jayneel Gandhi, Christopher J. Rossbach, and Onur Mutlu. Mosaic: Enabling application-transparent support for multiple page sizes in throughput processors. Operating Systems Review, 52(1):27–44, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alsberg:1972:EDF

[Als72] Peter A. Alsberg. Extensible data features in the operating system language OSL/2. Operating Systems Review, 6(1/2):31– 34, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andrews:1977:LFP

[AM77] Gregory R. Andrews and James R. McGraw. Language features for process interaction. Operating Systems Review, 11(2):114– 127, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Allchin:1985:SRA

[AM85] J. E. Allchin and M. S. McKendry. Synchronization and recovery of actions. Operating Systems Review, 19(1):32–45, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Atkinson:1987:DP

[AM87] Russell R. Atkinson and Edward M. McCreight. The dragon processor. Operating Systems Review, 21(4):65–69, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Azmandian:2011:VMM

[AMA^{+}11] Fatemeh Azmandian, Micha Moffie, Malak Alshawabkeh, Jennifer Dy, Javed Aslam, and David Kaeli. Virtual machine monitor-based lightweight intrusion detection. Operating Systems Review, 45(2):38–53, July 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Awad:2016:SSZ

[AMH^{+}16] Amro Awad, Pratyusa Manadhata, Stuart Haber, Yan Solihin, and William Horne. Silent shredder: Zero-cost shredding for secure non-volatile main memory controllers. Operating Systems Review, 50(2):263–276, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ananthanarayanan:1992:EID

[AMMR92] R. Ananthanarayanan, Sathis Menon, Ajay Mohindra, and Umakishore Ramachandran. Experiences in integrating distributed shared memory with virtual memory management. Operating Systems Review, 26(3):4–26, July 1992. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aviv:2012:ETE

[AMO^{+}12] Adam J. Aviv, Vin Mannino, Thanat Owlarn, Seth Shannin, Kevin Xu, and Boon Thau Loo. Experiences in teaching an educational user-level operating systems implementation project. Operating Systems Review, 46(2):80–86, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abernathy:1973:SDGa

[AMPS73a] David H. Abernathy, John S. Mancino, Charls R. Pearson, and Dona C. Swiger. Survey of design goals for operating systems. Operating Systems Review, 7(2):29–48, April 1973. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abernathy:1973:SDGb

[AMPS73b] David H. Abernathy, John S. Mancino, Charls R. Pearson, and Dona C. Swiger. Survey of design goals for operating systems. Operating Systems Review, 7(3):19–34, July 1973. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abernathy:1974:SDG

[AMPS74] David H. Abernathy, John S. Mancino, Charls R. Pearson, and Dona C. Swiger. Survey of design goals for operating systems. Operating Systems Review, 8(1):25–35, January 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aguilera:2007:SNP

[AMS^{+}07] Marcos K. Aguilera, Arif Merchant, Mehul Shah, Alistair Veitch, and Christos Karamanolis. Sinfonia: a new paradigm for building scalable distributed systems. Operating Systems Review, 41 (6):159–174, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amanton:2002:CCP

[AN02] Laurent Amanton and Mohamed Naïmi. The concept of causalphase ordering for overlapped broadcasts. Operating Systems Review, 36(3):67–81, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andrews:1981:NYS

[And81] Gregory R. Andrews. A note from your secretary-treasurer. Operating Systems Review, 15(4):1–2, October 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andrews:1983:RSM

[And83] Gregory R. Andrews. Report to the SIGOPS membership. Operating Systems Review, 17(1):2–3, January 1983. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Andler:1987:FR

[And87] Sten F. Andler. Financial report. Operating Systems Review, 21 (2):3–4, April 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:1995:PCS

[And95] Eric W. Anderson. The performance of the Container Shipping I/O system. Operating Systems Review, 29(5):229, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:2009:CRC

[And09] Thomas Anderson. Conference reviewing considered harmful. Operating Systems Review, 43(2):108–116, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anonymous:1975:ASI

[Ano75] Anonymous. Aims and scope for IEEE Transactions on Software Engineering. Operating Systems Review, 9(2):14–15, April 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anonymous:1978:E

[Ano78] Anonymous. Errata. Operating Systems Review, 12(2):11, April 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anonymous:1986:FR

[Ano86] Anonymous. Financial Report. Operating Systems Review, 20 (1):4, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Antonov:1990:RAO

[Ant90] Vadim G. Antonov. A regular architecture for operating system. Operating Systems Review, 24(3):22–39, July 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amiri:2000:AFP

[APGG00] Khalil Amiri, David Petrou, Greg Ganger, and Garth Gibson. Automatic function placement in distributed storage systems. Operating Systems Review, 34(2):36, April 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Apte:2007:APL

[AR07] Himani Apte and Meenali Rungta. Adding parity to the Linux ext3 file system. Operating Systems Review, 41(1):56–65, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arnaud:2010:ACI

[Arn10] Jean Arnaud. Automated control of Internet services. Operating Systems Review, 44(3):47–52, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abrossimov:1989:GVM

[ARS89] E. Abrossimov, M. Rozier, and M. Shapiro. Generic virtual memory management for operating system kernels. Operating Systems Review, 23(5):123–136, December 1989. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alt:2010:CSH

[AS10] Anne-Marie Alt and Daniel Simon. Control strategies for H.264 video decoding under resources constraints. Operating Systems Review, 44(3):53–58, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ajay:2017:GIL

[ASR^{+}17] Jerry Ajay, Chen Song, Aditya Singh Rathore, Chi Zhou, and Wenyao Xu. 3DGates: an instruction-level energy analysis and optimization of 3D printers. Operating Systems Review, 51(2): 419–433, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Anderson:2010:EM

[AT10] Eric Anderson and Joseph Tucek. Efficiency matters! Operating Systems Review, 44(1):40–45, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ananda:1992:SAR

[ATK92] A. L. Ananda, B. H. Tay, and E. K. Koh. A survey of asynchronous remote procedure calls. Operating Systems Review, 26(2):92–109, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Al-Theneyan:2001:EJU

[ATMZ01] Ahmed Al-Theneyan, Piyush Mehrotra, and Mohammed Zubair. Enhancing Jini for use across non-multicastable networks. Operating Systems Review, 35(2):21–30, April 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Azimi:2009:EOS

[ATSS09] Reza Azimi, David K. Tam, Livio Soares, and Michael Stumm. Enhancing operating system support for multicore processors by using hardware performance monitoring. Operating Systems Review, 43(2):56–65, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Albrecht:2006:PAM

[ATSV06] Jeannie Albrecht, Christopher Tuttle, Alex C. Snoeren, and Amin Vahdat. PlanetLab application management using plush. Operating Systems Review, 40(1):33–40, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Atwood:1984:UTO

[Atw84] J. W. Atwood. Use of Tunis in an operating systems design course. Operating Systems Review, 18(2):6–7, April 1984. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Acharya:1998:ADP

[AUS98] Anurag Acharya, Mustafa Uysal, and Joel Saltz. Active disks: programming model, algorithms and evaluation. Operating Systems Review, 32(5):81–91, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Appavoo:2008:PKB

[AUW08] Jonathan Appavoo, Volkmar Uhlig, and Amos Waterland. Project Kittyhawk: building a global-scale computer: Blue Gene/P as a generic computing platform. Operating Systems Review, 42(1):77–84, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Asmussen:2016:MHO

[AVN^{+}16] Nils Asmussen, Marcus Völp, Benedikt Nöthen, Hermann Härtig, and Gerhard Fettweis. M3: a hardware/operatingsystem co-design to tame heterogeneous manycores. Operating Systems Review, 50(2):189–203, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aksanli:2011:UGE

[AVZR11] Baris Aksanli, Jagannathan Venkatesh, Liuyi Zhang, and Tajana Rosing. Utilizing green energy prediction to schedule mixed batch and service jobs in data centers. Operating Systems Review, 45(3):53–57, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Agarwal:2017:TAT

[AW17]  Neha Agarwal and Thomas F. Wenisch.  Thermostat:

Application-transparent page management for two-tiered main memory. Operating Systems Review, 51(2):631–644, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Angstadt:2016:RPP

[AWS16] Kevin Angstadt, Westley Weimer, and Kevin Skadron. RAPID programming of pattern-recognition processors. Operating Systems Review, 50(2):593–605, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Adjie-Winoto:1999:DII

[AWSBL99] William Adjie-Winoto, Elliot Schwartz, Hari Balakrishnan, and Jeremy Lilley. The design and implementation of an intentional naming system. Operating Systems Review, 33(5):186–201, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Adjie-Winoto:2000:DII

[AWSBL00] William Adjie-Winoto, Elliot Schwartz, Hari Balakrishnan, and Jeremy Lilley. The design and implementation of an intentional naming system. Operating Systems Review, 34(2):22, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Amit:2017:H

[AWT17] Nadav Amit, Michael Wei, and Cheng-Chun Tu. Hypercallbacks. Operating Systems Review, 51(1):54–59, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Acharya:2008:SMC

[AWW08] Arup Acharya, Xiping Wang, and Charles Wright. SIP message classification: design and performance. Operating Systems Review, 42(1):100–101, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abe:2008:EAP

[AYK08] Yoshihisa Abe, Hiroshi Yamada, and Kenji Kono. Enforcing appropriate process execution for exploiting idle resources from outside operating systems. Operating Systems Review, 42(4): 27–40, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Aweke:2016:ASB

[AYQ^{+}16] Zelalem Birhanu Aweke, Salessawi Ferede Yitbarek, Rui Qiao, Reetuparna Das, Matthew Hicks, Yossi Oren, and Todd Austin. ANVIL: Software-based protection against nextgeneration rowhammer attacks. Operating Systems Review, 50 (2):743–755, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Alam:2018:DIY

[AZEE18] Hanna Alam, Tianhao Zhang, Mattan Erez, and Yoav Etsion. Do-it-yourself virtual memory translation. Operating Systems Review, 52(1):1–12, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bansal:2006:AGP

[BA06] Sorav Bansal and Alex Aiken. Automatic generation of peephole superoptimizers. Operating Systems Review, 40(5):394–403, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1990:FTC

[Bab90] Özalp Babaoğlu. Fault-tolerant computing based on Mach. Operating Systems Review, 24(1):27–39, January 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1991:RFA

[Bab91] Özalp Babaoğlu. Report on the fourth ACM SIGOPS European workshop fault tolerance support in distributed systems. Operating Systems Review, 25(1):19–43, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bacon:1981:ADS

[Bac81] Jean Bacon. An approach to distributed software systems. Operating Systems Review, 15(4):62–74, October 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bacon:1991:TRD

[Bac91] David F. Bacon. Transparent recovery in distributed systems (position paper). Operating Systems Review, 25(2):91–94, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bacon:1999:REA

[Bac99] Jean Bacon. Report on the Eighth ACM SIGOPS European Workshop. Operating Systems Review, 33(1):6–17, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bessani:2008:DBF

[BACF08] Alysson Neves Bessani, Eduardo Pelison Alchieri, Miguel Correia, and Joni Silva Fraga. DepSpace: a Byzantine fault-tolerant coordination service. Operating Systems Review, 42(4):163–176, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barreto:2011:ASF

[BAD^{+}11] Luciano Barreto, Aline Andrade, Adolfo Duran, Caique Lima, and Ademilson Lima. Abstract specification and formalization of an operating system kernel in Z. Operating Systems Review, 45(1):156–160, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Badrinath:1993:IMD

[BAI93] B. R. Badrinath, Arup Acharya, and Tomasz Imieliński. Impact of mobility on distributed computations. Operating Systems Review, 27(2):15–20, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baker:1995:GTP

[Bak95] Mary Baker. Going threadbare (panel session): sense or sedition? a debate on the threads abstraction. Operating Systems Review, 29(5):227, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balakrishnan:2024:TCW

[Bal24] Mahesh Balakrishnan. Taming consensus in the wild (with the shared log abstraction). Operating Systems Review, 58(1):1–6, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3689051.3689053.

Bershad:1989:LRP

[BALL89] B. Bershad, T. Anderson, E. Lazowska, and H. Levy. Lightweight remote procedure call. Operating Systems Review, 23(5):102–113, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bugnion:1996:CDP

[BAM^{+}96] Edouard Bugnion, Jennifer M. Anderson, Todd C. Mowry, Mendel Rosenblum, and Monica S. Lam. Compiler-directed page coloring for multiprocessors. Operating Systems Review, 30(5): 244–255, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babonneau:1977:AGS

[BAMM77] J. Y. Babonneau, M. S. Achard, G. Morisset, and M. B. Mounajjed. Automatic and general solution to the adaptation of programs in a paging environment. Operating Systems Review, 11 (5):109–116, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burrows:1989:LA

[BAN89] M. Burrows, M. Abadi, and R. Needham. A logic of authentication. Operating Systems Review, 23(5):1–13, December 1989.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burrows:1990:RN

[BAN90] Michael Burrows, Martín Abadi, and Roger Needham. Rejoinder to Nessett. Operating Systems Review, 24(2):39–40, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barnett:1979:GCV

[Bar79] Jeffrey A. Barnett. Garbage collection versus swapping. Operating Systems Review, 13(3):12–17, July 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bartlett:1981:NK

[Bar81] Joel F. Bartlett. A NonStop kernel. Operating Systems Review, 15(5):22–29, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baruchi:2014:SPT

[Bar14] Artur Baruchi. SOSP Professional Travel Scholarship: Reflections by recipient Artur Baruchi. Operating Systems Review, 48 (2):25, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baskett:1972:DCS

[Bas72] Forest Baskett. The dependence of computer system queues upon processing time distribution and central processor scheduling. Operating Systems Review, 6(1/2):109–113, June 1972. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baset:2012:CSP

[Bas12] Salman A. Baset. Cloud SLAs: present and future. Operating Systems Review, 46(2):57–66, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brundage:1975:CPD

[BB75] Robert E. Brundage and Alan P. Batson. Computational processor demands of Algol-60 programs. Operating Systems Review, 9(5):161–168, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balasubramanian:2017:SPR

[BBB^{+}17] Abhiram Balasubramanian, Marek S. Baranowski, Anton Burtsev, Aurojit Panda, Zvonimir Rakamari, and Leonid Ryzhyk. System programming in Rust: Beyond safety. Operating Systems Review, 51(1):94–99, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Benchaiba:2004:DME

[BBBAN04] M. Benchaïba, A. Bouabdallah, N. Badache, and M. Ahmed-Nacer. Distributed mutual exclusion algorithms in mobile ad hoc networks: an overview. Operating Systems Review, 38(1):74–89, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ball:2006:TSA

[BBC^{+}06] Thomas Ball, Ella Bounimova, Byron Cook, Vladimir Levin, Jakob Lichtenberg, Con McGarvey, Bohus Ondrusek, Sriram K. Rajamani, and Abdullah Ustuner. Thorough static analysis of device drivers. Operating Systems Review, 40(4):73–85, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barr:2002:NSL

[BBD^{+}02] Rimon Barr, John C. Bicket, Daniel S. Dantas, Bowei Du, T. W. Danny Kim, Bing Zhou, and Emin Gün Sirer. On the need for system-level support for ad hoc and sensor networks. Operating Systems Review, 36(2):1–5, April 2002. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barr:2010:VMV

[BBD^{+}10] Ken Barr, Prashanth Bungale, Stephen Deasy, Viktor Gyuris, Perry Hung, Craig Newell, Harvey Tuch, and Bruno Zoppis. The VMware mobile virtualization platform: is that a hypervisor in your pocket? Operating Systems Review, 44(4):124–135, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beschastnikh:2011:MTI

[BBE^{+}11] Ivan Beschastnikh, Yuriy Brun, Michael D. Ernst, Arvind Krishnamurthy, and Thomas E. Anderson. Mining temporal invariants from partially ordered logs. Operating Systems Review, 45 (3):39–46, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bakhshi:2007:FAT

[BBFH07] Rena Bakhshi, Francois Bonnet, Wan Fokkink, and Boudewijn Haverkort. Formal analysis techniques for gossiping protocols. Operating Systems Review, 41(5):28–36, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borg:1983:MSS

[BBG83] Anita Borg, Jim Baumbach, and Sam Glazer. A message system supporting fault tolerance. Operating Systems Review, 17(5):90– 99, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Butenuth:1996:COS

[BBH96] Roger Butenuth, Wolfgang Burke, and Hans-Ulrich Heiß. Cosy: an operating system for highly parallel computers. Operating Systems Review, 30(2):81–91, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bal:2000:DAS

[BBH^{+}00] Henri Bal, Raoul Bhoedjang, Rutger Hofman, Ceriel Jacobs, Thilo Kielmann, Jason Maassen, Rob van Nieuwpoort, John Romein, Luc Renambot, Tim Rühl, Ronald Veldema, Kees Verstoep, Aline Baggio, Gerco Ballintijn, Ihor Kuz, Guillaume Pierre, Maarten van Steen, Andy Tanenbaum, Gerben Doornbos, Desmond Germans, Hans Spoelder, Evert-Jan Baerends, Stan van Gisbergen, Hamideh Afsermanesh, Dick van Albada, Adam Belloum, David Dubbeldam, Zeger Hendrikse, Bob Hertzberger, Alfons Hoekstra, Kamil Iskra, Drona Kandhai, Dennis Koelma, Frank van der Linden, Benno Overeinder, Peter Sloot, Piero Spinnato, Dick Epema, Arjan van Gemund, Pieter Jonker, Andrei Radulescu, Cees van Reeuwijk, Henk Sips, Peter Knijnenburg, Michael Lew, Floris Sluiter, Lex Wolters, Hans Blom, Cees de Laat, and Aad van der Steen. The distributed ASCI Supercomputer project. Operating Systems Review, 34 (4):76–96, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhattiprolu:2008:VSC

[BBHL08] Sukadev Bhattiprolu, Eric W. Biederman, Serge Hallyn, and Daniel Lezcano. Virtual servers and checkpoint/restart in mainstream Linux. Operating Systems Review, 42(5):104–113, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bailey:1981:UDF

[BBM^{+}81] Kirk A. Bailey, Lee Boynton, Paul E. McKenney, Gary J. Oliver, and Dave Regan. User defined files. Operating Systems Review, 15(4):75–84, October 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bratanov:2009:VMW

[BBM09] Stanislav Bratanov, Roman Belenov, and Nikita Manovich. Virtual machines: a whole new world for performance analysis. Operating Systems Review, 43(2):46–55, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bobrow:1972:TPT

[BBMT72] Daniel G. Bobrow, Jerry D. Burchfiel, Daniel L. Murphy, and Raymond S. Tomlinson. TENEX: a paged time sharing system for the PDP-10. Operating Systems Review, 6(1/2):1–10, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barnett:1983:PLP

[BC83] Jeffrey A. Barnett and Alvin S. Cooperband. Priority is a limited property. Operating Systems Review, 17(3):9, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhandarkar:1991:PAC

[BC91a] Dileep Bhandarkar and Douglas W. Clark. Performance from architecture: comparing a RISC and a CISC with similar hardware organization. Operating Systems Review, 25(3S):310–319, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:1991:IPR

[BC91b] Kenneth Birman and Robert Cooper. The ISIS project: real experience with a fault tolerant programming system. Operating Systems Review, 25(2):103–107, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bovet:2001:RBO

[BC01] Daniel P. Bovet and Marco Cesati. A real bottom-up operating systems course. Operating Systems Review, 35(1):48–60, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bindal:2006:CSO

[BC06] Ruchir Bindal and Pei Cao. Can self-organizing P2P file distribution provide QoS guarantees? Operating Systems Review, 40(3):22–30, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Boutcher:2008:PTP

[BC08] David Boutcher and Abhishek Chandra. Practical techniques for purging deleted data using liveness information. Operating Systems Review, 42(5):85–94, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Boutcher:2010:DVM

[BC10] David Boutcher and Abhishek Chandra. Does virtualization make disk scheduling passé? Operating Systems Review, 44 (1):20–24, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Blair:1994:SIW

[BCC^{+}94] G. S. Blair, A. Campbell, G. Coulson, N. Davies, F. Garcia, and D. Shepherd. Summary of the 4th International Workshop on Network and Operating System Support for Digital Audio and Video (NOSSDAV’93). Operating Systems Review, 28(2): 22–33, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouchenak:2013:VCS

[BCC^{+}13] Sara Bouchenak, Gregory Chockler, Hana Chockler, Gabriela Gheorghe, Nuno Santos, and Alexander Shraer. Verifying cloud services: present and future. Operating Systems Review, 47(2): 6–19, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borriello:1987:RVC

[BCDN87] Gaetano Borriello, Andrew R. Cherenson, Peter B. Danzig, and Michael N. Nelson. RISCs vs. CISCs for Prolog: a case study.

Operating Systems Review, 21(4):136–145, October 1987. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bershad:1995:SEM

[BCE^{+}95] Brian N. Bershad, Craig Chambers, Susan Eggers, Chris Maeda, Dylan McNamee, Przemysław Pardyak, Stefan Savage, and Emin Gün Sirer. SPIN—an extensible microkernel for application-specific operating system services. Operating Systems Review, 29(1):74–77, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Berger:2008:TMS

[BCP^{+}08] Stefan Berger, Ramón Cáceres, Dimitrios Pendarakis, Reiner Sailer, Enriquillo Valdez, Ronald Perez, Wayne Schildhauer, and Deepa Srinivasan. TVDc: managing security in the trusted virtual datacenter. Operating Systems Review, 42(1):40–47, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bessani:2014:TPC

[BCR^{+}14] Alysson Bessani, Leucio A. Cutillo, Gianluca Ramunno, Norbert Schirmer, and Paolo Smiraglia. The TClouds platform: From the concept to the implementation of benchmark scenarios. Operating Systems Review, 48(2):13–22, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bortnikov:2010:BBS

[BCRS10] Vita Bortnikov, Gregory Chockler, Alexey Roytman, and Mike Spreitzer. Bulletin board: a scalable and robust eventually consistent shared memory over a peer-to-peer overlay. Operating Systems Review, 44(2):64–70, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Benitez:1991:CGS

[BD91] Manuel E. Benitez and Jack W. Davidson. Code generation for streaming: an access/execute mechanism. Operating Systems Review, 25(3S):132–141, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Budiu:2017:PPL

[BD17] Mihai Budiu and Chris Dodd. The P416 programming language. Operating Systems Review, 51(1):5–14, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Basu:2011:BDA

[BDDMR11] Sumit Basu, John Dunagan, Kevin Duh, and Kiran-Kumar Muniswamy-Reddy. BLR-D: applying bilinear logistic regression to factored diagnosis problems. Operating Systems Review, 45(3):31–38, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bobroff:2008:DJS

[BDF^{+}08] Norman Bobroff, Gargi Dasgupta, Liana Fong, Yanbin Liu, Balaji Viswanathan, Fabio Benedetti, and Jonathan Wagner. A distributed job scheduling and flow management system. Operating Systems Review, 42(1):63–70, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balegas:2015:TFI

[BDF^{+}15] Valter Balegas, Sérgio Duarte, Carla Ferreira, Rodrigo Rodrigues, Nuno Preguiça, Mahsa Najafzadeh, and Marc Shapiro. Towards fast invariant preservation in geo-replicated systems. Operating Systems Review, 49(1):121–125, January 2015. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolosky:2007:FPR

[BDH07] William J. Bolosky, John R. Douceur, and Jon Howell. The Farsite project: a retrospective. Operating Systems Review, 41(2): 17–26, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Butrico:2008:SEE

[BDK^{+}08] Maria Butrico, Dilma Da Silva, Orran Krieger, Michal Ostrowski, Bryan Rosenburg, Dan Tsafrir, Eric Van Hensbergen, Robert W. Wisniewski, and Jimi Xenidis. Specialized execution environments. Operating Systems Review, 42(1):106–107, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1997:GMV

[BDM97] Özalp Babaoğlu, Renzo Davoli, and Alberto Montresor. Group membership and view synchrony in partitionable asynchronous distributed systems: specifications. Operating Systems Review, 31(2):11–22, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1998:SSP

[BDMS98] Özalp Babaoğlu, Renzo Davoli, Alberto Montresor, and Roberto Segala. System support for partition-aware network applications. Operating Systems Review, 32(1):41–56, January 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bugnion:1997:DRC

[BDR97] Edouard Bugnion, Scott Devine, and Mendel Rosenblum. Disco: running commodity operating systems on scalable multiprocessors. Operating Systems Review, 31(5):143–156, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baldwin:2009:PSS

[BDS^{+}09] Adrian Baldwin, Chris Dalton, Simon Shiu, Krzysztof Kostienko, and Qasim Rajpoot. Providing secure services for a virtual infrastructure. Operating Systems Review, 43(1):44–51, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolosky:2000:MDC

[BDT00] William J. Bolosky, John R. Doucher, and Marvin Theimer. Mutually-distrusting cooperative file systems. Operating Systems Review, 34(2):29–30, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beckmann:1975:BR

[Bec75] Petr Beckmann. Book review. Operating Systems Review, 9(2): 8–9, April 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beck:1990:AMA

[Bec90] Bob Beck. AAMP: a multiprocessor approach for operating system and application migration. Operating Systems Review, 24(2):41–55, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bradlee:1991:IRA

[BEH91] David G. Bradlee, Susan J. Eggers, and Robert R. Henry. Integrating register allocation and instruction scheduling for RISCs.

Operating Systems Review, 25(3S):122–131, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Belpaire:1975:SSP

[Bel75] Gerald Belpaire. Synchronization: Is a synthesis of the problems possible? Operating Systems Review, 9(3):3–10, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burrows:2000:EFV

[BEL^{+}00] M. Burrows, U. Erlingson, S-T. A. Leung, M. T. Vandevoorde, C. A. Waldspurger, K. Walker, and W. E. Weihl. Efficient and flexible value sampling. Operating Systems Review, 34(5): 160–167, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Belson:2010:ASI

[Bel10] David Belson. Akamai State of the Internet Report, Q4 2009. Operating Systems Review, 44(3):27–37, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:1975:GMC

[BEW75] R. R. Brown, J. L. Elshoff, and M. R. Ward. The GM multiple console time sharing system. Operating Systems Review, 9(4):7– 17, October 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:1976:GMC

[BEW76] R. R. Brown, J. L. Elshoff, and M. R. Ward. The GM multiple console time sharing system. Operating Systems Review, 10(1): 17, January 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bisiani:1987:ASM

[BF87] Roberto Bisiani and Alessandro Forin. Architectural support for multilanguage parallel programming on heterogeneous systems. Operating Systems Review, 21(4):21–30, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bahmann:2008:EFK

[BF08] Helge Bahmann and Konrad Froitzheim. Extending futex for kernel to user notification. Operating Systems Review, 42(5):

18–26, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolosky:1997:DSM

[BFD97] William J. Bolosky, Robert P. Fitzgerald, and John R. Douceur. Distributed schedule management in the Tiger video fileserver. Operating Systems Review, 31(5):212–223, December 1997. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bagley:1975:SDS

[BFHW75] J. D. Bagley, E. R. Floto, S. C. Hsieh, and V. Watson. Sharing data and services in a virtual machine system. Operating Systems Review, 9(5):82–88, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolosky:1989:SET

[BFS89] W. Bolosky, R. Fitzgerald, and M. Scott. Simple but effective techniques for NUMA memory management. Operating Systems Review, 23(5):19–31, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barcellos:1994:HNO

[BFSG94] Antônio Marinho Pilla Barcellos, Valdir Rossi Belmonte Filho, João Frederico Lacava Schramm, and Cláudio Fernando Resin Geyer. The HetNOS network operating system: a tool for writing distributed applications. Operating Systems Review, 28 (4):34–47, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Behnam:2025:EEM

[BFZ^{+}25] Payman Behnam, Yaosheng Fu, Ritchie Zhao, Po-An Tsai, Zhididng Yu, and Alexey Tumanov. EMPIRIC: Exploring missing pieces in KV cache compression for reducing computation, storage, and latency in long-context LLM inference. Operating Systems Review, 59(2):46–54, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1987:SPM

[BGHL87] A. Birrell, J. Guttag, J. Horning, and R. Levin. Synchronization primitives for a multiprocessor: a formal specification. Operating Systems Review, 21(5):94–102, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Blasgen:1979:CP

[BGMP79] Mike Blasgen, Jim Gray, Mike Mitoma, and Tom Price. The convoy phenomenon. Operating Systems Review, 13(2):20–25, April 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouillot:2004:CMD

[BGS04] Nicolas Bouillot and Eric Gressier-Soudan. Consistency models for distributed interactive multimedia applications. Operating Systems Review, 38(4):20–32, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Belpaire:1975:FPR

[BH75] Gerald Belpaire and Nai-Ting Hsu. Formal properties of recursive Virtual Machine architectures. Operating Systems Review, 9(5):89–96, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bernstein:1981:PRT

[BH81] Arthur Bernstein and Paul K. Harter, Jr. Proving real-time properties of programs with temporal logic. Operating Systems Review, 15(5):1–11, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bowman:2021:TNG

[BH21] Benjamin Bowman and H. Howie Huang. Towards nextgeneration cybersecurity with graph AI. Operating Systems Review, 55(1):61–67, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3469379.3469386.

Bhattacharjee:2017:TTP

[Bha17] Abhishek Bhattacharjee. Translation-triggered prefetching. Operating Systems Review, 51(2):63–76, June 2017. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Banikazemi:2008:FMS

[BHB^{+}08] Mohammad Banikazemi, Jim Hafner, Wendy Belluomini, KK Rao, Dan Poff, and Bulent Abali. Flipstone: managing storage with fail-in-place and deferred maintenance service models. Operating Systems Review, 42(1):54–62, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:2019:CHI

[BHD19] Ken Birman, Bharath Hariharan, and Christopher De Sa. Cloud-hosted intelligence for real-time IoT applications. Operating Systems Review, 53(1):7–13, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bryce:1993:MPD

[BHJ^{+}93] C. Bryce, D. Hagimont, P. Joubert, C. Morin, G. Muller, and B. Rochat. Models and paradigms for distributed systems structuring: summary of sessions. Operating Systems Review, 27(2): 56–60, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baker:1991:MDF

[BHK^{+}91] Mary G. Baker, John H. Hartman, Michael D. Kupfer, Ken W. Shirriff, and John K. Ousterhout. Measurements of a distributed file system. Operating Systems Review, 25(5):198–212, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bacon:1994:EAC

[BHLM94] Jean Bacon, Richard Hayton, Sai Lai Lo, and Ken Moody. Extensible access control for a hierarchy of servers. Operating Systems Review, 28(3):4–15, July 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baskett:1977:TCD

[BHM77] Forest Baskett, John H. Howard, and John T. Montague. Task communication in DEMOS. Operating Systems Review, 11(5): 23–31, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bianchini:2017:IDE

[Bia17] Ricardo Bianchini. Improving datacenter efficiency. Operating Systems Review, 51(2):327, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:1985:RFT

[Bir85] Kenneth P. Birman. Replication and fault-tolerance in the ISIS system. Operating Systems Review, 19(5):79–86, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1991:PPS

[Bir91] Andrew Birrell. Position paper for SIGOPS workshop on fault tolerance support in distributed systems. Operating Systems Review, 25(1):83–84, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:1994:RCS

[Bir94] Ken Birman. A response to Cheriton and Skeen’s criticism of causal and totally ordered communication. Operating Systems Review, 28(1):11–21, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:2007:PLG

[Bir07] Ken Birman. The promise, and limitations, of gossip protocols. Operating Systems Review, 41(5):8–13, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bishop:1981:HTG

[Bis81] Matt Bishop. Hierarchical Take-Grant Protection systems. Operating Systems Review, 15(5):109–122, December 1981. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:2007:DHP

[BITW07] Andrew Birrell, Michael Isard, Chuck Thacker, and Ted Wobber. A design for high-performance flash disks. Operating Systems Review, 41(2):88–93, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beckman:2006:OSI

[BIYC06] Pete Beckman, Kamil Iskra, Kazutomo Yoshii, and Susan Coghlan. Operating system issues for petascale systems. Operating Systems Review, 40(2):29–33, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1981:CSB

[BJ81] Özalp Babaoğlu and William Joy. Converting a swap-based system to do paging in an architecture lacking page-referenced bits. Operating Systems Review, 15(5):78–86, December 1981. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:1987:EVS

[BJ87] K. Birman and T. Joseph. Exploiting virtual synchrony in distributed systems. Operating Systems Review, 21(5):123–138, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:2006:MCC

[BJK^{+}06] Özalp Babaoğlu, Márk Jelasity, Anne-Marie Kermarrec, Alberto Montresor, and Maarten van Steen. Managing clouds: a case for a fresh look at large unreliable dynamic networks. Operating Systems Review, 40(3):9–13, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birman:2015:BSP

[BJKT15] Ken Birman, Márk Jelasity, Robert Kleinberg, and Edward Tremel. Building a secure and privacy-preserving Smart Grid. Operating Systems Review, 49(1):131–136, January 2015. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brecht:2006:ENP

[BJL^{+}06] Tim Brecht, G. (John) Janakiraman, Brian Lynn, Vikram Saletore, and Yoshio Turner. Evaluating network processing efficiency with processor partitioning and asynchronous I/O. Operating Systems Review, 40(4):265–278, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Banatre:1991:STM

[BJM^{+}91] M. Banâtre, Ph. Joubert, Ch. Morin, G. Muller, B. Rochat, and P. Sanchez. Stable transactional memories and fault tolerant architectures. Operating Systems Review, 25(1):68–72, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Buzzard:1996:IHS

[BJM^{+}96] Greg Buzzard, David Jacobson, Milon Mackey, Scott Marovich, and John Wilkes. An implementation of the Hamlyn sendermanaged interface architecture. Operating Systems Review, 30 (SI):245–259, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1987:SEI

[BJW87] A. Birrell, M. Jones, and E. Wobber. A simple and efficient implementation of a small database. Operating Systems Review, 21(5):149–154, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bressoud:2008:CRT

[BK08] Thomas C. Bressoud and M. Frans Kaashoek. Chairs’ report on Twenty-First ACM Symposium on Operating Systems Principles. Operating Systems Review, 42(3):123–126, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouti:2012:SCB

[BK12] Adil Bouti and Jörg Keller. Securing cloud-based computations against malicious providers. Operating Systems Review, 46(2): 38–42, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Behnam:2025:TWS

[BKGT25] Payman Behnam, Alind Khare, Dhruv Garg, and Alexey Tumanov. Toward weight sharing paradigm for efficient AI: Training and inference serving. Operating Systems Review, 59(2): 34–45, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bornholt:2016:SCF

[BKL^{+}16] James Bornholt, Antoine Kaufmann, Jialin Li, Arvind Krishnamurthy, Emina Torlak, and Xi Wang. Specifying and checking file system crash-consistency models. Operating Systems Review, 50(2):83–98, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baratto:2005:TVD

[BKN05] Ricardo A. Baratto, Leonard N. Kim, and Jason Nieh. THINC: a virtual display architecture for thin-client computing. Operating Systems Review, 39(5):277–290, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bianchini:1996:HCL

[BKP^{+}96] R. Bianchini, L. I. Kontothanassis, R. Pinto, M. De Maria, M. Abud, and C. L. Amorim. Hiding communication latency and coherence overhead in software DSMs. Operating Systems Review, 30(5):198–209, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bessani:2012:LOW

[BKP^{+}12] Alysson Bessani, Rüdiger Kapitza, Dana Petcu, Paolo Romano, Spyridon V. Gogouvitis, Dimosthenis Kyriazis, and Roberto G. Cascella. A look to the old-world sky: EU-funded dependability cloud computing research. Operating Systems Review, 46(2): 43–56, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beck:1987:VAM

[BKT87] Bob Beck, Bob Kasten, and Shreekant Thakkar. VLSI assist for a multiprocessor. Operating Systems Review, 21(4):10–20, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bayer:1975:MME

[BL75] D. L. Bayer and H. Lycklama. MERT — a multi-environment real-time operating system. Operating Systems Review, 9(5):33– 42, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barkley:1989:LBS

[BL89] R. Barkley and T. Lee. A lazy buddy system bounded by two coalescing delays. Operating Systems Review, 23(5):167–176, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouabdallah:2000:DTB

[BL00] A. Bouabdallah and C. Laforest. A distributed token-based algorithm for the dynamic resource allocation problem. Operating Systems Review, 34(3):60–68, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baquero:2003:TPP

[BL03] Carlos Baquero and Nuno Lopes. Towards peer-to-peer content indexing. Operating Systems Review, 37(4):90–96, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Black:1983:ASC

[Bla83] Andrew P. Black. An asymmetric stream communication system. Operating Systems Review, 17(5):4–10, October 1983. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Black:1985:SDA

[Bla85] Andrew P. Black. Supporting distributed applications: experience with Eden. Operating Systems Review, 19(5):181–193, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Black:1991:UTO

[Bla91] Andrew P. Black. Understanding transactions in the operating system context. Operating Systems Review, 25(1):73–76, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Blackman:1995:BRW

[Bla95] Sally Blackman. Book review: The WEB Empowerment Book, Ralph Abraham, Frank Jas, and Willard Russell. Operating Systems Review, 29(4):2, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bornholt:2016:DBA

[BLC^{+}16] James Bornholt, Randolph Lopez, Douglas M. Carmean, Luis Ceze, Georg Seelig, and Karin Strauss. A DNA-based archival storage system. Operating Systems Review, 50(2):637–649, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bai:2017:VRE

[BLI17] Yuxin Bai, Victor W. Lee, and Engin Ipek. Voltage regulator efficiency aware power management. Operating Systems Review, 51(2):825–838, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barbalace:2017:BBH

[BLJ^{+}17] Antonio Barbalace, Robert Lyerly, Christopher Jelesnianski, Anthony Carno, Ho-Ren Chuang, Vincent Legout, and Binoy Ravindran. Breaking the boundaries in heterogeneous-ISA datacenters. Operating Systems Review, 51(2):645–659, June 2017.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1981:GED

[BLNS81] Andrew D. Birrell, Roy Levin, Roger M. Needham, and Michael D. Schroeder. Grapevine: An exercise in distributed computing. Operating Systems Review, 15(5):178–179, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bershad:1994:ACM

[BLRC94] Brian N. Bershad, Dennis Lee, Theodore H. Romer, and J. Bradley Chen. Avoiding conflict misses dynamically in large direct-mapped caches. Operating Systems Review, 28(5):158– 170, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burdorf:1990:NPT

[BM90] Christopher Burdorf and Jed Marti. Non-preemptive time warp scheduling algorithms. Operating Systems Review, 24(2):7–18, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bagrodia:1991:EIH

[BM91] Rajive Bagrodia and Sharad Mathur. Efficient implementation of high-level parallel programs. Operating Systems Review, 25 (3S):142–151, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baquero:1999:USC

[BM99] Carlos Baquero and Francisco Moura. Using structural characteristics for autonomous operation. Operating Systems Review, 33(4):90–96, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bond:2006:BBE

[BM06] Michael D. Bond and Kathryn S. McKinley. Bell: bit-encoding online memory leak detection. Operating Systems Review, 40 (5):61–72, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brito:2017:BSC

[BM17] Alisson Brito, Jr. and Leandro Becker Rivalino Matias. 6th Brazilian Symposium on Computing System Engineering. Operating Systems Review, 51(1):100, August 2017. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burke:2000:ASF

[BMA00] Jerome Burke, John McDonald, and Todd Austin. Architectural support for fast symmetric-key cryptography. Operating Systems Review, 34(5):178–189, December 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Berger:2000:HSM

[BMBW00] Emery D. Berger, Kathryn S. McKinley, Robert D. Blumofe, and Paul R. Wilson. Hoard: a scalable memory allocator for multithreaded applications. Operating Systems Review, 34(5): 117–128, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Benmohammed-Mahieddine:1994:PSI

[BMD94] K. Benmohammed-Mahieddine and P. M. Dew. A periodic symmetrically-initiated load balancing algorithm for distributed systems. Operating Systems Review, 28(1):66–79, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burtsev:2014:WSL

[BMER14] Anton Burtsev, Nikhil Mishrikoti, Eric Eide, and Robert Ricci. Weir: a streaming language for performance analysis. Operating Systems Review, 48(1):65–70, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Balkind:2016:OOS

[BMF^{+}16] Jonathan Balkind, Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Alexey Lavrov, Mohammad Shahrad, Adi Fuchs, Samuel Payne, Xiaohua Liang, Matthew Matl, and David Wentzlaff. OpenPiton: an open source manycore research framework. Operating Systems Review, 50(2):217–232, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bridges:2006:SSH

[BMK06] Patrick G. Bridges, Arthur B. MacCabe, and Orran Krieger. System software for high end computing. Operating Systems Review, 40(2):6–7, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bronevetsky:2004:ALC

[BMP^{+}04] Greg Bronevetsky, Daniel Marques, Keshav Pingali, Peter Szwed, and Martin Schulz. Application-level checkpointing for shared memory programs. Operating Systems Review, 38(5): 235–247, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brassil:2009:CSH

[BMR^{+}09] Jack Brassil, Rick McGeer, Raj Rajagopalan, Puneet Sharma, Praveen Yalagandula, Sujata Banerjee, David P. Reed, and Sung-Ju Lee. The CHART system: a high-performance, fair transport architecture based on explicit-rate signaling. Operating Systems Review, 43(1):26–35, January 2009. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bacon:1991:MSS

[BMTW91] Jean Bacon, Ken Moody, Sue Thomson, and Tim Wilson. A multi-service storage architecture. Operating Systems Review, 25(4):47–65, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bartoli:1993:WAS

[BMvdV93] Alberto Bartoli, Sape J. Mullender, and Martijn van der Valk. Wide-address spaces: exploring the design space. Operating Systems Review, 27(1):11–17, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borselius:2002:VTS

[BMW02a] Niklas Borselius, Chris J. Mitchell, and Aaron Wilson. On the value of threshold signatures. Operating Systems Review, 36 (4):30–35, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borselius:2002:PAU

[BMW02b] Niklas Borselius, Chris J. Mitchell, and Aaron Wilson. A pragmatic alternative to undetachable signatures. Operating Systems Review, 36(2):6–11, April 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1978:AGC

[BN78a] A. D. Birrell and R. M. Needham. An asynchronous garbage collector for the CAP filing system. Operating Systems Review, 12(2):31–33, April 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1978:CS

[BN78b] A. D. Birrell and R. M. Needham. Character streams. Operating Systems Review, 12(3):29–31, July 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1983:IRP

[BN83] Andrew D. Birrell and Bruce Jay Nelson. Implementing Remote procedure calls. Operating Systems Review, 17(5):3, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:2016:HBS

[BNE16] Fraser Brown, Andres Nötzli, and Dawson Engler. How to build static checking systems using orders of magnitude less code. Operating Systems Review, 50(2):143–157, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Birrell:1993:NO

[BNOW93] Andrew Birrell, Greg Nelson, Susan Owicki, and Edward Wobber. Network objects. Operating Systems Review, 27(5):217–230, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baker:1991:ASD

[BO91] Mary Baker and John Ousterhout. Availability in the Sprite distributed file system. Operating Systems Review, 25(2):95–98, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bull:1999:NMA

[BO99] John A. Bull and David J. Otway. A nested mutual authentication protocol. Operating Systems Review, 33(4):42–47, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barreto:2015:BSC

[BOB15] Raimundo Barreto, Rafael Obelheiro, and Leandro Becker. Brazilian Symposium on Computing System Engineering. Operating Systems Review, 49(2):1–2, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bochmann:1975:CPE

[Boc75] Gregor V. Bochmann. Communication protocols and error recovery procedures. Operating Systems Review, 9(3):45–50, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bodik:2011:OWM

[Bod11] Peter Bodík. Overview of the workshop on managing large-scale systems via the analysis of system logs and the application of machine learning techniques. Operating Systems Review, 45(3): 20–22, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Boettiger:2015:IDR

[Boe15] Carl Boettiger. An introduction to Docker for reproducible research. Operating Systems Review, 49(1):71–79, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borghoff:1992:DOD

[Bor92] Uwe M. Borghoff. Design of optimal distributed file systems: a framework for research. Operating Systems Review, 26(4):30– 61, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bordawekar:1998:CCF

[Bor98] Rajesh Bordawekar. A case for compositional file systems (extended abstract). Operating Systems Review, 32(3):72–80, July 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bosse:2006:VFA

[Bos06] Stefan Bosse. VAMNET: the functional approach to distributed programming. Operating Systems Review, 40(3):108–114, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouabdallah:1994:MEF

[Bou94] Abdelmadjid Bouabdallah. On mutual exclusion in faulty distributed systems. Operating Systems Review, 28(1):80–87, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Broner:1991:IRB

[BP91] Garbriel Broner and Patrick Powell. Intelligent I/O rule-based input/output processing for operating systems. Operating Systems Review, 25(3):10–26, July 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bisson:2012:DFF

[BPP12] Tim Bisson, Yuvraj Patel, and Shankar Pasupathy. Designing a fast file system crawler with incremental differencing. Operating Systems Review, 46(3):11–19, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barnes:2009:CPO

[BR09] Frederick R. M. Barnes and Carl G. Ritson. Checking processoriented operating system behaviour using CSP and refinement. Operating Systems Review, 43(4):45–49, December 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bouchenak:2010:SIW

[BR10] Sara Bouchenak and Eric Rutten. Summary of the 5th International Workshop on Feedback Control Implementation and Design in Computing Systems and Networks (FeBID 2010). Operating Systems Review, 44(3):38–40, July 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brereton:1983:DRI

[Bre83] Pearl Brereton. Detection and resolution of inconsistencies among distributed replicates of files. Operating Systems Review, 17(1):10–15, January 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bressoud:2008:SSN

[Bre08] Thomas C. Bressoud. Session scribe notes for Twenty-First ACM Symposium on Operating Systems Principles. Operating Systems Review, 42(3):136–151, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:1975:MCT

[Bro75] R. R. Brown. MCTS customer task environment. Operating Systems Review, 9(4):18–42, October 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:1976:MNP

[Bro76] Ralph R. Brown. MCTS nucleus: philosophy and praxis. Operating Systems Review, 10(1):39–60, January 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:2000:TBMa

[Bro00a] Aaron Brown. Towards benchmarks for maintainability, availability and growth/evolution. Operating Systems Review, 34(2): 34, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brown:2000:TBMb

[Bro00b] Aaron Brown. Towards benchmarks for maintainability, availability and growth/evolution(MAGE). Operating Systems Review, 34(2):41, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Beder:2000:AFT

[BRR^{+}00] D. M. Beder, A. Romanovsky, B. Randell, C. R. Snow, and R. J. Stroud. An application of fault tolerance patterns and coordinated atomic actions to a problem in railway scheduling. Operating Systems Review, 34(4):21–31, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brumfield:1986:GOS

[Bru86] Jeffrey A. Brumfield. A Guide To Operating Systems Literature. Operating Systems Review, 20(2):38–42, April 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Braunstein:1989:IEU

[BRW89] A. Braunstein, M. Riley, and J. Wilkes. Improving the efficiency of UNIX buffer caches. Operating Systems Review, 23(5):71–82, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Braban:1989:WSP

[BS89] Bruno Braban and Peter Schlenk. A well structured parallel file system for PM. Operating Systems Review, 23(2):25–38, April 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Babaoglu:1995:GCL

[BS95a] Özalp Babaoğlu and André Schiper. On group communication in large-scale distributed systems. Operating Systems Review, 29(1):62–67, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bressoud:1995:HBF

[BS95b] T. C. Bressoud and F. B. Schneider. Hypervisor-based fault tolerance. Operating Systems Review, 29(5):1–11, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Brustoloni:1996:EBS

[BS96] José Carlos Brustoloni and Peter Steenkiste. Effects of buffering semantics on I/O performance. Operating Systems Review, 30 (SI):277–291, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Butts:2002:DDI

[BS02] J. Adam Butts and Guri Sohi. Dynamic dead-instruction detection and elimination. Operating Systems Review, 36(5):199–210, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Borchert:2015:HLM

[BS15] Christoph Borchert and Olaf Spinczyk. Hardening an L4 microkernel against soft errors by aspect-oriented programming and whole-program analysis. Operating Systems Review, 49(2):37– 43, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bolosky:1991:NPT

[BSF^{+}91] William J. Bolosky, Michael L. Scott, Robert P. Fitzgerald, Robert J. Fowler, and Alan L. Cox. NUMA policies and their relation to memory architecture. Operating Systems Review, 25 (3S):212–221, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ballapuram:2008:EAS

[BSL08] Chinnakrishnan S. Ballapuram, Ahmad Sharif, and Hsien-Hsin S. Lee. Exploiting access semantics and program behavior to reduce snoop power in chip multiprocessors. Operating Systems Review, 42(2):60–69, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bairavasundaram:2012:RRS

[BSM^{+}12] Lakshmi N. Bairavasundaram, Gokul Soundararajan, Vipul Mathur, Kaladhar Voruganti, and Kiran Srinivasan. Responding rapidly to service level violations using virtual appliances. Operating Systems Review, 46(3):32–40, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Burcea:2008:PV

[BSMF08] Ioana Burcea, Stephen Somogyi, Andreas Moshovos, and Babak Falsafi. Predictor virtualization. Operating Systems Review, 42 (2):157–167, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bershad:1995:ESP

[BSP^{+}95] B. N. Bershad, S. Savage, P. Pardyak, E. G. Sirer, M. E. Fiuczynski, D. Becker, C. Chambers, and S. Eggers. Extensibility safety and performance in the SPIN operating system. Operating Systems Review, 29(5):267–283, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Baker:2006:FLR

[BSR^{+}06a] Mary Baker, Mehul Shah, David S. H. Rosenthal, Mema Roussopoulos, Petros Maniatis, TJ Giuli, and Prashanth Bungale. A fresh look at the reliability of long-term digital storage. Operating Systems Review, 40(4):221–234, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Binkert:2006:INI

[BSR06b] Nathan L. Binkert, Ali G. Saidi, and Steven K. Reinhardt. Integrated network interfaces for high-bandwidth TCP/IP. Operating Systems Review, 40(5):315–324, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhat:2015:HEE

[BSR^{+}15] Sharath K. Bhat, Ajithchandra Saya, Hemedra K. Rawat, Antonio Barbalace, and Binoy Ravindran. Harnessing energy efficiency of heterogeneous-ISA platforms. Operating Systems Review, 49(2):65–69, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhargava:2008:ATD

[BSSM08] Ravi Bhargava, Benjamin Serebrin, Francesco Spadini, and Srilatha Manne. Accelerating two-dimensional page walks for virtualized systems. Operating Systems Review, 42(2):26–35, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bush:1987:CSR

[BSUH87] William R. Bush, A. Dain Samples, David Ungar, and Paul N. Hilfinger. Compiling Smalltalk-80 to a RISC. Operating Systems Review, 21(4):112–116, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bakht:2011:SHM

[BTK11] Mehedi Bakht, Matt Trower, and Robin Kravets. Searchlight: helping mobile devices find their neighbors. Operating Systems Review, 45(3):71–76, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Basak:2010:VNS

[BTMS10] Debashis Basak, Rohit Toshniwal, Serge Maskalik, and Allwyn Sequeira. Virtualizing networking and security in the cloud. Operating Systems Review, 44(4):86–94, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Buckle:1977:RDT

[Buc77] Normand Buckle. Restricted data types, specification and enforcement of invariant properties of variables. Operating Systems Review, 11(2):68–76, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Budiu:2004:SC

[BVCG04] Mihai Budiu, Girish Venkataramani, Tiberiu Chelcea, and Seth Copen Goldstein. Spatial computation. Operating Systems Review, 38(5):14–26, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhoedjang:2000:EDA

[BVR^{+}00] Raoul A. F. Bhoedjang, Kees Verstoep, Tim Rühl, Henri E. Bal, and Rutger F. H. Hofman. Evaluating design alternatives for reliable communication on high-speed networks. Operating Systems Review, 34(5):71–81, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ballintijn:2000:CIP

[BvS00] Gerco Ballintijn and Maarten van Steen. Characterizing Internet performance to support wide-area application development. Operating Systems Review, 34(4):41–47, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Black:1995:ORH

[BW95] Andrew P. Black and Jonathan Walpole. Objects to the rescue! or httpd: the next generation operating system. Operating Systems Review, 29(1):91–95, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bozyigit:2001:ULP

[BW01] M. Bozyigit and M. Wasiq. User-level process checkpoint and restore for migration. Operating Systems Review, 35(2):86–96, April 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comment [RS02].

Basak:2012:MBD

[BWV^{+}12] Jayanta Basak, Kushal Wadhwani, Kaladhar Voruganti, Srinivasan Narayanamurthy, Vipul Mathur, and Siddhartha Nandi. Model building for dynamic multi-tenant provider environments. Operating Systems Review, 46(3):20–31, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhaskaran:2014:BCS

[BXS14] Meenakshi Sundaram Bhaskaran, Jian Xu, and Steven Swanson. Bankshot: caching slow storage in fast non-volatile memory. Operating Systems Review, 48(1):73–81, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ben-Yehuda:2008:AHS

[BY08] Muli Ben-Yehuda. 1st Annual Haifa Systems and Storage Conference (SYSTOR 2007): a message from the organizers. Operating Systems Review, 42(1):110, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ben-Yehuda:2008:OSF

[BYV08] Muli Ben-Yehuda and Eric Van Hensbergen. Open source as a foundation for systems research. Operating Systems Review, 42(1):2–4, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ben-Yehuda:2008:MGR

[BYVF08] Muli Ben-Yehuda, Eric Van Hensbergen, and Marc Fiuczynski. Minding the gap: R&D in the Linux kernel. Operating Systems Review, 42(5):1–3, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cabrera:1990:TSS

[Cab90] Luis-Felipe Cabrera. Technical summary of the second IEEE workshop on workstation operating systems. Operating Systems Review, 24(3):7–21, July 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chase:1989:ASP

[CAL^{+}89] J. Chase, F. Amador, E. Lazowska, H. Levy, and R. Littlefield. The Amber system: parallel programming on a network of multiprocessors. Operating Systems Review, 23(5):147–158, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Castro:1997:HHA

[CALM97] Miguel Castro, Atul Adya, Barbara Liskov, and Andrew C. Meyers. HAC: hybrid adaptive caching for distributed storage systems. Operating Systems Review, 31(5):102–115, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carlsen:1994:OPA

[Car94] Ulf Carlsen. Optimal privacy and authentication on a portable communications system. Operating Systems Review, 28(3):16– 23, July 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Casey:1991:LND

[Cas91] Liam Casey. Lessons from Norstar’s distributed call processing system. Operating Systems Review, 25(2):108–111, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Castori:1995:SRM

[Cas95] Pierre Castori. Semaphores revisited with MMS. Operating Systems Review, 29(3):49–63, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chase:2001:MES

[CAT^{+}01] Jeffrey S. Chase, Darrell C. Anderson, Prachi N. Thakar, Amin M. Vahdat, and Ronald P. Doyle. Managing energy and server resources in hosting centers. Operating Systems Review, 35(5):103–116, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Choffnes:2008:MPM

[CAW08] David Choffnes, Mark Astley, and Michael J. Ward. Migration policies for multi-core fair-share scheduling. Operating Systems Review, 42(1):92–93, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1993:IOS

[CB93] J. Bradley Chen and Brian N. Bershad. The impact of operating system structure on memory system performance. Operating Systems Review, 27(5):120–133, December 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Coulson:1995:APT

[CB95] Geoff Coulson and Gordon Blair. Architectural principles and techniques for distributed multimedia application support in operating systems. Operating Systems Review, 29(4):17–24, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:2017:EAT

[CB17] Guilherme Cox and Abhishek Bhattacharjee. Efficient address translation for architectures with multiple page sizes. Operating Systems Review, 51(2):435–448, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:2008:XEO

[CBC^{+}08] Russ Cox, Tom Bergan, Austin T. Clements, Frans Kaashoek, and Eddie Kohler. Xoc, an extension-oriented compiler for systems programming. Operating Systems Review, 42(2):244–254, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1998:UMA

[CBD^{+}98] Yuqun Chen, Angelos Bilas, Stefanos N. Damianakis, Cezary Dubnicki, and Kai Li. UTLB: a mechanism for address translation on network interfaces. Operating Systems Review, 32(5): 193–204, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chase:1992:OSA

[CBHLL92] Jeff Chase, Miche Baker-Harvey, Hank Levy, and Ed Lazowska. Opal: A Single Address Space System for 64-bit Architectures. Operating Systems Review, 26(2):9, April 1992. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carter:1991:IPM

[CBZ91] John B. Carter, John K. Bennett, and Willy Zwaenepoel. Implementation and performance of Munin. Operating Systems Review, 25(5):152–164, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Causot:1977:SDD

[CC77] Patrick Causot and Radhia Cousot. Static determination of dynamic properties of generalized type unions. Operating Systems Review, 11(2):77–94, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:1997:FTD

[CC97] Ye-In Chang and Yao-Jen Chang. A fault-tolerant dynamic triangular mesh protocol for distributed mutual exclusion. Operating Systems Review, 31(3):29–44, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:SES

[CC04] Ya-Fen Chang and Chin-Chen Chang. A secure and efficient strong-password authentication protocol. Operating Systems Review, 38(3):79–90, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2005:EAP

[CC05] Ya-Fen Chang and Chin-Chen Chang. An efficient authentication protocol for mobile satellite communication systems. Operating Systems Review, 39(1):70–84, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2021:WDF

[CC21] Rong Chen and Haibo Chen. Wukong: a distributed framework for fast and concurrent graph querying. Operating Systems Review, 55(1):77–83, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3469379.3469388.

Corbo:2006:SNE

[CCAP06] Jacomo Corbo, Antoni Calvó-Armengol, and David Parkes. A study of Nash equilibrium in contribution games for peer-to-peer networks. Operating Systems Review, 40(3):61–66, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Coons:2006:SPS

[CCB^{+}06] Katherine E. Coons, Xia Chen, Doug Burger, Kathryn S. McKinley, and Sundeep K. Kushwaha. A spatial path scheduling algorithm for EDGE architectures. Operating Systems Review, 40(5):129–140, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Costa:2005:VEE

[CCC^{+}05] Manuel Costa, Jon Crowcroft, Miguel Castro, Antony Rowstron, Lidong Zhou, Lintao Zhang, and Paul Barham. Vigilante: endto-end containment of Internet worms. Operating Systems Review, 39(5):133–147, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chou:2000:UML

[CCEH00] Andy Chou, Benjamin Chelf, Dawson Engler, and Mark Heinrich. Using meta-level compilation to check FLASH protocol code. Operating Systems Review, 34(5):59–70, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cena:1995:TRE

[CCG95] M. Cena, M. L. Crespo, and R. Gallard. Transparent remote execution in LAHNOS by means of a neural network device. Operating Systems Review, 29(1):17–28, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chow:1987:HMA

[CCH^{+}87] F. Chow, S. Correll, M. Himelstein, E. Killian, and L. Weber. How many addressing modes are enough? Operating Systems Review, 21(4):117–121, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2011:EAW

[CCHV11] Jie Chen, Ron C. Chiang, H. Howie Huang, and Guru Venkataramani. Energy-aware writes to non-volatile main memory. Operating Systems Review, 45(3):48–52, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:IDA

[CCK04a] Ya-Fan Chang, Chin-Chen Chang, and Chia-Lin Kao. An improvement on a deniable authentication protocol. Operating Systems Review, 38(3):65–74, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:SOT

[CCK04b] Ya-Fen Chang, Chin-Chen Chang, and Jui-Yi Kuo. A secure one-time password authentication scheme using Smart Cards without limiting login times. Operating Systems Review, 38 (4):80–90, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:1981:UMI

[CCLP81] George W. Cox, William M. Corwin, Konrad K. Lai, and Fred J. Pollack. A unified model and implementation for interprocess communication in a multiprocessor environment. Operating Systems Review, 15(5):125–126, December 1981. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1996:ABP

[CCM96] I-Cheng K. Chen, John T. Coffey, and Trevor N. Mudge. Analysis of branch prediction via data compression. Operating Systems Review, 30(5):128–137, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Castro:2008:BBR

[CCM08] Miguel Castro, Manuel Costa, and Jean-Philippe Martin. Better bug reporting with better privacy. Operating Systems Review, 42(2):319–328, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chick:2016:SKG

[CCS^{+}16] Oliver R. A. Chick, Lucian Carata, James Snee, Nikilesh Balakrishnan, and Ripduman Sohan. Shadow kernels: a general mechanism for kernel specialization in existing operating systems. Operating Systems Review, 50(1):3–8, January 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chiueh:1999:ISP

[cCVP99] Tzi cker Chiueh, Ganesh Venkitachalam, and Prashant Pradhan. Integrating segmentation and paging protection for safe, efficient and transparent software extensions. Operating Systems Review, 33(5):140–153, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chinya:2011:BDP

[CCW^{+}11] Gautham N. Chinya, Jamison D. Collins, Perry H. Wang, Hong Jiang, Guei-Yuan Lueh, Thomas A. Piazza, and Hong Wang. Bothnia: a dual-personality extension to the Intel integrated graphics driver. Operating Systems Review, 45(1):11–20, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chanda:2007:WTP

[CCZ07a] Anupam Chanda, Alan L. Cox, and Willy Zwaenepoel. Whodunit: transactional profiling for multi-tier applications. Operating Systems Review, 41(3):17–30, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Costa:2007:BSS

[CCZ^{+}07b] Manuel Costa, Miguel Castro, Lidong Zhou, Lintao Zhang, and Marcus Peinado. Bouncer: securing software by blocking bad input. Operating Systems Review, 41(6):117–130, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1995:LVM

[CD95a] D. R. Cheriton and K. J. Duda. Logged virtual memory. Operating Systems Review, 29(5):26–38, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1995:CMO

[CD95b] David R. Cheriton and Kenneth J. Duda. A caching model of operating system kernel functionality. Operating Systems Review, 29(1):83–86, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Castro:2002:SRS

[CDG^{+}02] Miguel Castro, Peter Druschel, Ayalvadi Ganesh, Antony Rowstron, and Dan S. Wallach. Secure routing for structured peerto-peer overlay networks. Operating Systems Review, 36(5S): 299–314, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chisnall:2017:CJS

[CDG^{+}17] David Chisnall, Brooks Davis, Khilan Gudka, David Brazdil, Alexandre Joannou, Jonathan Woodruff, A. Theodore Markettos, J. Edward Maste, Robert Norton, Stacey Son, Michael Roe, Simon W. Moore, Peter G. Neumann, Ben Laurie, and Robert N. M. Watson. CHERI JNI: Sinking the Java security model into the C. Operating Systems Review, 51(2):569–583, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandra:1994:SPM

[CDV^{+}94] Rohit Chandra, Scott Devine, Ben Verghese, Anoop Gupta, and Mendel Rosenblum. Scheduling and page migration for multiprocessor compute servers. Operating Systems Review, 28(5): 12–24, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cherupalli:2017:DAS

[CDY^{+}17] Hari Cherupalli, Henry Duwe, Weidong Ye, Rakesh Kumar, and John Sartori. Determining application-specific peak power and energy requirements for ultra-low power processors. Operating Systems Review, 51(2):3–16, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Canas:1988:PUO

[CE88] Daniel A. Cañas and Laura M. Esquivel. Portability and the UNIX operating system. Operating Systems Review, 22(2):6– 23, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1995:MPP

[CEC^{+}95] J. B. Chen, Y. Endo, K. Chan, D. Mazieres, A. Dias, M. Seltzer, and M. D. Smith. The measured performance of personal computer operating systems. Operating Systems Review, 29(5): 299–313, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cecchet:2000:DSM

[Cec00] Emmanuel Cecchet. Distributed shared memory for large computing clusters based on memory-mapped networks (poster session). Operating Systems Review, 34(2):35, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cerf:1975:FIC

[Cer75] Vinton Cerf. Formalisms for interprocess communication. Operating Systems Review, 9(3):43–44, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandra:2000:DQT

[CEV00] Surendar Chandra, Carla Schlatter Ellis, and Amin Vahdat. Differentiated QoS through quality aware transformation of Web content. Operating Systems Review, 34(2):36, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:1989:ICM

[CF89] A. Cox and R. Fowler. The implementation of a coherent memory abstraction on a NUMA multiprocessor: experiences with platinum. Operating Systems Review, 23(5):32–44, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chamberlin:1973:PAS

[CFL73] Donald D. Chamberlin, Samuel H. Fulier, and Leonard Y. Liu. A page allocation strategy for multiprogramming systems. Operating Systems Review, 7(4):66–72, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Citron:1998:AMM

[CFR98] Daniel Citron, Dror Feitelson, and Larry Rudolph. Accelerating multi-media processing by implementing memoing in multiplication and division units. Operating Systems Review, 32(5): 252–261, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carriero:1985:NLK

[CG85] Nicholas Carriero and David Gelernter. The S/Net’s Linda kernel (extended abstract). Operating Systems Review, 19(5):160, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cate:1991:CCC

[CG91] Vincent Cate and Thomas Gross. Combining the concepts of compression and caching for a two-level filesystem. Operating Systems Review, 25(3S):200–211, April 1991. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ciancarini:1993:LMM

[CG93] P. Ciancarini and N. Guerrini. Linda meets Minix. Operating Systems Review, 27(4):76–92, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calder:1994:RBC

[CG94] Brad Calder and Dirk Grunwald. Reducing branch costs via branch alignment. Operating Systems Review, 28(5):242–251, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2000:AGP

[CG00] Fay Chang and Garth Gibson. Automatic generation of I/O prefetching hints through speculative execution (poster session).

Operating Systems Review, 34(2):35, April 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chilimbi:2006:HIH

[CG06] Trishul M. Chilimbi and Vinod Ganapathy. HeapMD: identifying heap-based bugs using anomaly detection. Operating Systems Review, 40(5):219–228, December 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Costa:2007:EIC

[CGJ^{+}07] Paolo Costa, Vincent Gramoli, Márk Jelasity, Gian Paolo Jesi, Erwan Le Merrer, Alberto Montresor, and Leonardo Querzoni. Exploring the interdisciplinary connections of gossip-based systems. Operating Systems Review, 41(5):51–60, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2011:LBA

[CGKM11] Shimin Chen, Phillip B. Gibbons, Michael Kozuch, and Todd C. Mowry. Log-based architectures: using multicore to help software behave correctly. Operating Systems Review, 45(1):84–91, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2008:OVB

[CGL^{+}08] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualization-based approach to retrofitting protection in commodity operating systems. Operating Systems Review, 42(2):2–13, March 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1997:AUM

[CGM97] Liqun Chen, Dieter Gollmann, and Chris J. Mitchell. Authentication using minimally trusted servers. Operating Systems Review, 31(3):16–28, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cornilleau:1996:CCA

[CGS96a] T. Cornilleau and E. Gressier-Soudan. A combined-consistency approach: sequential & causal-consistency. Operating Systems Review, 30(4):33–44, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Costa:1996:LLL

[CGS^{+}96b] Manuel Costa, Paulo Guedes, Manuel Sequeira, Nuno Neves, and Miguel Castro. Lightweight logging for lazy release consistent distributed shared memory. Operating Systems Review, 30 (SI):59–73, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carr:1981:WSE

[CH81] Richard W. Carr and John L. Hennessy. WSCLOCK—a simple and effective algorithm for virtual memory management. Operating Systems Review, 15(5):87–95, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cooper:1998:CCM

[CH98] Keith D. Cooper and Timothy J. Harvey. Compiler-controlled memory. Operating Systems Review, 32(5):2–11, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chien:2007:SUL

[CH07] Hung-Yu Chien and Chen-Wei Huang. Security of ultralightweight RFID authentication protocols and its improvements. Operating Systems Review, 41(4):83–86, July 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carroll:2014:MMU

[CH14] Aaron Carroll and Gernot Heiser. Mobile multicores: use them or waste them. Operating Systems Review, 48(1):44–48, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chambers:1973:UCS

[Cha73] John M. Chambers. A user-controlled synchronization method. Operating Systems Review, 7(2):16–25, April 1973. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandras:1990:DMP

[Cha90] Rajan G. Chandras. Distributed message passing operating systems. Operating Systems Review, 24(1):7–17, January 1990. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chavez:1991:XTS

[Chá91] Jorge Buenabadd Chávez. XINIX time-sharing operating system. Operating Systems Review, 25(4):22–34, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:1996:DRS

[Cha96] Ye-In Chang. A dynamic request set based algorithm for mutual exclusion in distributed systems. Operating Systems Review, 30(2):52–62, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Connors:2000:HSD

[CHCmWH00] Daniel A. Connors, Hillery C. Hunter, Ben-Chung Cheng, and Wen mei W. Hwu. Hardware support for dynamic activation of compiler-directed computation reuse. Operating Systems Review, 34(5):222–233, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1975:RPS

[Che75a] Robert C. Chen. Representation of process synchronization. Operating Systems Review, 9(3):37–42, July 1975. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1975:ICS

[Che75b] T. C. Chen. Interprocess communication systems. Operating Systems Review, 9(3):63, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chesson:1975:NUS

[Che75c] Gregory L. Chesson. The network Unix system. Operating Systems Review, 9(5):60–66, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1984:EUR

[Che84] David R. Cheriton. An experiment using registers for fast message-based interprocess communication. Operating Systems Review, 18(4):12–20, October 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1985:PTP

[Che85] David R. Cheriton. Preliminary thoughts on problem-oriented shared memory: a decentralized approach to distributed systems. Operating Systems Review, 19(4):26–33, October 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:2000:Ta

[Che00a] David Cheriton. TRIAD. Operating Systems Review, 34(2):34, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:2000:Tb

[Che00b] David Cheriton. TRIAD. Operating Systems Review, 34(2):41, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2004:IAM

[Che04] Bi-Hui Chen. Improvement of authenticated multiple-key agreement protocol. Operating Systems Review, 38(3):35–41, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2017:BDA

[Che17] Yunji Chen. Big data analytics and intelligence at Alibaba Cloud. Operating Systems Review, 51(2):1, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Colin:2016:EIF

[CHLS16] Alexei Colin, Graham Harvey, Brandon Lucia, and Alanson P. Sample. An energy-interference-free hardware-software debugger for intermittent energy-harvesting systems. Operating Systems Review, 50(2):577–589, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Choen:1977:ITC

[Cho77] Ellis Choen. Information transmission in computational systems. Operating Systems Review, 11(5):133–139, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chu:1975:P

[Chu75] Wesley Chu. Performance. Operating Systems Review, 9(3):75, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cher:2004:SPM

[CHV04] Chen-Yong Cher, Antony L. Hosking, and T. N. Vijaykumar. Software prefetching for mark-sweep garbage collection: hardware analysis and software redesign. Operating Systems Review, 38(5):199–210, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2005:NMS

[CHY05] Ting-Yi Chang, Min-Shiang Hwang, and Wei-Pang Yang. A new multi-stage secret sharing scheme using one-way function. Operating Systems Review, 39(1):48–55, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chun:2010:ECH

[CII^{+}10] Byung-Gon Chun, Gianluca Iannaccone, Giuseppe Iannaccone, Randy Katz, Gunho Lee, and Luca Niccolini. An energy case for hybrid datacenters. Operating Systems Review, 44(1):76– 80, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chase:1993:DSA

[CIL93] Jeff Chase, Valérie Issarnay, and Hank Levy. Distribution in a single address space operating system. Operating Systems Review, 27(2):61–65, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calciu:2023:ULC

[CIP^{+}23] Irina Calciu, M. Talha Imran, Ivan Puddu, Sanidhya Kashyap, Hasan Al Maruf, Onur Mutlu, and Aasheesh Kolli. Using local cache coherence for disaggregated memory systems. Operating Systems Review, 57(1):21–28, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606557.3606561.

Cohen:1975:PHO

[CJ75] Ellis Cohen and David Jefferson. Protection in the Hydra Operating System. Operating Systems Review, 9(5):141–160, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2005:EDS

[CJ05] Yi-Hwa Chen and Jinn-Ke Jan. Enhancement of digital signature with message recovery using self-certified public keys and its variants. Operating Systems Review, 39(3):90–96, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cooksey:2002:SCD

[CJG02] Robert Cooksey, Stephan Jourdan, and Dirk Grunwald. A stateless, content-directed data prefetching mechanism. Operating Systems Review, 36(5):279–290, December 2002. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cosell:1975:OSC

[CJM^{+}75] B. P. Cosell, P. R. Johnson, J. H. Malman, R. E. Schantz, J. Sussman, R. H. Thomas, and D. C. Walden. An operational system for computer resource sharing. Operating Systems Review, 9(5):75–81, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Childers:2015:RPA

[CJM15] Bruce R. Childers, Alex K. Jones, and Daniel Mossé. A roadmap and plan of action for community-supported empirical evaluation in computer architecture. Operating Systems Review, 49 (1):108–117, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Campbell:1987:CCH

[CJR87] Roy Campbell, Garry Johnston, and Vincent Russo. Choices (class hierarchical open interface for custom embedded systems). Operating Systems Review, 21(3):9–17, July 1987. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chockler:2015:LWL

[CJRV15] Gregory Chockler, Flavio Junqueira, Rodrigo Rodrigues, and Ymir Vigfusson. LADIS’14: 8th Workshop on Large-Scale Distributed Systems and Middleware. Operating Systems Review, 49(1):118–120, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2009:CLC

[CJS^{+}09] Huacai Chen, Hai Jin, Zhiyuan Shao, Kan Hu, Ke Yu, and Kun Tian. ClientVisor: leverage COTS OS functionalities for power management in virtualized desktop environment. Operating Systems Review, 43(3):62–71, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheung:1986:SOT

[CK86] David Cheung and Tiko Kameda. Site optimal termination protocols for a distributed database under network partitioning. Operating Systems Review, 20(3):17–27, July 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chaiken:1991:LDS

[CKA91] David Chaiken, John Kubiatowicz, and Anant Agarwal. LimitLESS directories: A scalable cache coherence scheme. Operating Systems Review, 25(3S):224–234, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carter:1994:HSF

[CKD94] Nicholas P. Carter, Stephen W. Keckler, and William J. Dally. Hardware support for fast capability-based addressing. Operating Systems Review, 28(5):319–327, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cmelik:1991:AMS

[CKDK91] Robert F. Cmelik, Shing I. Kong, David R. Ditzel, and Edmund J. Kelly. An analysis of MIPS and SPARC instruction set utilization on the SPEC benchmarks. Operating Systems Review, 25(3S):290–302, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calder:1998:CCD

[CKJA98] Brad Calder, Chandra Krintz, Simmi John, and Todd Austin. Cache-conscious data placement. Operating Systems Review, 32 (5):139–149, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Crameri:2007:SDM

[CKK^{+}07] Olivier Crameri, Nikola Knezevic, Dejan Kostic, Ricardo Bianchini, and Willy Zwaenepoel. Staged deployment in Mirage, an integrated software upgrade testing and distribution system.

Operating Systems Review, 41(6):221–236, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Campbell:1999:OSA

[CKMV99] Andrew T. Campbell, Irene Katzela, Kazuho Miki, and John Vicente. Open Signaling for ATM, INTERNET and Mobile Networks (OPENSIG’98). Operating Systems Review, 33(2):15–28, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2016:DLD

[CKmWH16] Li-Wen Chang, Hee-Seok Kim, and Wen mei W. Hwu. DySel: Lightweight dynamic selection for kernel-based data-parallel programming model. Operating Systems Review, 50(2):667–680, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Coleman:2019:ADT

[CKN^{+}19] Cody Coleman, Daniel Kang, Deepak Narayanan, Luigi Nardi, Tian Zhao, Jian Zhang, Peter Bailis, Kunle Olukotun, Chris Ré, and Matei Zaharia. Analysis of DAWNBench, a time-to-accuracy machine learning performance benchmark. Operating Systems Review, 53(1):14–25, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Callahan:1991:SP

[CKP91] David Callahan, Ken Kennedy, and Allan Porterfield. Software prefetching. Operating Systems Review, 25(3S):40–52, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Craciunas:2008:RMT

[CKR08] Silviu S. Craciunas, Christoph M. Kirsch, and Harald Röck. I/O resource management through system call scheduling. Operating Systems Review, 42(5):44–54, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cargill:1987:CHS

[CL87] T. A. Cargill and B. N. Locanthi. Cheap hardware support for software debugging and profiling. Operating Systems Review, 21 (4):82–83, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheung:1995:EIO

[CL95] W. H. Cheung and Anthony H. S. Loong. Exploring issues of operating systems structuring: from microkernel to extensible systems. Operating Systems Review, 29(4):4–16, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Crampton:2001:AA

[CL01] Jason Crampton and George Loizou. Authorisation and antichains. Operating Systems Review, 35(3):6–15, July 2001. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:CGS

[CL04a] Chin-Chen Chang and Yeu-Pong Lai. A convertible group signature scheme. Operating Systems Review, 38(4):58–65, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:IDM

[CL04b] Chin-Chen Chang and Iuon-Chang Lin. An improvement of delegated multisignature scheme with document decomposition. Operating Systems Review, 38(4):52–57, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:2004:RFB

[CL04c] Chin-Chen Chang and Iuon-Chang Lin. Remarks on fingerprintbased remote user authentication scheme using Smart Cards. Operating Systems Review, 38(4):91–96, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Clark:1985:SSU

[Cla85] David D. Clark. The structuring of systems using upcalls. Operating Systems Review, 19(5):171–180, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Clark:1987:PPV

[Cla87] Douglas W. Clark. Pipelining and performance in the VAX 8800 processor. Operating Systems Review, 21(4):173–177, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chou:2005:XRR

[CLC05] Shihyu Chou, Eric Jui-Lin Lu, and Yi-Hui Chen. X-RDR: a rolebased delegation processor for Web-based information systems. Operating Systems Review, 39(1):4–21, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Criswell:2007:SVA

[CLDA07] John Criswell, Andrew Lenharth, Dinakar Dhurjati, and Vikram Adve. Secure virtual architecture: a safe execution environment for commodity operating systems. Operating Systems Review, 41(6):351–366, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chong:2007:SWA

[CLM^{+}07] Stephen Chong, Jed Liu, Andrew C. Myers, Xin Qi, K. Vikram, Lantian Zheng, and Xin Zheng. Secure Web application via automatic partitioning. Operating Systems Review, 41(6):31– 44, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandra:1994:WTS

[CLR94] Satish Chandra, James R. Larus, and Anne Rogers. Where is time spent in message-passing and shared-memory programs? Operating Systems Review, 28(5):61–73, December 1994. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cantin:2006:SP

[CLS06] Jason F. Cantin, Mikko H. Lipasti, and James E. Smith. Stealth prefetching. Operating Systems Review, 40(5):274–282, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Coffman:1975:SSR

[CM75] E. G. Coffman, Jr. and I. Mitrani. Selecting a scheduling rule that meets pre-specified response time demands. Operating Systems Review, 9(5):187–191, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chang:1987:SAP

[CM87] A. Chang and M. Mergen. 801 Storage: architecture and programming. Operating Systems Review, 21(5):109–110, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Clulow:2006:SCG

[CM06] Jolyon Clulow and Tyler Moore. Suicide for the common good: a new strategy for credential revocation in self-organizing systems. Operating Systems Review, 40(3):18–21, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Correia:2013:DIC

[CM13] Miguel Correia and Neeraj Mittal. Dependability issues in cloud computing: extended papers from the 1st International Workshop on Dependability Issues in Cloud Computing — DISCCO. Operating Systems Review, 47(2):20–22, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Correia:2014:CCD

[CM14] Miguel Correia and Neeraj Mittal. Cloud computing dependability: Report and extended papers of the Second International Workshop on Dependability Issues in Cloud Computing. Operating Systems Review, 48(2):1–2, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chakravorty:2006:HCS

[CMK^{+}06] Sayantan Chakravorty, Celso L. Mendes, Laxmikant V. Kalé, Terry Jones, Andrew Tauferner, Todd Inglett, and José Moreira. HPC-Colony: services and interfaces for very large systems. Operating Systems Review, 40(2):43–49, April 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chung:2006:TTM

[CMM^{+}06] JaeWoong Chung, Chi Cao Minh, Austen McDonald, Travis Skare, Hassan Chafi, Brian D. Carlstrom, Christos Kozyrakis, and Kunle Olukotun. Tradeoffs in transactional memory virtualization. Operating Systems Review, 40(5):371–381, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1977:TPR

[CMMS77] David R. Cheriton, Michael A. Malcolm, Lawrence S. Melen, and Gary R. Sager. Thoth, a portable real-time operating system (extended abstract). Operating Systems Review, 11(5):171, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:2002:PMB

[CMN02] Landon P. Cox, Christopher D. Murray, and Brian D. Noble. Pastiche: making backup cheap and easy. Operating Systems Review, 36(5S):285–298, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chun:2007:AAO

[CMSK07] Byung-Gon Chun, Petros Maniatis, Scott Shenker, and John Kubiatowicz. Attested append-only memory: making adversaries stick to their word. Operating Systems Review, 41(6): 189–204, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carr:1994:COI

[CMT94] Steve Carr, Kathryn S. McKinley, and Chau-Wen Tseng. Compiler optimizations for improving data locality. Operating Systems Review, 28(5):252–262, December 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandy:2007:RTP

[CN07] John A. Chandy and Sumit Narayan. Reliability tradeoffs in personal storage systems. Operating Systems Review, 41(1):37– 41, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1996:RFC

[CNC^{+}96] Peter M. Chen, Wee Teck Ng, Subhachandra Chandra, Christopher Aycock, Gurushankar Rajamani, and David Lowell. The Rio file cache: surviving operating system crashes. Operating Systems Review, 30(5):74–83, December 1996. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chanson:1989:BMG

[CNL89] S. T. Chanson, G. W. Neufeld, and L. Liang. A bibliography on multicast and group communications. Operating Systems Review, 23(4):20–25, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Colwell:1987:VAT

[CNO^{+}87] Robert P. Colwell, Robert P. Nix, John J. O’Donnell, David B. Papworth, and Paul K. Rodman. A VLIW architecture for a trace scheduling compiler. Operating Systems Review, 21(4): 180–192, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chuang:2006:UPB

[CNV^{+}06] Weihaw Chuang, Satish Narayanasamy, Ganesh Venkatesh, Jack Sampson, Michael Van Biesbrouck, Gilles Pokam, Brad Calder, and Osvaldo Colavin. Unbounded page-based transactional memory. Operating Systems Review, 40(5):347–358, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Collier:1973:AIS

[Col73] William W. Collier. Asynchronous interactions on shared data. Operating Systems Review, 7(2):6–15, April 1973. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cook:1978:CUC

[Coo78] Douglas Cook. The cost of using the CAP computer’s protection facilities. Operating Systems Review, 12(2):26–30, April 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cooper:1985:RDP

[Coo85] Eric C. Cooper. Replicated distributed programs. Operating Systems Review, 19(5):63–78, December 1985. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cooper:1986:RPC

[Coo86] Eric C. Cooper. Replicated procedure call. Operating Systems Review, 20(1):44–56, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cooper:1994:ECT

[Coo94] Robert Cooper. Experience with causally and totally ordered communication support: a cautionary tale. Operating Systems Review, 28(1):28–31, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Colaco:2008:UFW

[COS^{+}08] Eduardo M. Colaço, Marcelo Iury S. Oliveira, Alexandro S. Soares, Francisco Brasileiro, and Dalton S. Guerrero. Using a file working set model to speed up the recovery of Peer-to-Peer backup systems. Operating Systems Review, 42(6):64–70, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Costa:2013:BGB

[Cos13] Paolo Costa. Bridging the gap between applications and networks in data centers. Operating Systems Review, 47(1):3–8, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Carretero:1996:PPF

[CPdM^{+}96] J. Carretero, F. Pérez, P. de Miguel, F. García, and L. Alonso. ParFiSys: a parallel file system for MPP. Operating Systems Review, 30(2):74–80, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cansado:2010:CFD

[CPM10] Jacinto C. A. Cansado, João H. S. Pereira, and Edson T. Midorikawa. Considering the frequency dimension into on demand adaptive algorithms. Operating Systems Review, 44(1):110–115, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Choi:2008:ABP

[CPT08] Bumyong Choi, Leo Porter, and Dean M. Tullsen. Accurate branch prediction for short threads. Operating Systems Review, 42(2):125–134, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Camargos:2007:SMH

[CPW07] Lásaro Camargos, Fernando Pedone, and Marcin Wieloch. Sprint: a middleware for high-performance transaction processing. Operating Systems Review, 41(3):385–398, June 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Coffman:1972:SSP

[CR72] E. G. Coffman and Thomas A. Ryan. A study of storage partitioning using a mathematical model of locality. Operating Systems Review, 6(1/2):122–129, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chandy:1975:SPO

[CR75] K. M. Chandy and P. F. Reynolds. Scheduling partially ordered tasks with probabilistic execution times. Operating Systems Review, 9(5):169–177, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Couceiro:2012:WDT

[CR12] Maria Couceiro and Paolo Romano. Where does transactional memory research stand and what challenges lie ahead?: WTM 2012, EuroTM Workshop on Transactional Memory. Operating Systems Review, 46(2):87–92, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Craft:1983:RMD

[Cra83] Daniel H. Craft. Resource management in a decentralized system. Operating Systems Review, 17(5):11–19, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chapin:1995:HFC

[CRD^{+}95] J. Chapin, M. Rosenblum, S. Devine, T. Lahiri, D. Teodosiu, and A. Gupta. Hive: fault containment for shared-memory multiprocessors. Operating Systems Review, 29(5):12–25, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cristian:1991:FTA

[Cri91] Flaviu Cristian. Fault-tolerance in the Advanced Automation System. Operating Systems Review, 25(2):117–121, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Crichlow:1994:COP

[Cri94] Joel M. Crichlow. Combining optimism and pessimism to produce high availability in distributed transaction processing. Operating Systems Review, 28(3):43–64, July 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Casey:1977:DSD

[CS77] L. Casey and N. Shelness. A domain structure for distributed computer systems. Operating Systems Review, 11(5):101–108, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1993:ULC

[CS93] David R. Cheriton and Dale Skeen. Understanding the limitations of causally and totally ordered communication. Operating Systems Review, 27(5):44–57, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Creak:2000:TOS

[CS00] G. Alan Creak and Robert Sheehan. A top-down operating systems course. Operating Systems Review, 34(3):69–80, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Casey:2008:IFD

[CS08] Eoghan Casey and Gerasimos J. Stellatos. The impact of full disk encryption on digital forensics. Operating Systems Review, 42(3):93–98, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calciu:2017:BBC

[CSBA17a] Irina Calciu, Siddhartha Sen, Mahesh Balakrishnan, and Marcos K. Aguilera. Black-box concurrent data structures for NUMA architectures. Operating Systems Review, 51(2):207– 221, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calciu:2017:HIC

[CSBA17b] Irina Calciu, Siddhartha Sen, Mahesh Balakrishnan, and Marcos K. Aguilera. How to implement any concurrent data structure for modern servers. Operating Systems Review, 51(1):24– 32, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Churchill:2017:SLS

[CSBA17c] Berkeley Churchill, Rahul Sharma, JF Bastien, and Alex Aiken. Sound loop superoptimization for Google Native Client. Operating Systems Review, 51(2):313–326, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheng:2008:TCI

[CSJZ08] Bin Cheng, Lex Stein, Hai Jin, and Zheng Zhang. Towards cinematic Internet video-on-demand. Operating Systems Review, 42(4):109–122, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Culler:1991:FGP

[CSS^{+}91] David E. Culler, Anurag Sah, Klaus E. Schauser, Thorsten von Eicken, and John Wawrzynek. Fine-grain parallelism with minimal hardware support: a compiler-controlled threaded abstract machine. Operating Systems Review, 25(3S):164–175, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Calvelli:1993:ARS

[CV93] Claudio Calvelli and Vijay Varadharajan. Authentication and revocation in SPM extended abstract. Operating Systems Review, 27(4):42–57, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chiueh:2000:ISP

[CVP00] Tzi-Cker Chiueh, Ganesh Venkitachalam, and Prashant Pradhan. Integrating segmentation and paging protection for safe, efficient and transparent software extensions. Operating Systems Review, 34(2):20, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cachin:2014:WRH

[CvR14] Christian Cachin and Robbert van Renesse. Workshop report: HotDep 2013 — The 9th Workshop on Hot Topics in Dependable Systems. Operating Systems Review, 48(1):19–20, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cao:1992:AMR

[CW92] Jiannong Cao and K. C. Wang. An abstract model of rollback recovery control in distributed systems. Operating Systems Review, 26(4):62–76, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Crandall:2006:TSD

[CWdO^{+}06] Jedidiah R. Crandall, Gary Wassermann, Daniela A. S. de Oliveira, Zhendong Su, S. Felix Wu, and Frederic T. Chong. Temporal search: detecting hidden malware timebombs with virtual machines. Operating Systems Review, 40(5):25–36, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheng:2005:MAS

[CWL05] Chien-Fu Cheng, Shu-Ching Wang, and Tyne Liang. Multiagent schema of Mobile IP protocol for mobile environment. Operating Systems Review, 39(4):46–65, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chakraborty:2006:CSE

[CWS06] Koushik Chakraborty, Philip M. Wells, and Gurindar S. Sohi. Computation spreading: employing hardware migration to specialize CMP cores on-the-fly. Operating Systems Review, 40(5): 283–292, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chuan:2005:LBN

[CXMX05] Zhan Chuan, Lu Xianliang, Hou Mengshu, and Zhou Xu. A LVQ-based neural network anti-spam email approach. Operating Systems Review, 39(1):34–39, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chou:2001:ESO

[CYC^{+}01] Andy Chou, Junfeng Yang, Benjamin Chelf, Seth Hallem, and Dawson Engler. An empirical study of operating systems errors. Operating Systems Review, 35(5):73–88, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2017:PPQ

[CYG^{+}17] Quan Chen, Hailong Yang, Minyi Guo, Ram Srivatsa Kannan, Jason Mars, and Lingjia Tang. Prophet: Precise QoS prediction on non-preemptive accelerators to improve utilization in warehouse-scale computers. Operating Systems Review, 51(2): 17–32, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2016:BQA

[CYMT16] Quan Chen, Hailong Yang, Jason Mars, and Lingjia Tang. Baymax: QoS awareness and increased utilization for nonpreemptive accelerators in warehouse scale computers. Operating Systems Review, 50(2):681–696, June 2016. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cheriton:1983:DVK

[CZ83] David R. Cheriton and Willy Zwaenepoel. The distributed V kernel and its performance for diskless workstations. Operating Systems Review, 17(5):129–140, October 1983. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ciortea:2009:CST

[CZB^{+}09] Liviu Ciortea, Cristian Zamfir, Stefan Bucur, Vitaly Chipounov, and George Candea. Cloud9: a software testing service. Operating Systems Review, 43(4):5–10, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cohen:2005:CIC

[CZG^{+}05] Ira Cohen, Steve Zhang, Moises Goldszmidt, Julie Symons, Terence Kelly, and Armando Fox. Capturing, indexing, clustering, and retrieving system history. Operating Systems Review, 39(5): 105–118, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dalal:1975:MSS

[Dal75] Yogen K. Dalal. More on selecting sequence numbers. Operating Systems Review, 9(3):25–36, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dasser:1992:TTO

[Das92] Mahmoud Dasser. TOMP a total ordering multicast protocol. Operating Systems Review, 26(1):32–40, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Danthine:1975:CPN

[DB75] André A. S. Danthine and Joseph Bremer. Communication protocols in a network context. Operating Systems Review, 9(3): 87–92, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Davcev:1985:CRC

[DB85] Dančo Davčev and Walter A. Burkhard. Consistency and recovery control for replicated files. Operating Systems Review, 19 (5):87–96, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Druschel:1996:LRP

[DB96] Peter Druschel and Gaurav Banga. Lazy receiver processing (LRP): a network subsystem architecture for server systems. Operating Systems Review, 30(SI):261–275, October 1996. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dwyer:1997:MAF

[DB97] Dane Dwyer and Vaduvur Bharghavan. A mobility-aware file system for partially connected operation. Operating Systems Review, 31(1):24–30, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Douceur:1999:PBR

[DB99] John R. Douceur and William J. Bolosky. Progress-based regulation of low-importance processes. Operating Systems Review, 33(5):247–260, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Douceur:2000:PBR

[DB00a] John R. Douceur and William J. Bolosky. Process-based regulation of low-importance processes. Operating Systems Review, 34(2):26–27, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duesterwald:2000:SPH

[DB00b] Evelyn Duesterwald and Vasanth Bala. Software profiling for hot path prediction: less is more. Operating Systems Review, 34(5):202–211, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dutta:2011:WPA

[DB11] Prabal Dutta and Ricardo Bianchini. Workshop on power aware computing and systems (HotPower’11). Operating Systems Review, 45(3):47, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deBruijn:2008:PFL

[dBB08] Willem de Bruijn and Herbert Bos. PipesFS: fast Linux I/O in the Unix tradition. Operating Systems Review, 42(5):55–63, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Douglis:2006:MSC

[DBH^{+}06] Fred Douglis, Michael Branson, Kirsten Hildrum, Bin Rong, and Fan Ye. Multi-site cooperative data stream analysis. Operating Systems Review, 40(3):31–37, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Devietti:2008:HAS

[DBMZ08] Joe Devietti, Colin Blundell, Milo M. K. Martin, and Steve Zdancewic. Hardbound: architectural support for spatial safety of the C programming language. Operating Systems Review, 42 (2):103–114, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denehy:2004:DSA

[DBP^{+}04] Timothy E. Denehy, John Bent, Florentina I. Popovici, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Deconstructing storage arrays. Operating Systems Review, 38(5): 59–71, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dagand:2009:FFP

[DBR09] Pierre-Evariste Dagand, Andrew Baumann, and Timothy Roscoe. Filet-o-fish: practical and dependable domain-specific languages for OS development. Operating Systems Review, 43 (4):35–39, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Draves:1991:UCI

[DBRD91] Richard P. Draves, Brian N. Bershad, Richard F. Rashid, and Randall W. Dean. Using continuations to implement thread management and communication in operating systems. Operating Systems Review, 25(5):122–136, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duda:1999:BVT

[DC99] Kenneth J. Duda and David R. Cheriton. Borrowed-virtual-time (BVT) scheduling: supporting latency-sensitive threads in a general-purpose scheduler. Operating Systems Review, 33(5): 261–276, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duda:2000:BVT

[DC00] Kenneth J. Duda and David R. Cheriton. Borrowed-virtual-time (BVT) scheduling: supporting latency-sensitive threads in a general-purpose scheduler. Operating Systems Review, 34(2): 27–28, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dwarkadas:1996:ICT

[DCZ96] Sandhya Dwarkadas, Alan L. Cox, and Willy Zwaenepoel. An integrated compile-time/run-time software distributed shared memory system. Operating Systems Review, 30(5):186–197, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1980:MCS

[DD80] Peter J. Denning and T. Don Dennis. On minimizing contention at semaphores. Operating Systems Review, 14(2):9–16, April 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Desnoyers:2012:LMC

[DD12] Mathieu Desnoyers and Michel R. Dagenais. Lockless multicore high-throughput buffering scheme for kernel tracing. Operating Systems Review, 46(3):65–81, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Didona:2016:PAM

[DDK^{+}16] Diego Didona, Nuno Diegues, Anne-Marie Kermarrec, Rachid Guerraoui, Ricardo Neves, and Paolo Romano. ProteusTM: Abstraction meets performance in transactional memory. Operating Systems Review, 50(2):757–771, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Drumond:2018:AAC

[DDM^{+}18] Mario Drumond, Alexandros Daglis, Nooshin Mirzadeh, Dmitrii Ustiugov, Javier Picorel, Babak Falsafi, Boris Grot, and Dionisios Pnevmatikatos. Algorithm/architecture co-design for nearmemory processing. Operating Systems Review, 52(1):109–122, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Delabrida:2016:BWG

[DDOL16] Saul Delabrida, Thiago D’Angelo, Ricardo A. R. Oliveira, and Antonio A. F. Loureiro. Building wearables for geology: an operating system approach. Operating Systems Review, 50(1):31– 45, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dan:1999:QAM

[DDYM99] Pei Dan, Wang Dongsheng, Zhang Youhui, and Shen Meiming. Quasi-asynchronous migration: a novel migration protocol for PVM tasks. Operating Systems Review, 33(2):5–14, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dellar:1980:RBS

[Del80] Carl Dellar. Removing backing store administration from the CAP operating system. Operating Systems Review, 14(4):41– 49, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1974:ITD

[Den74a] Peter J. Denning. Is it not time to define “structured programming”? Operating Systems Review, 8(1):6–7, January 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1974:SPL

[Den74b] Peter J. Denning. Is “structured programming” any longer the right term? Operating Systems Review, 8(4):4–6, October 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1978:NYS

[Den78] Dorothy E. Denning. A note from your secretary-treasurer. Operating Systems Review, 12(3):2–4, July 1978. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1979:NYS

[Den79] Dorothy E. Denning. A note from your secretary-treasurer. Operating Systems Review, 13(3):1–2, July 1979. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1980:NYS

[Den80] Dorothy E. Denning. A note from your secretary-treasurer. Operating Systems Review, 14(3):2–4, July 1980. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Deng:2007:AR

[Den07] Yuhui Deng. Author response. Operating Systems Review, 41 (4):78, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Desnoyers:2010:EEN

[Des10] Peter Desnoyers. Empirical evaluation of NAND flash memory performance. Operating Systems Review, 44(1):50–54, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Damron:2006:HTM

[DFL06] Peter Damron, Alexandra Fedorova, and Yossi Lev. Hybrid transactional memory. Operating Systems Review, 40(5):336– 346, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duchien:2000:POR

[DFS00] Laurence Duchien, Gérard Florin, and Lionel Seinturier. Partial order relations in distributed object environments. Operating Systems Review, 34(4):56–75, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deGois:2010:OPD

[dGdB10] Lourival A. de Gois and Walter C. da Borelli. Optimization of procedures for discovery and information of idle resources in distributed systems. Operating Systems Review, 44(1):103– 109, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Demers:1988:EAR

[DGH^{+}88] Alan Demers, Dan Greene, Carl Houser, Wes Irish, John Larson, Scott Shenker, Howard Sturgis, Dan Swinehart, and Doug Terry. Epidemic algorithms for replicated database maintenance. Operating Systems Review, 22(1):8–32, January 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Day:1973:AOS

[DH73] Paul Day and John Hines. ARGOS: An operating system for a computer utility supporting interactive instrument control. Operating Systems Review, 7(4):28–37, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ding:1995:ULP

[DH95] Yun Ding and Patrick Horster. Undetectable on-line password guessing attacks. Operating Systems Review, 29(4):77–86, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ding:1996:WKA

[DH96] Yun Ding and Patrick Horster. Why the Kuperee authentication system fails. Operating Systems Review, 30(2):42–51, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DePauw:2010:VAT

[DH10] Wim De Pauw and Stephen Heisig. Visual and algorithmic tooling for system trace analysis: a case study. Operating Systems Review, 44(1):97–102, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DeCandia:2007:DAH

[DHJ^{+}07] Giuseppe DeCandia, Deniz Hastorun, Madan Jampani, Gunavardhan Kakulapati, Avinash Lakshman, Alex Pilchin, Swaminathan Sivasubramanian, Peter Vosshall, and Werner Vogels. Dynamo: Amazon’s highly available key-value store. Operating Systems Review, 41(6):205–220, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Debnath:2015:RHT

[DHK^{+}15] Biplob Debnath, Alireza Haghdoost, Asim Kadav, Mohammed G. Khatib, and Cristian Ungureanu. Revisiting hash table design for phase change memory. Operating Systems Review, 49(2):18–26, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Daniels:1991:SLS

[DHRS91] Dean Daniels, Roger Haskin, Jon Reinke, and Wayne Sawdon. Shared logging services for fault-tolerant distributed computing. Operating Systems Review, 25(1):65–68, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dijkstra:2005:MRO

[Dij05] Edsger W. Dijkstra. My recollections of operating system design. Operating Systems Review, 39(2):4–40, April 2005. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dimitoglou:1998:DMT

[Dim98] George Dimitoglou. Deadlocks and methods for their detection, prevention and recovery in modern operating systems. Operating Systems Review, 32(3):51–54, July 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dodge:2005:SIL

[DIN05] Catherine Dodge, Cynthia Irvine, and Thuy Nguyen. A study of initialization in Linux and OpenBSD. Operating Systems Review, 39(2):79–93, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dion:1980:CFS

[Dio80] Jeremy Dion. The Cambridge File Server. Operating Systems Review, 14(4):26–35, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Du:2019:WPC

[DIS19] Yifan Du, Valérie Issarny, and Françoise Sailhan. When the power of the crowd meets the intelligence of the middleware: The mobile phone sensing case. Operating Systems Review, 53(1):85–90, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deJonge:1993:LDN

[dJKH93] Wiebren de Jonge, M. Frans Kaashoek, and Wilson C. Hsieh. The Logical Disk: a new approach to improving file systems. Operating Systems Review, 27(5):15–28, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). 14th ACM Symposium on Operating Systems Principles, Ashville, NC, USA.

Dhawan:2017:CCA

[DJS^{+}17] Medhavi Dhawan, Gurprit Johal, Jim Stabile, Vjekoslav Brajkovic, James Chang, Kapil Goyal, Kevin James, Zeeshan Lokhandwala, Anny Martinez Manzanilla, Roger Michoud, Maithem Munshed, Srinivas Neginhal, Konstantin Spirov, Michael Wei, Scott Fritchie, Chris Rossbach, Ittai Abraham, and Dahlia Malkhi. Consistent clustered applications with Corfu. Operating Systems Review, 51(1):78–82, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1975:SPL

[DK75] Peter J. Denning and Kevin C. Kahn. A study of program locality and lifetime functions. Operating Systems Review, 9(5): 207–216, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Desnoyers:2015:ITW

[DK15] Peter Desnoyers and Gokul Kandiraju. INFLOW 2015: The Third Workshop on Interactions of NVM/FLash with Operating systems and Workload: INFLOW’15 message from the Chairs. Operating Systems Review, 49(2):17, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Delimitrou:2016:HRE

[DK16] Christina Delimitrou and Christos Kozyrakis. HCloud: Resource-efficient provisioning in shared cloud systems. Operating Systems Review, 50(2):473–488, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Delimitrou:2017:BKW

[DK17] Christina Delimitrou and Christos Kozyrakis. Bolt: I know what you did last summer … in the cloud. Operating Systems Review, 51(2):599–613, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dunlap:2002:REI

[DKC^{+}02] George W. Dunlap, Samuel T. King, Sukru Cinar, Murtaza A. Basrai, and Peter M. Chen. ReVirt: enabling intrusion analysis through virtual-machine logging and replay. Operating Systems Review, 36(5S):211–224, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dabek:2001:WAC

[DKK^{+}01] Frank Dabek, M. Frans Kaashoek, David Karger, Robert Morris, and Ion Stoica. Wide-area cooperative storage with CFS. Operating Systems Review, 35(5):202–215, December 2001. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dalton:2010:TP

[DKK10] Michael Dalton, Hari Kannan, and Christos Kozyrakis. Tainting is not pointless. Operating Systems Review, 44(2):88–92, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DaSilva:2006:KIO

[DKW^{+}06] Dilma Da Silva, Orran Krieger, Robert W. Wisniewski, Amos Waterland, David Tam, and Andrew Baumann. K42: an infrastructure for operating system research. Operating Systems Review, 40(2):34–42, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dayal:2009:MOB

[DKW^{+}09] Umeshwar Dayal, Harumi Kuno, Janet L. Wiener, Kevin Wilkinson, Archana Ganapathi, and Stefan Krompass. Managing operational business intelligence workloads. Operating Systems Review, 43(1):92–98, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dietrich:2015:DVS

[DL15] Christian Dietrich and Daniel Lohmann. The dataref versuchung: Saving time through better internal repeatability. Operating Systems Review, 49(1):51–60, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dall:2018:AVP

[DLLN18] Christoffer Dall, Shih-Wei Li, Jin Tack Lim, and Jason Nieh. ARM virtualization: Performance and architectural implications. Operating Systems Review, 52(1):45–56, July 2018. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deLara:2000:PCBa

[dLWZ00a] Eyal de Lara, Dan Wallach, and Willy Zwaenepoel. Puppeteer: component-based adaptation for mobile computing. Operating Systems Review, 34(2):33, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deLara:2000:PCBb

[dLWZ00b] Eyal de Lara, Dan Wallach, and Willy Zwaenepoel. Puppeteer: component-based adaptation for mobile computing (poster session). Operating Systems Review, 34(2):40, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Danzig:1990:HRT

[DM90] Peter B. Danzig and Stephen Melvin. High resolution timing with low resolution clocks and microsecond resolution timer for Sun workstations. Operating Systems Review, 24(1):23–26, January 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ditzel:1987:DTS

[DMB87] David R. Ditzel, Hubert R. McLellan, and Alan D. Berenbaum. Design tradeoffs to support the C programming language in the CRISP microprocessor. Operating Systems Review, 21(4):158– 163, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Desnoyers:2013:MCS

[DMD13] Mathieu Desnoyers, Paul E. McKenney, and Michel R. Dagenais. Multi-core systems modeling for formal verification of parallel algorithms. Operating Systems Review, 47(2):51–65, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dumitras:2010:RSA

[DNT10] Tudor Dumitraş, Iulian Neamtiu, and Eli Tilevich. Report on the Second ACM Workshop on Hot Topics in Software Upgrades (HotSWUp’09). Operating Systems Review, 44(4):146–152, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL http://www.hotswup.org/2009/.

Dabbous:2009:ORW

[DO09] Walid Dabbous and Maximilian Ott. Overview of the ROADS’09 workshop. Operating Systems Review, 43(4):50–53, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oliveira:2012:SMC

[dOL12] Frederico Alvares de Oliveira, Jr. and Thomas Ledoux. Selfmanagement of cloud applications and infrastructure for energy optimization. Operating Systems Review, 46(2):10–18, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deOliveira:2012:MAM

[dORF12] Augusto Born de Oliveira, Ahmad Saif Ur Rehman, and Sebastian Fischmeister. mTags: augmenting microkernel messages with lightweight metadata. Operating Systems Review, 46(2): 67–79, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DosReis:1988:NBA

[Dos88] Anthony J. Dos Reis. A note on Ben-Ari’s concurrent programming system. Operating Systems Review, 22(3):41–42, July 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deOliveira:2008:BWO

[dOS08] Rômulo Silva de Oliveira and Alexandre Sztajnberg. Brazilian workshop on operating systems. Operating Systems Review, 42 (6):50–51, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Douglis:1993:RCD

[Dou93] Fred Douglis. On the role of compression in distributed systems. Operating Systems Review, 27(2):88–93, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Douceur:2009:PRV

[Dou09] John R. Douceur. Paper rating vs. paper ranking. Operating Systems Review, 43(2):117–121, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Druschel:1993:FHB

[DP93] Peter Druschel and Larry L. Peterson. Fbufs: a high-bandwidth cross-domain transfer facility. Operating Systems Review, 27(5): 189–202, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dalton:2009:TVP

[DPW^{+}09] Chris I. Dalton, David Plaquin, Wolfgang Weidner, Dirk Kuhlmann, Boris Balacheff, and Richard Brown. Trusted virtual platforms: a key enabler for converged client devices. Operating Systems Review, 43(1):36–43, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Damm:1989:RTO

[DRSK89] A. Damm, J. Reisinger, W. Schwabl, and H. Kopetz. The realtime operating system of MARS. Operating Systems Review, 23(3):141–157, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Doekemeijer:2024:ZRA

[DRTT24] Krijn Doekemeijer, Zebin Ren, Nick Tehrany, and Animesh Trivedi. ZWAL: Rethinking write-ahead logs for ZNS SSDs with zone appends. Operating Systems Review, 58(1):53–60, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3689051.3689060.

Druschel:1992:MPO

[Dru92] Peter Druschel. Modularity and protection are orthogonal, or “Why μ-kernel Architectures are Flawed”. Operating Systems Review, 26(2):22, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1972:PWS

[DS72] Peter J. Denning and Stuart C. Schwartz. Properties of the working-set model. Operating Systems Review, 6(1/2):130–140, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1973:DSP

[DS73] Peter J. Denning and Jeffrey R. Spirn. Dynamic storage partitioning. Operating Systems Review, 7(4):73–79, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Denning:1980:EVO

[DS80] Peter J. Denning and Harold S. Stone. An exchange of views on operating systems courses. Operating Systems Review, 14 (4):71–82, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Davis:1990:NSP

[DS90] Don Davis and Ralph Swick. Network security via private-key certificates. Operating Systems Review, 24(4):64–67, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Davari:1992:SUP

[DS92] Sadegh Davari and Lui Sha. Sources of unbounded priority inversions in real-time systems and a comparative study of possible solutions. Operating Systems Review, 26(2):110–120, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DaSilva:2006:PPA

[DS06] Jeff Da Silva and J. Gregory Steffan. A probabilistic pointer analysis for speculative optimizations. Operating Systems Review, 40(5):416–425, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dowty:2009:GVV

[DS09] Micah Dowty and Jeremy Sugerman. GPU virtualization on VMware’s hosted I/O architecture. Operating Systems Review, 43(3):73–82, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Desikan:2004:SSR

[DSBK04] Rajagopalan Desikan, Simha Sethumadhavan, Doug Burger, and Stephen W. Keckler. Scalable selective re-execution for EDGE architectures. Operating Systems Review, 38(5):120–132, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

daSilva:2011:VBW

[dSBP11] Dilma da Silva, Luciano Barreto, and Paulo César A. Pereira. VII Brazilian Workshop on Operating Systems. Operating Systems Review, 45(1):155, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deSa:2013:ARB

[dSFdAM13] Alírio Santos de Sá, Allan Edgard Silva Freitas, and Raimundo José de Araújo Macêdo. Adaptive request batching for Byzantine replication. Operating Systems Review, 47(1):35–42, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Das:2005:HKM

[DSGP05] Manik Lal Das, Ashutosh Saxena, Ved P. Gulati, and Deepak B. Phatak. Hierarchical key management scheme using polynomial interpolation. Operating Systems Review, 39(1):40–47, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

dosSantos:2016:EAF

[dSM16] Caio Augusto R. dos Santos and Rivalino Matias, Jr. Exploratory analysis on failure causes in a mass-market operating system. Operating Systems Review, 50(1):18–30, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dovrolis:2001:HHI

[DTR01] Constantinos Dovrolis, Brad Thayer, and Parameswaran Ramanathan. HIP: hybrid interrupt-polling for the network interface. Operating Systems Review, 35(4):50–60, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dube:2000:SHC

[Dub00] Rohit Dube. Scalable hierarchical coarse-grained timers. Operating Systems Review, 34(1):11–20, January 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duchamp:1989:ATM

[Duc89] D. Duchamp. Analysis of transaction management performance. Operating Systems Review, 23(5):177–190, December 1989. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Duchamp:1992:SSW

[Duc92] Daniel Duchamp. Systems Software for Wireless Mobile Computing. Operating Systems Review, 26(2):10, April 1992. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dunstan:1991:SFS

[Dun91] Neil Dunstan. Semaphores for fair scheduling monitor conditions. Operating Systems Review, 25(3):27–31, July 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Davidson:1987:EIS

[DV87] Jack W. Davidson and Richard A. Vaughan. The effect of instruction set complexity on program size and memory performance. Operating Systems Review, 21(4):60–64, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deVivo:1996:BRD

[dV96] Marco de Vivo. Book review: Distributed Operating Systems by Andrew S. Tanenbaum. Operating Systems Review, 30(1): 3, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

deVivo:1998:ISA

[dVdVI98] Marco de Vivo, Gabriela O. de Vivo, and Germinal Isern. Internet security attacks at the basic levels. Operating Systems Review, 32(2):4–15, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dawkins:2012:SRI

[DVS12] Scott Dawkins, Kaladhar Voruganti, and John D. Strunk. Systems research and innovation in data ONTAP. Operating Systems Review, 46(3):1–3, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Deng:2007:HSG

[DW07a] Yuhui Deng and Frank Wang. A heterogeneous storage grid enabled by grid service. Operating Systems Review, 41(1):7– 13, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Deng:2007:OCS

[DW07b] Yuhui Deng and Frank Wang. Opportunities and challenges of storage Grid enabled by Grid service. Operating Systems Review, 41(4):79–82, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

DaSilva:2008:I

[DW08] Dilma Da Silva and Robert W. Wisniewski. Introduction. Operating Systems Review, 42(1):1, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dawu:2001:TES

[DY01] Gu Dawu and Wang Yi. On the techniques of enhancing the security of block ciphers. Operating Systems Review, 35(4):94– 96, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dolev:2010:STR

[DY10] Shlomi Dolev and Reuven Yagel. Stabilizing trust and reputation for self-stabilizing efficient hosts in spite of Byzantine guests. Operating Systems Review, 44(3):65–74, July 2010. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dean:1995:MDS

[DZ95] Dawson Dean and Richard Zippel. Matching data storage to application needs. Operating Systems Review, 29(1):68–73, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dong:2011:RNF

[DZP^{+}11] Yuan Dong, Haiyang Zhu, Jinzhan Peng, Fang Wang, Michael P. Mesnier, Dawei Wang, and Sun C. Chan. RFS: a network file system for mobile devices and the cloud. Operating Systems Review, 45(1):101–111, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Easton:1972:PSL

[Eas72] William B. Easton. Process synchronization without long-term interlock. Operating Systems Review, 6(1/2):95–100, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Emmerich:2007:IRM

[EAS07] Wolfgang Emmerich, Mikio Aoyama, and Joe Sventek. The impact of research on middleware technology. Operating Systems Review, 41(1):89–112, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elyasi:2017:EIR

[EAS^{+}17] Nima Elyasi, Mohammad Arjomand, Anand Sivasubramaniam, Mahmut T. Kandemir, Chita R. Das, and Myoungsoo Jung. Exploiting intra-request slack to improve SSD performance. Operating Systems Review, 51(2):375–388, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ekanadham:1978:SNT

[EB78] K. Ekanadham and A. J. Bernstein. Some new transitions in hierarchical level structures. Operating Systems Review, 12(4):34– 38, October 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

ElSayed:2016:UFS

[EBP16] Salem El Sayed, Matthias Bolten, and Dirk Pleiter. Using file system counters in modelling parallel I/O architectures. Operating Systems Review, 50(3):37–46, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eisenhauer:2001:MTC

[EBS01] Greg Eisenhauer, Fabián E. Bustamante, and Karsten Schwan. A middleware toolkit for client-initiated service specialization. Operating Systems Review, 35(2):7–20, April 2001. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Engler:2001:BDB

[ECH^{+}01] Dawson Engler, David Yu Chen, Seth Hallem, Andy Chou, and Benjamin Chelf. Bugs as deviant behavior: a general approach to inferring errors in systems code. Operating Systems Review, 35(5):57–72, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eruno:1973:SIT

[ECS73] J. Eruno, E. G. Coffman, Jr., and R. Sethi. Scheduling independent tasks to reduce mean finishing-time (extended abstract). Operating Systems Review, 7(4):102–103, October 1973. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Editors:2013:EES

[Edi13] Editors. An energy-efficient self-provisioning approach for cloud resources management. Operating Systems Review, 47(3):2–9, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elnikety:2006:TUD

[EDP06] Sameh Elnikety, Steven Dropsho, and Fernando Pedone. Tashkent: uniting durability with transaction ordering for highperformance scalable database replication. Operating Systems Review, 40(4):117–130, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elnikety:2007:TMA

[EDZ07] Sameh Elnikety, Steven Dropsho, and Willy Zwaenepoel. Tashkent+: memory-aware load balancing and update filtering in replicated databases. Operating Systems Review, 41(3):399– 412, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eyerman:2006:PCA

[EEKS06] Stijn Eyerman, Lieven Eeckhout, Tejas Karkhanis, and James E. Smith. A performance counter architecture for computing accurate CPI components. Operating Systems Review, 40(5): 175–184, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eichert:2002:CVA

[EENV02] Stuart Eichert, Osman N. Ertugay, Dan Nessett, and Suresh Vobbilisetty. Commercially viable active networking. Operating Systems Review, 36(1):8–22, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ellard:2012:GCV

[EER12] Daniel Ellard, Craig Everhart, and Theresa Raj. Glitz: crossvendor federated file systems. Operating Systems Review, 46 (3):4–10, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eugster:2007:APG

[EFL07] Patrick Eugster, Pascal Felber, and Fabrice Le Fessant. The ‘art’ of programming gossip-based systems. Operating Systems Review, 41(5):37–42, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elson:2002:FGN

[EGE02] Jeremy Elson, Lewis Girod, and Deborah Estrin. Fine-grained network time synchronization using reference broadcasts. Operating Systems Review, 36(5S):147–163, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elson:2007:MIW

[EHD07] Jeremy Elson, Jon Howell, and John R. Douceur. MapCruncher: integrating the world’s geographic information. Operating Systems Review, 41(2):50–59, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eide:2015:FSI

[Eid15] Eric Eide. Foreword: Special issue on repeatability and sharing of experimental artifacts. Operating Systems Review, 49(1):1– 2, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ezzati-Jivan:2013:FCS

[EJD13] Naser Ezzati-Jivan and Michel R. Dagenais. A framework to compute statistics of system parameters from very large trace files. Operating Systems Review, 47(1):43–54, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Efstathopoulos:2008:MFG

[EK08] Petros Efstathopoulos and Eddie Kohler. Manageable finegrained information flow. Operating Systems Review, 42(4):301– 313, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Endo:2014:CCA

[EKF^{+}14] Hiroshi Endo, Hiroyoshi Kodama, Hiroyuki Fukuda, Toshio Sugimoto, Takashi Horie, and Masao Kondo. Cooperative control architecture of fan-less servers and fresh-air cooling in container servers for low power operation. Operating Systems Review, 48 (1):34–38, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ekanayake:2004:ULP

[EKM04] Virantha Ekanayake, Clinton Kelly IV, and Rajit Manohar. An ultra low-power processor for sensor networks. Operating Systems Review, 38(5):27–36, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Engler:1995:EOS

[EKO95a] D. R. Engler, M. F. Kaashoek, and J. O’Toole, Jr. Exokernel: an operating system architecture for application-level resource management. Operating Systems Review, 29(5):251–266, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Engler:1995:OSK

[EKO95b] Dawson R. Engler, M. Frans Kaashoek, and James W. O’Toole, Jr. The operating system kernel as a secure programmable machine. Operating Systems Review, 29(1):78–82, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Efstathopoulos:2005:LEP

[EKV^{+}05] Petros Efstathopoulos, Maxwell Krohn, Steve VanDeBogart, Cliff Frey, David Ziegler, Eddie Kohler, David Mazières, Frans Kaashoek, and Robert Morris. Labels and event processes in the asbestos operating system. Operating Systems Review, 39 (5):17–30, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Esquivel:1995:QOC

[ELG95] S. Esquivel, G. Leguizamon, and R. Gallard. A quasi-optimal cluster allocation strategy for parallel program execution in distributed systems using genetic algorithms. Operating Systems Review, 29(2):82–96, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ellis:1973:PDC

[Ell73] Clarence A. Ellis. On the probability of deadlock in computer systems. Operating Systems Review, 7(4):88–95, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ellis:1977:CCD

[Ell77] Clarence A. Ellis. Consistency and correctness of duplicate database systems. Operating Systems Review, 11(5):67–84, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Edwards:2015:CRC

[ELR15] Sarah Edwards, Xuan Liu, and Niky Riga. Creating repeatable computer science and networking experiments on shared, public testbeds. Operating Systems Review, 49(1):90–99, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Edwards:1989:ERM

[EM89] D. Edwards and M. Mckendry. Exploiting read-mostly workloads in the FileNet file system. Operating Systems Review, 23 (5):58–70, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Erlingsson:2006:AHE

[EM06] Úlfar Erlingsson and John MacCormick. Ad hoc extensibility and access control. Operating Systems Review, 40(3):93–101, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eide:2009:PFW

[EMS09] Eric Eide, Gilles Muller, and Olaf Spinczyk. PLOS 2009: Fifth Workshop on Programming Languages and Operating Systems. Operating Systems Review, 43(4):31–34, December 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eide:2011:SPS

[EMSPS11] Eric Eide, Gilles Muller, Wolfgang Schröder-Preikschat, and Olaf Spinczyk. Summary of PLOS 2011: the Sixth Workshop on Programming Languages and Operating Systems. Operating Systems Review, 45(3):1–4, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

ElHajj:2016:SPM

[EMZ^{+}16] Izzat El Hajj, Alexander Merritt, Gerd Zellweger, Dejan Milojicic, Reto Achermann, Paolo Faraboschi, Wen mei Hwu, Timothy Roscoe, and Karsten Schwan. SpaceJMP: Programming with multiple virtual address spaces. Operating Systems Review, 50(2):353–368, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Erlichson:1996:SAP

[ENCH96] Andrew Erlichson, Neal Nuckolls, Greg Chesson, and John Hennessy. SoftFLASH: analyzing the performance of clustered distributed virtual shared memory. Operating Systems Review, 30 (5):210–220, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Enslow:1975:OAE

[Ens75] Philip H. Enslow. OSCL (1) activity in Europe. Operating Systems Review, 9(2):16–17, April 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Engler:1994:DER

[EP94] Dawson R. Engler and Todd A. Proebsting. DCG: an efficient, retargetable dynamic code generation system. Operating Systems Review, 28(5):263–272, December 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Erbsen:2020:SHL

[EPG^{+}20] Andres Erbsen, Jade Philipoom, Jason Gross, Robert Sloan, and Adam Chlipala. Simple high-level code for cryptographic arithmetic: With proofs, without compromises. Operating Systems Review, 54(1):23–30, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3421473.3421477.

Eriksen:2014:YSF

[Eri14] Marius Eriksen. Your server as a function. Operating Systems Review, 48(1):51–57, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ezhilchelvan:2010:LPR

[ES10] Paul Ezhilchelvan and Santosh Shrivastava. Learning from the past for resolving dilemmas of asynchrony. Operating Systems Review, 44(2):58–63, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Engelmann:2006:MAR

[ESB^{+}06] Christian Engelmann, Stephen L. Scott, David E. Bernholdt, Narasimha R. Gottumukkala, Chokchai Leangsuksun, Jyothish Varma, Chao Wang, Frank Mueller, Aniruddha G. Shet, and P. Sadayappan. MOLAR: adaptive runtime support for high-end computing operating and runtime systems. Operating Systems Review, 40(2):63–72, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eskicioglu:1996:CBD

[Esk96] M. Rasit Eskicioğlu. A comprehensive bibliography of distributed shared memory. Operating Systems Review, 30(1):71– 96, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Estrin:2002:KAS

[Est02] Deborah Estrin. Keynote address: Sensor network research: emerging challenges for architecture, systems, and languages. Operating Systems Review, 36(5):1–4, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Eggert:2005:ISP

[ET05] Lars Eggert and Joseph D. Touch. Idletime scheduling with preemption intervals. Operating Systems Review, 39(5):249–262, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Etsion:2007:FGK

[ETKF07] Yoav Etsion, Dan Tsafrir, Scott Kirkpatrick, and Dror G. Feitelson. Fine grained kernel logging with KLogger: experience and insights. Operating Systems Review, 41(3):259–272, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Emck:1989:EOS

[EVvdW89] J. H. Emck, J. H. Voskamp, and A. J. van der Wal. EPEP: an operating system designed for experiment-control. Operating Systems Review, 23(4):33–44, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Elshoff:1976:MOS

[EW76] James L. Elshoff and Mitchel R. Ward. The MCTS operating system. Operating Systems Review, 10(1):18–38, January 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Endo:1996:ULE

[EWCS96] Yasuhiro Endo, Zheng Wang, J. Bradley Chen, and Margo Seltzer. Using latency to evaluate interactive system performance. Operating Systems Review, 30(SI):185–199, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fabry:1973:CCB

[Fab73] R. S. Fabry. The case for capability based computers (extended abstract). Operating Systems Review, 7(4):120, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Faber:1998:OTW

[Fab98] Theodore Faber. Optimizing throughout in a workstation-based network file system over a high bandwidth local area network. Operating Systems Review, 32(1):29–40, January 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fahndrich:2006:LSF

[FAH^{+}06] Manuel Fähndrich, Mark Aiken, Chris Hawblitzel, Orion Hodson, Galen Hunt, James R. Larus, and Steven Levi. Language support for fast and reliable message-based communication in singularity OS. Operating Systems Review, 40(4):177–190, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ford:1997:FOS

[FBB^{+}97] Bryan Ford, Godmar Back, Greg Benson, Jay Lepreau, Albert Lin, and Olin Shivers. The Flux OSKit: a substrate for kernel and language research. Operating Systems Review, 31(5):38–51, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Frohlich:2012:BSC

[FBL^{+}12] Antonio Augusto Fröhlich, Leandro Buss Becker, George Lima, Stefan Petters, Dilma M. da Silva, and Edna N. Silva Barros. Brazilian Symposium on Computing System Engineering. Operating Systems Review, 46(1):52, January 2012. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Finlayson:1987:LFE

[FC87] R. Finlayson and D. Cheriton. Log files: an extended file service exploiting write-once storage. Operating Systems Review, 21(5): 139–148, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fournier:2010:ABD

[FD10] Pierre-Marc Fournier and Michel R. Dagenais. Analyzing blocking to debug performance problems on multi-core systems. Operating Systems Review, 44(2):77–87, April 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Freitas:2014:PET

[FdAM14] Allan Edgard Silva Freitas and Raimundo José de Araújo Macêdo. A performance evaluation tool for hybrid and dynamic distributed systems. Operating Systems Review, 48(1):11–18, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feak:1983:SIS

[Fêa83] Viiveke Fêak, editor. Security, IFIP/Sec’83: proceedings of the First Security Conference, Stockholm, Sweden, 16–19 May 1983. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86669-8 (Elsevier). LCCN QA76.9.A25 S4 1983.

Feitelson:2015:RRC

[Fei15] Dror G. Feitelson. From repeatability to reproducibility and corroboration. Operating Systems Review, 49(1):3–11, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feske:2007:CSC

[Fes07] Norman Feske. A case study on the cost and benefit of dynamic RPC marshalling for low-level system components. Operating Systems Review, 41(4):40–48, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Forman:2009:EDL

[FES09] George Forman, Kave Eshghi, and Jaap Suermondt. Efficient detection of large-scale redundancy in enterprise file systems. Operating Systems Review, 43(1):84–91, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Faibish:2008:SVU

[FFBG08] Sorin Faibish, Stephen Fridella, Peter Bixby, and Uday Gupta. Storage virtualization using a block-device file system. Operating Systems Review, 42(1):119–126, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fernandess:2007:GTF

[FFM07] Yaacov Fernandess, Antonio Fernández, and Maxime Monod. A generic theoretical framework for modeling gossip-based algorithms. Operating Systems Review, 41(5):19–27, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Freudenthal:1991:PCF

[FG91] Eric Freudenthal and Allan Gottlieb. Process coordination with fetch-and-increment. Operating Systems Review, 25(3S):260– 268, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fox:1996:ANC

[FGBA96] Armando Fox, Steven D. Gribble, Eric A. Brewer, and Elan Amir. Adapting to network and client variability via on-demand dynamic distillation. Operating Systems Review, 30(5):160–170, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fox:1997:CBS

[FGC^{+}97] Armando Fox, Steven D. Gribble, Yatin Chawathe, Eric A. Brewer, and Paul Gauthier. Cluster-based scalable network services. Operating Systems Review, 31(5):78–91, December 1997.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Friedman:2007:GMB

[FGR^{+}07] Roy Friedman, Daniela Gavidia, Luis Rodrigues, Aline Carneiro Viana, and Spyros Voulgaris. Gossiping on MANETs: the beauty and the beast. Operating Systems Review, 41(5):67–74, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Francez:1985:SCA

[FH85] Nissim Francez and Brent Hailpern. Script: a communication abstraction mechanism. Operating Systems Review, 19(2):53– 67, April 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ford:1995:UAI

[FHL95] Bryan Ford, Mike Hibler, and Jay Lepreau. Using annotated interface definitions to optimize RPC. Operating Systems Review, 29(5):232, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ford:1996:MMR

[FHL^{+}96] Bryan Ford, Mike Hibler, Jay Lepreau, Patrick Tullmann, Godmar Back, and Stephen Clawson. Microkernels meet recursive virtual machines. Operating Systems Review, 30(SI):137–151, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fall:2011:REG

[FIM^{+}11] Kevin Fall, Gianluca Iannaccone, Maziar Manesh, Sylvia Ratnasamy, Katerina Argyraki, Mihai Dobrescu, and Norbert Egi. RouteBricks: enabling general purpose network infrastructure. Operating Systems Review, 45(1):112–125, January 2011. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Finlayson:1992:SCV

[Fin92] Ross Finlayson. Structuring and Communication in the Vanguard OS Kernel. Operating Systems Review, 26(2):30, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fiuczynski:2006:POH

[Fiu06] Marc E. Fiuczynski. PlanetLab: overview, history, and future directions. Operating Systems Review, 40(1):6–10, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fu:1998:VSS

[FJLC98] Chao-Ying Fu, Matthew D. Jennings, Sergei Y. Larin, and Thomas M. Conte. Value speculation scheduling for high performance processors. Operating Systems Review, 32(5):262–271, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fink:2017:VMD

[FKZ17] Bryan Fink, Eric Knauft, and Gene Zhang. vSAN: Modern distributed storage. Operating Systems Review, 51(1):33–37, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fischer:1977:EIO

[FL77] Charles N. Fischer and Richard J. LeBlanc. Efficient implementation and optimization of run-time checking in PASCAL. Operating Systems Review, 11(2):19–24, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fleisch:1981:APS

[Fle81] Brett D. Fleisch. An architecture for pup services on a distributed operating system. Operating Systems Review, 15(1):26– 44, January 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fleisch:1983:OSP

[Fle83] Brett D. Fleisch. Operating systems: a perspective on future trends. Operating Systems Review, 17(2):14–17, April 1983. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fleisch:2007:PDC

[Fle07] Brett D. Fleisch. Program director’s column: can nuggets make a difference? Operating Systems Review, 41(1):3–4, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Franklin:2008:RDV

[FLM^{+}08] Jason Franklin, Mark Luk, Jonathan M. McCune, Arvind Seshadri, Adrian Perrig, and Leendert van Doorn. Remote detection of virtual machine monitors with fuzzy benchmarking. Operating Systems Review, 42(3):83–92, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feiertag:1977:PMS

[FLR77] R. J. Feiertag, K. N. Levitt, and L. Robinson. Proving multilevel security of a system design. Operating Systems Review, 11(5): 57–65, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feustel:1998:DUI

[FM98] Edward A. Feustel and Terry Mayfield. The DGSA: unmet information security challenges for operating system designers. Operating Systems Review, 32(1):3–22, January 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flautner:2002:VAP

[FM02] Krisztián Flautner and Trevor Mudge. Vertigo: automatic performance-setting for Linux. Operating Systems Review, 36 (5S):105–116, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Frost:2007:GFS

[FMK^{+}07] Christopher Frost, Mike Mammarella, Eddie Kohler, Andrew de los Reyes, Shant Hovsepian, Andrew Matsuoka, and Lei Zhang. Generalized file system dependencies. Operating Systems Review, 41(6):307–320, December 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feeley:1995:IGM

[FMP^{+}95] M. J. Feeley, W. E. Morgan, E. P. Pighin, A. R. Karlin, H. M. Levy, and C. A. Thekkath. Implementing global memory management in a workstation cluster. Operating Systems Review, 29(5):201–212, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Factor:2007:NPA

[FNRC^{+}07] Michael Factor, Dalit Naor, Simona Rabinovici-Cohen, Leeat Ramati, Petra Reshef, and Julian Satran. The need for preservation aware storage: a position paper. Operating Systems Review, 41(1):19–23, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feiertag:1972:MIO

[FO72] R. J. Feiertag and E. I. Organick. The Multics input/output system. Operating Systems Review, 6(1/2):35–38, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fridrich:1981:FFS

[FO81] M. Fridrich and W. Older. The Felix File Server. Operating Systems Review, 15(5):37–44, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fogel:1974:VPA

[Fog74] Marc H. Fogel. The VMOS paging algorithm: a practical implementation of the working set model. Operating Systems Review, 8(1):8–17, January 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fontao:1972:CAA

[Fon72] Rafael O. Fontao. A concurrent algorithm for avoiding deadlocks in multiprocess multiple resource systems. Operating Systems Review, 6(1/2):72–79, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fossum:1987:PXF

[Fos87] Timothy V. Fossum. PC-XINU features and installation. Operating Systems Review, 21(3):30–33, July 1987. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Foulk:1974:YAA

[Fou74] Clinton R. Foulk. Yet another attempt to define “structured programming”. Operating Systems Review, 8(3):4–5, July 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fleisch:1989:MCD

[FP89] B. Fleisch and G. Popek. Mirage: a coherent distributed shared memory design. Operating Systems Review, 23(5):211–223, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Furht:1989:PRI

[FPG89] Borko Furht, J. Parker, and D. Grostick. Performance of REAL/ IX^{T M}-fully preemptive real time UNIX. Operating Systems Review, 23(4):45–52, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fitzgerald:1985:IVM

[FR85] Robert Fitzgerald and Richard F. Rashid. The integration of virtual memory management and interprocess communication in accent (abstract only). Operating Systems Review, 19(5):13– 14, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fromentin:1994:LSD

[FR94] Eddy Fromentin and Michel Raynal. Local states in distributed computations: a few relations and formulas. Operating Systems Review, 28(2):65–72, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Francis:1980:SOS

[Fra80] N. D. Francis. Simulation of operating systems: a functional flowchart. Operating Systems Review, 14(3):16–21, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Franky:1995:DPS

[Fra95] Maria Consuelo Franky. DGDBM: programming support for distributed transactions over replicated files. Operating Systems Review, 29(3):64–74, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flatt:2000:CPC

[FRL00] Matthew Flatt, Alastair Reid, and Jay Lepreau. CpU: practical components for systems software. Operating Systems Review, 34 (2):32, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fong:1995:TFS

[FS95] Liana L. Fong and Mark S. Squillante. Time-function scheduling: a general approach to controllable resource. Operating Systems Review, 29(5):230, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ford:1996:CIS

[FS96] Bryan Ford and Sai Susarla. CPU inheritance scheduling. Operating Systems Review, 30(SI):91–105, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flinn:1999:EAA

[FS99] Jason Flinn and M. Satyanarayanan. Energy-aware adaptation for mobile applications. Operating Systems Review, 33(5):48–63, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flinn:2000:EAA

[FS00] Jason Flinn and M. Satyanarayanan. Energy-aware adaptation for mobile applications. Operating Systems Review, 34(2):13– 14, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fetzer:2008:SED

[FS08a] Christof Fetzer and Martin Süßkraut. Switchblade: enforcing dynamic personalized system call models. Operating Systems Review, 42(4):273–286, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fong:2008:DVS

[FS08b] Liana Fong and Małgorzata Steinder. Duality of virtualization: simplification and complexity. Operating Systems Review, 42 (1):96–97, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fuller:1973:SMT

[Ful73] Samuel H. Fuller. Summary of minimal-total-processing-time drum and disk scheduling disciplines. Operating Systems Review, 7(4):54–57, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Flautner:2000:TLP

[FURM00] Kristián Flautner, Rich Uhlig, Steve Reinhardt, and Trevor Mudge. Thread-level parallelism and interactive performance of desktop applications. Operating Systems Review, 34(5):129–138, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Freiling:2006:IIC

[FV06] Felix C. Freiling and Hagen Völzer. Illustrating the impossibility of crash-tolerant consensus in asynchronous systems. Operating Systems Review, 40(2):105–109, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ferles:2020:SRA

[FVDS20] Kostas Ferles, Jacob Van Geffen, Isil Dillig, and Yannis Smaragdakis. Symbolic reasoning for automatic signal placement. Operating Systems Review, 54(1):64–76, August 2020. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL

https://dl.acm.org/doi/10.1145/3421473.3421482.

Fogel:1972:EID

[FW72] Marc Fogel and Joseph Winograd. EINSTEIN: an internal driver in a time-sharing environment. Operating Systems Review, 6(3):6–14, October 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Friedman:1977:AAP

[FW77] Daniel P. Friedman and David S. Wise. Aspects of applicative programming for file systems (preliminary version). Operating Systems Review, 11(2):41–55, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ferraiuolo:2017:VPH

[FXZ^{+}17] Andrew Ferraiuolo, Rui Xu, Danfeng Zhang, Andrew C. Myers, and G. Edward Suh. Verification of a practical hardware security architecture through static information flow analysis. Operating Systems Review, 51(2):555–568, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fan:2016:CSG

[FZL16] Songchun Fan, Seyed Majid Zahedi, and Benjamin C. Lee. The computational sprinting game. Operating Systems Review, 50 (2):561–575, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fingler:2023:DGA

[FZY^{+}23] Henrique Fingler, Zhiting Zhu, Esther Yoon, Zhipeng Jia, Emmett Witchel, and Christopher J. Rossbach. Disaggregated GPU acceleration for serverless applications. Operating Systems Review, 57(1):10–20, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3606557.3606560.

Govindan:1991:SIM

[GA91] Ramesh Govindan and David P. Anderson. Scheduling and IPC mechanisms for continuous media. Operating Systems Review, 25(5):68–80, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gritzalis:1998:SIS

[GA98] Stefanos Gritzalis and George Aggelis. Security issues surrounding programming languages for mobile code: JAVA vs. Safe-Tcl. Operating Systems Review, 32(2):16–32, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gulati:2008:TDS

[GA08] Ajay Gulati and Irfan Ahmad. Towards distributed storage resource management using flow control. Operating Systems Review, 42(6):10–16, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gaines:1972:OSB

[Gai72] R. Stockton Gaines. An operating system based on the concept of a supervisory computer. Operating Systems Review, 6(1/2): 17–23, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gaines:1975:P

[Gai75] R. Stockton Gaines. Protection. Operating Systems Review, 9(3):57–58, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gait:1978:EEP

[Gai78] Jason Gait. Easy entry: the password encryption problem. Operating Systems Review, 12(3):54–60, July 1978. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goel:2002:STS

[GAK^{+}02] Ashvin Goel, Luca Abeni, Charles Krasic, Jim Snow, and Jonathan Walpole. Supporting time-sensitive applications on a commodity OS. Operating Systems Review, 36(5S):165–180, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gannon:1977:EED

[Gan77] J. D. Gannon. An experimental evaluation of data types on programming reliability. Operating Systems Review, 11(2):141, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gantenbein:1992:ABD

[Gan92] Rex E. Gantenbein. An annotated bibliography of dependable distributed computing. Operating Systems Review, 26(2):60–81, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ganti:2008:PAL

[Gan08] Ashwin Ganti. Plan 9 authentication in Linux. Operating Systems Review, 42(5):27–33, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garfinkel:2007:CDV

[Gar07] Simson L. Garfinkel. Complete delete vs. time machine computing. Operating Systems Review, 41(1):42–44, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ghita:2013:TAP

[GAT13] Denisa Ghita, Katerina Argyraki, and Patrick Thiran. Toward accurate and practical network tomography. Operating Systems Review, 47(1):22–26, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goscinski:1990:RML

[GB90] Andrzej Goscinski and Mirion Bearman. Resource management in large distributed systems. Operating Systems Review, 24(4): 7–25, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gopalakrishnan:1993:PPA

[GB93] R. Gopalakrishnan and Andreas D. Bovopoulos. A protocol processing architecture for networked multimedia computers. Operating Systems Review, 27(3):19–33, July 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2001:RFR

[GB01] B. Gupta and S. K. Banerjee. A Roll-Forward Recovery Scheme for Solving the Problem of Coasting Forward for Distributed Systems. Operating Systems Review, 35(3):55–66, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gifford:1985:ALS

[GBBL85] David K. Gifford, Robert W. Baldwin, Stephen T. Berlin, and John M. Lucassen. An architecture for large scale information systems. Operating Systems Review, 19(5):161–170, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gribble:2000:PDD

[GBCH00] Steven D. Gribble, Eric A. Brewer, David Culler, and Joseph M. Hellerstein. Persistent distributed data structures to simplify cluster-based Internet services. Operating Systems Review, 34 (2):37–38, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guerra:2010:EPS

[GBG^{+}10] Jorge Guerra, Wendy Belluomini, Joseph Glider, Karan Gupta, and Himabindu Pucha. Energy proportionality for storage: impact and feasibility. Operating Systems Review, 44(1):35–39, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goldszmidt:2010:TAP

[GBZP10] Moises Goldszmidt, Mihai Budiu, Yue Zhang, and Michael Pechuk. Toward automatic policy refinement in repair services for large distributed systems. Operating Systems Review, 44(2): 47–51, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gray:1989:LEF

[GC89] C. Gray and D. Cheriton. Leases: an efficient fault-tolerant mechanism for distributed file cache consistency. Operating Systems Review, 23(5):202–210, December 1989. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Greenwald:1996:SBN

[GC96] Michael Greenwald and David Cheriton. The synergy between non-blocking synchronization and operating system structure. Operating Systems Review, 30(SI):123–136, October 1996. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guerrero:2005:ECB

[GC05] Jorge Herrerías Guerrero and Roberto Gómez Cárdenas. An example of communication between security tools: IPTables — Snort. Operating Systems Review, 39(3):34–43, July 2005. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ge:2008:PDQ

[GC08] Ping Ge and Hailong Cai. Providing differentiated QoS for peer-to-peer file sharing systems. Operating Systems Review, 42(6):17–23, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goel:2012:RTP

[GC12] Atul Goel and Peter Corbett. RAID triple parity. Operating Systems Review, 46(3):41–49, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ge:2017:GGC

[GCJ17] Xinyang Ge, Weidong Cui, and Trent Jaeger. GRIFFIN: Guarding control flows using Intel processor trace. Operating Systems Review, 51(2):585–598, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gallagher:1994:DMD

[GCM^{+}94] David M. Gallagher, William Y. Chen, Scott A. Mahlke, John C. Gyllenhaal, and Wen mei W. Hwu. Dynamic memory disambiguation using the memory conflict buffer. Operating Systems Review, 28(5):183–193, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gummaraju:2008:SPG

[GCTR08] Jayanth Gummaraju, Joel Coburn, Yoshio Turner, and Mendel Rosenblum. Streamware: programming general-purpose multicore processors using streams. Operating Systems Review, 42 (2):297–307, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gueye:2013:CMA

[GDRT13] Soguy M. K. Gueye, Noel De Palma, Eric Rutten, and Alain Tchana. Coordinating multiple administration loops using discrete control. Operating Systems Review, 47(3):18–25, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gertz:1972:SRH

[Ger72] Jeffrey L. Gertz. Storage reallocation in hierarchical associative memories. Operating Systems Review, 6(1/2):58–63, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gerber:1977:PSC

[Ger77] A. J. Gerber. Process synchronization by counter variables. Operating Systems Review, 11(4):6–17, October 1977. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gracioli:2015:DER

[GF15] Giovani Gracioli and Antônio Augusto Fröhlich. On the design and evaluation of a real-time operating system for cachecoherent multicore architectures. Operating Systems Review, 49 (2):2–16, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goel:2008:RSS

[GFPcF08] Ashvin Goel, Kamran Farhadi, Kenneth Po, and Wu chang Feng. Reconstructing system state for intrusion analysis. Operating Systems Review, 42(3):21–28, April 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garey:1973:BSL

[GG73] M. R. Garey and R. L. Grehem. Bounds on scheduling with limited resources. Operating Systems Review, 7(4):104–111, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gopal:1991:SFT

[GG91] Gita Gopal and Nancy D. Griffeth. Software fault tolerance in telecommunications systems. Operating Systems Review, 25 (2):112–116, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gharachorloo:1991:PEM

[GGH91] Kourosh Gharachorloo, Anoop Gupta, and John Hennessy. Performance evaluation of memory consistency models for sharedmemory multiprocessors. Operating Systems Review, 25(3S): 245–257, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goldsack:2009:SCM

[GGL^{+}09] Patrick Goldsack, Julio Guijarro, Steve Loughran, Alistair Coles, Andrew Farrell, Antonio Lain, Paul Murray, and Peter Toft. The SmartFrog configuration management framework. Operating Systems Review, 43(1):16–25, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goyal:1996:HCS

[GGV96] Pawan Goyal, Xingang Guo, and Harrick M. Vin. A hierarchical CPU scheduler for multimedia operating systems. Operating Systems Review, 30(SI):107–121, October 1996. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guttag:1977:SEA

[GHM77] Jhon V. Guttag, Ellis Horowitz, and David R. Musser. Some extensions to algebraic specifications. Operating Systems Review, 11(2):63–67, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guy:1992:FRF

[GHP92] Richard G. Guy, John S. Heidemann, and Thomas W. Page, Jr. The Ficus Replicated File System. Operating Systems Review, 26(2):26, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Grace:2008:EOO

[GHP^{+}08] Paul Grace, Danny Hughes, Barry Porter, Gordon S. Blair, Geoff Coulson, and Francois Taiani. Experiences with open overlays: a middleware approach to network heterogeneity. Operating Systems Review, 42(4):123–136, May 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ghodsi:2007:ESB

[GHW07] Ali Ghodsi, Seif Haridi, and Hakim Weatherspoon. Exploiting the synergy between gossiping and structured overlays. Operating Systems Review, 41(5):61–66, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gifford:1981:CSI

[Gif81] David K. Gifford. Cryptographic sealing for information secrecy and authentication. Operating Systems Review, 15(5):123–124, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gilbert:1978:MSL

[Gil78] D. C. Gilbert. Modeling spin locks with queuing networks. Operating Systems Review, 12(1):29–42, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Girling:1982:ORH

[Gir82] C. Gray Girling. Object representation on a heterogeneous network. Operating Systems Review, 16(4):49–59, October 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gifford:1991:SFS

[GJSO91] David K. Gifford, Pierre Jouvelot, Mark A. Sheldon, and James W. O’Toole, Jr. Semantic file systems. Operating Systems Review, 25(5):16–25, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guangchun:2003:ABL

[GJXJ03a] Luo Guangchun, Zhang Jun, Lu Xianliang, and Lu Jun. Active block layout: a high performance disk layout mechanism. Operating Systems Review, 37(1):5–13, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guangchun:2003:HNC

[GJXJ03b] Luo Guangchun, Zhang Jun, Lu Xianliang, and Lu Jun. HCCM: a novel cache consistence mechanism. Operating Systems Review, 37(2):25–36, April 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goldenberg:1991:VVI

[GKD91] Ruth E. Goldenberg, Lawrence J. Kenah, and Denise E. Dumas. VAX/VMS internals and data structures: version 5.2. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1991. ISBN 1-55558-059-9. xxvi + 1427 pp. LCCN QA76.76.O63 G638 1991.

Ghemawat:1995:UMO

[GKL95] Sanjay Ghemawat, M. Frans Kaashoek, and Barbara Liskov. Using a modified object buffer to improve the write performance of an object-oriented database. Operating Systems Review, 29 (5):235, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gibson:2000:FVS

[GKO^{+}00] Jeff Gibson, Robert Kunz, David Ofelt, Mark Horowitz, John Hennessy, and Mark Heinrich. FLASH vs. (simulated) FLASH: closing the simulation loop. Operating Systems Review, 34(5): 49–58, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2011:ASP

[GKS11] Vishakha Gupta, Rob Knauerhase, and Karsten Schwan. Attaining system performance points: revisiting the end-to-end argument in system design for heterogeneous many-core systems. Operating Systems Review, 45(1):3–10, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guerraoui:2007:SBS

[GKV07] Rachid Guerraoui, Michal Kapalka, and Jan Vitek. STM-Bench7: a benchmark for software transactional memory. Operating Systems Review, 41(3):315–324, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gu:1999:EJT

[GLC99] Yan Gu, B. S. Lee, and Wentong Cai. Evaluation of Java thread performance on two different multithreaded kernels. Operating Systems Review, 33(1):34–46, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gan:2022:EPC

[GLD^{+}22] Yu Gan, Mingyu Liang, Sundar Dev, David Lo, and Christina Delimitrou. Enabling practical cloud performance debugging with unsupervised learning. Operating Systems Review, 56(1): 34–41, June 2022. CODEN OSRED8. ISSN 0163-5980 (print),

1943-586X (electronic). URL https://dl.acm.org/doi/10.
1145/3544497.3544503.

Guerrero:1993:IEA

[GLG93] R. Guerrero, L. Leguizamon, and R. Gallard. Implementation and evaluation of alternative process schedulers in MINIX. Operating Systems Review, 27(1):79–100, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2004:DDD

[GLL04] B. Gupta, Z. Liu, and Z. Liang. On designing direct dependency: based fast recovery algorithms for distributed systems. Operating Systems Review, 38(1):58–73, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ghosh:1998:PMA

[GMM98] Somnath Ghosh, Margaret Martonosi, and Sharad Malik. Precise miss analysis for program transformations with caches of arbitrary associativity. Operating Systems Review, 32(5):228–239, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Geschke:1977:EEM

[GMS77] Charles M. Geschke, James H. Morris, Jr., and Edwin H. Satterthwaite. Early experience with Mesa. Operating Systems Review, 11(2):138, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gangwani:2016:CBS

[GMT16] Tanmay Gangwani, Adam Morrison, and Josep Torrellas. CAS-PAR: Breaking serialization in lock-free multicore synchronization. Operating Systems Review, 50(2):789–804, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garnett:1980:AGC

[GN80] N. H. Garnett and R. M. Needham. An asynchronous garbage collector for the Cambridge File Server. Operating Systems Review, 14(4):36–40, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Grunwald:1996:WPO

[GN96] Dirk Grunwald and Richard Neves. Whole-program optimization for time and space efficient threads. Operating Systems Review, 30(5):50–59, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gibson:1998:CEH

[GNA^{+}98] Garth A. Gibson, David F. Nagle, Khalil Amiri, Jeff Butler, Fay W. Chang, Howard Gobioff, Charles Hardin, Erik Riedel, David Rochberg, and Jim Zelenka. A cost-effective, highbandwidth storage architecture. Operating Systems Review, 32 (5):92–103, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Grider:2009:CGF

[GNB^{+}09] Gary Grider, James Nunez, John Bent, Steve Poole, Rob Ross, and Evan Felix. Coordinating government funding of file system and I/O research through the high end computing university research activity. Operating Systems Review, 43(1):2–7, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gong:1989:SCB

[Gon89] Li Gong. On security in capability-based systems. Operating Systems Review, 23(2):56–60, April 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gong:1992:SRD

[Gon92] Li Gong. A security risk of depending on synchronized clocks. Operating Systems Review, 26(1):49–53, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goodman:1987:CMV

[Goo87] James R. Goodman. Coherency for multiprocessor virtual address caches. Operating Systems Review, 21(4):72–81, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gorski:1978:MRA

[Gór78] Janusz Górski. A modular representation of the access control system. Operating Systems Review, 12(3):61–77, July 1978. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gordon:1987:WMG

[Gor87] Robert L. Gordon. Window management, graphics, and operating systems. Operating Systems Review, 21(3):5–8, July 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gordon:2006:SSP

[Gor06] Minor Gordon. Small-scale peer-to-peer overlays. Operating Systems Review, 40(3):45–48, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gopalakrishnan:1995:RTU

[GP95] R. Gopalakrishnan and Guru M. Parulkar. A real-time upcall facility for protocol processing with QoS guarantees. Operating Systems Review, 29(5):231, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gil:2005:TCS

[GP05] Marisa Gil and Ruben Pinilla. Thread coloring: a scheduler proposal from user to hardware threads. Operating Systems Review, 39(2):54–70, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Goel:2005:TIR

[GPF^{+}05] Ashvin Goel, Kenneth Po, Kamran Farhadi, Zheng Li, and Eyal de Lara. The Taser intrusion recovery system. Operating Systems Review, 39(5):163–176, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gunawi:2007:IFS

[GPK^{+}07] Haryadi S. Gunawi, Vijayan Prabhakaran, Swetha Krishnan, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Improving file system reliability with I/O shepherding. Operating Systems Review, 41(6):293–306, December 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Georgiou:1987:ECI

[GPR87] C. J. Georgiou, S. L. Palmer, and P. L. Rosenfeld. An experimental coprocessor for implementing persistent objects on an IBM 4381. Operating Systems Review, 21(4):84–87, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gomaa:2004:HRL

[GPV04] Mohamed Gomaa, Michael D. Powell, and T. N. Vijaykumar. Heat-and-run: leveraging SMT and CMP to manage power density through the operating system. Operating Systems Review, 38(5):260–270, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gao:2017:TSE

[GPY^{+}17] Mingyu Gao, Jing Pu, Xuan Yang, Mark Horowitz, and Christos Kozyrakis. TETRIS: Scalable and efficient neural network acceleration with 3D memory. Operating Systems Review, 51 (2):751–764, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gray:2014:SPT

[Gra14] Cary Gray. SOSP Professional Travel Scholarship: Reflections by recipient Cary Gray. Operating Systems Review, 48(2):24, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ganapathy:2008:DIM

[GRB^{+}08] Vinod Ganapathy, Matthew J. Renzelmann, Arini Balakrishnan, Michael M. Swift, and Somesh Jha. The design and implementation of microdrivers. Operating Systems Review, 42 (2):168–178, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Greenberg:1972:ADS

[Gre72] Mark L. Greenberg. An algorithm for drum storage management in time-sharing systems. Operating Systems Review, 6(1/2):141– 148, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gaines:1978:SSP

[GS78] R. Stockton Gaines and Norman Z. Shapiro. Some security principles and their application to computer security. Operating Systems Review, 12(3):19–28, July 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gopinath:1989:CWO

[GS89] P. Gopinath and K. Schwan. CHAOS: why one cannot have only an operating system for real-time applications. Operating Systems Review, 23(3):106–125, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ghosh:1990:CST

[GS90] H. Ghosh and S. Sreedhar. A comparative study of two simple network file access models. Operating Systems Review, 24(4):68– 77, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gwertzman:1995:ARA

[GS95] James S. Gwertzman and Margo Seltzer. Autonomous replication across wide-area internetworks. Operating Systems Review, 29(5):234, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gorti:2013:RAD

[GS13] Naga Pavan Kumar Gorti and Arun K. Somani. Reliability aware dynamic voltage and frequency scaling for improved microprocessor lifetime. Operating Systems Review, 47(3):10–17, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ghanbari:2010:SDQ

[GSA10] Saeed Ghanbari, Gokul Soundararajan, and Cristiana Amza. SelfTalk for Dena: query language and runtime support for evaluating system behavior. Operating Systems Review, 44(1):30– 34, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gschwind:1994:FAU

[Gsc94] Michael K. Gschwind. FTP access as a user-defined file system. Operating Systems Review, 28(2):73–80, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guo:2016:HDI

[GSCM16] Qing Guo, Karin Strauss, Luis Ceze, and Henrique S. Malvar. High-density image storage using approximate memory cells. Operating Systems Review, 50(2):413–426, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ganger:2000:IMB

[GSGN00] Greg Ganger, Steve Schlosser, John Griffin, and David Nagle. Incorporating MEMS-based storage into computer systems. Operating Systems Review, 34(2):30, April 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2008:MSP

[GSM08] Karan Gupta, Prasenjit Sarkar, and Lesley Mbogo. MIRAGE: storage provisioning in large data centers using balanced component utilizations. Operating Systems Review, 42(1):104–105, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gharachorloo:2000:ADA

[GSSV00] Kourosh Gharachorloo, Madhu Sharma, Simon Steely, and Stephen Van Doren. Architecture and design of AlphaServer GS320. Operating Systems Review, 34(5):13–24, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2017:HCS

[GSW^{+}17] Abhishek Gupta, Rick Spillane, Wenguang Wang, Maxime Austruy, Vahid Fereydouny, and Christos Karamanolis. Hybrid cloud storage: Bridging the gap between compute clusters and cloud storage. Operating Systems Review, 51(1):48–53, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gordon:2006:ECG

[GTA06] Michael I. Gordon, William Thies, and Saman Amarasinghe. Exploiting coarse-grained task, data, and pipeline parallelism in stream programs. Operating Systems Review, 40(5):151–162, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Govil:1999:CDR

[GTHR99] Kinshuk Govil, Dan Teodosiu, Yongqiang Huang, and Mendel Rosenblum. Cellular Disco: resource management using virtual clusters on shared-memory multiprocessors. Operating Systems Review, 33(5):154–169, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Govil:2000:CDR

[GTHR00] Kingshuk Govil, Dan Teodosiu, Yongqiang Huang, and Mendel Rosenblum. Cellular disco: resource management using virtual clusters on shared-memory multiprocessors. Operating Systems Review, 34(2):21, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gordon:2002:SCC

[GTK^{+}02] Michael I. Gordon, William Thies, Michal Karczmarek, Jasper Lin, Ali S. Meli, Andrew A. Lamb, Chris Leger, Jeremy Wong, Henry Hoffmann, David Maze, and Saman Amarasinghe. A stream compiler for communication-exposed architectures. Operating Systems Review, 36(5):291–303, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gidra:2011:ASG

[GTSS11] Lokesh Gidra, Gaël Thomas, Julien Sopena, and Marc Shapiro. Assessing the scalability of garbage collectors on many cores. Operating Systems Review, 45(3):15–19, December 2011. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guerra:2008:CAB

[GUB^{+}08] Jorge Guerra, Luis Useche, Medha Bhadkamkar, Ricardo Koller, and Raju Rangaswami. The case for active block layer extensions. Operating Systems Review, 42(6):3–9, October 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guenther:1987:REU

[Gue87] G. R. Guenther. Running 7th edition UNIX programs on a VAX in compatibility mode. Operating Systems Review, 21(1):30–33, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guenther:1988:ECS

[Gue88] Grant R. Guenther. Extended control services in operating system interfaces. Operating Systems Review, 22(4):20–24, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2001:DBA

[Gup01] Vijay Gupta. A distributed backoff algorithm to support realtime traffic on Ethernet. Operating Systems Review, 35(3):43– 66, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gupta:2005:TAI

[Gup05] Vijay Shivshanker Gupta. Trust and accountability issues in scalable invalidation-based Web cache consistency. Operating Systems Review, 39(4):23–36, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gurumurthi:2007:SDS

[Gur07] Sudhanva Gurumurthi. Should disks be speed demons or brainiacs? Operating Systems Review, 41(1):33–36, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Geva:2004:CFI

[GW04] Mordechai Geva and Yair Wiseman. A common framework for inter-process communication on a cluster. Operating Systems Review, 38(4):33–44, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gwinn:1994:SMT

[Gwi94] Joe Gwinn. Some measurements of timeline gaps in VAX/VMS. Operating Systems Review, 28(2):92–96, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gwinn:2005:QSV

[Gwi05] Joseph M. Gwinn. Quality-of-service versus Realtime. Operating Systems Review, 39(4):18–22, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ge:2008:DSS

[GWSY08] Yi Ge, Chen Wang, Xiaowei Shen, and Honesty Young. A database scale-out solution for emerging write-intensive commercial workloads. Operating Systems Review, 42(1):102–103, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guangchun:2003:MND

[GXJJ03] Luo Guangchun, Lu Xianliang, Li Jiong, and Zhang Jun. MA-DIDS: a novel distributed IDS based on mobile agent. Operating Systems Review, 37(1):46–53, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gan:2019:LDL

[GZH^{+}19] Yu Gan, Yanqi Zhang, Kelvin Hu, Dailun Cheng, Yuan He, Meghna Pancholi, and Christina Delimitrou. Leveraging deep learning to improve performance predictability in cloud microservices with Seer. Operating Systems Review, 53(1):34–39, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Habermann:1972:SCP

[Hab72] A. Nico Habermann. Synchronization of communicating processes. Operating Systems Review, 6(1/2):80–85, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2017:PSA

[HABZ17] Jian Huang, Michael Allen-Bond, and Xuechen Zhang. Pallas: Semantic-aware checking for finding deep bugs in fast path. Operating Systems Review, 51(2):709–722, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hac:1985:DFS

[Hac85] Anna Hac. Distributed file systems — a survey. Operating Systems Review, 19(1):15–18, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haddon:1977:NMC

[Had77] Bruce K. Haddon. Nested monitor calls. Operating Systems Review, 11(4):18–23, October 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haddon:1983:RTA

[Had83] Bruce K. Haddon. Review of “Technical aspects of data communication”: (second edition) by John E. McNamara. Digital Press, Educational Services, Digital Equipment Corporation, Bedford, Massachusetts, 1982. Operating Systems Review, 17(3):7, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [McN77, McN82, McN88].

Haddon:1984:BRS

[Had84] Bruce K. Haddon. Book review of “Security, IFIP/Sec’83: proceedings of the first security conference” North-Holland Publishing Co. 1983. Operating Systems Review, 18(3):14, July 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Fêa83].

Haddon:1985:RIS

[Had85] Bruce K. Haddon. Review of “Information systems design methodologies: a feature analysis”. Edited by T. W. Olle, H. G. Sol, and C. J. Tully. North-Holland Publishing Co. 1983. Operating Systems Review, 19(4):4–5, October 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [OSV82, OST83, OSV86].

Haddon:1993:BRG

[Had93] Bruce K. Haddon. Book review: Global Software: Developing Applications for the International Market by Dave Taylor: (Springer-Verlag, New York 1992). Operating Systems Review, 27(1):5–6, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haddon:2001:ISS

[Had01] Bruce K. Haddon. IEEE storage system standards. Operating Systems Review, 35(1):8–16, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haeberlen:2010:CAC

[Hae10] Andreas Haeberlen. A case for the accountable cloud. Operating Systems Review, 44(2):52–57, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hunt:2007:SPI

[HAF^{+}07] Galen Hunt, Mark Aiken, Manuel Fähndrich, Chris Hawblitzel, Orion Hodson, James Larus, Steven Levi, Bjarne Steensgaard, David Tarditi, and Ted Wobber. Sealing OS processes to improve dependability and safety. Operating Systems Review, 41 (3):341–354, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hagmann:1987:RCF

[Hag87] R. Hagmann. Reimplementing the Cedar file system using logging and group commit. Operating Systems Review, 21(5): 155–162, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Halvorsen:2000:NLFb

[Hal00a] Pål Halvorsen. Network level farming: speeding up a multimedia storage server. Operating Systems Review, 34(2):41, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Halvorsen:2000:NLFa

[Hal00b] Pål Halvorsen. Network level framing: speeding up a multimedia storage server. Operating Systems Review, 34(2):34, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hansen:1972:STS

[Han72] Per Brinch Hansen. Short-term scheduling in multiprogramming systems. Operating Systems Review, 6(1/2):101–105, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hansen:1983:UPC

[Han83] Per Brinch Hansen. Using personal computers in operating system courses. Operating Systems Review, 17(3):41–44, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harper:1982:MEW

[Har82] M. E. Harper. Mutual exclusion within both software- and hardware-driven kernel primitives. Operating Systems Review, 16(4):60–68, October 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hardy:1985:KA

[Har85] Norman Hardy. KeyKOS architecture. Operating Systems Review, 19(4):8–25, October 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harris:1988:IOS

[Har88] David L. Harris. An input/output subsystem for the Hawk operating system kernel. Operating Systems Review, 22(2):32–44, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hatkanagalekar:1994:NSI

[Hat94] Pradeep Hatkanagalekar. A note on structured interrupts. Operating Systems Review, 28(2):88–91, April 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hoch:1980:ICP

[HB80] Charles Hoch and J. C. Browne. An implementation of capabilities on the PDP-11/45. Operating Systems Review, 14(3): 22–32, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hales:2006:TAS

[HB06] David Hales and Özalp Babaoğlu. Towards automatic social bootstrapping of peer-to-peer protocols. Operating Systems Review, 40(3):56–60, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harji:2013:OTL

[HBB13] Ashif S. Harji, Peter A. Buhr, and Tim Brecht. Our troubles with Linux Kernel upgrades and why you should care. Operating Systems Review, 47(2):66–72, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harchol-Balter:1995:EPL

[HBD95] Mor Harchol-Balter and Allen B. Downey. Exploiting process lifetime distributions for dynamic load balancing. Operating Systems Review, 29(5):236, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herder:2006:MHR

[HBG^{+}06] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. MINIX 3: a highly reliable, selfrepairing operating system. Operating Systems Review, 40(3): 80–89, July 2006. CODEN OSRED8. ISSN 0163-5980 (print),

1943-586X (electronic). URL http://www.minix3.org/doc/
OSR-2006.pdf;   http://www.minix3.org/docs/jorrit-herder/
osr-jul06.pdf.

Huang:2006:PMA

[HBP06] Mark Huang, Andy Bavier, and Larry Peterson. PlanetFlow: maintaining accountability for network services. Operating Systems Review, 40(1):89–94, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harty:1992:ACP

[HC92] Kieran Harty and David R. Cheriton. Application-controlled Physical Memory using External Page-Cache Management. Operating Systems Review, 26(2):19, April 1992. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hu:1995:YCE

[HC95] Ping Hu and Bruce Christianson. Is your computing environment secure?: security problems with interrupt handling mechanisms. Operating Systems Review, 29(4):87–96, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hauswirth:2004:LOM

[HC04] Matthias Hauswirth and Trishul M. Chilimbi. Low-overhead memory leak detection using adaptive statistical profiling. Operating Systems Review, 38(5):156–164, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huh:2004:CDM

[HCBS04] Jaehyuk Huh, Jichuan Chang, Doug Burger, and Gurindar S. Sohi. Coherence decoupling: making use of incoherence. Operating Systems Review, 38(5):97–106, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heath:2006:MFT

[HCG^{+}06] Taliver Heath, Ana Paula Centeno, Pradeep George, Luiz Ramos, and Yogesh Jaluria. Mercury and Freon: temperature emulation and management for server systems. Operating Systems Review, 40(5):106–116, December 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hansen:2007:ETT

[HCJ07] Jacob Gorm Hansen, Eske Christiansen, and Eric Jul. Evil twins: two models for TCB reduction in HPC clusters. Operating Systems Review, 41(4):20–29, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:2008:MVE

[HCK08] Shan He, Renan G. Cattelan, and Darko Kirovski. Modeling viral economies for digital media. Operating Systems Review, 42(4):149–162, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hammond:2004:PTC

[HCW^{+}04] Lance Hammond, Brian D. Carlstrom, Vicky Wong, Ben Hertzberg, Mike Chen, Christos Kozyrakis, and Kunle Olukotun. Programming with transactional coherence and consistency (TCC). Operating Systems Review, 38(5):1–13, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:1997:MDA

[HCZ97] Yanxiang He, Donald H. Cooley, and Jianping Zhang. A model for a distributed OS automatic generation system. Operating Systems Review, 31(4):78–84, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:1998:PMM

[HCZ98] Yanxiang He, Donald H. Cooley, and Jianping Zhang. Planning management of multiagent-based distributed open: computing environment model. Operating Systems Review, 32(1):57–64, January 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hagimont:2012:SAE

[HD12] Daniel Hagimont and Noel De Palma. A simulator to assess energy saving strategies and policies in HPC workloads. Operating Systems Review, 46(2):2–9, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hines:2009:PCL

[HDG09] Michael R. Hines, Umesh Deshpande, and Kartik Gopalan. Post-copy live migration of virtual machines. Operating Systems Review, 43(3):14–26, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hoang:2021:CCS

[HDGP21] Loc Hoang, Roshan Dathathri, Gurbinder Gill, and Keshav Pingali. CuSP: a customizable streaming edge partitioner for distributed graph analytics. Operating Systems Review, 55(1):47– 60, June 2021. CODEN OSRED8. ISSN 0163-5980 (print),

1943-586X (electronic). URL https://dl.acm.org/doi/10.
1145/3469379.3469385.

Hayashi:1994:AAS

[HDH^{+}94] Kenichi Hayashi, Tsunehisa Doi, Takeshi Horie, Yoichi Koyanagi, Osamu Shiraki, Nobutaka Imamura, Toshiyuki Shimizu, Hiroaki Ishihata, and Tatsuya Shindo. AP1000+: architectural support of PUT/GET interface for parallelizing compiler. Operating Systems Review, 28(5):196–207, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:2002:FBC

[HDL^{+}02] Yanxiang He, Zhuomin Du, Xuhui Li, Donald H. Cooley, and Jing He. A field-based collaboration strategy in MADCE. Operating Systems Review, 36(1):88–96, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harry:1995:DVF

[HdRC95] Michael Harry, Juan Miguel del Rosario, and Alok Choudhary. The design of VIP-FS: a virtual, parallel file system for high performance parallel and distributed computing. Operating Systems Review, 29(3):35–48, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heimbigner:1978:WDD

[Hei78] Dennis Heimbigner. Writing device drivers in Concurrent Pascal. Operating Systems Review, 12(4):16–33, October 1978. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heiser:2007:TTC

[HEK^{+}07] Gernot Heiser, Kevin Elphinstone, Ihor Kuz, Gerwin Klein, and Stefan M. Petters. Towards trustworthy computing systems: taking microkernels to the next level. Operating Systems Review, 41(4):3–11, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Honig:2011:SES

[HEKSP11] Timo Hönig, Christopher Eibel, Rüdiger Kapitza, and Wolfgang Schröder-Preikschat. SEEP: exploiting symbolic execution for energy-aware programming. Operating Systems Review, 45(3): 58–62, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hemmendinger:1988:CIG

[Hem88] David Hemmendinger. A correct implementation of general semaphores. Operating Systems Review, 22(3):42–44, July 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hemmendinger:1989:CCU

[Hem89] David Hemmendinger. Comments on “A Correct and Unrestrictive Implementation of General Semaphores”. Operating Systems Review, 23(1):7–8, January 1989. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Kea88, Hem88, TT00].

Herriot:1977:TIP

[Her77] Robert G. Herriot. Towards the ideal programming language. Operating Systems Review, 11(2):56–62, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herbert:1978:NPA

[Her78] A. J. Herbert. A new protection architecture for the Cambridge capability computer. Operating Systems Review, 12(1):24–28, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herlihy:1986:CHA

[Her86] Maurice Herlihy. Comparing how atomicity mechanisms support replication. Operating Systems Review, 20(3):31–39, July 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herlihy:1987:OCC

[Her87] Maurice Herlihy. Optimistic concurrency control for abstract data types. Operating Systems Review, 21(2):33–44, April 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herlihy:1992:MIH

[Her92a] Maurice Herlihy. A Methodology for Implementing Highly Concurrent Data Objects. Operating Systems Review, 26(2):12, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herrtwich:1992:SSI

[Her92b] Ralf Guido Herrtwich. Summary of the Second International Workshop on Network and Operating System Support for Digital Audio and Video. Operating Systems Review, 26(2):32–59, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herbert:2007:WHP

[Her07] Andrew Herbert. What happened to Pastry. Operating Systems Review, 41(2):10–16, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herrod:2010:SRD

[Her10] Stephen Alan Herrod. Systems research and development at VMware. Operating Systems Review, 44(4):1–2, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heuring:1997:BRE

[Heu97] Vincent P. Heuring. Book review: Essential Java Fast, John Cowell. Operating Systems Review, 31(4):2, October 1997. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:2008:DOB

[HF08] Shuibing He and Dan Feng. Design of an object-based storage device based on I/O processor. Operating Systems Review, 42 (6):30–35, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ho:2006:PTB

[HFC^{+}06] Alex Ho, Michael Fetterman, Christopher Clark, Andrew Warfield, and Steven Hand. Practical taint-based protection using demand emulation. Operating Systems Review, 40(4):29– 41, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hayes:1987:ADE

[HFWZ87] John R. Hayes, Martin E. Fraeman, Robert L. Williams, and Thomas Zaremba. An architecture for the direct execution of the Forth programming language. Operating Systems Review, 21(4):42–49, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hebbard:1980:PAM

[HGB^{+}80] B. Hebbard, P. Grosso, T. Baldridge, C. Chan, D. Fishman, P. Goshgarian, T. Hilton, J. Hoshen, K. Hoult, G. Huntley, M. Stolarchuk, and L. Warner. A penetration analysis of the Michigan Terminal System. Operating Systems Review, 14(1): 7–20, January 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heinlein:1994:IMP

[HGDG94] John Heinlein, Kourosh Gharachorloo, Scott Dresser, and Anoop Gupta. Integration of message passing and shared memory in the Stanford FLASH multiprocessor. Operating Systems Review, 28(5):38–50, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hendricks:2007:LOB

[HGR07] James Hendricks, Gregory R. Ganger, and Michael K. Reiter. Low-overhead Byzantine fault-tolerant storage. Operating Systems Review, 41(6):73–86, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hardy:1988:CDW

[HH88] Norm Hardy and Norman Hardy. The Confused Deputy: (or why capabilities might have been invented). Operating Systems Review, 22(4):36–38, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Holmes:1989:DPH

[HH89] V. P. Holmes and D. L. Harris. A designer’s perspective of the Hawk multiprocessor operating system kernel. Operating Systems Review, 23(3):158–172, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huebner:2008:ROS

[HH08] Ewa Huebner and Frans Henskens. The role of operating systems in computer forensics. Operating Systems Review, 42(3):1–3, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hartig:1997:PKB

[HHLS97] Hermann Härtig, Michael Hohmuth, Jochen Liedtke, and Sebastian Schönberg. The performance of μ-kernel-based systems.

Operating Systems Review, 31(5):66–77, December 1997. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2005:FDD

[HHS05] Hai Huang, Wanda Hung, and Kang G. Shin. FS2: dynamic data replication in free disk space for improving disk performance and energy consumption. Operating Systems Review, 39 (5):263–276, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hillsberg:1981:GTS

[Hil81] Bruce Light Hillsberg. Generic terminal support. Operating Systems Review, 15(2):10–15, April 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hilzer:1992:SPC

[Hil92] Ralph C. Hilzer, Jr. Synchronization of the producer/consumer problem using semaphores, monitors, and the Ada rendezvous. Operating Systems Review, 26(3):31–39, July 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hills:1993:SI

[Hil93] Ted Hills. Structured interrupts. Operating Systems Review, 27 (1):51–68, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hills:1994:RNS

[Hil94] Ted Hills. Response to a note on structured interrupts. Operating Systems Review, 28(4):31–33, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hansen:2010:SVM

[HJ10] Jacob Gorm Hansen and Eric Jul. Scalable virtual machine storage using local disks. Operating Systems Review, 44(4):71– 79, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Han:2016:IMD

[HJrCH16] Jaeung Han, Seungheun Jeon, Young ri Choi, and Jaehyuk Huh. Interference management for distributed parallel applications in consolidated clusters. Operating Systems Review, 50(2):443–456, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hauser:1993:UTI

[HJT^{+}93] Carl Hauser, Christian Jacobi, Marvin Theimer, Brent Welch, and Mark Weiser. Using threads in interactive systems: a case study. Operating Systems Review, 27(5):94–105, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Halfmann:1999:ESP

[HK99] Udo Halfmann and Winfried E. Kühnhauser. Embedding security policies into a distributed computing environment. Operating Systems Review, 33(2):51–64, April 1999. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Howell:2000:RDS

[HK00] Jon Howell and David Kotz. Restricted delegation: seamlessly spanning administrative boundaries. Operating Systems Review, 34(2):38–39, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haeberlen:2007:PPA

[HKD07] Andreas Haeberlen, Petr Kouznetsov, and Peter Druschel. Peer-Review: practical accountability for distributed systems. Operating Systems Review, 41(6):175–188, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Handurukande:2006:PSB

[HKL^{+}06] S. B. Handurukande, A.-M. Kermarrec, F. Le Fessant, L. Massoulié, and S. Patarin. Peer sharing behaviour in the eDonkey network, and implications for the design of server-less file sharing systems. Operating Systems Review, 40(4):359–371, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Howard:1987:SPD

[HKM^{+}87] J. Howard, M. Kazar, S. Menees, D. Nichols, M. Satyanarayanan, Robert N. Sidebotham, and M. West. Scale and performance in a distributed file system. Operating Systems Review, 21(5):1–2, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heinrich:1994:PIF

[HKO^{+}94] Mark Heinrich, Jeffrey Kuskin, David Ofelt, John Heinlein, Joel Baxter, Jaswinder Pal Singh, Richard Simoni, Kourosh Gharachorloo, David Nakahira, Mark Horowitz, Anoop Gupta, Mendel Rosenblum, and John Hennessy. The performance impact of flexibility in the Stanford FLASH multiprocessor. Operating Systems Review, 28(5):274–285, December 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heiser:2016:RAP

[HKPvR16] Gernot Heiser, Kenji Kono, KyoungSoo Park, and Robbert van Renesse. Report on the Asia–Pacific Systems Workshop 2015 (APSys’15). Operating Systems Review, 50(1):1–2, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hopper:1979:AMM

[HKU79] K. Hopper, H. J. Kugler, and C. Unger. Abstract machines modelling network control systems. Operating Systems Review, 13(1):10–24, January 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hong:1992:MPG

[HL92] Zhao Hong and Huatian Li. A mechanism of process group for application reliability in distributed systems. Operating Systems Review, 26(1):66–77, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hua:1996:DCM

[HL96] Ji Hua and Xie Li. A distributed computing model based on multiserver. Operating Systems Review, 30(4):3–11, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hwang:2005:TAU

[HL05] Kuo-Feng Hwang and I-En Liao. Two attacks on a user friendly remote authentication scheme with Smart Cards. Operating Systems Review, 39(2):94–96, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hunt:2007:SRS

[HL07] Galen C. Hunt and James R. Larus. Singularity: rethinking the software stack. Operating Systems Review, 41(2):37–49, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hou:1997:MBR

[HLFZ97] Jianmin Hou, Xuandong Li, Xiaocong Fan, and Guoliang Zheng. A message-based real-time model by object-oriented technique. Operating Systems Review, 31(3):45–51, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Han:2002:DMA

[HLL^{+}02] Hong Han, Xian Liang Lu, Jun Lu, Chen Bo, and Ren Li Yong. Data mining aided signature discovery in network-based intrusion detection system. Operating Systems Review, 36(4):7–13, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hwang:2004:KAS

[HLL04] Min-Shiang Hwang, Li-Hua Li, and Cheng-Chi Lee. A key authentication scheme with non-repudiation. Operating Systems Review, 38(3):75–78, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haring:1998:IWP

[HLR98] Gunter Haring, Christoph Lindemann, and Martin Reiser. International workshop performance evaluation. Operating Systems Review, 32(2):2–3, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hong:1990:AOM

[HM90] Zhao Hong and Wayne McCoy. An associated object model for distributed systems. Operating Systems Review, 24(4):34– 51, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hayter:1991:DAN

[HM91] Mark Hayter and Derek McAuley. The desk area network. Operating Systems Review, 25(4):14–21, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hosking:1993:PTA

[HM93] Antony L. Hosking and J. Eliot B. Moss. Protection traps and alternatives for memory management of an object-oriented language. Operating Systems Review, 27(5):106–119, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heiser:2020:TPT

[HMK20] Gernot Heiser, Toby Murray, and Gerwin Klein. Towards provable timing-channel prevention. Operating Systems Review, 54 (1):1–7, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.org/

doi/10.1145/3421473.3421475.

Howard:2017:RPF

[HMS17] Heidi Howard, Dahlia Malkhi, and Sasha Spiegelman. Revisiting the Paxos Foundations: a look at summer internship work at VMware Research. Operating Systems Review, 51(1):67–71, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haskin:1987:RMQ

[HMSC87] R. Haskin, Y. Malachi, W. Sawdon, and G. Chan. Recovery management in QuickSilver. Operating Systems Review, 21(5): 107–108, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herbert:1981:SCS

[HN81] A. J. Herbert and R. M. Needham. Sequencing computation steps in a network. Operating Systems Review, 15(5):59–63, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hay:2008:FEV

[HN08] Brian Hay and Kara Nance. Forensics examination of volatile system data using virtual introspection. Operating Systems Review, 42(3):74–82, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hayden:2012:RTW

[HN12] Christopher M. Hayden and Iulian Neamtiu. Report on the Third Workshop on Hot Topics in Software Upgrades (HotSWUp’11). Operating Systems Review, 46(1):93–99, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hsiao:2017:ASI

[HNK^{+}17] Chun-Hung Hsiao, Satish Narayanasamy, Essam Muhammad Idris Khan, Cristiano L. Pereira, and Gilles A. Pokam. AsyncClock: Scalable inference of asynchronous event causality. Operating Systems Review, 51(2):193–205, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hall:1991:PCA

[HO91] C. Brian Hall and Kevin O’Brien. Performance characteristics of architectural features of the IBM RISC System/6000. Operating Systems Review, 25(3S):303–309, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hartman:1993:ZSN

[HO93] John H. Hartman and John K. Ousterhout. The Zebra striped network file system. Operating Systems Review, 27(5):29–43, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hofri:1990:PME

[Hof90] Micha Hofri. Proof of a mutual exclusion algorithm—a classic example. Operating Systems Review, 24(1):18–22, January 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hofri:2007:STC

[Hof07] Micha Hofri. Service transparency considered harmful: letter to the editor. Operating Systems Review, 41(4):77, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hogan:1988:PIS

[Hog88] Carole B. Hogan. Protection imperfect: the security of some computing environments. Operating Systems Review, 22(3):7– 27, July 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See note [Wel88].

Hohmuth:2007:I

[Hoh07] Michael Hohmuth. Introduction. Operating Systems Review, 41 (4):1–2, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Holt:1972:SDP

[Hol72] Richard C. Holt. Some deadlock properties of computer systems. Operating Systems Review, 6(1/2):64–71, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Holt:1982:TUL

[Hol82] R. C. Holt. Tunis: a Unix look-alike written in concurrent Euclid (abstract). Operating Systems Review, 16(1):4–5, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Holt:1988:DMT

[Hol88] R. C. Holt. Device management in TURNING PLUS. Operating Systems Review, 22(1):33–41, January 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hopper:1990:PES

[Hop90] Andy Hopper. Pandora — an experimental system for multimedia applications. Operating Systems Review, 24(2):19–34, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Howry:1972:MSP

[How72] Sam Howry. A multiprogramming system for process control. Operating Systems Review, 6(1/2):24–30, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Howard:1982:RPR

[How82] John H. Howard. Reply to “on proof rules for monitors”. Operating Systems Review, 16(4):8–9, October 1982. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [AB82].

Haro:1993:MEO

[HP93] Christophe Haro and Christian Proust. A multitasking executive for operating systems courses. Operating Systems Review, 27(3):

97–107, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heidemann:1995:PCC

[HP95] J. Heidemann and G. Popek. Performance of cache coherence in stackable filing. Operating Systems Review, 29(5):127–141, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Halvorsen:2000:IPO

[HPG00] Pål Halvorsen, Thomas Plagemann, and Vera Goebel. The IN-STANCE project: operating system enhancements to support multimedia servers. Operating Systems Review, 34(2):36, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hamilton:1993:SFB

[HPM93] Graham Hamilton, Michael L. Powell, and James G. Mitchell. Subcontract: a flexible base for distributed programming. Operating Systems Review, 27(5):69–79, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hillyer:1992:BFM

[HR92] Bruce K. Hillyer and Bethany S. Robinson. The BBFS Filesystem Model. Operating Systems Review, 26(2):18, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harrison:1975:POS

[HRU75] Michael A. Harrison, Walter L. Ruzzo, and Jeffrey D. Ullman. On protection in operating systems. Operating Systems Review, 9(5):14–24, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ha:2008:CNT

[HRX08] Sangtae Ha, Injong Rhee, and Lisong Xu. CUBIC: a new TCPfriendly high-speed TCP variant. Operating Systems Review, 42(5):64–74, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haldar:1988:ESM

[HS88] S. Haldar and D. Subramanian. An efficient solution to the mutual exclusion problem using unfair and weak semaphore.

Operating Systems Review, 22(2):60–66, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Haldar:1991:FPS

[HS91] S. Haldar and D. K. Subramanian. Fairness in processor scheduling in time sharing systems. Operating Systems Review, 25(1): 4–18, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:1996:IBR

[HS96] Andrew S. Huang and John Paul Shen. The intrinsic bandwidth requirements of ordinary programs. Operating Systems Review, 30(5):105–114, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hasabnis:2016:LAI

[HS16] Niranjan Hasabnis and R. Sekar. Lifting assembly to intermediate representation: a novel approach leveraging compilers. Operating Systems Review, 50(2):311–324, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hsieh:1989:FCI

[Hsi89] C. Samuel Hsieh. Further comments on implementation of general semaphores. Operating Systems Review, 23(1):9–10, January 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comment [Kea88, Hem89, TT00].

Heidemann:2001:BEW

[HSI^{+}01] John Heidemann, Fabio Silva, Chalermek Intanagonwiwat, Ramesh Govindan, Deborah Estrin, and Deepak Ganesan. Building efficient wireless sensor networks with low-level naming. Operating Systems Review, 35(5):146–159, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Helme:1997:SFF

[HSK97] Arne Helme and Tage Stabell-Kulø. Security functions for a file repository. Operating Systems Review, 31(2):3–8, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hu:2017:TFC

[HSL17] Yang Hu, Mingcong Song, and Tao Li. Towards “full containerization” in containerized network function virtualization. Operating Systems Review, 51(2):467–481, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Howard:2015:RRD

[HSMC15] Heidi Howard, Malte Schwarzkopf, Anil Madhavapeddy, and Jon Crowcroft. Raft refloated: Do we have consensus? Operating Systems Review, 49(1):12–21, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2001:VBC

[HSPC01] Z. Huang, C. Sun, M. Purvis, and S. Cranefield. View-based consistency and false sharing effect in distributed shared memory. Operating Systems Review, 35(2):51–60, April 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hoke:2006:ICM

[HSS^{+}06] Evan Hoke, Jimeng Sun, John D. Strunk, Gregory R. Ganger, and Christos Faloutsos. InteMon: continuous mining of sensor data in large-scale self-infrastructures. Operating Systems Review, 40(3):38–44, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hill:2000:SAD

[HSW^{+}00] Jason Hill, Robert Szewczyk, Alec Woo, Seth Hollar, David Culler, and Kristofer Pister. System architecture directions for networked sensors. Operating Systems Review, 34(5):93–104, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hoque:2015:SDB

[HT15] Mohammad A. Hoque and Sasu Tarkoma. Sudden drop in the battery level?: Understanding Smartphone state of charge anomaly. Operating Systems Review, 49(2):70–74, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Meyer:2008:PVD

[hTMAC^{+}08] Dut h T. Meyer, Gitika Aggarwal, Brendan Cully, Geoffrey Lefebvre, Michael J. Feeley, Norman C. Hutchinson, and Andrew Warfield. Parallax: virtual disks for virtual machines. Operating Systems Review, 42(4):41–54, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herlihy:2001:OMD

[HTW01] Maurice Herlihy, Srikanta Tirthapura, and Roger Wattenhofer. Ordered Multicast and Distributed Swap. Operating Systems Review, 35(1):85–96, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Heiser:2006:VMM

[HUL06] Gernot Heiser, Volkmar Uhlig, and Joshua LeVasseur. Are virtual-machine monitors microkernels done right? Operating Systems Review, 40(1):95–99, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Halder:1992:COW

[HV92] S. Halder and K. Vidyasankar. Counterexamples to a one writer multireader atomic variable construction of Burns and Peterson. Operating Systems Review, 26(1):78–87, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hamberg:2008:UMC

[HV08] Roelof Hamberg and Frits Vaandrager. Using model checkers in an introductory course on operating systems. Operating Systems Review, 42(6):101–111, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hawblitzel:2002:LFJ

[HvE02] Chris Hawblitzel and Thorsten von Eicken. Luna: a flexible Java protection system. Operating Systems Review, 36(5S):391–403, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hammond:1998:DSS

[HWO98] Lance Hammond, Mark Willey, and Kunle Olukotun. Data speculation support for a chip multiprocessor. Operating Systems Review, 32(5):58–69, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2001:HRD

[HXL01] Tao Huang, Teng Xu, and Xianliang Lu. A high resolution disk I/O trace system. Operating Systems Review, 35(4):82– 87, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hu:2010:GBU

[HYM10] Wenjin Hu, Tao Yang, and Jeanna N. Matthews. The good, the bad and the ugly of consumer cloud storage. Operating Systems Review, 44(3):110–115, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hwang:2003:ASM

[HYS03] Min-Shiang Hwang, Chao-Chen Yang, and Cheng-Yeh Shiu. An authentication scheme for mobile satellite communication systems. Operating Systems Review, 37(4):42–47, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hazelwood:2009:COA

[HZ09] Kim Hazelwood and Mohamed Zahran. Challenges and opportunities at all levels: interactions among operating systems, compilers, and multicore processors. Operating Systems Review, 43 (2):3–4, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

He:1997:SSD

[HZCC97] Yanxiang He, Jianping Zhang, Donald H. Cooley, and Li Chen. Semantics subsystem in distributed OS formalization generating system. Operating Systems Review, 31(4):85–92, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Isard:2007:DDD

[IBY^{+}07] Michael Isard, Mihai Budiu, Yuan Yu, Andrew Birrell, and Dennis Fetterly. Dryad: distributed data-parallel programs from sequential building blocks. Operating Systems Review, 41(3): 59–72, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Iyer:2001:ASD

[ID01] Sitaram Iyer and Peter Druschel. Anticipatory scheduling: a disk scheduling framework to overcome deceptive idleness in synchronous I/O. Operating Systems Review, 35(5):117–130, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ibrahim:2025:ECA

[ID25] Shadi Ibrahim and Jad Darrous. Erasure coding aware block placement for data-intensive applications. Operating Systems Review, 59(2):62–69, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Izraelevitz:2016:FAP

[IKK16] Joseph Izraelevitz, Terence Kelly, and Aasheesh Kolli. Failureatomic persistent memory updates via JUSTDO logging. Operating Systems Review, 50(2):427–442, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Inouye:1992:EVA

[IKWS92] Jon Inouye, Ravindranath Konuru, Jonathan Walpole, and Bart Sears. The effects of virtually addressed caches on virtual memory design and performance. Operating Systems Review, 26 (4):14–29, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ipek:2006:EEA

[ÏMC^{+}06] Engin Ïpek, Sally A. McKee, Rich Caruana, Bronis R. de Supinski, and Martin Schulz. Efficiently exploring architectural design spaces via predictive modeling. Operating Systems Review, 40 (5):195–206, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Isard:2007:AAD

[Isa07] Michael Isard. Autopilot: automatic data center management. Operating Systems Review, 41(2):60–67, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Isaacs:2008:RSS

[Isa08] Rebecca Isaacs. Report on the 2007 SOSP shadow program committee. Operating Systems Review, 42(3):127–131, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Iskra:2000:IDE

[IvdLH^{+}00] K. A. Iskra, F. van der Linden, Z. W. Hendrikse, B. J. Overeinder, G. D. van Albada, and P. M. A. Sloot. The implementation of dynamite: an environment for migrating PVM tasks. Operating Systems Review, 34(3):40–55, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:2006:GMB

[JADAD06] Stephen T. Jones, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Geiger: monitoring the buffer cache in a virtual machine environment. Operating Systems Review, 40(5):14–24, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Janson:1975:DLE

[Jan75] Philippe A. Janson. Dynamic linking and environment initialization in a multi-domain process. Operating Systems Review, 9 (5):43–50, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Janson:1981:UTE

[Jan81] P. A. Janson. Using type-extension to organize virtual-memory mechanisms. Operating Systems Review, 15(4):6–38, October 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Johansen:2006:FSS

[JAvR06] Håvard Johansen, André Allavena, and Robbert van Renesse. Fireflies: Scalable support for intrusion-tolerant network overlays. Operating Systems Review, 40(4):3–13, October 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jann:2008:EEP

[JBDP08] Joefon Jann, R. Sarma Burugula, Niteesh Dubey, and Pratap Pattnaik. End-to-end performance of commercial applications in the face of changing hardware. Operating Systems Review, 42 (1):13–20, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jefferson:1987:DST

[JBW^{+}87] D. Jefferson, B. Beckman, F. Wieland, L. Blume, and M. Diloreto. Distributed simulation and the Time Warp Operating System. Operating Systems Review, 21(5):77–93, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jeong:2019:SSG

[JCY^{+}19] Eunji Jeong, Sungwoo Cho, Gyeong-In Yu, Joo Seong Jeong, Dong-Jin Shin, Taebum Kim, and Byung-Gon Chun. Speculative symbolic graph execution of imperative deep learning programs. Operating Systems Review, 53(1):26–33, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joseph:1995:RTM

[JdLT^{+}95] A. D. Joseph, A. F. de Lespinasse, J. A. Tauber, D. K. Gifford, and M. F. Kaashoek. Rover: a toolkit for mobile information access. Operating Systems Review, 29(5):156–171, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jeffay:1992:NOS

[Jef92] Kevin Jeffay. Network and Operating System Support for Digital Audio and Video. Operating Systems Review, 26(2):16, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jamrozik:1996:RNL

[JFV^{+}96] Hervé A. Jamrozik, Michael J. Feeley, Geoffrey M. Voelker, James Evans II, Anna R. Karlin, Henry M. Levy, and Mary K. Vernon. Reducing network latency using subpages in a global memory environment. Operating Systems Review, 30(5):258– 267, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1993:HAV

[JH93] Alan Jones and Andrew Hopper. Handling audio and video streams in a distributed environment. Operating Systems Review, 27(5):231–243, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Javed:2011:PHN

[JHC^{+}11] Umar Javed, Dongsu Han, Ramon Caceres, Jeffrey Pang, Srinivasan Seshan, and Alexander Varshavsky. Predicting handoffs in 3G networks. Operating Systems Review, 45(3):65–70, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jeon:2016:TTD

[JHK^{+}16] Myeongjae Jeon, Yuxiong He, Hwanju Kim, Sameh Elnikety, Scott Rixner, and Alan L. Cox. TPC: Target-driven parallelism combining prediction and correction to reduce tail latency in interactive services. Operating Systems Review, 50(2):129–141, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jambor:2007:ILL

[JHT^{+}07] Martin Jambor, Tomas Hruby, Jan Taus, Kuba Krchak, and Viliam Holub. Implementation of a Linux log-structured file system with a garbage collector. Operating Systems Review, 41 (1):24–32, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jin:1999:FRF

[Jin99] Hai Jin. On-the-fly reconstruction of the failed disk in RAID. Operating Systems Review, 33(3):32–42, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joshi:2005:DPP

[JKDC05] Ashlesha Joshi, Samuel T. King, George W. Dunlap, and Peter M. Chen. Detecting past and present intrusions through vulnerability-specific predicates. Operating Systems Review, 39 (5):91–104, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jul:2000:HPC

[JKH^{+}00] Eric Jul, Povl Koch, Jørgen S. Hansen, Michael Svendsen, Kim Henriksen, Kenn Nielsen, and Mads Dydensborg. Highperformance cluster-based Internet servers. Operating Systems Review, 34(2):38, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jiekak:2013:RCS

[JKL^{+}13] Steve Jiekak, Anne-Marie Kermarrec, Nicolas Le Scouarnec, Gilles Straub, and Alexandre Van Kempen. Regenerating codes: a system perspective. Operating Systems Review, 47(2):23–32, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Junqueira:2013:DB

[JKR13] Flavio P. Junqueira, Ivan Kelly, and Benjamin Reed. Durability with BookKeeper. Operating Systems Review, 47(1):9–15, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jacobsen:2015:LCD

[JKS^{+}15] Charles Jacobsen, Muktesh Khole, Sarah Spall, Scotty Bauer, and Anton Burtsev. Lightweight capability domains: Towards decomposing the Linux kernel. Operating Systems Review, 49 (2):44–50, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Johnson:1995:CHP

[JKW95] K. L. Johnson, M. F. Kaashoek, and D. A. Wallach. CRL: highperformance all-software distributed shared memory. Operating Systems Review, 29(5):213–226, December 1995. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1975:ESP

[JL75] Anita K. Jones and Richard J. Lipton. The enforcement of security policies for computation. Operating Systems Review, 9 (5):197–206, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jul:1987:FGM

[JLHB87] E. Jul, H. Levy, N. Hutchinson, and A. Black. Fine-grained mobility in the emerald system. Operating Systems Review, 21 (5):105–106, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:2005:AMS

[JLR^{+}05] Cliff Jones, David Lomet, Alexander Romanovsky, Gerhard Weikum, Alan Fekete, Marie-Claude Gaudel, Henry F. Korth, Rogerio de Lemos, Eliot Moss, Ravi Rajwar, Krithi Ramamritham, Brian Randell, and Luis Rodrigues. The atomic manifesto: a story in four quarks. Operating Systems Review, 39(2): 41–46, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jun:1990:MIT

[JLZx90] Chen Jun, Xie Li, and Sun Zhong-xiu. A model for intelligent task scheduling in a large distributed system. Operating Systems Review, 24(4):26–33, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jia:1995:OSK

[JM95] Xiaohua Jia and Mamoru Maekawa. Operating system kernel automatic construction. Operating Systems Review, 29(3):91– 96, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jacob:1998:LSM

[JM98] Bruce L. Jacob and Trevor N. Mudge. A look at several memory management units, TLB-refill mechanisms, and page table organizations. Operating Systems Review, 32(5):295–306, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joao:2008:IPO

[JMK^{+}08] Jose A. Joao, Onur Mutlu, Hyesoon Kim, Rishi Agarwal, and Yale N. Patt. Improving the performance of object-oriented languages with dynamic predication of indirect jumps. Operating Systems Review, 42(2):80–90, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Johnson:1991:CRB

[Joh91] Douglas Johnson. The case for a read barrier. Operating Systems Review, 25(3S):279–287, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1980:CAR

[Jon80] Anita K. Jones. Capability architecture revisited. Operating Systems Review, 14(3):33–35, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1992:TIU

[Jon92] Michael B. Jones. A Toolkit for Interposing User Code at the System Interface. Operating Systems Review, 26(2):21, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1993:IAT

[Jon93] Michael B. Jones. Interposition agents: transparently interposing user code at the system interface. Operating Systems Review, 27(5):80–93, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Juang:2002:EEC

[JOW^{+}02] Philo Juang, Hidekazu Oki, Yong Wang, Margaret Martonosi, Li Shiuan Peh, and Daniel Rubenstein. Energy-efficient computing for wildlife tracking: design tradeoffs and early experiences with ZebraNet. Operating Systems Review, 36(5):96–107, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joseph:1978:MNM

[JP78] M. Joseph and V. R. Prasad. More on nested monitor calls. Operating Systems Review, 12(2):21–25, April 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joshi:2005:MVO

[JR05] Rushikesh K. Joshi and Subash Rajaa. Modeling VP operation: the Diwali Festival Problem. Operating Systems Review, 39(2): 51–53, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:1997:CRT

[JRR97] Michael B. Jones, Daniela Roşu, and Marcel-Cãtãlin Roşu. CPU reservations and time constraints: efficient, predictable scheduling of independent activities. Operating Systems Review, 31(5): 198–211, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joukov:2008:GME

[JS08] Nikolai Joukov and Josef Sipek. GreenFS: making enterprise computers greener by protecting them better. Operating Systems Review, 42(4):69–80, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jevdjic:2017:ASC

[JSCM17] Djordje Jevdjic, Karin Strauss, Luis Ceze, and Henrique S. Malvar. Approximate storage of compressed and encrypted videos.

Operating Systems Review, 51(2):361–373, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jain:2008:ALI

[JSDG08] Shvetank Jain, Fareha Shafique, Vladan Djeric, and Ashvin Goel. Application-level isolation and recovery with solitude. Operating Systems Review, 42(4):95–107, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jia:2015:SOC

[JSS^{+}15] Qin Jia, Zhiming Shen, Weijia Song, Robbert van Renesse, and Hakim Weatherspoon. Supercloud: Opportunities and challenges. Operating Systems Review, 49(1):137–141, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jalote:1990:WIA

[JT90] Pankaj Jalote and Satish K. Tripathi. Workshop on integrated approach for fault tolerance-current state and future requirements. Operating Systems Review, 24(1):40–57, January 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ji:2000:SOI

[JTG^{+}00] Li Ji, Li Tianning, Chen Guihai, Xie Li, and C. L. Wang. Strategies optimization and integration in DSM. Operating Systems Review, 34(3):29–39, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jamshidi:2021:DDP

[JV21] Kasra Jamshidi and Keval Vora. A deeper dive into patternaware subgraph exploration with PEREGRINE. Operating Systems Review, 55(1):1–10, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3469379.3469381.

Ju:1996:SPT

[JW96] Jiubin Ju and Yong Wang. Scheduling PVM tasks. Operating Systems Review, 30(3):22–31, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ji:2001:CAF

[JW01] Dongyao Ji and Yuming Wang. Comments on “An approach to the formal verification of the two-party cryptographic protocols” by Zhang, Li and Xiao. Operating Systems Review, 35(1):6– 7, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [ZLX99].

Jia:2024:KIB

[JW24] Zhipeng Jia and Emmett Witchel. The key ideas behind Boki’s shared logs. Operating Systems Review, 58(1):7–14, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3689051.
3689054.

Jun:2002:SNF

[JXG^{+}02] Lu Jun, Lu Xianliang, Luo Guangchun, Han Hong, and Zhou Xu. STFS: a novel file system for efficient small writes. Operating Systems Review, 36(4):50–54, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jiang:2021:VES

[JXG21] Xiaolin Jiang, Chengshuo Xu, and Rajiv Gupta. VRGQ: Evaluating a stream of iterative graph queries via value reuse. Operating Systems Review, 55(1):11–20, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3469379.3469382.

Jun:2002:CAW

[JXHQ02] Lu Jun, Lu Xianliang, Han Hong, and Wei Qingsong. A cooperative asynchronous write mechanism for NAS. Operating Systems Review, 36(3):25–32, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jiang:2022:MMF

[JXQ^{+}22] Yuting Jiang, Yifan Xiong, Lei Qu, Cheng Luo Luo, Chen Tian, Peng Cheng, and Yongqiang Xiong. Moneo: Monitoring finegrained metrics nonintrusively in AI infrastructure. Operating Systems Review, 56(1):18–25, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3544497.3544501.

Ju:1993:PCU

[JXT93] Jiubin Ju, Gaochao Xu, and Jie Tao. Parallel computing using idle workstations. Operating Systems Review, 27(3):87–96, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ju:1995:IDL

[JXY95] Jiubin Ju, Gaochao Xu, and Kun Yang. An intelligent dynamic load balancer for workstation clusters. Operating Systems Review, 29(1):7–16, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joye:1998:IBS

[JY98] Marc Joye and Sung-Ming Yen. ID-based secret-key cryptography. Operating Systems Review, 32(4):33–39, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Johnson:1991:TOR

[JZ91] David B. Johnson and Willy Zwaenepoel. Transparent optimistic rollback recovery. Operating Systems Review, 25(2):99– 102, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jin:2002:DPO

[JZZW02] Chao Jin, Weimin Zheng, Feng Zhou, and Yinghui Wu. A distributed persistent object store for scalable service. Operating Systems Review, 36(4):36–49, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:BRLb

[Kad95a] Basim Kadhim. Book review: Linux Universe, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(4):3, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:BRLa

[Kad95b] Basim Kadhim. Book review: Linux: Unleashing the Workstation in Your PC, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(2):2–3, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotla:2007:ZSB

[KAD^{+}07] Ramakrishna Kotla, Lorenzo Alvisi, Mike Dahlin, Allen Clement, and Edmund Wong. Zyzzyva: speculative Byzantine fault tolerance. Operating Systems Review, 41(6):45–58, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kahn:1972:ASC

[Kah72] Gilles Kahn. An approach to systems correctness. Operating Systems Review, 6(1/2):86–94, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kahn:1985:FRS

[Kah85] Kevin C. Kahn. Financial report to the SIGOPS membership. Operating Systems Review, 19(2):2, April 1985. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kain:1975:HEP

[Kai75] Richard Y. Kain. How to evaluate page replacement algorithms. Operating Systems Review, 9(5):1–5, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Koyano:2013:SML

[KAI^{+}13] Sou Koyano, Shingo Ata, Hisashi Iwamoto, Yuji Yano, Yasuto Kuroda, Kazunari Inoue, and Ikuo Oka. A study on micro level traffic prediction for energy-aware routers. Operating Systems Review, 47(3):26–33, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kamga:2013:CFE

[Kam13] Christine Mayap Kamga. CPU frequency emulation based on DVFS. Operating Systems Review, 47(3):34–41, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kant:1983:ELC

[Kan83] Krishna Kant. Efficient local checkpointing for software fault tolerance. Operating Systems Review, 17(2):11–13, April 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krieger:2006:KBC

[KAR^{+}06] Orran Krieger, Marc Auslander, Bryan Rosenburg, Robert W. Wisniewski, Jimi Xenidis, Dilma Da Silva, Michal Ostrowski, Jonathan Appavoo, Maria Butrico, Mark Mergen, Amos Waterland, and Volkmar Uhlig. K42: building a complete operating system. Operating Systems Review, 40(4):133–145, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Konishi:2006:LIL

[KAS^{+}06] Ryusuke Konishi, Yoshiji Amagai, Koji Sato, Hisashi Hifumi, Seiji Kihara, and Satoshi Moriai. The Linux implementation of a log-structured file system. Operating Systems Review, 40(3): 102–107, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kenah:1984:VVI

[KB84] Lawrence J. Kenah and Simon F. Bate. VAX/VMS internals and data structures. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1984. ISBN 0-932376-52-5 (paperback). xix + 795 pp. LCCN QA76.76.O63 K46 1984.

Keeton:2006:RRR

[KBB^{+}06] Kimberly Keeton, Dirk Beyer, Ernesto Brau, Arif Merchant, Cipriano Santos, and Alex Zhang. On the road to recovery: restoring data after disasters. Operating Systems Review, 40(4): 235–248, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kulkarni:1994:OSC

[KBC94] Dinesh C. Kulkarni, Arindam Banerji, and David L. Cohn. Operating systems and cost management. Operating Systems Review, 28(1):5–10, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kubiatowicz:2000:OAG

[KBC^{+}00] John Kubiatowicz, David Bindel, Yan Chen, Steven Czerwinski, Patrick Eaton, Dennis Geels, Ramakrishna Gummadi, Sean Rhea, Hakim Weatherspoon, Chris Wells, and Ben Zhao. OceanStore: an architecture for global-scale persistent storage. Operating Systems Review, 34(5):190–201, December 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2002:ANU

[KBK02] Changkyu Kim, Doug Burger, and Stephen W. Keckler. An adaptive, non-uniform cache structure for wire-delay dominated on-chip caches. Operating Systems Review, 36(5):211–222, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadav:2010:DRR

[KBPM10] Asim Kadav, Mahesh Balakrishnan, Vijayan Prabhakaran, and Dahlia Malkhi. Differential RAID: rethinking RAID for SSD reliability. Operating Systems Review, 44(1):55–59, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karamcheti:1994:SOM

[KC94] Vijay Karamcheti and Andrew A. Chien. Software overhead in messaging layers: where does the time go? Operating Systems Review, 28(5):51–60, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kao:1995:ESA

[KC95] I.-Lung Kao and Randy Chow. An efficient and secure authentication protocol using uncertified keys. Operating Systems Review, 29(3):14–21, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ku:2005:CFR

[KC05] Wei-Chi Ku and Shuai-Min Chen. Cryptanalysis of a flexible remote user authentication scheme using Smart Cards. Operating Systems Review, 39(1):90–96, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ku:2005:WYR

[KCC05] Wei-Chi Ku, Min-Hung Chiang, and Shen-Tien Chang. Weaknesses of Yoon–Ryu–Yoo’s hash-based password authentication scheme. Operating Systems Review, 39(1):85–89, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kahn:1981:IMO

[KCD^{+}81] Kevin C. Kahn, William M. Corwin, T. Don Dennis, Herman D’Hooge, David E. Hubka, Linda A. Hutchins, John T. Montague, and Fred J. Pollack. iMAX: A multiprocessor operating system for an object-based computer. Operating Systems Review, 15(5):127–136, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ku:2003:WLL

[KCL03] Wei-Chi Ku, Chien-Ming Chen, and Hui-Lung Lee. Weaknesses of Lee–Li–Hwang’s hash-based password authentication scheme. Operating Systems Review, 37(4):19–25, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krintz:1998:OET

[KCLZ98] Chandra Krintz, Brad Calder, Han Bok Lee, and Benjamin G. Zorn. Overlapping execution with transfer using non-strict execution for mobile programs. Operating Systems Review, 32(5): 159–169, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kwon:2016:SPT

[KDL^{+}16] Youngjin Kwon, Alan M. Dunn, Michael Z. Lee, Owen S. Hofmann, Yuanzhong Xu, and Emmett Witchel. Sego: Pervasive trusted metadata for efficiently verified untrusted system services. Operating Systems Review, 50(2):277–290, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kirovski:2002:ETS

[KDP02] Darko Kirovski, Milenko Drinić, and Miodrag Potkonjak. Enabling trusted software integrity. Operating Systems Review, 36 (5):108–120, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kgil:2006:PUS

[KDS^{+}06] Taeho Kgil, Shaun D’Souza, Ali Saidi, Nathan Binkert, Ronald Dreslinski, Trevor Mudge, Steven Reinhardt, and Krisztian Flautner. PicoServer: using 3D stacking technology to enable a compact energy efficient chip multiprocessor. Operating Systems Review, 40(5):117–128, December 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kleiman:1995:IT

[KE95] Steve Kleiman and Joe Eykholt. Interrupts as threads. Operating Systems Review, 29(2):21–26, April 1995. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kearns:1988:CUI

[Kea88] Phil Kearns. A correct and unrestrictive implementation of general semaphores. Operating Systems Review, 22(4):46–48, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comments [Hem89, Hsi89, TT00].

Keedy:1979:SOS

[Kee79] J. L. Keedy. On structuring operating systems with monitors. Operating Systems Review, 13(1):5–9, January 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krishnan:2019:AIR

[KEF^{+}19] Sanjay Krishnan, Aaron J. Elmore, Michael Franklin, John Paparrizos, Zechao Shang, Adam Dziedzic, and Rui Liu. Artificial intelligence in resource-constrained and shared environments. Operating Systems Review, 53(1):1–6, July 2019. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaashoek:1997:APF

[KEG^{+}97] M. Frans Kaashoek, Dawson R. Engler, Gregory R. Ganger, Hector M. Briceño, Russell Hunt, David Mazières, Thomas Pinckney, Robert Grimm, John Jannotti, and Kenneth Mackenzie. Application performance and flexibility on exokernel systems. Operating Systems Review, 31(5):52–65, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Keppel:1991:PIF

[Kep91] David Keppel. A portable interface for on-the-fly instruction space modification. Operating Systems Review, 25(3S):86–95, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2007:LPM

[KEP07] Dohun Kim, Jugwan Eom, and Chanik Park. L4oprof: a performance-monitoring-unit-based software-profiling framework for the L4 microkernel. Operating Systems Review, 41(4):

69–76, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kumar:2009:TBP

[KF09] Viren Kumar and Alexandra Fedorova. Towards better performance per watt in virtual environments on asymmetric single-ISA multi-core systems. Operating Systems Review, 43(3):105– 109, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotz:1999:MAF

[KG99] David Kotz and Robert S. Gray. Mobile agents and the future of the Internet. Operating Systems Review, 33(3):7–13, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kenah:1988:VVI

[KGB88] Lawrence J. Kenah, Ruth E. Goldenberg, and Simon F. Bate. VAX/VMS internals and data structures: version 4.4. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1988. ISBN 1-55558-008-4 (paperback). xvii + 979 pp. LCCN QA76.76.O63 K47 1988.

Kunal:2009:HDS

[KGGK09] K. Kunal, K. George, M. Gautam, and V. Kamakoti. HTM design spaces: complete decoupling from caches and achieving highly concurrent transactions. Operating Systems Review, 43(2):98–99, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kannan:2018:HDH

[KGGS18] Sudarsun Kannan, Ada Gavrilovska, Vishal Gupta, and Karsten Schwan. HeteroOS: OS design for heterogeneous memory management in datacenters. Operating Systems Review, 52(1):13–26, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krishna:2006:CSM

[KGS06] Arvind S. Krishna, Aniruddha S. Gokhale, and Douglas C. Schmidt. Context-specific middleware specialization techniques for optimizing software product-line architectures. Operating Systems Review, 40(4):205–218, October 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kang:2017:NCI

[KHG^{+}17] Yiping Kang, Johann Hauswald, Cao Gao, Austin Rovinski, Trevor Mudge, Jason Mars, and Lingjia Tang. Neurosurgeon: Collaborative intelligence between the cloud and mobile edge. Operating Systems Review, 51(2):615–629, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Klues:2007:ICC

[KHL^{+}07] Kevin Klues, Vlado Handziski, Chenyang Lu, Adam Wolisz, David Culler, David Gay, and Philip Levis. Integrating concurrency control and energy management in device drivers. Operating Systems Review, 41(6):251–264, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kieburtz:1987:RAS

[Kie87] Richard B. Kieburtz. A RISC architecture for symbolic computation. Operating Systems Review, 21(4):146–155, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kilicote:2000:PPA

[Kil00] Han Kilicote. PASIS: perpetually available and secure information systems (poster session). Operating Systems Review, 34 (2):35, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Koskinen:2008:BIE

[KJ08] Eric Koskinen and John Jannotti. BorderPatrol: isolating events for black-box tracing. Operating Systems Review, 42(4):191–203, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kavulya:2011:PEC

[KJH^{+}11] Soila P. Kavulya, Kaustubh Joshi, Matti Hiltunen, Scott Daniels, Rajeev Gandhi, and Priya Narasimhan. Practical experiences with chronics discovery in large telecommunications systems. Operating Systems Review, 45(3):23–30, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kharbutli:2006:CEP

[KJS^{+}06] Mazen Kharbutli, Xiaowei Jiang, Yan Solihin, Guru Venkataramani, and Milos Prvulovic. Comprehensively and efficiently protecting the heap. Operating Systems Review, 40(5):207–218, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kavi:1984:AQ

[KK84] Krishna M. Kavi and K. Krishnamohan. Architecture quality. Operating Systems Review, 18(1):11–19, January 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2016:NEN

[KKB^{+}16] Wook-Hee Kim, Jinwoong Kim, Woongki Baek, Beomseok Nam, and Youjip Won. NVWAL: Exploiting NVRAM in write-ahead logging. Operating Systems Review, 50(2):385–398, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kumar:2002:CSR

[KKC02] Arun Kumar, Neeran Karnik, and Girish Chafle. Context sensitivity in role-based access control. Operating Systems Review, 36(3):53–66, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krepska:2011:HPP

[KKFB11] Elzbieta Krepska, Thilo Kielmann, Wan Fokkin, and Henri Bal. HipG: parallel processing of large-scale graphs. Operating Systems Review, 45(2):3–13, July 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2017:TAA

[KKK^{+}17] Channoh Kim, Jaehyeok Kim, Sungmin Kim, Dooyoung Kim, Namho Kim, Gitae Na, Young H. Oh, Hyeon Gyu Cho, and Jae W. Lee. Typed architectures: Architectural support for lightweight scripting. Operating Systems Review, 51(2):77–90, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kawahito:2006:NIR

[KKM^{+}06] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. A new idiom recognition framework for exploiting hardware-assist instructions. Operating Systems Review, 40(5):382–393, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kawahito:2000:ENP

[KKN00] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective null pointer check elimination utilizing hardware trap. Operating Systems Review, 34(5):139–149, December 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kandlur:1989:HDR

[KKS89] D. D. Kandlur, D. L. Kiskis, and K. G. Shin. HARTOS: a distributed real-time operating system. Operating Systems Review, 23(3):72–89, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kwon:2016:LCI

[KKS^{+}16] Yonghwi Kwon, Dohyeong Kim, William Nick Sumner, Kyungtae Kim, Brendan Saltaformaggio, Xiangyu Zhang, and Dongyan Xu. LDX: Causality inference by lightweight dual execution. Operating Systems Review, 50(2):503–515, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kang:1998:ASN

[KL98] Sung-Il Kang and Heung-Kyu Lee. Analysis and solution of nonpreemptive policies for scheduling readers and writers. Operating Systems Review, 32(3):30–50, July 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comment [Küh99].

Kumar:2002:UMC

[KL02] Sanjeev Kumar and Kai Li. Using model checking to debug device firmware. Operating Systems Review, 36(5S):61–74, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kiciman:2007:APR

[KL07] Emre Kiciman and Benjamin Livshits. AjaxScope: a platform for remotely monitoring the client-side behavior of Web 2.0 applications. Operating Systems Review, 41(6):17–30, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Klimovic:2017:RRF

[KLK17] Ana Klimovic, Heiner Litz, and Christos Kozyrakis. ReFlex: Remote flash ≈ local flash. Operating Systems Review, 51(2): 345–359, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karlin:1991:ESC

[KLMO91] Anna R. Karlin, Kai Li, Mark S. Manasse, and Susan Owicki. Empirical studies of competitive spinning for a shared-memory multiprocessor. Operating Systems Review, 25(5):41–55, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Klossner:1980:PBO

[Klo80] Andrew Klossner. A parallel between operating system and human government. Operating Systems Review, 14(2):28–31, April 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kronenberg:1985:VEA

[KLS85] Nancy P. Kronenberg, Henry M. Levy, and William D. Strecker. VAXclusters (extended abstract): a closely-coupled distributed system. Operating Systems Review, 19(5):1, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karmon:2008:GPE

[KLS08] Kfir Karmon, Liran Liss, and Assaf Schuster. GWiQ-P: an efficient decentralized Grid-wide quota enforcement protocol. Operating Systems Review, 42(1):111–118, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kansal:2010:SLC

[KLS^{+}10] Aman Kansal, Jie Liu, Abhishek Singh, Ripal Nathuji, and Tarek Abdelzaher. Semantic-less coordination of power management and application performance. Operating Systems Review, 44(1):66–70, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2003:IBP

[KLY03] Hyun-Sung Kim, Sung-Woon Lee, and Kee-Young Yoo. IDbased password authentication scheme using smart cards and fingerprints. Operating Systems Review, 37(4):32–41, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Keller:2014:FSD

[KMA^{+}14] Gabriele Keller, Toby Murray, Sidney Amani, Liam O’Connor, Zilin Chen, Leonid Ryzhyk, Gerwin Klein, and Gernot Heiser. File systems deserve verification too! Operating Systems Review, 48(1):58–64, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kohler:2002:PLO

[KMC02] Eddie Kohler, Robert Morris, and Benjie Chen. Programming language optimizations for modular router configurations. Operating Systems Review, 36(5):251–263, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krieger:2010:EMC

[KMK10] Orran Krieger, Phil McGachey, and Arkady Kanevsky. Enabling a marketplace of clouds: VMware’s vCloud director. Operating Systems Review, 44(4):103–114, December 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kashyap:2016:OSA

[KMK16] Sanidhya Kashyap, Changwoo Min, and Taesoo Kim. Opportunistic spinlocks: Achieving virtual machine scalability in the clouds. Operating Systems Review, 50(1):9–16, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kuperman:2016:PR

[KMN^{+}16] Yossi Kuperman, Eyal Moscovici, Joel Nider, Razya Ladelsky, Abel Gordon, and Dan Tsafrir. Paravirtual remote I/O. Operating Systems Review, 50(2):49–65, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Keeton:2010:LFV

[KMSV10] Kimberly Keeton, Charles B. Morrey III, Craig A. N. Soules, and Alistair Veitch. LazyBase: freshness vs. performance in information management. Operating Systems Review, 44(1):15– 19, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Khalidi:1993:EFS

[KN93] Yousef A. Khalidi and Michael N. Nelson. Extensible file systems in spring. Operating Systems Review, 27(5):1–14, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotz:1996:FPP

[KN96] David Kotz and Nils Nieuwejaar. Flexibility and performance of parallel file systems. Operating Systems Review, 30(2):63–73, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Knott:1974:PCP

[Kno74] Gary D. Knott. A proposal for certain process management and intercommunication primitives. Operating Systems Review, 8(4):7–44, October 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Knott:1975:PCP

[Kno75] Gary D. Knott. A proposal for certain process management and intercommunication primitives. Operating Systems Review, 9(1):19–41, January 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Koren:2006:SLK

[Kor06] Oded Koren. A study of the Linux kernel evolution. Operating Systems Review, 40(2):110–112, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kosaraju:1973:LDS

[Kos73] S. Rao Kosaraju. Limitations of Dijkstra’s Semaphore Primitives and Petri nets. Operating Systems Review, 7(4):122–126, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotulski:1988:CIP

[Kot88] Leszek Kotulski. Comments on implementation of P and V primitives with help of binary semaphores. Operating Systems Review, 22(2):53–59, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kuenning:1997:AHM

[KP97] Geoffrey H. Kuenning and Gerald J. Popek. Automated hoarding for mobile computers. Operating Systems Review, 31(5): 264–275, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kavka:1993:EDM

[KPG93] C. Kavka, M. Printista, and R. Gallard. Extending device management in Minix. Operating Systems Review, 27(2):35–43, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kang:1999:PEN

[KPL99] Sung-Il Kang, Kihyun Pyun, and Heung-Kyu Lee. Performance evaluation of non-preemptive policies for scheduling readers and writers. Operating Systems Review, 33(3):43–61, July 1999. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kulkarni:2008:OPB

[KPR^{+}08] Milind Kulkarni, Keshav Pingali, Ganesh Ramanarayanan, Bruce Walter, Kavita Bala, and L. Paul Chew. Optimistic parallelism benefits from data partitioning. Operating Systems Review, 42(2):233–243, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kliot:2009:LFC

[KPS09] Gabriel Kliot, Erez Petrank, and Bjarne Steensgaard. A lockfree, concurrent, and incremental stack scanning mechanism for garbage collectors. Operating Systems Review, 43(3):3–13, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaufmann:2016:HPP

[KPS^{+}16a] Antoine Kaufmann, SImon Peter, Naveen Kr. Sharma, Thomas Anderson, and Arvind Krishnamurthy. High performance packet processing with FlexNIC. Operating Systems Review, 50(2):67– 81, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kolli:2016:HPT

[KPS^{+}16b] Aasheesh Kolli, Steven Pelley, Ali Saidi, Peter M. Chen, and Thomas F. Wenisch. High-performance transactions for persistent memories. Operating Systems Review, 50(2):399–411, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karges:1997:DIP

[KRS97] Jonathan Karges, Otto Ritter, and Sándor Suhai. Design and implementation of a parallel pipe. Operating Systems Review, 31(2):60–94, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kruijer:1982:PMC

[Kru82] H. S. M. Kruijer. Processor management in a concurrent Pascal kernel. Operating Systems Review, 16(2):7–17, April 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kant:1982:GCR

[KS82] Krishna Kant and Abraham Silberschatz. On the generalized critical region construct. Operating Systems Review, 16(3):4–16, July 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kepecs:1985:SSO

[KS85] Jonathan Kepecs and Marvin Solomon. SODA: a simplified operating system for distributed applications. Operating Systems Review, 19(4):45–56, October 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kistler:1991:DOC

[KS91a] James J. Kistler and M. Satyanarayanan. Disconnected operation in the Coda file system. Operating Systems Review, 25 (5):213–225, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kistler:1991:TDO

[KS91b] James Jay Kistler and M. Satyanarayanan. Transparent disconnected operation for fault-tolerance. Operating Systems Review, 25(1):77–80, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kifer:1992:OEO

[KS92] Michael Kifer and Scott A. Smolka. OSP: an environment for operating system projects. Operating Systems Review, 26(4):98– 100, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Klostermeyer:1995:RDP

[KS95] William F. Klostermeyer and Kankanahalli Srinivas. Reducing disk power consumption in a portable computer. Operating Systems Review, 29(2):27–32, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kwon:1999:CSR

[KS99] Taekyoung Kwon and Jooseok Song. Clarifying straight replays and forced delays. Operating Systems Review, 33(1):47–52, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadav:2009:LMD

[KS09] Asim Kadav and Michael M. Swift. Live migration of directaccess devices. Operating Systems Review, 43(3):95–104, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotra:2017:HSC

[KSCK17] Jagadish B. Kotra, Narges Shahidi, Zeshan A. Chishti, and Mahmut T. Kandemir. Hardware-software co-design to mitigate DRAM refresh overheads: a case for refresh-aware process scheduling. Operating Systems Review, 51(2):723–736, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2014:PCM

[KSDC14] Hyojun Kim, Sangeetha Seshadri, Clement L. Dickey, and Lawrence Chiu. Phase change memory in enterprise storage systems: silver bullet or snake oil? Operating Systems Review, 48(1):82–89, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Knezevic:2009:TCE

[KSK09] Nikola Knežević, Simon Schubert, and Dejan Kostić. Towards a cost-effective networking testbed. Operating Systems Review, 43 (4):66–71, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Koelbel:1990:WEB

[KSL90] Chuck Koelbel, Gene Spafford, and George Leach. Workshop on experiences with building distributed and multiprocessor systems. Operating Systems Review, 24(2):2–6, April 1990. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kehne:1992:NBP

[KSL92] A. Kehne, J. Schönwälder, and H. Langendörfer. A nonce-based protocol for multiple authentications. Operating Systems Review, 26(4):84–89, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kong:2008:PTD

[KSLA08] Jiantao Kong, Karsten Schwan, Min Lee, and Mustaque Ahamad. Protectit: trusted distributed services operating on sensitive data. Operating Systems Review, 42(4):137–147, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karcher:2009:ATS

[KSP09] Thomas Karcher, Christoph Schaefer, and Victor Pankratius. Auto-tuning support for manycore applications: perspectives for operating systems and compilers. Operating Systems Review, 43(2):96–97, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krause:1973:TSA

[KSS73] K. L. Krause, V. Y. Shen, and H. D. Schwetman. A taskscheduling algorithm for a multiprogramming computer system. Operating Systems Review, 7(4):112–118, October 1973. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krishnamurthy:1996:EAS

[KSS^{+}96] Arvind Krishnamurthy, Klaus E. Schauser, Chris J. Scheiman, Randolph Y. Wang, David E. Culler, and Katherine Yelick. Evaluation of architectural support for global address-based communication in large-scale parallel machines. Operating Systems Review, 30(5):37–48, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaashoek:1991:FTU

[KT91a] M. Frans Kaashoek and Andrew S. Tanenbaum. Fault tolerance using group communication. Operating Systems Review, 25(2): 71–74, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Katevenis:1991:RBP

[KT91b] Manolis Katevenis and Nestoras Tzartzanis. Reducing the branch penalty by rearranging instructions in a double-width memory. Operating Systems Review, 25(3S):15–27, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kamga:2012:ESE

[KTB12] Christine Mayap Kamga, Giang Son Tran, and Laurent Broto. Extended scheduler for efficient frequency scaling in virtualized systems. Operating Systems Review, 46(2):28–35, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ku:2003:TSA

[KTC03] Wei-Chi Ku, Hao-Chuan Tsai, and Shuai-Min Chen. Two simple attacks on Lin–Shen–Hwang’s strong-password authentication protocol. Operating Systems Review, 37(4):26–31, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2017:KPC

[KTG^{+}17] Jinchun Kim, Elvira Teran, Paul V. Gratz, Daniel A. Jiménez, Seth H. Pugsley, and Chris Wilkerson. Kill the program counter: Reconstructing program behavior in the processor cache hierarchy. Operating Systems Review, 51(2):737–749, June 2017. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaashoek:1989:ERB

[KTH89] M. Frans Kaashoek, A. S. Tanenbaum, and S. F. Hummel. An efficient reliable broadcast protocol. Operating Systems Review, 23(4):5–19, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kimbrel:1996:TDC

[KTP^{+}96] Tracy Kimbrel, Andrew Tomkins, R. Hugo Patterson, Brian Bershad, Pei Cao, Edward W. Felten, Garth A. Gibson, Anna R. Karlin, and Kai Li. A trace-driven comparison of algorithms for parallel prefetching and caching. Operating Systems Review, 30 (SI):19–34, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ku:2004:HBS

[Ku04] Wei-Chi Ku. A hash-based strong-password authentication scheme without using Smart Cards. Operating Systems Review, 38(1):29–34, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kuhnhauser:1998:CIA

[Küh98] Winfried E. Kühnhauser. A classification of interdomain actions. Operating Systems Review, 32(4):47–61, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kuhnhauser:1999:CKH

[Küh99] Winfried E. Kühnhauser. A comment on S. Kang’s and H. Lee’s paper on “Analysis and solution of non-preemptive policies for scheduling readers and writers” (OSR 32(2)). Operating Systems Review, 33(2):4, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [KL98].

Kuhnhauser:2004:RKO

[Küh04] Winfried E. Kühnhauser. Root Kits: an operating systems viewpoint. Operating Systems Review, 38(1):12–23, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kutti:1984:WDK

[Kut84] Swamy Kutti. Why a distributed kernel? Operating Systems Review, 18(4):5–11, October 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kaashoek:1992:FIP

[KvRvST92] M. Frans Kaashoek, Robbert van Renesse, Hans van Staveren, and Andrew S. Tanenbaum. FLIP; an Internetwork Protocol for supporting distributed systems. Operating Systems Review, 26 (2):29, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kermarrec:2007:GDS

[KvS07] Anne-Marie Kermarrec and Maarten van Steen. Gossiping in distributed systems. Operating Systems Review, 41(5):2–7, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kun:2000:SMA

[KXD00] Yang Kun, Guo Xin, and Liu Dayou. Security in mobile agent system: problems and approaches. Operating Systems Review, 34(1):21–28, January 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kanev:2017:MAM

[KXWB17] Svilen Kanev, Sam Likun Xi, Gu-Yeon Wei, and David Brooks. Mallacc: Accelerating memory allocation. Operating Systems Review, 51(2):33–45, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2002:DEC

[KY02] Dongkeun Kim and Donald Yeung. Design and evaluation of compiler algorithms for pre-execution. Operating Systems Review, 36(5):159–170, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Krohn:2007:IFC

[KYB^{+}07] Maxwell Krohn, Alexander Yip, Micah Brodsky, Natan Cliffer, M. Frans Kaashoek, Eddie Kohler, and Robert Morris. Information flow control for standard OS abstractions. Operating Systems Review, 41(6):321–334, December 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kwon:2017:IHP

[KYP^{+}17] Youngjin Kwon, Hangchen Yu, Simon Peter, Christopher J. Rossbach, and Emmett Witchel. Ingens: Huge page support for the OS and hypervisor. Operating Systems Review, 51(1):83– 93, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Khazraee:2017:MNO

[KZVT17] Moein Khazraee, Lu Zhang, Luis Vega, and Michael Bedford Taylor. Moonwalk: NRE optimization in ASIC clouds. Operating Systems Review, 51(2):511–526, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lim:1994:RSA

[LA94] Beng-Hong Lim and Anant Agarwal. Reactive synchronization algorithms for multiprocessors. Operating Systems Review, 28 (5):25–35, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lepreau:2000:HCE

[LAAW00] Jay Lepreau, Chris Alfeld, David Andersen, and Kristin Wright. A highly configurable emulation facility for distributed systems and networks. Operating Systems Review, 34(2):30, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lorch:2006:SWM

[LAB^{+}06] Jacob R. Lorch, Atul Adya, William J. Bolosky, Ronnie Chaiken, John R. Douceur, and Jon Howell. The SMART way to migrate replicated stateful services. Operating Systems Review, 40(4):103–115, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lampson:1991:ADS

[LABW91] Butler Lampson, Martín Abadi, Michael Burrows, and Edward Wobber. Authentication in distributed systems: theory and practice. Operating Systems Review, 25(5):165–182, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lacoste:2000:TSP

[Lac00] Marc Lacoste. Towards a secure platform for distributed mobile object computing. Operating Systems Review, 34(2):56–73, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lakhotia:1985:IE

[Lak85] Arun Lakhotia. Implication and equivalence I/O. Operating Systems Review, 19(1):46–52, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lampson:1974:P

[Lam74] Butler W. Lampson. Protection. Operating Systems Review, 8 (1):18–24, January 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lampson:1975:SIS

[Lam75] Butler Lampson. Synchronization: Introduction by the session chairman. Operating Systems Review, 9(3):1–2, July 1975. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lampson:1983:HCS

[Lam83] Butler W. Lampson. Hints for computer system design. Operating Systems Review, 17(5):33–48, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lamport:1985:SPU

[Lam85] Leslie Lamport. Solved problems, unsolved problems and nonproblems in concurrency. Operating Systems Review, 19(4):34– 44, October 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lampson:2000:CSR

[Lam00] Butler Lampson. Computer systems research (invited talk) (summary only): past and future. Operating Systems Review, 34 (2):8–9, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Landau:1989:SSC

[Lan89] Charles R. Landau. Security in a secure capability-based system. Operating Systems Review, 23(4):2–4, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

LaRowe:1992:OSR

[LaR92] Rick LaRowe. Operating Systems Research Related to the Galactica Net Architecture. Operating Systems Review, 26(2): 13, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lauer:1981:ODO

[Lau81] Hugh C. Lauer. Observations on the development of an operating system. Operating Systems Review, 15(5):30–36, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lau:1984:TPN

[Lau84] Francis C. M. Lau. Two-part names and process termination. Operating Systems Review, 18(3):28–30, July 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lazowska:1992:ASOa

[Laz92a] Edward D. Lazowska. 13th ACM Symposium on Operating Systems Principles: panel session presentations. Operating Systems Review, 26(1):3–17, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lazowska:1992:ASOb

[Laz92b] Edward D. Lazowska. 13th ACM Symposium on Operating Systems Principles: “work in progress”; abstracts. Operating Systems Review, 26(2):7, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lunn:1981:ARL

[LB81] K. Lunn and K. H. Bennett. An algorithm for resource location in a loosely linked distributed computer system. Operating Systems Review, 15(2):16–20, April 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

LeLann:1991:RAB

[LB91] G. Le Lann and G. Bres. Reliable atomic broadcast in distributed systems with omission faults. Operating Systems Review, 25(2):80–86, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2006:AER

[LB06] Benjamin C. Lee and David M. Brooks. Accurate and efficient regression modeling for microarchitectural performance and power prediction. Operating Systems Review, 40(5):185– 194, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2008:ETL

[LB08] Benjamin C. Lee and David Brooks. Efficiency trends and limits from comprehensive microarchitectural adaptivity. Operating Systems Review, 42(2):36–47, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1991:TYE

[LBB^{+}91] Jochen Liedtke, Ulrich Bartling, Uwe Beyer, Dietmar Heinrichs, Rudolf Ruland, and Gyula Szalay. Two years of experience with a μ-kernel based OS. Operating Systems Review, 25(2):51–62, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1998:STS

[LBF^{+}98] Walter Lee, Rajeev Barua, Matthew Frank, Devabhaktuni Srikrishna, Jonathan Babb, Vivek Sarkar, and Saman Amarasinghe. Space-time scheduling of instruction-level parallelism on a raw machine. Operating Systems Review, 32(5):46–57, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liyong:2003:NPP

[LBJ03] Ren Liyong, Chen Bo, and Wu Jing. A novel packet-pair-based inferring bandwidth congestion control mechanism for layered multicast. Operating Systems Review, 37(4):63–69, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Laadan:2007:DPV

[LBP^{+}07] Oren Laadan, Ricardo A. Baratto, Dan B. Phung, Shaya Potter, and Jason Nieh. DejaView: a personal virtual computer recorder. Operating Systems Review, 41(6):279–292, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2006:MEM

[LBvH06] Xin Li, Marian Boldt, and Reinhard von Hanxleden. Mapping Esterel onto a multi-threaded embedded processor. Operating Systems Review, 40(5):303–314, December 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Low:1993:FGO

[LC93] Marie Rose Low and Bruce Christianson. Fine grained object protection in UNIX. Operating Systems Review, 27(1):33–50, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lowell:1997:FTR

[LC97] David E. Lowell and Peter M. Chen. Free transactions with Rio Vista. Operating Systems Review, 31(5):92–101, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levis:2002:MTV

[LC02] Philip Levis and David Culler. Maté: a tiny virtual machine for sensor networks. Operating Systems Review, 36(5):85–95, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2004:SOT

[LC04a] Min-Hui Lin and Chin-Chen Chang. A secure one-time password authentication scheme with low-computation for mobile communications. Operating Systems Review, 38(2):76–84, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ling:2004:MCF

[LC04b] Yibei Ling and Wai Chen. Measuring cache freshness by additive age. Operating Systems Review, 38(3):12–17, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levin:1975:PMS

[LCC^{+}75] R. Levin, E. Cohen, W. Corwin, F. Pollack, and W. Wulf. Policy/mechanism separation in Hydra. Operating Systems Review, 9(5):132–140, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2017:SSA

[LCCZ17] Kaiwei Li, Jianfei Chen, Wenguang Chen, and Jun Zhu. SaberLDA: Sparsity-aware learning of topic models on GPUs. Operating Systems Review, 51(2):497–509, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Luderer:1981:DUS

[LCH^{+}81] G. W. R. Luderer, H. Che, J. P. Haggerty, P. A. Kirslis, and W. T. Marshall. A distributed UNIX system based on a virtual circuit switch. Operating Systems Review, 15(5):160–168, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loo:2005:IDO

[LCH^{+}05] Boon Thau Loo, Tyson Condie, Joseph M. Hellerstein, Petros Maniatis, Timothy Roscoe, and Ion Stoica. Implementing declarative overlays. Operating Systems Review, 39(5):75–90, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liskov:1987:IA

[LCJS87] B. Liskov, D. Curtis, P. Johnson, and R. Scheifer. Implementation of Argus. Operating Systems Review, 21(5):111–122, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lagar-Cavilla:2011:TBS

[LCJV^{+}11] H. Andrés Lagar-Cavilla, Kaustubh Joshi, Alexander Varshavsky, Jeffrey Bickford, and Darwin Parra. Traffic backfilling: subsidizing lunch for delay-tolerant applications in UMTS networks. Operating Systems Review, 45(3):77–81, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Las-Casas:2024:LAA

[LCKFA24] Pedro Las-Casas, Alok Gautum Kumbhare, Rodrigo Fonseca, and Sharad Agarwal. LLexus: an AI agent system for incident management. Operating Systems Review, 58(1):23–36, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3689051.3689056.

Lin:2016:SKT

[LCL^{+}16] Xiaofeng Lin, Yu Chen, Xiaodong Li, Junjie Mao, Jiaquan He, Wei Xu, and Yuanchun Shi. Scalable kernel TCP design and implementation for short-lived connections. Operating Systems Review, 50(2):339–352, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2001:CNP

[LCTK01] Bu-Sung Lee, Wen-Tong Cai, Stephen J. Turner, and Jit-Beng Koh. Comparison of network protocol and architecture for distributed virtual simulation environment. Operating Systems Review, 35(3):30–42, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Linderman:2008:MPM

[LCWM08] Michael D. Linderman, Jamison D. Collins, Hong Wang, and Teresa H. Meng. Merge: a programming model for heterogeneous multi-core systems. Operating Systems Review, 42(2):287– 296, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1996:GPT

[LE96] Jochen Liedtke and Kevin Elphinstone. Guarded page tables on Mips R4600 or an exercise in architecture-dependent micro optimization. Operating Systems Review, 30(1):4–15, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Le:1998:OET

[Le98] Bich C. Le. An out-of-order execution technique for runtime binary translators. Operating Systems Review, 32(5):151–158, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lepreau:2000:SSS

[LE00] Jay Lepreau and Eric Eide. Session summaries from the 17th Symposium on Operating Systems Principle (SOSP’99). Operating Systems Review, 34(2):4–5, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ledru:1997:APE

[Led97] Pascal Ledru. Adaptive parallelism: an early experiment with Java remote method invocation. Operating Systems Review, 31 (4):24–29, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1999:BBB

[Lee99] Jong-Hyeon Lee. The big brother ballot. Operating Systems Review, 33(3):19–25, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lantz:1986:TUD

[LEH86] Keith A. Lantz, Judy L. Edighoffer, and Bruce L. Hitson. Towards a universal directory service. Operating Systems Review, 20(2):43–53, April 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leibfried:1989:DDR

[Lei89] T. F. Leibfried. A deadlock detection and recovery algorithm using the formalism of a directed graph matrix. Operating Systems Review, 23(2):45–55, April 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

LaRowe:1991:RNM

[LEK91] Richard P. LaRowe, Jr., Carla Schlatter Ellis, and Laurence S. Kaplan. The robustness of NUMA memory management. Operating Systems Review, 25(5):137–151, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lesokhin:2017:PFS

[LER^{+}17] Ilya Lesokhin, Haggai Eran, Shachar Raindel, Guy Shapiro, Sagi Grimberg, Liran Liss, Muli Ben-Yehuda, Nadav Amit, and Dan Tsafrir. Page fault support for network controllers. Operating Systems Review, 51(2):449–466, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leschke:2004:ASF

[Les04] Tim Leschke. Achieving speed and flexibility by separating management from protection: embracing the Exokernel operating system. Operating Systems Review, 38(4):5–19, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levy:1988:SFR

[Lev88] Hank Levy. SIGOPS Financial Report August 1, 1988. Operating Systems Review, 22(4):1–2, October 1988. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levy:1990:NSD

[Lev90] Hank Levy. New SIGOPS dues structure. Operating Systems Review, 24(3):1, July 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levine:2003:DD

[Lev03a] Gertrude Neuman Levine. Defining deadlock. Operating Systems Review, 37(1):54–64, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levine:2003:DDF

[Lev03b] Gertrude Neuman Levine. Defining deadlock with fungible resources. Operating Systems Review, 37(3):5–11, July 2003. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levine:2005:CDP

[Lev05] Gertrude Neuman Levine. The classification of deadlock prevention and avoidance is erroneous. Operating Systems Review, 39(2):47–50, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levin:2007:PCR

[Lev07] Roy Levin. A perspective on computing research management. Operating Systems Review, 41(2):3–9, April 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ludwich:2013:FVC

[LF13] Mateus Krepsky Ludwich and Antônio Augusto Fröhlich. On the formal verification of component-based embedded operating systems. Operating Systems Review, 47(1):28–34, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lacage:2009:NUI

[LFH^{+}09] Mathieu Lacage, Martin Ferrari, Mads Hansen, Thierry Turletti, and Walid Dabbous. NEPI: using independent simulators, emulators, and testbeds for easy experimentation. Operating Systems Review, 43(4):60–65, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liaw:2004:SPA

[LFW04] Horng-Twu Liaw, Shiou-Wei Fan, and Wei-Chen Wu. A simple password authentication using a polynomial. Operating Systems Review, 38(4):74–79, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lou:2010:MDD

[LFWL10] Jian-Guang Lou, Qiang Fu, Yi Wang, and Jiang Li. Mining dependency in distributed systems through unstructured logs analysis. Operating Systems Review, 44(1):91–96, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lebeck:2000:PAP

[LFZE00] Alvin R. Lebeck, Xiaobo Fan, Heng Zeng, and Carla Ellis. Power aware page allocation. Operating Systems Review, 34(5): 105–116, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lumb:2004:DSD

[LG04] Christopher R. Lumb and Richard Golding. D-SPTF: decentralized request distribution in brick-based storage systems. Operating Systems Review, 38(5):37–47, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liskov:1991:RHF

[LGG^{+}91] Barbara Liskov, Sanjay Ghemawat, Robert Gruber, Paul Johnson, and Liuba Shrira. Replication in the Harp file system. Operating Systems Review, 25(5):226–238, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Laudon:1994:IMT

[LGH94] James Laudon, Anoop Gupta, and Mark Horowitz. Interleaving: a multithreading technique targeting multiprocessors and workstations. Operating Systems Review, 28(5):308–318, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liskov:1991:RUF

[LGJS91] Barbara Liskov, Robert Gruber, Paul Johnson, and Liuba Shrira. A replicated Unix file system (extended abstract). Operating Systems Review, 25(1):60–64, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leite:2014:BSC

[LGMF14] Julius Leite, Raphael Guerra, Rivalino Matias, Jr., and Antônio Augusto Fröhlich. Brazilian Symposium on Computer System Engineering, November 4–8 2013, Niterói, Brazil. Operating Systems Review, 48(1):1, 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liang:2007:RDM

[LGN07] Jin Liang, Indranil Gupta, and Klara Nahrstedt. Reliable ondemand management operations for large-scale distributed applications. Operating Systems Review, 41(5):82–88, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lomas:1989:RRP

[LGSN89] T. Lomas, L. Gong, J. Saltzer, and R. Needhamn. Reducing risks from poorly chosen keys. Operating Systems Review, 23 (5):14–18, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:ICK

[LH04] Tian-Fu Lee and Tzonelih Hwang. Improved conference key distribution protocol based on a symmetric balanced incomplete block design. Operating Systems Review, 38(3):58–64, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:SAA

[LHL04] Cheng-Chi Lee, Min-Shiang Hwang, and I-En Liao. A server assisted authentication protocol for detecting error vectors. Operating Systems Review, 38(2):93–96, April 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levinthal:1987:PCG

[LHPL87] Adam Levinthal, Pat Hanrahan, Mike Paquette, and Jim Lawson. Parallel computers for graphics applications. Operating Systems Review, 21(4):193–198, October 1987. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lindsay:1983:CCR

[LHWY83] Bruce G. Lindsay, Laura M. Haas, Paul F. Wilms, and Robert A. Yost. Computation & communication in R: a distributed database manager. Operating Systems Review, 17(5):1–2, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2002:FRU

[LHY02] Cheng-Chi Lee, Min-Shiang Hwang, and Wei-Peng Yang. A flexible remote user authentication scheme using Smart Cards. Operating Systems Review, 36(3):46–52, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liebl:1993:ADS

[Lie93a] Armin Liebl. Authentication in distributed systems: a bibliography. Operating Systems Review, 27(4):31–41, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1993:IIK

[Lie93b] Jochen Liedtke. Improving IPC by kernel design. Operating Systems Review, 27(5):175–188, December 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1994:SNIa

[Lie94a] Jochen Liedtke. A short note on implementing “new” machine instructions by software for efficient test of page accessibility. Operating Systems Review, 28(1):61–65, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1994:SNIb

[Lie94b] Jochen Liedtke. A short note on implementing thread exclusiveness and address space locking. Operating Systems Review, 28(3):38–42, July 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1995:MKC

[Lie95a] J. Liedtke. On micro-kernel construction. Operating Systems Review, 29(5):237–250, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1995:ASS

[Lie95b] Jochen Liedtke. Address space sparsity and fine granularity. Operating Systems Review, 29(1):87–90, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1995:SNS

[Lie95c] Jochen Liedtke. A short note a small virtually-addressed control blocks. Operating Systems Review, 29(3):31–34, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liedtke:1996:SNC

[Lie96] Jochen Liedtke. A short note on cheap fine-grained time measurement. Operating Systems Review, 30(2):92–94, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lightner:1994:BRT

[Lig94] Michael Lightner. Book review: TEX in Practice, Volumes 1– 4 by Stephan von Bechtolsheim: (Springer-Verlag, New York 1993). Operating Systems Review, 28(2):6–8, April 1994. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lightner:1995:BRT

[Lig95] Michael Lightner. Book review: TEX in Practice, Stephan von Bechtolsheim. Operating Systems Review, 29(1):2–6, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lindsay:1981:BLS

[Lin81] D. C. Lindsay. On binding layers of software. Operating Systems Review, 15(2):33–37, April 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lions:1978:OSC

[Lio78] J. Lions. An operating system case study. Operating Systems Review, 12(3):46–53, July 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lipner:1975:CCP

[Lip75] Steven B. Lipner. A comment on the confinement problem. Operating Systems Review, 9(5):192–196, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liskov:1972:DVO

[Lis72] Barbara H. Liskov. The design of the Venus Operating System. Operating Systems Review, 6(1/2):11–16, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lister:1977:PNM

[Lis77] Andrew Lister. The problem of nested monitor calls. Operating Systems Review, 11(3):5–7, July 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Litant:1987:BRC

[Lit87] Thomas F. Litant. Book review: Computer Security: the Practical Issues in a Troubled World, (Elsevier Science Publishers, Amsterdam 1985). Operating Systems Review, 21(1):3–5, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Litman:1988:DDO

[Lit88] Ami Litman. The DUNIX distributed operating system. Operating Systems Review, 22(1):42–51, January 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liaqat:2016:SEE

[LJdL^{+}16] Daniyal Liaqat, Silviu Jingoi, Eyal de Lara, Ashvin Goel, Wilson To, Kevin Lee, Italo De Moraes Garcia, and Manuel Saldana. Sidewinder: an energy efficient and developer friendly heterogeneous architecture for continuous mobile sensing. Operating Systems Review, 50(2):205–215, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2002:UIO

[LJS^{+}02] Tao Li, Lizy Kurian John, Anand Sivasubramaniam, N. Vijaykrishnan, and Juan Rubio. Understanding and improving operating system effects in control flow prediction. Operating Systems Review, 36(5):68–80, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lv:2006:FTC

[LJW^{+}06] Qin Lv, William Josephson, Zhe Wang, Moses Charikar, and Kai Li. Ferret: a toolkit for content-based similarity search of feature-rich data. Operating Systems Review, 40(4):317–330, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:1997:GCM

[LJX97a] Qun Li, Hua Ji, and Li Xie. Group consistency model which separates the intra-group consistency maintenance from the intergroup consistency maintenance in large scale DSM systems. Operating Systems Review, 31(2):23–35, April 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:1997:BPF

[LJX97b] Qun Li, Jie Jing, and Li Xie. BFXM: a parallel file system model based on the mechanism of distributed shared memory. Operating Systems Review, 31(4):30–40, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:SSP

[LJY04] Jung-Seuk Lee, Jun-Cheol Jeon, and Kee-Young Yoo. A security scheme for protecting security policies in firewall. Operating Systems Review, 38(2):69–72, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1991:PCP

[LK91] Edward K. Lee and Randy H. Katz. Performance consequences of parity placement in disk arrays. Operating Systems Review, 25(3S):190–199, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ludwig:2001:FSE

[LK01] Stefan Ludwig and Winfried Kalfa. File system encryption with integrated user management. Operating Systems Review, 35 (4):88–93, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Libster:2008:PIM

[LK08] Eugene Libster and Jesse D. Kornblum. A proposal for an integrated memory acquisition mechanism. Operating Systems Review, 42(3):14–20, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leverich:2010:EEH

[LK10] Jacob Leverich and Christos Kozyrakis. On the energy (in)efficiency of Hadoop clusters. Operating Systems Review, 44(1):61–65, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1991:FPP

[LKB91] Roland L. Lee, Alex Y. Kwok, and Fayé A. Briggs. The floatingpoint performance of a superscalar SPARC processor. Operating Systems Review, 25(3S):28–37, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2003:PSAa

[LKKY03a] Sung-Woon Lee, Woo-Hun Kim, Hyun-Sung Kim, and Kee-Young Yoo. Parallizable simple authenticated key agreement protocol. Operating Systems Review, 37(2):13–18, April 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2003:PSAb

[LKKY03b] Sung-Woon Lee, Woo-Hun Kim, Hyun-Sung Kim, and Kee-Young Yoo. Parallizable simple authenticated key agreement protocol. Operating Systems Review, 37(3):17–22, July 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:1999:BRH

[LKvR^{+}99] Xiaoming Liu, Christoph Kreitz, Robbert van Renesse, Jason Hickey, Mark Hayden, Kenneth Birman, and Robert Constable. Building reliable, high-performance communication systems from components. Operating Systems Review, 33(5):80–92, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2000:BRH

[LKvR^{+}00] Xiaoming Liu, Christoph Kreitz, Robbert van Renesse, Jason Hickey, Mark Hayden, Kenneth Birman, and Robert Constable. Building reliable, high-performance communication systems from components. Operating Systems Review, 34(2):16–17, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:CUA

[LKY04] Sung-Woon Lee, Hyun-Sung Kim, and Kee-Young Yoo. Cryptanalysis of a user authentication scheme using hash functions. Operating Systems Review, 38(1):24–28, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:1998:SLB

[LL98] Sanglu Lu and Xie Li. A scalable loading balancing system for NOWs. Operating Systems Review, 32(3):55–63, July 1998. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lepak:2002:TSS

[LL02] Kevin M. Lepak and Mikko H. Lipasti. Temporally silent stores. Operating Systems Review, 36(5):30–41, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:CMK

[LL04] Narn-Yih Lee and Ming-Feng Lee. Comments on multiparty key exchange scheme. Operating Systems Review, 38(4):70–73, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2016:MTP

[LL16] Felix Xiaozhu Lin and Xu Liu. memif: Towards programming heterogeneous memory asynchronously. Operating Systems Review, 50(2):369–383, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lazowska:1981:AES

[LLA^{+}81] Edward D. Lazowska, Henry M. Levy, Guy T. Almes, Michael J. Fischer, Robert J. Fowler, and Stephen C. Vestal. The architecture of the Eden system. Operating Systems Review, 15(5): 148–159, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2004:PDE

[LLD^{+}04] Xiaodong Li, Zhenmin Li, Francis David, Pin Zhou, Yuanyuan Zhou, Sarita Adve, and Sanjeev Kumar. Performance directed energy management for main memory and disks. Operating Systems Review, 38(5):271–283, December 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2002:RUA

[LLH02] Cheng-Chi Lee, Li-Hua Li, and Min-Shiang Hwang. A remote user authentication scheme using hash functions. Operating Systems Review, 36(4):23–29, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2004:CES

[LLH04] Li-Hua Li, Chi-Yu Liu, and Min-Shiang Hwang. Cryptanalysis of an efficient secure group signature scheme. Operating Systems Review, 38(4):66–69, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1996:SSA

[LLK96] Jongwoon Lee, Sungyoung Lee, and Hyungill Kim. Scheduling soft aperiodic tasks in adaptable fixed-priority systems. Operating Systems Review, 30(4):17–28, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2017:DAD

[LLL^{+}17] Haopeng Liu, Guangpu Li, Jeffrey F. Lukman, Jiaxin Li, Shan Lu, Haryadi S. Gunawi, and Chen Tian. DCatch: Automatically detecting distributed concurrency bugs in cloud systems. Operating Systems Review, 51(2):677–691, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leesatapornwongsa:2016:TTN

[LLLG16] Tanakorn Leesatapornwongsa, Jeffrey F. Lukman, Shan Lu, and Haryadi S. Gunawi. TaxDC: a taxonomy of non-deterministic concurrency bugs in datacenter distributed systems. Operating Systems Review, 50(2):517–530, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2017:ITN

[LLN^{+}17] Ming Liu, Liang Luo, Jacob Nelson, Luis Ceze, Arvind Krishnamurthy, and Kishore Atreya. IncBricks: Toward in-network computation with an in-network cache. Operating Systems Review, 51(2):795–809, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ladin:1991:LRE

[LLS91] Rivka Ladin, Barbara Liskov, and Liuba Shrira. Lazy replication: exploiting the semantics of distributed services (extended abstract). Operating Systems Review, 25(1):49–55, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2008:PBH

[LLS^{+}08] Tiancheng Liu, Ying Li, Andrew Schofield, Matt Hogstrom, Kewei Sun, and Ying Chen. Partition-based heap memory management in an application server. Operating Systems Review, 42(1):98, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Logan:2024:EDS

[LLSK24] Luke Logan, Jay Lofstead, Xian-He Sun, and Anthony Kougkas. An evaluation of DAOS for simulation and deep learning HPC-Workloads. Operating Systems Review, 58(1):37–44, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3689051.
3689058.

Luo:2005:RCS

[LLY05] Guangchun Luo, Xianliang Lu, and Ting Yang. The research on consistency of space/time of IDS. Operating Systems Review, 39(3):44–51, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Luk:1996:CBP

[LM96] Chi-Keung Luk and Todd C. Mowry. Compiler-based prefetching for recursive data structures. Operating Systems Review, 30 (5):222–233, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Low:1997:JAS

[LM97] Marie Rose Low and James A. Malcolm. A joint authorisation scheme. Operating Systems Review, 31(1):88–96, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lakshman:2010:CDS

[LM10] Avinash Lakshman and Prashant Malik. Cassandra: a decentralized structured storage system. Operating Systems Review, 44(2):35–40, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lachenmann:2007:RML

[LMG^{+}07] Andreas Lachenmann, Pedro José Marrón, Matthias Gauger, Daniel Minder, Olga Saukh, and Kurt Rothermel. Removing the memory limitations of sensor networks with flash-based virtual memory. Operating Systems Review, 41(3):131–144, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ling:2000:AOT

[LML00] Yibei Ling, Tracy Mullen, and Xiaola Lin. Analysis of optimal thread pool size. Operating Systems Review, 34(2):42–55, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leslie:1993:POS

[LMM93] Ian M. Leslie, Derek McAuley, and Sape J. Mullender. Pegasus—operating system support for distributed multimedia systems. Operating Systems Review, 27(1):69–78, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Laden:2012:ADF

[LMV12] Guy Laden, Roie Melamed, and Ymir Vigfusson. Adaptive and dynamic funnel replication in clouds. Operating Systems Review, 46(1):40–46, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lauer:1979:DOS

[LN79] Hugh C. Lauer and Roger M. Needham. On the duality of operating system structures. Operating Systems Review, 13(2): 3–19, April 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lvin:2008:ATA

[LNBZ08] Vitaliy B. Lvin, Gene Novark, Emery D. Berger, and Benjamin G. Zorn. Archipelago: trading address space for reliability and security. Operating Systems Review, 42(2):115–124, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loepere:1985:RCC

[Loe85] Keith Loepere. Resolving covert channels within a B2 class secure system. Operating Systems Review, 19(3):9–28, July 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loepere:1989:CCL

[Loe89] Keith Loepere. The covert channel limiter revisited. Operating Systems Review, 23(2):39–44, April 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loepere:2005:STM

[Loe05] Keith Loepere. Stackable thread mechanisms. Operating Systems Review, 39(4):4–17, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lohr:1977:BCP

[Löh77] Klaus-Peter Löhr. Beyond concurrent Pascal. Operating Systems Review, 11(5):173–180, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lomet:1977:PSS

[Lom77] D. B. Lomet. Process structuring, synchronization, and recovery using atomic actions. Operating Systems Review, 11(2):128–137, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ledlie:2009:CTB

[LOM^{+}09] Jonathan Ledlie, Billy Odero, Einat Minkov, Imre Kiss, and Joseph Polifroni. Crowd translator: on building localized speech recognizers through micropayments. Operating Systems Review, 43(4):84–89, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Long:1993:NBM

[Lon93] Darrell D. E. Long. A note on bit-mapped free sector management. Operating Systems Review, 27(2):7–9, April 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lorin:1986:EAO

[Lor86] Harold Lorin. An expanded approach to objects. Operating Systems Review, 20(1):6–11, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Love:1977:EIE

[Lov77] Tom Love. An experimental investigation of the effect of program structure on program understanding. Operating Systems Review, 11(2):105–113, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Litiu:2001:DMC

[LP01] Radu Litiu and Atul Prakash. DACIA: a mobile component framework for building adaptive distributed applications. Operating Systems Review, 35(2):31–42, April 2001. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2007:MAI

[LPH^{+}07] Shan Lu, Soyeon Park, Chongfeng Hu, Xiao Ma, Weihang Jiang, Zhenmin Li, Raluca A. Popa, and Yuanyuan Zhou. MUVI: automatically inferring multi-variable access correlations and detecting related semantic and concurrency bugs. Operating Systems Review, 41(6):103–116, December 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2017:EML

[LPM17] Cha V. Li, Vinicius Petrucci, and Daniel Mossé. Exploring machine learning for thread characterization on heterogeneous multiprocessors. Operating Systems Review, 51(1):113–123, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Letia:2010:CCC

[LPS10] Mihai Letia, Nuno Preguiça, and Marc Shapiro. Consistency without concurrency control in large, dynamic systems. Operating Systems Review, 44(2):29–34, April 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2008:LMC

[LPSZ08] Shan Lu, Soyeon Park, Eunsoo Seo, and Yuanyuan Zhou. Learning from mistakes: a comprehensive study on real world concurrency bug characteristics. Operating Systems Review, 42(2):329– 339, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2008:UPH

[LRS^{+}08] Man-Lap Li, Pradeep Ramachandran, Swarup Kumar Sahoo, Sarita V. Adve, Vikram S. Adve, and Yuanyuan Zhou. Understanding the propagation of hard errors to software and implications for resilient system design. Operating Systems Review, 42 (2):265–276, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Larus:1994:LMS

[LRV94] James R. Larus, Brad Richards, and Guhan Viswanathan. LCM: memory system support for parallel language implementation. Operating Systems Review, 28(5):208–218, December 1994. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lam:1991:CPO

[LRW91] Monica D. Lam, Edward E. Rothberg, and Michael E. Wolf. The cache performance and optimizations of blocked algorithms. Operating Systems Review, 25(3S):63–74, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lam:1975:ALA

[LS75] Shui Lam and Ravi Sethi. Analysis of a level algorithm for preemptive scheduling. Operating Systems Review, 9(5):178–186, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lamport:1986:BCS

[LS86] Leslie Lamport and P. M. Melliar Smith. Byzantine clock synchronization. Operating Systems Review, 20(3):10–16, July 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lorence:1990:ITM

[LS90] Mark J. Lorence and M. Satyanarayanan. IPwatch: a tool for monitoring network locality. Operating Systems Review, 24 (1):58–80, January 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:1994:IOT

[LS94] Qi Lu and M. Satyanarayanan. Isolation-only transactions for mobile computing. Operating Systems Review, 28(2):81–87, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Loo:2009:IWN

[LS09] Boon Thau Loo and Stefan Saroiu. 5th International Workshop on Networking Meets Databases (NetDB 2009). Operating Systems Review, 43(4):17–18, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2000:FCRa

[LSA^{+}00a] Chenyang Lu, John A. Stankovic, Tarek Abdelzaher, Sang H. Son, and Gang Tao. Feedback control real-time scheduling: support for performance guarantees in unpredictable environments. Operating Systems Review, 34(2):33, April 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2000:FCRb

[LSA^{+}00b] Chenyang Lu, John A. Stankovic, Tarek Abdelzaher, Sang H. Son, and Gang Tao. Feedback control real-time scheduling: support for performance guarantees in unpredictable environments (poster session). Operating Systems Review, 34(2):40, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liskov:1977:AMC

[LSAS77] Barbara Liskov, Alan Snyder, Russell Atkinson, and Craig Schaffert. Abstraction mechanisms in CLU. Operating Systems Review, 11(2):140, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2000:TPE

[LSH00] Chun-Li Lin, Hung-Min Sun, and Tzonelih Hwang. Three-party encrypted key exchange: attacks and a solution. Operating Systems Review, 34(4):12–20, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2001:EPD

[LSH01] Chun-Li Lin, Hung-Min Sun, and Tzonelih Hwang. Efficient and practical DHEKE protocols. Operating Systems Review, 35 (1):41–47, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2003:SEOa

[LSH03a] Chih-Wei Lin, Jau-Ji Shen, and Min-Shiang Hwang. Security enhancement for Optimal Strong-Password Authentication protocol. Operating Systems Review, 37(2):7–12, April 2003. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2003:SEOb

[LSH03b] Chih-Wei Lin, Jau-Ji Shen, and Min-Shiang Hwang. Security enhancement for Optimal Strong-Password Authentication Protocol. Operating Systems Review, 37(3):12–16, July 2003. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2008:LLA

[LSKK08] Sungjin Lee, Dongkun Shin, Young-Jin Kim, and Jihong Kim. LAST: locality-aware sector translation for NAND flash memory-based storage systems. Operating Systems Review, 42 (6):36–42, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2017:LAC

[LSL^{+}17] Ang Li, Shuaiwen Leon Song, Weifeng Liu, Xu Liu, Akash Kumar, and Henk Corporaal. Locality-aware CTA clustering for modern GPUs. Operating Systems Review, 51(2):297–311, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lustig:2016:CVM

[LSMB16] Daniel Lustig, Geet Sethi, Margaret Martonosi, and Abhishek Bhattacharjee. COATCheck: Verifying memory ordering at the hardware-OS interface. Operating Systems Review, 50(2):233– 247, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2007:CPC

[LSP07] Chuanpeng Li, Kai Shen, and Athanasios E. Papathanasiou. Competitive prefetching for concurrent sequential I/O. Operating Systems Review, 41(3):189–202, June 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lowell:2004:DVM

[LSS04] David E. Lowell, Yasushi Saito, and Eileen J. Samberg. Devirtualizable virtual machines enabling general, single-node, online maintenance. Operating Systems Review, 38(5):211–223, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lohmann:2006:QAA

[LST^{+}06] Daniel Lohmann, Fabian Scheler, Reinhard Tartler, Olaf Spinczyk, and Wolfgang Schröder-Preikschat. A quantitative analysis of aspects in the eCos kernel. Operating Systems Review, 40(4):191–204, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lecuyer:2019:PAQ

[LSV^{+}19] Mathias L‘ecuyer, Riley Spahn, Kiran Vodrahalli, Roxana Geambasu, and Daniel Hsu. Privacy accounting and quality control in the Sage differentially private ML platform. Operating Systems Review, 53(1):75–84, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:1996:PDV

[LT96] Edward K. Lee and Chandramohan A. Thekkath. Petal: distributed virtual disks. Operating Systems Review, 30(5):84–92, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lubowich:2011:PDL

[LT11] Yuval Lubowich and Gadi Taubenfeld. On the performance of distributed lock-based synchronization? Operating Systems Review, 45(2):28–37, July 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Levi:1989:MHR

[LTCA89] S. Levi, S. K. Tripathi, S. D. Carson, and A. K. Agrawala. The MARUTI hard real-time operating system. Operating Systems Review, 23(3):90–105, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2006:ADA

[LTQZ06] Shan Lu, Joseph Tucek, Feng Qin, and Yuanyuan Zhou. AVIO: detecting atomicity violations via access interleaving invariants.

Operating Systems Review, 40(5):37–48, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lux:1995:AOM

[Lux95] Wolfgang Lux. Adaptable object migration: concept and implementation. Operating Systems Review, 29(2):54–69, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2001:CNI

[LW01] Yuhong Li and Lars Wolf. Collection of network information in active networks. Operating Systems Review, 35(4):39–49, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:IAK

[LW04] Narn-Yih Lee and Chien-Nan Wu. Improved authentication key exchange protocol without using one-way hash function. Operating Systems Review, 38(2):85–92, April 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lin:2005:PFA

[LWMX05] Zhiqiang Lin, Chao Wang, Bing Mao, and Li Xie. A policy flexible architecture for secure operating system. Operating Systems Review, 39(3):24–33, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lustig:2017:ASC

[LWPG17] Daniel Lustig, Andrew Wright, Alexandros Papakonstantinou, and Olivier Giroux. Automated synthesis of comprehensive memory model litmus test suites. Operating Systems Review, 51(2):661–675, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lober:2009:DIL

[LWQ09] William B. Lober, Stephen Wagner, and Christina Quiles. Development and implementation of a loosely coupled, multi-site, networked and replicated electronic medical record in Haiti. Operating Systems Review, 43(4):79–83, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lipasti:1996:VLL

[LWS96] Mikko H. Lipasti, Christopher B. Wilkerson, and John Paul Shen. Value locality and load value prediction. Operating Systems Review, 30(5):138–147, December 1996. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2004:SCB

[LWY^{+}04] Bu-Sung Lee, Wing-Keong Woo, Chai-Kiat Yeo, Teck-Meng Lim, Bee-Hwa Lim, Yuxiong He, and Jie Song. Secure communications between bandwidth brokers. Operating Systems Review, 38(1):43–57, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lu:2000:SSN

[LX00] Sanglu Lu and Li Xie. Scalable scheduling on a network of workstations. Operating Systems Review, 34(2):74–83, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liang:2019:CLS

[LXYZ19] Chieh-Jan Mike Liang, Hui Xue, Mao Yang, and Lidong Zhou. The case for learning-and-system co-design. Operating Systems Review, 53(1):68–74, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Luo:2003:PBO

[LZ03] Lei Luo and Ming-Yuan Zhu. Partitioning based operating system: a formal model. Operating Systems Review, 37(3):23–35, July 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2017:DBD

[LZC^{+}17] Mengxing Liu, Mingxing Zhang, Kang Chen, Xuehai Qian, Yongwei Wu, Weimin Zheng, and Jinglei Ren. DudeTM: Building durable transactions with decoupling for persistent memory. Operating Systems Review, 51(2):329–343, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Li:2022:IFT

[LZH^{+}22] Yichen Li, Xu Zhang, Shilin He, Zhuangbin Chen, Yu Kang, Jinyang Liu, Liqun Li, Yingnong Dang, Feng Gao, Zhangwei Xu, Saravan Rajmohan, Qingwei Lin, Dongmei Zhang, and Michael R. Lyu. An intelligent framework for timely, accurate, and comprehensive cloud incident detection. Operating Systems Review, 56(1):1–7, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3544497.3544499.

Liu:2003:ODS

[LZJ03] Jing Liu, Mingyang Zheng, and Jiubin Ju. Offering different services by server clusters. Operating Systems Review, 37(1):14– 22, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maegaard:1979:ROS

[MA79] Henrik Maegaard and Aksel Andreasen. REPOS: An operating system for the PDP-11. Operating Systems Review, 13(3):6–11, July 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Minet:1991:ABO

[MA91] Pascale Minet and Emmanuelle Anceaume. Atomic broadcast in one phase. Operating Systems Review, 25(2):87–90, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:2006:SBI

[MA06] Jason E. Miller and Anant Agarwal. Software-based instruction caching for embedded processors. Operating Systems Review, 40(5):293–302, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McDougall:2010:VPP

[MA10] Richard McDougall and Jennifer Anderson. Virtualization performance: perspectives and challenges ahead. Operating Systems Review, 44(4):40–56, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mesnier:2011:DSS

[MA11] Michael P. Mesnier and Jason B. Akers. Differentiated storage services. Operating Systems Review, 45(1):45–53, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

MacLaren:1977:EHP

[Mac77] M. Donald MacLaren. Exception handling in PL/I. Operating Systems Review, 11(2):101–104, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Madsen:1981:CSSa

[Mad81a] Johannes Madsen. A computer system supporting data abstraction. Operating Systems Review, 15(1):45–72, January 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Madsen:1981:CSSb

[Mad81b] Johannes Madsen. A computer system supporting data abstraction. Operating Systems Review, 15(2):38–78, April 1981. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maffeis:1994:DIC

[Maf94] Silvano Maffeis. Design and implementation of a configurable mixed-media file system. Operating Systems Review, 28(4):4– 10, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mahoney:1994:OOF

[Mah94] Bill Mahoney. An “open” oriented file system. Operating Systems Review, 28(1):48–54, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mahmood:1998:TAD

[Mah98] Amjad Mahmood. Task allocation in distributed computing systems. Operating Systems Review, 32(3):21–29, July 1998. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maas:2016:THL

[MAHK16] Martin Maas, Krste Asanović, Tim Harris, and John Kubiatowicz. Taurus: a holistic language runtime system for coordinating distributed managed-language applications. Operating Systems Review, 50(2):457–471, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muller:2007:VMS

[MAK07] René Müller, Gustavo Alonso, and Donald Kossmann. A virtual machine for sensor networks. Operating Systems Review, 41(3):

145–158, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Malrait:2010:QOC

[Mal10] Luc Malrait. QoS-oriented control of server systems. Operating Systems Review, 44(3):59–64, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mao:2009:DDN

[Mao09] Yun Mao. On the declarativity of declarative networking. Operating Systems Review, 43(4):19–24, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Markatos:1997:VWS

[Mar97] Evangelos P. Markatos. Visualizing working sets. Operating Systems Review, 31(4):3–11, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Masuda:1977:EPL

[Mas77] Takashi Masuda. Effect of program localities on memory management strategies. Operating Systems Review, 11(5):117–124, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Massalin:1987:SLS

[Mas87] Henry Massalin. Superoptimizer: a look at the smallest program. Operating Systems Review, 21(4):122–126, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mysore:2006:IC

[MAS^{+}06] Shashidhar Mysore, Banit Agrawal, Navin Srivastava, Sheng-Chih Lin, Kaustav Banerjee, and Tim Sherwood. Introspective 3D chips. Operating Systems Review, 40(5):264–273, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Matthews:2004:CRR

[Mat04] Jeanna Neefe Matthews. The case for repeated research in operating systems. Operating Systems Review, 38(2):5–7, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Matthews:2006:OSR

[Mat06] Jeanna N. Matthews. Operating systems review: looking back and looking forward. Operating Systems Review, 40(1):1–2, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Matthews:2007:OSR

[Mat07] Jeanna N. Matthews. Operating systems review: year in review. Operating Systems Review, 41(1):1–2, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Matthews:2010:WPO

[Mat10] Jeanna Neefe Matthews. Workshop proceedings and other publications in operating system review. Operating Systems Review, 44(1):1, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Myers:1980:HIC

[MB80] G. J. Myers and B. R. S. Buckingham. A hardware implementation of capability-based addressing. Operating Systems Review, 14(4):13–25, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:1991:ECS

[MB91] Jeffrey C. Mogul and Anita Borg. The effect of context switches on cache performance. Operating Systems Review, 25(3S):75– 84, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maeda:1993:PSD

[MB93] Chris Maeda and Brian N. Bershad. Protocol service decomposition for high-performance networking. Operating Systems Review, 27(5):244–255, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Merkel:2006:BPC

[MB06] Andreas Merkel and Frank Bellosa. Balancing power consumption in multiprocessor systems. Operating Systems Review, 40 (4):403–414, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Merkel:2008:TAV

[MB08] Andreas Merkel and Frank Bellosa. Task activity vectors: a new metric for temperature-aware scheduling. Operating Systems Review, 42(4):1–12, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Malkhi:2012:PCF

[MBD^{+}12] Dahlia Malkhi, Mahesh Balakrishnan, John D. Davis, Vijayan Prabhakaran, and Ted Wobber. From Paxos to CORFU: a flashspeed shared log. Operating Systems Review, 46(1):47–51, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Moravan:2006:SNT

[MBM^{+}06] Michelle J. Moravan, Jayaram Bobba, Kevin E. Moore, Luke Yen, Mark D. Hill, Ben Liblit, Michael M. Swift, and David A. Wood. Supporting nested transactional memory in logTM. Operating Systems Review, 40(5):359–370, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mukkara:2016:WID

[MBS16] Anurag Mukkara, Nathan Beckmann, and Daniel Sanchez. Whirlpool: Improving dynamic cache management with static data classification. Operating Systems Review, 50(2):113–127, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Min:1991:ECB

[MC91] Sang L. Min and Jong-Deok Choi. An efficient cache-based access anomaly detection scheme. Operating Systems Review, 25 (3S):235–244, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:1996:CKU

[MC96] Chris J. Mitchell and Liqun Chen. Comments on the S/KEY user authentication scheme. Operating Systems Review, 30(4): 12–16, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maniatis:2011:STP

[MC11] Petros Maniatis and Byung-Gon Chun. Small trusted primitives for dependable systems. Operating Systems Review, 45(1):126– 141, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mishra:2006:TES

[MCC^{+}06] Mahim Mishra, Timothy J. Callahan, Tiberiu Chelcea, Girish Venkataramani, Seth C. Goldstein, and Mihai Budiu. Tartan: evaluating spatial computation for whole program execution. Operating Systems Review, 40(5):163–174, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mcdaniel:1977:MEA

[Mcd77] Gene Mcdaniel. Metric (extended abstract): A kernel instrumentation system for distributed environments. Operating Systems Review, 11(5):93–99, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McDonald:2000:DFM

[McD00] Ian McDonald. Distributed, flexible memory management in an operating system supporting quality of service. Operating Systems Review, 34(2):39, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mallik:2008:PMU

[MCD^{+}08] Arindam Mallik, Jack Cosgrove, Robert P. Dick, Gokhan Memik, and Peter Dinda. PICSEL: measuring user-perceived performance to control dynamic frequency scaling. Operating Systems Review, 42(2):70–79, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mohomed:2006:UUA

[MCdL06] Iqbal Mohomed, Jim Chengming Cai, and Eyal de Lara. URICA: Usage-awaRe Interactive Content Adaptation for mobile devices. Operating Systems Review, 40(4):345–358, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Misra:2017:ELT

[MCGL17] Pulkit A. Misra, Jeffrey S. Chase, Johannes Gehrke, and Alvin R. Lebeck. Enabling lightweight transactions with precision time. Operating Systems Review, 51(2):779–794, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muthitacharoen:2001:LBN

[MCM01] Athicha Muthitacharoen, Benjie Chen, and David Mazières. A low-bandwidth network file system. Operating Systems Review, 35(5):174–187, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:2007:ESR

[MCM07] Barton P. Miller, Gregory Cooksey, and Fredrick Moore. An empirical study of the robustness of MacOS applications using random testing. Operating Systems Review, 41(1):78–86, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McNamara:1977:TAD

[McN77] John E. McNamara. Technical aspects of data communication. Digital Press, Educational Services Dept., Bedford, MA, USA, 1977. ISBN 0-932376-01-0. xi + 387 pp. LCCN TK5105 .M33 1978.

McNamara:1982:TAD

[McN82] John E. McNamara. Technical aspects of data communication. Digital Press, Bedford, MA, USA, second edition, 1982. ISBN 0-932376-18-5. xi + 330 pp. LCCN TK5105 .M4 1982. US$32.00.

McNamara:1988:TAD

[McN88] John E. McNamara. Technical aspects of data communication. Digital Press, Rockport, MA, third edition, 1988. ISBN 1-55558-007-6. xi + 383 pp. LCCN TK5105 .M33 1988. US$42.00.

McMahan:2017:ASF

[MCN^{+}17] Joseph McMahan, Michael Christensen, Lawton Nichols, Jared Roesch, Sung-Yee Guo, Ben Hardekopf, and Timothy Sherwood. An architecture supporting formal and compositional binary analysis. Operating Systems Review, 51(2):177–191, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Moreto:2009:FQF

[MCR^{+}09] Miquel Moreto, Francisco J. Cazorla, Alex Ramirez, Rizos Sakellariou, and Mateo Valero. FlexDCP: a QoS framework for CMP architectures. Operating Systems Review, 43(2):86–96, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mellor-Crummey:1991:SC

[MCS91] John M. Mellor-Crummey and Michael L. Scott. Synchronization without contention. Operating Systems Review, 25(3S):269– 278, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mao:2016:RFR

[MCXS16] Junjie Mao, Yu Chen, Qixue Xiao, and Yuanchun Shi. RID: Finding reference count bugs with inconsistent path pair checking. Operating Systems Review, 50(2):531–544, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:1981:CTN

[MD81] James G. Mitchell and Jeremy Dion. A comparison of two network-based file servers. Operating Systems Review, 15(5): 45–46, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Montresor:2001:MDN

[MDB01] Alberto Montresor, Renzo Davoli, and Özalp Babaoğlu. Middleware for dependable network services in partitionable distributed systems. Operating Systems Review, 35(1):73–96, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mowry:1996:ACI

[MDK96] Todd C. Mowry, Angela K. Demke, and Orran Krieger. Automatic compiler-inserted I/O prefetching for out-of-core applications. Operating Systems Review, 30(SI):3–17, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maynard:1994:CCC

[MDO94] Ann Marie Grizzaffi Maynard, Colette M. Donnelly, and Bret R. Olszewski. Contrasting characteristics and cache performance of technical and multi-user commercial workloads. Operating Systems Review, 28(5):145–156, December 1994. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mattson:2000:CS

[MDR^{+}00] Peter Mattson, William J. Dally, Scott Rixner, Ujval J. Kapasi, and John D. Owens. Communication scheduling. Operating Systems Review, 34(5):82–92, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mickens:2009:SDW

[MdS09] James Mickens and Dilma da Silva. SOSP diversity workshop. Operating Systems Review, 43(4):90–91, December 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Monteiro:2008:AVM

[ME08] Steena D. S. Monteiro and Robert F. Erbacher. An authentication and validation mechanism for analyzing syslogs forensically. Operating Systems Review, 42(3):41–50, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mukherjee:1994:MII

[MEG94] Bodhisattwa Mukherjee, Greg Eisenhauer, and Kaushik Ghosh. A machine independent interface for lightweight threads. Operating Systems Review, 28(1):33–47, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mummert:1995:EWC

[MES95] L. B. Mummert, M. R. Ebling, and M. Satyanarayanan. Exploiting weak connectivity for mobile file access. Operating Systems Review, 29(5):143–155, December 1995. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Metzner:1982:SOS

[Met82] J. R. Metzner. Structuring operating systems literature for the graduate course. Operating Systems Review, 16(4):10–25, October 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Merlin:1975:RMS

[MF75] Philip M. Merlin and David J. Farber. Recoverability of modular systems. Operating Systems Review, 9(3):51–56, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McKenney:2020:RUL

[MFBWW20] Paul E. McKenney, Joel Fernandes, Silas Boyd-Wickizer, and Jonathan Walpole. RCU usage in the Linux kernel: Eighteen years later. Operating Systems Review, 54(1):47–63, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3421473.3421481.

Muck:2012:ICH

[MFGSP12] Tiago Rogerio Muck, Antonio Augusto Frohlich, Michael Gernoth, and Wolfgang Schroder-Preikschat. Implementing OS components in hardware using AOP. Operating Systems Review, 46(1):64–72, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Madden:2002:TTA

[MFHH02] Samuel Madden, Michael J. Franklin, Joseph M. Hellerstein, and Wei Hong. TAG: a Tiny AGgregation service for ad-hoc sensor networks. Operating Systems Review, 36(5S):131–146, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mashtizadeh:2017:TPD

[MGT^{+}17] Ali José Mashtizadeh, Tal Garfinkel, David Terei, David Mazieres, and Mendel Rosenblum. Towards practical defaulton multi-core record/replay. Operating Systems Review, 51(2): 693–708, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Madhavapeddy:2007:MCF

[MHD^{+}07] Anil Madhavapeddy, Alex Ho, Tim Deegan, David Scott, and Ripduman Sohan. Melange: creating a ‘functional’ Internet. Operating Systems Review, 41(3):101–114, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mills:1977:CFM

[Mil77] Philip M. Mills. Control functions for a multiprocessor architecture. Operating Systems Review, 11(1):26–40, January 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1978:UPO

[Mil78] Richard Miller. UNIX: a portable operating system? Operating Systems Review, 12(3):32–37, July 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1990:PDM

[Mil90] Frank W. Miller. Predictive deadline multi-processing. Operating Systems Review, 24(4):52–63, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1992:PMP

[Mil92] Frank W. Miller. The performance of a mixed priority realtime scheduling algorithm. Operating Systems Review, 26(4):5– 13, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:1996:JBF

[Mit96] James G. Mitchell. JavaOS: back to the future. Operating Systems Review, 30(SI):1, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:2000:MSN

[Mit00] Chris J. Mitchell. Making serial number based authentication robust against loss of state. Operating Systems Review, 34(3): 56–59, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maggio:1991:FSC

[MK91] Maria D. Maggio and David W. Krumme. A flexible system call interface for interprocessor communication in a distributed memory multicomputer. Operating Systems Review, 25(2):4–21, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Morris:1999:CMR

[MKJK99] Robert Morris, Eddie Kohler, John Jannotti, and M. Frans Kaashoek. The Click modular router. Operating Systems Review, 33(5):217–231, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Morris:2000:CMR

[MKJK00] Robert Morris, Eddie Kohler, John Jannotti, and M. Frans Kaashoek. The click modular router. Operating Systems Review, 34(2):24–25, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mazieres:1999:SKM

[MKKW99] David Mazières, Michael Kaminsky, M. Frans Kaashoek, and Emmett Witchel. Separating key management from file system security. Operating Systems Review, 33(5):124–139, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mazieres:2000:SKM

[MKKW00] David Mazières, Michael Kaminsky, M. Frans Kaashoek, and Emmett Witchel. Separating key management from file system security. Operating Systems Review, 34(2):19–20, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mai:2019:THP

[MKL^{+}19] Luo Mai, Alexandros Koliousis, Guo Li, Andrei-Octavian Brabete, and Peter Pietzuch. Taming hyper-parameters in deep learning systems. Operating Systems Review, 53(1):52–58, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McDonald:2008:SID

[MKY08] J. Todd McDonald, Yong C. Kim, and Alec Yasinsac. Software issues in digital forensics. Operating Systems Review, 42(3):29– 40, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mohan:1985:ECP

[ML85] C. Mohan and B. Lindsay. Efficient commit protocols for the tree of processes model of distributed transactions. Operating Systems Review, 19(2):40–52, April 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Myers:1997:DMI

[ML97] Andrew C. Myers and Barbara Liskov. A decentralized model for information flow control. Operating Systems Review, 31(5): 129–142, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mamrak:1983:PRD

[MLB83] Sandra A. Mamrak, Dennis Leinbaugh, and Toby S. Berk. A progress report on the Desperanto research project: software support for distributed processing. Operating Systems Review, 17(1):17–29, January 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Meyrowitz:1981:BAD

[MM81] Norman Meyrowitz and Margaret Moser. BRUWIN: An adaptable design strategy for window manager/virtual terminal systems. Operating Systems Review, 15(5):180–189, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mull:1991:EST

[MM91] Allison J. Mull and P. Tobin Maginnis. Evolutionary steps toward a distributed operating system: theory and implementation. Operating Systems Review, 25(4):4–13, October 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1992:SAO

[MM92] Barton P. Miller and Charles McDowell. Summary of ACM/ ONR workshop on parallel and distributed debugging. Operating Systems Review, 26(1):18–31, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1993:SAO

[MM93] Barton P. Miller and Charles McDowell. Summary of ACM/ ONR workshop on parallel and distributed debugging. Operating Systems Review, 27(4):8–23, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mysore:2008:UVF

[MMAS08] Shashidhar Mysore, Bita Mazloom, Banit Agrawal, and Timothy Sherwood. Understanding and visualizing full systems with data flow tomography. Operating Systems Review, 42(2):211– 221, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Moore:1996:CSM

[MMB96] A. W. Moore, A. J. McGregor, and J. W. Breen. A comparison of system monitoring methods, passive network monitoring and kernel instrumentation. Operating Systems Review, 30(1):16– 38, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muthitacharoen:2002:IRW

[MMGC02] Athicha Muthitacharoen, Robert Morris, Thomer M. Gil, and Benjie Chen. Ivy: a read/write peer-to-peer file system. Operating Systems Review, 36(5S):31–44, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McHugh:2008:PNF

[MMN08] John McHugh, Ron McLeod, and Vagishwari Nagaonkar. Passive network forensics: behavioural classification of network hosts based on connection patterns. Operating Systems Review, 42(3):99–111, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mueller:1983:NTM

[MMP83] Erik T. Mueller, Johanna D. Moore, and Gerald J. Popek. A nested transaction mechanism for LOCUS. Operating Systems Review, 17(5):71–89, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Markuze:2016:TIP

[MMT16] Alex Markuze, Adam Morrison, and Dan Tsafrir. True IOMMU protection from DMA attacks: When copy is faster than zero copy. Operating Systems Review, 50(2):249–262, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McKenney:2010:WGM

[MMTW10] Paul E. McKenney, Maged M. Michael, Josh Triplett, and Jonathan Walpole. Why the grass may not be greener on the other side: a comparison of locking vs. transactional memory. Operating Systems Review, 44(3):93–101, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Malkhi:2007:PRS

[MNP07] Dahlia Malkhi, Lev Novik, and Chris Purcell. P2P replica synchronization with vector sets. Operating Systems Review, 41(2): 68–74, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marzullo:1985:MTD

[MO85] Keith Marzullo and Susan Owicki. Maintaining the time in a distributed system. Operating Systems Review, 19(3):44–54, July 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:2006:EMB

[Mog06] Jeffrey C. Mogul. Emergent (mis)behavior vs. complex software systems. Operating Systems Review, 40(4):293–304, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:2008:PSH

[Mog08] Jeffrey C. Mogul. Policies for the SIGOPS Hall of Fame Award. Operating Systems Review, 42(3):132–135, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:2009:WWO

[Mog09] Jeffrey C. Mogul. WOWCS: the Workshop on Organizing Workshops, Conferences, and Symposia for Computer Systems. Operating Systems Review, 43(2):106–107, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mohan:1978:SRO

[Moh78] C. Mohan. Survey of recent operating systems research, designs and implementations. Operating Systems Review, 12(1):53–89, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Montgomery:1977:MSM

[Mon77] Warren A. Montgomery. Measurements of sharing in Multics. Operating Systems Review, 11(5):85–90, November 1977. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mooney:1982:UUI

[Moo82] James D. Mooney. USIM: a user interface manager. Operating Systems Review, 16(1):32–40, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mooney:1992:CAO

[Moo92] James D. Mooney. The CTRON approach to operating system support for software portability. Operating Systems Review, 26 (4):90–97, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mosberger:1993:MCM

[Mos93] David Mosberger. Memory consistency models. Operating Systems Review, 27(1):18–26, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mountassir:1996:DCD

[Mou96] H. Mountassir. Decidability of a class of dual communicating finite state machines. Operating Systems Review, 30(3):59–66, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Manning:1975:STP

[MP75] Eric Manning and R. W. Peebles. Segment transfer protocols for a homogeneous computer network. Operating Systems Review, 9(3):65–73, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1981:XOS

[MP81] Barton Miller and David Presotto. XOS: an operating system for the X-tree architecture. Operating Systems Review, 15(2): 21–32, April 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Malkawi:1985:CDM

[MP85] Mohammad Malkawi and Janek Patel. Compiler directed memory management policy for numerical programs. Operating Systems Review, 19(5):97–106, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Massalin:1989:TIO

[MP89] H. Massalin and C. Pu. Threads and input/output in the synthesis kernel. Operating Systems Review, 23(5):191–201, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muller:1991:HPM

[MP91] Keith Muller and Joseph Pasquale. A high performance multistructured file system design. Operating Systems Review, 25 (5):56–67, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Massalin:1992:LFMa

[MP92a] Henry Massalin and Calton Pu. A Lock-Free Multiprocessor OS Kernel. Operating Systems Review, 26(2):8, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Massalin:1992:LFMb

[MP92b] Henry Massalin and Calton Pu. A Lock-Free Multiprocessor OS Kernel. Operating Systems Review, 26(2):108, April 1992. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mosberger:1996:MPE

[MP96] David Mosberger and Larry L. Peterson. Making paths explicit in the Scout operating system. Operating Systems Review, 30 (SI):153–167, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Musuvathi:2002:CPA

[MPC^{+}02] Madanlal Musuvathi, David Y. W. Park, Andy Chou, Dawson R. Engler, and David L. Dill. CMC: a pragmatic approach to model checking real code. Operating Systems Review, 36(5S): 75–88, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Midorikawa:2008:ARB

[MPC08] Edson T. Midorikawa, Ricardo L. Piantola, and Hugo H. Cassettari. On adaptive replacement based on LRU with working area restriction algorithm. Operating Systems Review, 42(6):81– 92, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muir:2006:POP

[MPF^{+}06] Steve Muir, Larry Peterson, Marc Fiuczynski, Justin Cappos, and John Hartman. Privileged operations in the PlanetLab virtualised environment. Operating Systems Review, 40(1):75–88, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mislove:2006:EBO

[MPHD06] Alan Mislove, Ansley Post, Andreas Haeberlen, and Peter Druschel. Experiences in building and operating ePOST, a reliable peer-to-peer application. Operating Systems Review, 40(4):147– 159, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muller:2006:SPC

[MPLH06] Gilles Muller, Yoann Padioleau, Julia L. Lawall, and René Rydhof Hansen. Semantic patches considered helpful. Operating Systems Review, 40(3):90–92, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McCune:2008:HLC

[MPP^{+}08a] Jonathan M. McCune, Bryan Parno, Adrian Perrig, Michael K. Reiter, and Arvind Seshadri. How low can you go?: recommendations for hardware-supported minimal TCB code execution. Operating Systems Review, 42(2):14–25, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McCune:2008:FEI

[MPP^{+}08b] Jonathan M. McCune, Bryan J. Parno, Adrian Perrig, Michael K. Reiter, and Hiroshi Isozaki. Flicker: an execution infrastructure for TCB minimization. Operating Systems Review, 42(4):315–328, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Magenheimer:1987:IMD

[MPPZ87] Daniel J. Magenheimer, Liz Peters, Karl Pettis, and Dan Zuras. Integer multiplication and division on the HP precision architecture. Operating Systems Review, 21(4):90–99, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mayes:1995:ULT

[MQW95] K. R. Mayes, S. Quick, and B. C. Warboys. User-level threads on a general hardware interface. Operating Systems Review, 29 (4):57–62, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Makris:2007:DAU

[MR07] Kristis Makris and Kyung Dong Ryu. Dynamic and adaptive updates of non-quiescent subsystems in commodity operating system kernels. Operating Systems Review, 41(3):327–340, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:1987:PFE

[MRA87] J. Mogul, R. Rashid, and M. Accetta. The packer filter: an efficient mechanism for user-level network code. Operating Systems Review, 21(5):39–51, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Matthews:1997:IPL

[MRC^{+}97] Jeanna Neefe Matthews, Drew Roselli, Adam M. Costello, Randolph Y. Wang, and Thomas E. Anderson. Improving the performance of log-structured file systems with adaptive methods. Operating Systems Review, 31(5):238–251, December 1997. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Muralidharan:2016:AAC

[MRH^{+}16] Saurav Muralidharan, Amit Roy, Mary Hall, Michael Garland, and Piyush Rai. Architecture-adaptive code variant tuning. Operating Systems Review, 50(2):325–338, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mawhirter:2021:GHP

[MRH^{+}21] Daniel Mawhirter, Sam Reinehr, Connor Holmes, Tongping Liu, and Bo Wu. GraphZero: a high-performance subgraph matching system. Operating Systems Review, 55(1):21–37, June 2021. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

URL https://dl.acm.org/doi/10.1145/3469379.3469383.

Muniswamy-Reddy:2009:PFC

[MRS09] Kiran-Kumar Muniswamy-Reddy and Margo Seltzer. Provenance as first class cloud data. Operating Systems Review, 43 (4):11–16, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McCue:1991:SFT

[MS91a] Daniel L. McCue and Santosh K. Shrivastava. Structuring faulttolerant object systems for portability. Operating Systems Review, 25(2):118–121, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Moffett:1991:CDA

[MS91b] Jonathan D. Moffett and Morris S. Sloman. Content-dependent access control. Operating Systems Review, 25(2):63–70, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mummert:1994:VGC

[MS94] L. Mummert and M. Satyanarayanan. Variable granularity cache coherence. Operating Systems Review, 28(1):55–60, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:2000:SDS

[MS00] Donald Miller and Alan Skousen. The Sombrero distributed single address space operating system. Operating Systems Review, 34(2):37, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Martin:2000:TSA

[MSA^{+}00] Milo M. K. Martin, Daniel J. Sorin, Anatassia Ailamaki, Alaa R. Alameldeen, Ross M. Dickson, Carl J. Mauer, Kevin E. Moore, Manoj Plakal, Mark D. Hill, and David A. Wood. Timestamp snooping: an approach for extending SMPs. Operating Systems Review, 34(5):25–36, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mangione-Smith:1991:VRD

[MSAD91] William Mangione-Smith, Santosh G. Abraham, and Edward S. Davidson. Vector register design for polycyclic vector scheduling. Operating Systems Review, 25(3S):154–163, April 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mukherjee:2002:CSA

[MSB^{+}02] Shubhendu S. Mukherjee, Federico Silla, Peter Bannon, Joel Emer, Steve Lang, and David Webb. A comparative study of arbitration algorithms for the Alpha 21364 pipelined router. Operating Systems Review, 36(5):223–234, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Minnich:2006:RWK

[MSC^{+}06] Ronald G. Minnich, Matthew J. Sottile, Sung-Eun Choi, Erik Hendriks, and Jim McKie. Right-weight kernels: an off-the-shelf alternative to custom light-weight kernels. Operating Systems Review, 40(2):22–28, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mohan:1985:MDT

[MSF85] C. Mohan, R. Strong, and S. Finkelstein. Method for distributed transaction commit and recovery using Byzantine Agreement within clusters of processors. Operating Systems Review, 19(3): 29–43, July 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marsh:1991:FCU

[MSLM91] Brian D. Marsh, Michael L. Scott, Thomas J. LeBlanc, and Evangelos P. Markatos. First-class user-level threads. Operating Systems Review, 25(5):110–121, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Machanick:1998:HST

[MSP98] Philip Machanick, Pierre Salverda, and Lance Pompe. Hardwaresoftware trade-offs in a direct Rambus implementation of the RAMpage memory hierarchy. Operating Systems Review, 32(5): 105–114, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mercaldi:2006:IST

[MSP^{+}06] Martha Mercaldi, Steven Swanson, Andrew Petersen, Andrew Putnam, Andrew Schwerin, Mark Oskin, and Susan J. Eggers. Instruction scheduling for a tiled dataflow architecture. Operating Systems Review, 40(5):141–150, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Melliar-Smith:1977:SRR

[MSR77] P. M. Melliar-Smith and B. Randell. Software reliability: The role of programmed exception handling. Operating Systems Review, 11(2):95–100, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mullender:1985:DFS

[MT85] Sape J. Mullender and Andrew S. Tanenbaum. A distributed file service based on optimistic concurrency control. Operating Systems Review, 19(5):51–62, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McKinley:1996:QAL

[MT96] Kathryn S. McKinley and Olivier Temam. A quantitative analysis of loop nest locality. Operating Systems Review, 30(5):

94–104, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Martinez:2002:SSA

[MT02] José F. Martínez and Josep Torrellas. Speculative synchronization: applying thread-level speculation to explicitly parallel applications. Operating Systems Review, 36(5):18–29, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Merrifield:2017:PIE

[MT17] Timothy Merrifield and H. Reza Taheri. Performance implications of extended page tables on virtualized x86 processors. Operating Systems Review, 51(1):38–47, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mergen:2006:VHP

[MUKX06] Mark F. Mergen, Volkmar Uhlig, Orran Krieger, and Jimi Xenidis. Virtualization for high-performance computing. Operating Systems Review, 40(2):8–11, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mullender:1987:RSE

[Mul87] Sape Mullender. Report on the Second European SIGOPS Workshop “Making Distributed Systems Work”. Operating Systems Review, 21(1):49–84, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mullender:1986:DMM

[MV86] Sape J. Mullender and Paul M. B. Vitanyi. Distributed matchmaking for processes in computer networks. Operating Systems Review, 20(2):54–64, April 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Madhyastha:2006:OOA

[MVKA06] Harsha V. Madhyastha, Arun Venkataramani, Arvind Krishnamurthy, and Thomas Anderson. Oasis: an overlay-aware network stack. Operating Systems Review, 40(1):41–48, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Malkhi:2013:WRL

[MvR13] Dahlia Malkhi and Robbert van Renesse. Workshop report on LADIS 2012. Operating Systems Review, 47(1):1–2, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McQuillan:1975:SCH

[MW75] John M. McQuillan and David C. Walden. Some considerations for a high performance message-based interprocess communication system. Operating Systems Review, 9(3):77–86, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marzullo:1991:MRT

[MW91] Keith Marzullo and Mark Wood. Making real-time reactive systems reliable. Operating Systems Review, 25(1):45–48, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Marzullo:1992:TBR

[MW92] Keith Marzullo and Mark D. Wood. A Toolkit for Building Reactive Systems or A Minute on Meta. Operating Systems Review, 26(2):23, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McKenney:2008:ITL

[MW08] Paul E. McKenney and Jonathan Walpole. Introducing technology into the Linux kernel: a case study. Operating Systems Review, 42(5):4–17, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mogul:2009:CSR

[MW09] Jeffrey C. Mogul and Jay J. Wylie. Computer systems research at HP Labs. Operating Systems Review, 43(1):8–9, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mengshu:2005:TMP

[MXXC05] Hou Mengshu, Lu Xianliang, Zhou Xu, and Zhan Chuan. A trust model of P2P system based on confirmation theory. Operating Systems Review, 39(1):56–62, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:1998:FPH

[MY98] Chris J. Mitchell and Chan Yeob Yeun. Fixing a problem in the Helsinki protocol. Operating Systems Review, 32(4):21–24, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Maggi:2008:SIF

[MZI08] Federico Maggi, Stefano Zanero, and Vincenzo Iozzo. Seeing the invisible: forensic uses of anomaly detection and machine learning. Operating Systems Review, 42(3):51–58, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mao:2002:LRM

[MZWZ02] Yun Mao, Youhui Zhang, Dongsheng Wang, and Weimin Zheng. LND: a reliable multi-tier storage device in NOW. Operating Systems Review, 36(1):70–80, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Naimi:1993:DAK

[Nai93] Mohamed Naimi. Distributed algorithm for K-entries to critical section based on the directed graphs. Operating Systems Review, 27(4):67–75, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Naimi:1996:DME

[Nai96] Mohamed Naimi. Distributed mutual exclusion on hypercubes. Operating Systems Review, 30(3):46–51, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nicoara:2008:CSE

[NAR08] Angela Nicoara, Gustavo Alonso, and Timothy Roscoe. Controlled, systematic, and efficient code replacement for running Java programs. Operating Systems Review, 42(4):233–246, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Natarajan:1980:AAT

[Nat80] N. Natarajan. Atomic actions and timestamps. Operating Systems Review, 14(2):25–27, April 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1977:CFS

[NB77] R. M. Needham and A. D. Birrell. The CAP filing system. Operating Systems Review, 11(5):11–16, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nehmer:1991:FTA

[NB91] J. Nehmer and T. Becker. A fault tolerance approach for distributed ISDN control systems. Operating Systems Review, 25 (2):126–129, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nicholas:2000:OTD

[NB00] Tyrone Nicholas and Jerzy A. Barchanski. Overview of TOS: a distributed educational operating system in Java. Operating Systems Review, 34(1):2–10, January 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nellans:2009:EMC

[NBB09] David Nellans, Rajeev Balasubramonian, and Erik Brunvand. OS execution on multi-cores: is out-sourcing worthwhile? Operating Systems Review, 43(2):104–105, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nelson:2020:NSS

[NBK^{+}20] Luke Nelson, James Bornholt, Arvind Krishnamurthy, Emina Torlak, and Xi Wang. Noninterference specifications for secure systems. Operating Systems Review, 54(1):31–39, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3421473.3421478.

Nicol:1987:OSD

[NBW87] John R. Nicol, Gordon S. Blair, and Jonathan Walpole. Operating system design: towards a holistic approach? Operating Systems Review, 21(1):11–19, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nostro:2014:ITA

[NCBB14] Nicola Nostro, Andrea Ceccarelli, Andrea Bondavalli, and Francesco Brancati. Insider threat assessment: a model-based methodology. Operating Systems Review, 48(2):3–12, July 2014.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nightingale:2005:SED

[NCF05] Edmund B. Nightingale, Peter M. Chen, and Jason Flinn. Speculative execution in a distributed file system. Operating Systems Review, 39(5):191–205, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nassiffe:2012:OQS

[NCL12] Ríad Nassiffe, Eduardo Camponogara, and George Lima. Optimizing quality of service in real-time systems under energy constraints. Operating Systems Review, 46(1):82–92, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1972:HDF

[Nee72] R. M. Needham. Handling difficult faults in operating systems. Operating Systems Review, 6(1/2):55–57, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1977:CPI

[Nee77] R. M. Needham. The CAP project — an interim evaluation. Operating Systems Review, 11(5):17–22, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1979:ACA

[Nee79] Roger M. Needham. Adding capability access to conventional file servers. Operating Systems Review, 13(1):3–4, January 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Negishi:2000:TCS

[Neg00] Yasushi Negishi. Tuplink: a communication system for PDAs and micro-devices. Operating Systems Review, 34(2):37, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nessett:1982:IPD

[Nes82] D. M. Nessett. Identifier protection in a distributed operating system. Operating Systems Review, 16(1):26–31, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nessett:1990:CBA

[Nes90] Dan M. Nessett. A critique of the Burrows, Abadi and Needham logic. Operating Systems Review, 24(2):35–38, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Neuman:1989:NCL

[Neu89] B. C. Neuman. The need for closure in large distributed systems. Operating Systems Review, 23(4):28–30, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Neuman:1992:VSP

[Neu92] B. Clifford Neuman. The Virtual System Project. Operating Systems Review, 26(2):17, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Neugebauer:2000:ULP

[Neu00] Rolf Neugebauer. A Unix-like personality supporting quality-of-service. Operating Systems Review, 34(2):39, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Newton:1979:DPD

[New79] Glen Newton. Deadlock prevention, detection, and resolution: an annotated bibliography. Operating Systems Review, 13(2): 33–44, April 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ng:1999:CST

[Ng99] Siaw-Lynn Ng. Comments on “On the Security of Three-Party Cryptographic Protocols” by Xu, Zhang, Zhu. Operating Systems Review, 33(3):5–6, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [XZZ98].

Nagarajan:2009:RMM

[NG09] Vijay Nagarajan and Rajiv Gupta. Runtime monitoring on multicores via OASES. Operating Systems Review, 43(2):15– 24, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nalli:2017:APM

[NHH^{+}17] Sanketh Nalli, Swapnil Haria, Mark D. Hill, Michael M. Swift, Haris Volos, and Kimberly Keeton. An analysis of persistent memory use with WHISPER. Operating Systems Review, 51 (2):135–148, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1983:HCS

[NHM83] R. M. Needham, A. J. Herbert, and J. G. Mitchell. How to connect stable memory to a computer. Operating Systems Review, 17(1):16, January 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nichols:1987:UIW

[Nic87] D. Nichols. Using idle workstations in a shared computing environment. Operating Systems Review, 21(5):5–12, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Navarro:2002:PTO

[NIDC02] Juan Navarro, Sitararn Iyer, Peter Druschel, and Alan Cox. Practical, transparent operating system support for superpages. Operating Systems Review, 36(5S):89–104, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nieh:1995:SPS

[NL95] Jason Nieh and Monica S. Lam. SMART: a processor scheduler for multimedia applications. Operating Systems Review, 29 (5):233, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Necula:1996:SKE

[NL96] George C. Necula and Peter Lee. Safe kernel extensions without run-time checking. Operating Systems Review, 30(SI):229–243, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nieh:1997:DIE

[NL97] Jason Nieh and Monica S. Lam. The design, implementation and evaluation of SMART: a scheduler for multimedia applications. Operating Systems Review, 31(5):184–197, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nelson:1995:HAS

[NLO95] M. N. Nelson, M. Linton, and S. Owicki. A highly available scalable ITV system. Operating Systems Review, 29(5):54–67, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nanda:2000:MPR

[NMS^{+}00] Ashwini Nanda, Kwok-Ken Mak, Krishnan Sugarvanam, Ramendra K. Sahoo, Vijayaraghavan Soundarararjan, and T. Basil Smith. MemorIES3: a programmable, real-time hardware emulation tool for multiprocessor server design. Operating Systems Review, 34(5):37–48, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

NBS:1975:PPS

[NN75] National Bureau of Standards and National Science Foundation. A preliminary prospectus for a software engineering handbook. Operating Systems Review, 9(2):10–13, April 1975. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nguyen:2017:WSP

[NP17] Donald Nguyen and Keshav Pingali. What scalable programs need from transactional memory. Operating Systems Review, 51 (2):105–118, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nakao:2006:SRO

[NPB06] Akihiro Nakao, Larry Peterson, and Andy Bavier. Scalable routing overlay networks. Operating Systems Review, 40(1):49–61, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Narayanasamy:2006:RSM

[NPC06] Satish Narayanasamy, Cristiano Pereira, and Brad Calder. Recording shared memory dependencies using strata. Operating Systems Review, 40(5):229–240, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nightingale:2008:PSC

[NPCF08] Edmund B. Nightingale, Daniel Peek, Peter M. Chen, and Jason Flinn. Parallelizing security checks on commodity hardware.

Operating Systems Review, 42(2):308–318, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Noureddine:2013:REM

[NRS13] Adel Noureddine, Romain Rouvoy, and Lionel Seinturier. A review of energy measurement approaches. Operating Systems Review, 47(3):42–49, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1987:AR

[NS87] R. M. Needham and M. D. Schroeder. Authentication revisited. Operating Systems Review, 21(1):7, January 1987. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Neuman:1993:NUT

[NS93] B. Clifford Neuman and Stuart G. Stubblebine. A note on the use of timestamps as nonces. Operating Systems Review, 27(2): 10–14, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nathuji:2007:VCP

[NS07] Ripal Nathuji and Karsten Schwan. VirtualPower: coordinated power management in virtualized enterprise systems. Operating Systems Review, 41(6):265–278, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nowatzki:2016:ABS

[NS16] Tony Nowatzki and Karthikeyan Sankaralingam. Analyzing behavior specialized acceleration. Operating Systems Review, 50 (2):697–711, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Narang:2011:PDM

[NSKS11] Ankur Narang, Abhinav Srivastava, Naga Praveen Kumar Katta, and Rudrapatna K. Shyamasundar. Performance driven multi-objective distributed scheduling for parallel computations. Operating Systems Review, 45(2):14–27, July 2011. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Noble:1997:AAA

[NSN^{+}97] Brian D. Noble, M. Satyanarayanan, Dushyanth Narayanan, James Eric Tilton, Jason Flinn, and Kevin R. Walker. Agile application-aware adaptation for mobility. Operating Systems Review, 31(5):276–287, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nygren:2010:ANP

[NSS10] Erik Nygren, Ramesh K. Sitaraman, and Jennifer Sun. The Akamai network: a platform for high-performance Internet applications. Operating Systems Review, 44(3):2–19, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nygren:2010:NSR

[NSW10] Erik Nygren, Ramesh K. Sitaraman, and Joel Wein. Networked systems research at Akamai. Operating Systems Review, 44(3): 1, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Naas:2021:EUK

[NTC^{+}21] Mohammed Islam Naas, François Trahay, Alexis Colin, Pierre Olivier, Stéphane Rubini, Frank Singhoff, and Jalil Boukhobza. EZIOTracer: Unifying kernel and user space I/O tracing for data-intensive applications. Operating Systems Review, 55(1): 88–98, June 2021. CODEN OSRED8. ISSN 0163-5980 (print),

1943-586X (electronic). URL https://dl.acm.org/doi/10.
1145/3469379.3469391.

Nicolas:2022:SSL

[NTHAB22] Louis-Marie Nicolas, Luis Thomas, Yassine Hadjadj-Aoul, and Jalil Boukhobza. SLRL: a simple least remaining lifetime file eviction policy for HPC multi-tier storage systems. Operating Systems Review, 56(1):70–76, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3544497.3544509.

Nutt:1974:ICS

[Nut74] Gary J. Nutt. An implementation of a computer simulation system. Operating Systems Review, 8(3):6–7, July 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nutt:1994:BRC

[Nut94a] Gary J. Nutt. Book review: Coloured Petri Nets: Basic Concepts, Analysis Methods and Practical Use (volume 1) by Kurt Jensen: (Springer-Verlag, 1992). Operating Systems Review, 28 (1):1–2, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nuttall:1994:BSS

[Nut94b] Mark Nuttall. A brief survey of systems providing process or object migration facilities. Operating Systems Review, 28(4):64– 80, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nieh:2006:ETO

[NV06] Jason Nieh and Chris Vaill. Experiences teaching operating systems using virtual platforms and Linux. Operating Systems Review, 40(2):100–104, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Needham:1977:CCC

[NW77] R. M. Needham and R. D. H. Walker. The Cambridge CAP computer and its protection system. Operating Systems Review, 11(5):1–10, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nelson:1987:CSN

[NWO87] M. Nelson, B. Welch, and J. Ousterhout. Caching in the Sprite network file system. Operating Systems Review, 21(5):3– 4, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nijim:2005:PAA

[NXQ05] Mais Nijim, Tao Xie, and Xiao Qin. Performance analysis of an admission controller for CPU- and I/O-intensive applications in self-managing computer systems. Operating Systems Review, 39 (4):37–45, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ottoni:2008:COG

[OA08] Guilherme Ottoni and David I. August. Communication optimizations for global multi-threaded instruction scheduling. Operating Systems Review, 42(2):222–232, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ousterhout:2009:CRS

[OAE^{+}09] John Ousterhout, Parag Agrawal, David Erickson, Christos Kozyrakis, Jacob Leverich, David Mazières, Subhasish Mitra, Aravind Narayanan, Guru Parulkar, Mendel Rosenblum, Stephen M. Rumble, Eric Stratmann, and Ryan Stutsman. The case for RAMClouds: scalable high-performance storage entirely in DRAM. Operating Systems Review, 43(4):92–105, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olagunju:1986:EPI

[OB86] Amos O. Olagunju and Elvis Borders. Emulators; prospective instruments for instruction in systems programming. Operating Systems Review, 20(4):16–24, October 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ostrowski:2010:SAL

[OB10] Krzysztof Ostrowski and Ken Birman. Storing and accessing live mashup content in the cloud. Operating Systems Review, 44(2): 7–11, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ouarnoughi:2016:ICP

[OBSR16] Hamza Ouarnoughi, Jalil Boukhobza, Frank Singhoff, and Stéphane Rubini. Integrating I/Os in Cloudsim for performance and energy estimation. Operating Systems Review, 50(3):27–36, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ong:2000:IVM

[OCF00] Joon Suan Ong, Yvonne Coady, and Michael J. Feeley. Integrating virtual memory with user-level network communication. Operating Systems Review, 34(2):36–37, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oh:2014:IPL

[OCLN14] Yongseok Oh, Jongmoo Choi, Donghee Lee, and Sam H. Noh. Improving performance and lifetime of the SSD RAID-based host cache through a log-structured approach. Operating Systems Review, 48(1):90–97, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ousterhout:1989:BBC

[OD89] John Ousterhout and Fred Douglis. Beating the I/O bottleneck: a case for log-structured file systems. Operating Systems Review, 23(1):11–28, January 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ousterhout:1985:TDA

[ODH^{+}85] John K. Ousterhout, Hervé Da Costa, David Harrison, John A. Kunze, Mike Kupfer, and James G. Thompson. A trace-driven analysis of the UNIX 4.2 BSD file system. Operating Systems Review, 19(5):15–24, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oestreicher:1991:SRG

[Oes91] Dan Oestreicher. A simple reliable globally-ordered broadcast service. Operating Systems Review, 25(4):66–76, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oestreicher:2001:ECJ

[Oes01] Dan Oestreicher. Experience with a commercial Java implementation of group communication using reliable multicast. Operating Systems Review, 35(4):21–31, October 2001. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oyamada:2016:BSC

[OFB16] Márcio Oyamada, Antônio Augusto Fröhlich, and Leandro Becker. 5th Brazilian Symposium on Computing System Engineering. Operating Systems Review, 50(1):17, January 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ozkasap:2006:EBA

[ÖGA06] Öznur Özkasap, Zülküf Genç, and Emre Atsan. Epidemic-based approaches for reliable multicast in mobile ad hoc networks. Operating Systems Review, 40(3):73–79, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olson:2017:CGM

[OHW17] Lena E. Olson, Mark D. Hill, and David A. Wood. Crossing guard: Mediating host-accelerator coherence interactions. Operating Systems Review, 51(2):163–176, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ogata:2002:BFO

[OKN02] Kazunori Ogata, Hideaki Komatsu, and Toshio Nakatani. Bytecode fetch optimization for a Java interpreter. Operating Systems Review, 36(5):58–67, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oplinger:2002:ESR

[OL02] Jeffrey Oplinger and Monica S. Lam. Enhancing software reliability with speculative threads. Operating Systems Review, 36 (5):184–196, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oliver:1990:PDD

[Oli90] Roger Oliver. Protection in a distributed document processing system. Operating Systems Review, 24(2):56–65, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ong:2002:UVC

[OLLY02] T. M. Ong, T. M. Lim, B. S. Lee, and C. K. Yeo. Unicorn: voluntary computing over Internet. Operating Systems Review, 36(2):36–51, April 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oki:1985:ROS

[OLS85] Brian M. Oki, Barbara H. Liskov, and Robert W. Scheifler. Reliable object storage to support atomic actions. Operating Systems Review, 19(5):147–159, December 1985. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olszewski:2007:JIN

[OMCB07] Marek Olszewski, Keir Mierle, Adam Czajkowski, and Angela Demke Brown. JIT instrumentation: a novel approach to dynamically instrument operating systems. Operating Systems Review, 41(3):3–16, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

OToole:1993:CCG

[ONG93] James O’Toole, Scott Nettles, and David Gifford. Concurrent compacting garbage collection of a persistent heap. Operating Systems Review, 27(5):161–174, December 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olukotun:1996:CSC

[ONH^{+}96] Kunle Olukotun, Basem A. Nayfeh, Lance Hammond, Ken Wilson, and Kunyung Chang. The case for a single-chip multiprocessor. Operating Systems Review, 30(5):2–11, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Opderbeck:1975:ECP

[Opd75] Holger Opderbeck. On the efficiency of control procedures for computer communication networks. Operating Systems Review, 9(3):93–96, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oki:1993:IBA

[OPSS93] Brian Oki, Manfred Pfluegl, Alex Siegel, and Dale Skeen. The Information Bus: an architecture for extensible distributed systems. Operating Systems Review, 27(5):58–68, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Otway:1987:ETM

[OR87] Dave Otway and Owen Rees. Efficient and timely mutual authentication. Operating Systems Review, 21(1):8–10, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Oestreicher:1980:SES

[OS80] Dan Oestreicher and J. I. Strauss. A set of efficient semaphoring instructions. Operating Systems Review, 14(3):36–45, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

OShea:1992:RRP

[O’S92] G. O’Shea. Redundant rights in protection systems. Operating Systems Review, 26(3):27–30, July 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Osman:2002:DIZ

[OSSN02] Steven Osman, Dinesh Subhraveti, Gong Su, and Jason Nieh. The design and implementation of Zap: a system for migrating computing environments. Operating Systems Review, 36(5S):

361–376, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olle:1983:ISD

[OST83] T. William Olle, H. G. (Henk G.) Sol, and C. J. (Colin J.) Tully, editors. Information systems design methodologies: a feature analysis: Proceedings of the IFIP WG 8.1 Working Conference on Feature Analysis of Information Systems Design Methodologies, York, U.K., 5–7 July, 1983. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86705-8 (U.S.). LCCN QA76.9.S88 I35 1983.

Olle:1982:ISD

[OSV82] T. William Olle, H. G. (Henk G.) Sol, and A. A. Verrijn Stuart, editors. Information systems design methodologies: a comparative review: proceedings of the IFIP WG 8.1 Working Conference on Comparative Review of Information Systems Design Methodologies, Noordwijkerhout, The Netherlands, 10–14 May 1982. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86407-5. LCCN Z699.A1 I37 1982.

Olle:1986:ISD

[OSV86] T. William Olle, H. G. (Henk G.) Sol, and A. A. Verrijn Stuart, editors. Information systems design methodologies: improving the practice: proceedings of the IFIP WG 8.1 Working Conference on Comparative Review of Information Systems Design Methodologies, Improving the Practice, Noordwijkerhout, The Netherlands, 5–7 May, 1986. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70014-5 (U.S.). LCCN QA76.9.S88 I345 1986.

Oikawa:1995:RDU

[OT95] Shuichi Oikawa and Hideyuki Tokuda. Reflection of developing user-level real-time thread packages. Operating Systems Review, 29(4):63–76, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ott:2018:SDI

[Ott18] David E. Ott. Software defined infrastructure: Rethinking cybersecurity with a more capable toolset. Operating Systems Review, 52(1):129–133, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ousterhout:1981:MDO

[Ous81] John K. Ousterhout. Medusa, a distributed operating system, volume 1 of Computer science. Data bases and distributed systems. UMI Research Press, Ann Arbor, Mich., 1981. ISBN 0-8357-1201-X. xii + 139 pp. LCCN QA76.6 .O92 1981.

Ong:2006:KLS

[OVS^{+}06] Hong Ong, Jeffrey Vetter, R. Scott Studham, Collin McCurdy, Bruce Walker, and Alan Cox. Kernel-level single system image for petascale computing. Operating Systems Review, 40(2):50– 54, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Owen:1984:DCI

[Owe84] Kenneth Owen. Data communications: IFIP’s international “network” of experts. Operating Systems Review, 18(1):20–26, January 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pu:1995:OIS

[PAB^{+}95] C. Pu, T. Autrey, A. Black, C. Consel, C. Cowan, J. Inouye, L. Kethana, J. Walpole, and K. Zhang. Optimistic incremental specialization: streamlining a commercial operating system. Operating Systems Review, 29(5):314–321, December 1995. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pai:1998:LAR

[PAB^{+}98] Vivek S. Pai, Mohit Aron, Gaurov Banga, Michael Svendsen, Peter Druschel, Willy Zwaenepoel, and Erich Nahum. Localityaware request distribution in cluster-based network servers. Operating Systems Review, 32(5):205–216, December 1998. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2016:ATC

[PAM^{+}16] Jongse Park, Emmanuel Amaro, Divya Mahajan, Bradley Thwaites, and Hadi Esmaeilzadeh. AxGames: Towards crowdsourcing quality target determination in approximate computing. Operating Systems Review, 50(2):623–636, June 2016. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parnas:1978:NPN

[Par78] David L. Parnas. The non-problem of nested monitor calls. Operating Systems Review, 12(1):12–18, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pasquale:1992:PS

[Pas92] Joseph Pasquale. Project Sequoia 2000. Operating Systems Review, 26(2):14, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:2002:MSE

[Pat02a] D. Patiyoot. Migration /evolution of security towards wireless ATM. Operating Systems Review, 36(1):23–30, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:2002:SIW

[Pat02b] Danai Patiyoot. Security issues for wireless ATM networks. Operating Systems Review, 36(1):31–57, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Payne:1977:FTE

[Pay77] A. J. Payne. A formalised technique for expressing system exercisers. Operating Systems Review, 11(3):8–12, July 1977. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pardyak:1996:DBE

[PB96] Przemysław Pardyak and Brian N. Bershad. Dynamic binding for an extensible system. Operating Systems Review, 30(SI):201– 212, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Portokalidis:2008:EID

[PB08] Georgios Portokalidis and Herbert Bos. Eudaemon: involuntary and on-demand emulation against zero-day exploits. Operating Systems Review, 42(4):287–299, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Princehouse:2009:CPG

[PB09] Lonnie Princehouse and Ken Birman. Code-partitioning gossip. Operating Systems Review, 43(4):40–44, December 2009.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Prabhakaran:2005:IFS

[PBA^{+}05] Vijayan Prabhakaran, Lakshmi N. Bairavasundaram, Nitin Agrawal, Haryadi S. Gunawi, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. IRON file systems. Operating Systems Review, 39(5):206–220, December 2005. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pariag:2007:CPW

[PBH^{+}07] David Pariag, Tim Brecht, Ashif Harji, Peter Buhr, Amol Shukla, and David R. Cheriton. Comparing the performance of Web server architectures. Operating Systems Review, 41(3): 231–243, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peisert:2008:CFF

[PBM08] Sean Peisert, Matt Bishop, and Keith Marzullo. Computer forensics in forensis. Operating Systems Review, 42(3):112–122, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Purandare:2022:AWC

[PBM22] Devashish R. Purandare, Daniel Bittman, and Ethan L. Miller. Analysis and workload characterization of the CERN EOS storage system. Operating Systems Review, 56(1):55–61, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3544497.
3544507.

Peter:2008:SES

[PBR^{+}08] Simon Peter, Andrew Baumann, Timothy Roscoe, Paul Barham, and Rebecca Isaacs. 30 seconds is not enough!: a study of operating system timer usage. Operating Systems Review, 42 (4):205–218, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pelleg:2008:VBD

[PBYH^{+}08] Dan Pelleg, Muli Ben-Yehuda, Rick Harper, Lisa Spainhower, and Tokunbo Adeshiyan. Vigilant: out-of-band detection of failures in virtual machines. Operating Systems Review, 42(1):26– 31, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pruitt:1975:ART

[PC75] J. L. Pruitt and W. W. Case. Architecture of a real time operating system. Operating Systems Review, 9(5):51–59, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Powell:1991:FTD

[PCD91] David Powell, Marc Chérèque, and David Drackley. Faulttolerance in Delta-4. Operating Systems Review, 25(2):122–125, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pillai:2014:TEP

[PCH^{+}14] Thanumalayan Sankaranarayana Pillai, Vijay Chidambaram, Joo-Young Hwang, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Towards efficient, portable application-level consistency. Operating Systems Review, 48(1):26–31, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pei:2000:SKC

[PCP00] Pengjun Pei, Guohua Cui, and Kun Peng. On a session key compromise problem in [KC95] protocol. Operating Systems Review, 34(3):16–18, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peacock:1989:DAL

[Pea89] J. Kent Peacock. Deadlock avoidance in loosely-coupled multiprocessors with finite buffer pools. Operating Systems Review, 23(2):20–24, April 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Philbin:1996:TSC

[PEA^{+}96] James Philbin, Jan Edler, Otto J. Anshus, Craig C. Douglas, and Kai Li. Thread scheduling for cache locality. Operating Systems Review, 30(5):60–71, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Penry:2009:MDS

[Pen09] David A. Penry. Multicore diversity: a software developer’s nightmare. Operating Systems Review, 43(2):100–101, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Perl:1992:PAC

[Per92] Sharon E. Perl. Performance Assertion Checking. Operating Systems Review, 26(2):25, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peterson:1976:RCF

[Pet76] James L. Peterson. Referee coordination for the fifth Symposium on Operating Systems Principles. Operating Systems Review, 10(1):7–16, January 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peterson:1993:LON

[Pet93] Larry L. Peterson. Life on the OS/network boundary. Operating Systems Review, 27(2):94–98, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peng:2002:MPC

[PFGD02] Bi Peng, Xie Fei, Yang Guangwen, and Wang Dingxing. A multiprotocol cross-domain communication model for metacomputing systems. Operating Systems Review, 36(2):52–63, April 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Plehn:2022:DAC

[PFK^{+}22] Julius Plehn, Anna Fuchs, Michael Kuhn, Jakob Lüttgau, and Thomas Ludwig. Data-aware compression for HPC using machine learning. Operating Systems Review, 56(1):62–69, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3544497.3544508.

Popek:1973:FRV

[PG73] Gerald J. Popek and Robert P. Goldberg. Formal requirements for virtualizable third generation architectures. Operating Systems Review, 7(4):121, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Panadiwal:1996:HPA

[PG96] Rajmohan Panadiwal and Andrzej M. Goscinski. A high performance and adaptive commit protocol for a distributed environment. Operating Systems Review, 30(3):52–58, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pinilla:2003:JPI

[PG03a] Ruben Pinilla and Marisa Gil. JVM: platform independent vs. performance dependent. Operating Systems Review, 37(2):44– 56, April 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pinilla:2003:UJT

[PG03b] Ruben Pinilla and Marisa Gil. ULT: a Java threads model for platform independent execution. Operating Systems Review, 37 (4):48–62, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patel:2006:BGA

[PG06] Jay A. Patel and Indranil Gupta. Bridging the gap: augmenting centralized systems with P2P technologies. Operating Systems Review, 40(3):14–17, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Prasad:2016:PMR

[PG16] Aravinda Prasad and K. Gopinath. Prudent memory reclamation in procrastination-based synchronization. Operating Systems Review, 50(2):99–112, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patterson:1995:IPC

[PGG^{+}95] R. H. Patterson, G. A. Gibson, E. Ginting, D. Stodolsky, and J. Zelenka. Informed prefetching and caching. Operating Systems Review, 29(5):79–95, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patterson:1993:SRR

[PGS93] R. Hugo Patterson, Garth A. Gibson, and M. Satyanarayanan. A status report on research in transparent informed prefetching. Operating Systems Review, 27(2):21–34, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pattabiraman:2008:SPC

[PGZ08] Karthik Pattabiraman, Vinod Grover, and Benjamin G. Zorn. Samurai: protecting critical data in unsafe languages. Operating Systems Review, 42(4):219–232, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1977:NDE

[PHL^{+}77] G. J. Popek, J. J. Horning, B. W. Lampson, J. G. Mitchell, and R. L. London. Notes on the design of Euclid. Operating Systems Review, 11(2):11–18, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peterson:1989:RXK

[PHOA89] L. Peterson, N. Hutchinson, S. O’Malley, and M. Abbott. RPC in the x-Kernel: evaluating new design techniques. Operating Systems Review, 23(5):91–101, December 1989. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peir:1996:ICP

[PHYO96] Jih-Kwon Peir, Windsor W. Hsu, Honesty Young, and Shauchi Ong. Improving cache performance with balanced tag and data paths. Operating Systems Review, 30(5):268–278, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Piotrowski:1989:FNE

[Pio89] Walter G. Piotrowski. Are file names enough? Operating Systems Review, 23(4):26–27, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patwardhan:2006:DTS

[PJDL06] Jaidev P. Patwardhan, Vijeta Johri, Chris Dwyer, and Alvin R. Lebeck. A defect tolerant self-organizing nanoscale SIMD architecture. Operating Systems Review, 40(5):241–251, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1975:PVM

[PK75] Gerald J. Popek and Charles S. Kline. The PDP-11 virtual machine architecture: A case study. Operating Systems Review, 9(5):97–105, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Perkovic:1996:ODR

[PK96] Dejan Perkovic and Peter J. Keleher. Online data-race detection via coherency guarantees. Operating Systems Review, 30(SI):47– 57, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Povzner:2008:EGD

[PKB^{+}08] Anna Povzner, Tim Kaldewey, Scott Brandt, Richard Golding, Theodore M. Wong, and Carlos Maltzahn. Efficient guaranteed disk request scheduling with Fahrrad. Operating Systems Review, 42(4):13–25, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Prabhakar:2016:GCH

[PKB^{+}16] Raghu Prabhakar, David Koeplinger, Kevin J. Brown, HyoukJoong Lee, Christopher De Sa, Christos Kozyrakis, and Kunle Olukotun. Generating configurable hardware from parallel patterns. Operating Systems Review, 50(2):651–665, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Povzner:2009:AAE

[PKM^{+}09] Anna Povzner, Kimberly Keeton, Arif Merchant, Charles B. Morrey III, Mustafa Uysal, and Marcos K. Aguilera. Autograph: automatically extracting workflow file signatures. Operating Systems Review, 43(1):76–83, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pollack:1981:IOF

[PKW81] Fred J. Pollack, Kevin C. Kahn, and Roy M. Wilkinson. The iMAX-432 object filing system. Operating Systems Review, 15 (5):137–147, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Petri:1995:LBF

[PL95] S. Petri and H. Langendörfer. Load balancing and fault tolerance in workstation clusters migrating groups of communicating processes. Operating Systems Review, 29(4):25–36, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2001:SEK

[PL01] Chang-Seop Park and Dong-Hoon Lee. Secure and efficient key management for dynamic multicast groups. Operating Systems Review, 35(4):32–38, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peir:1998:CDM

[PLH98] Jih-Kwon Peir, Yongjoon Lee, and Windsor W. Hsu. Capturing dynamic memory reference behavior with adaptive cache topology. Operating Systems Review, 32(5):240–250, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Padioleau:2008:DAC

[PLHM08] Yoann Padioleau, Julia Lawall, René Rydhof Hansen, and Gilles Muller. Documenting and automating collateral evolutions in Linux device drivers. Operating Systems Review, 42(4):247–260, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Padioleau:2006:UCE

[PLM06] Yoann Padioleau, Julia L. Lawall, and Gilles Muller. Understanding collateral evolution in Linux device drivers. Operating Systems Review, 40(4):59–71, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Powell:1983:PMD

[PM83] Michael L. Powell and Barton P. Miller. Process migration in DEMOS/MP. Operating Systems Review, 17(5):110–119, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pascal:2003:PID

[PM03] Patricia Pascal and Thierry Monteil. PAPER: influence of deterministic customers in time sharing scheduler. Operating Systems Review, 37(1):34–45, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pradhan:2000:SHP

[PN00] Prashant Pradhan and Anindya Neogi. Suez: high-performance real-time IP router. Operating Systems Review, 34(2):39, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Petrini:2006:SSF

[PNT06] Fabrizio Petrini, Jarek Nieplocha, and Vinod Tipparaju. SFT: scalable fault tolerance. Operating Systems Review, 40(2):55– 62, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ponder:1997:OUD

[Pon97] Carl Ponder. Organizing UNIX directories as lattices. Operating Systems Review, 31(4):72–77, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Poole:1973:WTT

[Poo73] Peter C. Poole. When is a test not a test? Operating Systems Review, 7(2):4–5, April 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1975:DSC

[Pop75] Gerald J. Popek. On data secure computer networks. Operating Systems Review, 9(3):59–62, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Porter:2010:DSC

[Por10] George Porter. Decoupling storage and computation in Hadoop with SuperDataNodes. Operating Systems Review, 44(2):41–46, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Potier:1977:ADP

[Pot77] Dominique Potier. Analysis of demand paging policies with swapped working sets. Operating Systems Review, 11(5):125– 131, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Powell:1977:DFS

[Pow77] Michael L. Powell. The DEMOS file system. Operating Systems Review, 11(5):33–42, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Power:1989:DSE

[Pow89] June Power. Distributed system evolution-some observations. Operating Systems Review, 23(4):31–32, October 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Powell:1983:PRB

[PP83] Michael L. Powell and David L. Presotto. Publishing: a reliable broadcast communication mechanism. Operating Systems Review, 17(5):100–109, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2006:CMS

[PP06] KyoungSoo Park and Vivek S. Pai. CoMon: a mostly-scalable monitoring system for PlanetLab. Operating Systems Review, 40(1):65–74, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parikh:2009:NWN

[PP09] Tapan Parikh and Vivek Pai. NSDR 2009 3rd Workshop on Networked Systems for Developing Regions. Operating Systems Review, 43(4):72, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2017:DRM

[PPM17] Jason Jong Kyu Park, Yongjun Park, and Scott Mahlke. Dynamic resource management for efficient utilization of multitasking GPUs. Operating Systems Review, 51(2):527–540, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pfitscher:2014:COD

[PPO14] Ricardo J. Pfitscher, Mauricio A. Pillon, and Rafael R. Obelheiro. Customer-oriented diagnosis of memory provisioning for IaaS clouds. Operating Systems Review, 48(1):2–10, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pettit:2018:BPH

[PPS^{+}18] Justin Pettit, Ben Pfaff, Joe Stringer, Cheng-Chun Tu, Brenden Blanco, and Alex Tessmer. Bringing platform harmony to VMware NSX. Operating Systems Review, 52(1):123–128, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pike:1993:UNS

[PPT^{+}93] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, and Phil Winterbottom. The use of name spaces in Plan 9. Operating Systems Review, 27(2):72–76, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peinl:1983:SMD

[PR83] Peter Peinl and Andreas Reuter. Synchronizing multiple database processes in a tightly coupled multiprocessor environment. Operating Systems Review, 17(1):30–37, January 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Peterson:2006:DPP

[PR06] Larry Peterson and Timothy Roscoe. The design principles of PlanetLab. Operating Systems Review, 40(1):11–16, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Paiva:2015:DPD

[PR15] João Paiva and Luís Rodrigues. On data placement in distributed systems. Operating Systems Review, 49(1):126–130, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pratt:1986:AMM

[Pra86] S. J. Pratt. The alchemy model: a model for homogeneous and heterogeneous distributed computing system. Operating Systems Review, 20(2):25–37, April 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pratt:1987:CTA

[Pra87] S. J. Pratt. Catastrophe theory: its application in operating system design. Operating Systems Review, 21(2):23–32, April 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pai:1996:EMC

[PRAH96] Vijay S. Pai, Parthasarathy Ranganathan, Sarita V. Adve, and Tracy Harton. An evaluation of memory consistency models for shared-memory systems with ILP processors. Operating Systems Review, 30(5):12–23, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Poirier:2010:AOS

[PRD10] Benjamin Poirier, Robert Roy, and Michel Dagenais. Accurate offline synchronization of distributed traces using kernel-level events. Operating Systems Review, 44(3):75–87, July 2010. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Perl:1996:SWN

[PS96] Sharon E. Perl and Richard L. Sites. Studies of Windows NT performance using dynamic execution traces. Operating Systems Review, 30(SI):169–183, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1998:TAP

[PS98] D. Patiyoot and S. J. Shepherd. Techniques for authentication protocols and key distribution on wireless ATM networks. Operating Systems Review, 32(4):25–32, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1999:MES

[PS99a] D. Patiyoot and S. J. Shepherd. Modelling and evaluation of security induced delay in wireless ATM networks. Operating Systems Review, 33(3):26–31, July 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1999:WSS

[PS99b] D. Patiyoot and S. J. Shepherd. WASS: a security services for wireless ATM networks. Operating Systems Review, 33(4):36– 41, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1999:SIA

[PS99c] Danai Patiyoot and S. J. Shepard. Security issues in ATM networks. Operating Systems Review, 33(4):22–35, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1999:CST

[PS99d] Danai Patiyoot and S. J. Shepherd. Cryptographic security techniques for wireless networks. Operating Systems Review, 33(2):

36–50, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patiyoot:1999:WWA

[PS99e] Danai Patiyoot and S. J. Shepherd. WASS: wireless ATM security system. Operating Systems Review, 33(2):29–35, April 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pillai:2001:RTD

[PS01] Padmanabhan Pillai and Kang G. Shin. Real-time dynamic voltage scaling for low-power embedded operating systems. Operating Systems Review, 35(5):89–102, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Paik:2009:SLC

[PS09] Michael Paik and Lakshminarayanan Subramanian. Signet: lowcost auditable transactions using SIMs and mobile phones. Operating Systems Review, 43(4):73–78, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Portokalidis:2006:AEF

[PSB06] Georgios Portokalidis, Asia Slowinska, and Herbert Bos. Argos: an emulator for fingerprinting zero-day attacks for advertised honeypots with automatic signature generation. Operating Systems Review, 40(4):15–27, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Payne:2007:LAS

[PSC^{+}07] Bryan D. Payne, Reiner Sailer, Ramón Cáceres, Ron Perez, and Wenke Lee. A layered approach to simplified access control in virtualized systems. Operating Systems Review, 41(4):12–19, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Parashar:2006:SSB

[PSG06] Angshuman Parashar, Anand Sivasubramaniam, and Sudhanva Gurumurthi. SlicK: slice-based locality exploitation for efficient redundant multithreading. Operating Systems Review, 40(5): 95–105, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Patrick:2008:CEO

[PSK08] Christina M. Patrick, SeungWoo Son, and Mahmut Kandemir. Comparative evaluation of overlap strategies with study of I/O overlap in MPI-IO. Operating Systems Review, 42(6):43–49, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Papagiannis:2016:IOS

[PSMB16] Anastasios Papagiannis, Giorgos Saloustros, Manolis Marazakis, and Angelos Bilas. Iris: an optimized I/O stack for lowlatency storage devices. Operating Systems Review, 50(3):3–11, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Petersen:1997:FUP

[PST^{+}97] Karin Petersen, Mike J. Spreitzer, Douglas B. Terry, Marvin M. Theimer, and Alan J. Demers. Flexible update propagation for weakly consistent replication. Operating Systems Review, 31(5): 288–301, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Padala:2007:ACV

[PSZ^{+}07] Pradeep Padala, Kang G. Shin, Xiaoyun Zhu, Mustafa Uysal, Zhikui Wang, Sharad Singhal, Arif Merchant, and Kenneth Salem. Adaptive control of virtualized resources in utility computing environments. Operating Systems Review, 41(3):289–302, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Phothilimthana:2016:SS

[PTBD16] Phitchaya Mangpo Phothilimthana, Aditya Thakur, Rastislav Bodik, and Dinakar Dhurjati. Scaling up superoptimization. Operating Systems Review, 50(2):297–310, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pu:1993:RLS

[Pu93] Calton Pu. Relaxing the limitations of serializable transactions in distributed systems. Operating Systems Review, 27(2):66–71, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Paradinas:1995:NDI

[PV95] Pierre Paradinas and Jean-Jacques Vandewalle. New directions for integrated circuit cards operating systems. Operating Systems Review, 29(1):56–61, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Powers:2017:BBG

[PVB17] Bobby Powers, John Vilk, and Emery D. Berger. Browsix: Bridging the gap between Unix and the browser. Operating Systems Review, 51(2):253–266, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Perl:1993:PAC

[PW93] Sharon E. Perl and William E. Weihl. Performance assertion checking. Operating Systems Review, 27(5):134–145, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pfitzmann:1998:HBF

[PW98] Birgit Pfitzmann and Michael Waidner. How to break frauddetectable key recovery. Operating Systems Review, 32(1):23– 28, January 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Popek:1981:LNT

[PWC^{+}81] G. Popek, B. Walker, J. Chow, D. Edwards, C. Kline, G. Rudisin, and G. Thiel. LOCUS a network transparent, high reliability distributed system. Operating Systems Review, 15(5): 169–177, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pei:2019:BER

[PWT^{+}19] Kexin Pei, Shiqi Wang, Yuchi Tian, Justin Whitehouse, Carl Vondrick, Yinzhi Cao, Baishakhi Ray, Suman Jana, and Junfeng Yang. Bringing engineering rigor to deep learning. Operating Systems Review, 53(1):59–67, July 2019. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Qie:2002:DPU

[QPP02] Xiaohu Qie, Ruoming Pang, and Larry Peterson. Defensive programming: using an annotation toolkit to build DoS-resistant software. Operating Systems Review, 36(5S):45–60, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Qin:2005:RTB

[QTSZ05] Feng Qin, Joseph Tucek, Jagadeesan Sundaresan, and Yuanyuan Zhou. Rx: treating bugs as allergies—a safe method to survive software failures. Operating Systems Review, 39(5):235–248, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rauchwerger:2006:SMW

[RA06] Lawrence Rauchwerger and Nancy M. Amato. SmartApps: middle-ware for adaptive applications on reconfigurable platforms. Operating Systems Review, 40(2):73–82, April 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rellermeyer:2007:CSP

[RA07] Jan S. Rellermeyer and Gustavo Alonso. Concierge: a service platform for resource-constrained devices. Operating Systems Review, 41(3):245–258, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Robinson:2007:HAM

[RAF07] Jeffrey Choi Robinson and Jim Alves-Foss. A high assurance MLS file server. Operating Systems Review, 41(1):45–53, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ramachandran:2000:SIU

[Ram00] Umakishore Ramachandran. System infrastructure for ubiquitous presence. Operating Systems Review, 34(2):31, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rana:1982:THD

[Ran82] S. P. Rana. Triple-handed dining philosophers. Operating Systems Review, 16(1):6–9, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rattan:1987:MMU

[Rat87] I. Rattan. Memory management units for microcomputer operating systems. Operating Systems Review, 21(1):34–38, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rattner:2011:RI

[Rat11] Justin Rattner. Research at Intel. Operating Systems Review, 45(1):1–2, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rodrigues:2012:SWL

[RAVC12] Luis Rodrigues, Divy Agrawal, Ymir Vigfusson, and Gregory Chockler. Summary of the 5th Workshop on Large-Scale Distributed Systems and Middleware (LADIS 2011). Operating Systems Review, 46(1):1–3, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raynal:1991:STD

[Ray91] Michel Raynal. A simple taxonomy for distributed mutual exclusion algorithms. Operating Systems Review, 25(2):47–50, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raynal:1992:ALC

[Ray92] Michel Raynal. About logical clocks for distributed systems. Operating Systems Review, 26(1):41–48, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Russell:1975:ERP

[RB75] David L. Russell and Thomas H. Bredt. Error resynchronization in producer-consumer systems. Operating Systems Review, 9(5): 106–113, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reddy:1993:NBT

[RB93] P. Krishna Reddy and S. Bhalla. A non-blocking transaction data flow graph based approach for replicated data. Operating Systems Review, 27(3):46–54, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rolon:2024:BMP

[RB24] Sebastián Rolón and Oana Balmau. Is bare-metal I/O performance with user-defined storage drives inside VMs possible? Benchmarking libvfio-user vs. common storage virtualization configurations. Operating Systems Review, 58(1):45–52, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3689051.3689059.

Rosenblum:1995:IAT

[RBH^{+}95] M. Rosenblum, E. Bugnion, S. A. Herrod, E. Witchel, and A. Gupta. The impact of architectural trends on operating system performance. Operating Systems Review, 29(5):285–298, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Riviere:2007:CGC

[RBLP07] Étienne Rivière, Roberto Baldoni, Harry Li, and José Pereira. Compositional gossip: a conceptual architecture for designing gossip-based applications. Operating Systems Review, 41(5):43– 50, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rao:2001:PBP

[RCC01] Herman Chung-Hwa Rao, Yih-Farn Chen, and Ming-Feng Chen. A proxy-based personal Web archiving service. Operating Systems Review, 35(1):61–72, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rodrigues:2001:BUA

[RCL01] Rodrigo Rodrigues, Miguel Castro, and Barbara Liskov. BASE: using abstraction to improve fault tolerance. Operating Systems Review, 35(5):15–28, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Repantis:2010:SMI

[RCSW10] Thomas Repantis, Jeff Cohen, Scott Smith, and Joel Wein. Scaling a monitoring infrastructure for the Akamai network. Operating Systems Review, 44(3):20–26, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ruckert:1987:LSC

[RD87] Rogert G. Ruckert and John D. Dean. Launching a successful CPME program in a multi-vendor environment. Operating Systems Review, 21(1):39–48, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Robinson:1997:NPR

[RD97] John T. Robinson and Murthy V. Devarakonda. Note on a problem with Reed and Long’s FBR results. Operating Systems Review, 31(1):3–4, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rowstron:2001:SMC

[RD01] Antony Rowstron and Peter Druschel. Storage management and caching in PAST, a large-scale, persistent peer-to-peer storage utility. Operating Systems Review, 35(5):188–201, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rusu:2012:GSF

[RD12] Florin Rusu and Alin Dobra. GLADE: a scalable framework for efficient analytics. Operating Systems Review, 46(1):12–18, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reinhardt:1985:DFA

[Rei85] Steve Reinhardt. A data-flow approach to multitasking on CRAY X-MP computers. Operating Systems Review, 19(5): 107–114, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reiter:1992:ISG

[Rei92] Michael Reiter. Integrating security in a group oriented distributed system. Operating Systems Review, 26(2):27, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Redstone:2000:AOS

[REL00] Joshua A. Redstone, Susan J. Eggers, and Henry M. Levy. An analysis of operating system behavior on a simultaneous multithreaded architecture. Operating Systems Review, 34(5): 245–256, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ranganathan:1998:ESD

[RF98] Narayan Ranganathan and Manoj Franklin. An empirical study of decentralized ILP execution models. Operating Systems Review, 32(5):272–281, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reis:2017:SAC

[RF17] João Gabriel Reis and Antônio Augusto Fröhlich. OS support for adaptive components in self-aware systems. Operating Systems Review, 51(1):101–112, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rajwar:2002:TLF

[RG02] Ravi Rajwar and James R. Goodman. Transactional lock-free execution of lock-based programs. Operating Systems Review, 36(5):5–17, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ranganathan:1998:PDW

[RGAB98] Parthasarathy Ranganathan, Kourosh Gharachorloo, Sarita V. Adve, and Luiz André Barroso. Performance of database workloads on shared-memory systems with out-of-order processors. Operating Systems Review, 32(5):307–318, December 1998. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Richmond:1997:NPM

[RH97] Michael Richmond and Michael Hitchens. A new process migration algorithm. Operating Systems Review, 31(1):31–42, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ruiz:2015:RSA

[RHMR15] Cristian Ruiz, Salem Harrache, Michael Mercier, and Olivier Richard. Reconstructable software appliances with Kameleon. Operating Systems Review, 49(1):80–89, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rossbach:2007:TUM

[RHP^{+}07] Christopher J. Rossbach, Owen S. Hofmann, Donald E. Porter, Hany E. Ramadan, Bhandari Aditya, and Emmett Witchel. TxLinux: using and managing hardware transactional memory in an operating system. Operating Systems Review, 41(6): 87–102, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rajbhandari:2017:OCM

[RHR^{+}17] Samyam Rajbhandari, Yuxiong He, Olatunji Ruwase, Michael Carbin, and Trishul Chilimbi. Optimizing CNNs on multicores for scalability, performance and goodput. Operating Systems Review, 51(2):267–280, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rickert:1988:NBC

[Ric88] Neil W. Rickert. Non-Byzantine clock synchronization—a programming experiment. Operating Systems Review, 22(1):73–78, January 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Riedel:2007:I

[Rie07] Erik Riedel. Introduction. Operating Systems Review, 41(1): 5–6, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rippert:2003:PFO

[Rip03] Christophe Rippert. Protection in flexible operating system architectures. Operating Systems Review, 37(4):8–18, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rizzo:1997:VFA

[Riz97] Luigi Rizzo. A very fast algorithm for RAM compression. Operating Systems Review, 31(2):36–45, April 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Riekstin:2014:NME

[RJK^{+}14] Ana Carolina Riekstin, Sean James, Aman Kansal, Jie Liu, and Eric Peterson. No more electrical infrastructure: towards fuel cell powered data centers. Operating Systems Review, 48(1):39– 43, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reed:1977:SES

[RK77] David P. Reed and Rajendra K. Kanodia. Synchronization with eventcounts and sequencers (extended abstract). Operating Systems Review, 11(5):91, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reid:1983:FSS

[RK83] Loretta Guarino Reid and Philip L. Karlton. A file system supporting cooperation between programs. Operating Systems Review, 17(5):20–29, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raghavan:2011:RRP

[RK11] Karthik Raghavan and V. Kamakoti. ROSY: recovering processor and memory systems from hard errors. Operating Systems Review, 45(3):82–84, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reddy:2011:BFH

[RKBH11] Dheeraj Reddy, David Koufaty, Paul Brett, and Scott Hahn. Bridging functional heterogeneity in multicore architectures. Operating Systems Review, 45(1):21–33, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raghunath:2011:DDF

[RKV11] Arun Raghunath, John Keys, and Mona Vij. Direct data flows. Operating Systems Review, 45(1):54–61, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reed:1996:ACA

[RL96] Benjamin Reed and Darrell D. E. Long. Analysis of caching algorithms for distributed file systems. Operating Systems Review, 30(3):12–21, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Regnier:2008:EIH

[RLB08] Paul Regnier, George Lima, and Luciano Barreto. Evaluation of interrupt handling timeliness in real-time Linux operating systems. Operating Systems Review, 42(6):52–63, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ren:2017:SDH

[RLD^{+}17] Ao Ren, Zhe Li, Caiwen Ding, Qinru Qiu, Yanzhi Wang, Ji Li, Xuehai Qian, and Bo Yuan. SC–DCNN: Highly-scalable deep convolutional neural network using stochastic computing. Operating Systems Review, 51(2):405–418, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romer:1996:SPI

[RLV^{+}96] Theodore H. Romer, Dennis Lee, Geoffrey M. Voelker, Alec Wolman, Wayne A. Wong, Jean-Loup Baer, Brian N. Bershad, and Henry M. Levy. The structure and performance of interpreters. Operating Systems Review, 30(5):150–159, December 1996. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Roth:1998:DBP

[RMS98] Amir Roth, Andreas Moshovos, and Gurindar S. Sohi. Dependence based prefetching for linked data structures. Operating Systems Review, 32(5):115–126, December 1998. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ram:2001:CCM

[RMSB01] D. Janaki Ram, M. Uma Mahesh, N. S. K. Chandra Sekhar, and Chitra Babu. Causal consistency in mobile environment. Operating Systems Review, 35(1):34–40, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Richardson:1983:TFM

[RN83] M. F. Richardson and R. M. Needham. The TRIPOS filing machine, a front end to a file server. Operating Systems Review, 17(5):120–128, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raja:1993:SDP

[RN93] Prasad Raja and Guevara Noubir. Static and dynamic polling mechanisms for Fieldbus networks. Operating Systems Review, 27(3):34–45, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ryutov:2000:RESa

[RN00a] Tatyana Ryutov and Clifford Neuman. Representation and evaluation of security policies. Operating Systems Review, 34(2):34, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ryutov:2000:RESb

[RN00b] Tatyana Ryutov and Clifford Neuman. Representation and evaluation of security policies (poster session). Operating Systems Review, 34(2):41, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rosenblum:1991:DIL

[RO91] Mendel Rosenblum and John K. Ousterhout. The design and implementation of a log-structured file system. Operating Systems Review, 25(5):1–15, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Robinson:1996:ASS

[Rob96] John T. Robinson. Analysis of steady-state segment storage utilizations in a log-structured file system with least-utilized segment cleaning. Operating Systems Review, 30(4):29–32, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Robbins:1998:IEI

[Rob98] Steven Robbins. Introducing empirical investigation in undergraduate operating systems. Operating Systems Review, 32(4): 77–80, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Robbins:2008:TPA

[Rob08] Steven Robbins. A three pronged approach to teaching undergraduate operating systems. Operating Systems Review, 42 (6):93–100, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rakotoarivelo:2009:OCM

[ROJS09] Thierry Rakotoarivelo, Maximilian Ott, Guillaume Jourjon, and Ivan Seskar. OMF: a control and management framework for networking testbeds. Operating Systems Review, 43(4):54–59, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ricci:2006:LRA

[ROLV06] Robert Ricci, David Oppenheimer, Jay Lepreau, and Amin Vahdat. Lessons from resource allocators for large-scale multiuser testbeds. Operating Systems Review, 40(1):25–32, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romanovsky:1993:FTS

[Rom93] Alexander B. Romanovsky. Fault tolerance: synchronization of redundancy. Operating Systems Review, 27(4):58–66, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romanovsky:1995:SDW

[Rom95] A. B. Romanovsky. Software diversity as a way to wellstructured concurrent software. Operating Systems Review, 29 (3):85–90, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romanovsky:1997:CGS

[Rom97] Alexander B. Romanovsky. Conversational group service. Operating Systems Review, 31(1):54–63, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rose:1978:PEU

[Ros78] Greg Rose. Performance evaluation under Unix and a study of PDP-11 instruction usage. Operating Systems Review, 12(3): 38–45, July 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rosenburg:1989:LST

[Ros89] B. Rosenburg. Low-synchronization translation lookaside buffer consistency in large-scale shared-memory multiprocessors. Operating Systems Review, 23(5):137–146, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Roscoe:1994:LNS

[Ros94] Timothy Roscoe. Linkage in the Nemesis single address space operating system. Operating Systems Review, 28(4):48–55, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Roscoe:1995:CIS

[Ros95] Timothy Roscoe. CLANGER: an interpreted systems programming language. Operating Systems Review, 29(2):13–20, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rosenblum:2006:IVC

[Ros06] Mendel Rosenblum. Impact of virtualization on computer architecture and operating systems. Operating Systems Review, 40(5):1, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Routh:1984:PAA

[Rou84] Richard LeRoy Routh. A proposal for an architectural approach which apparently solves all known software-based internal computer security problems. Operating Systems Review, 18(3):31– 39, July 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ramamurthy:2007:PDE

[RP07] Pratap Ramamurthy and Ramanathan Palaniappan. Performancedirected energy management using BOS. Operating Systems Review, 41(1):66–77, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rodriquez:1997:NCM

[RPM97] Santiago Rodríquez, Antonio Pérez, and Rafael Méndez. A new checkpoint mechanism for real time operating systems. Operating Systems Review, 31(4):55–62, October 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ruan:2008:DCS

[RPNT08] Yaoping Ruan, Vivek S. Pai, Erich Nahum, and John M. Tracey. Do commodity SMT processors need more OS research? Operating Systems Review, 42(1):21–25, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rodriguez-Rosell:1972:EDH

[RR72] Juan Rodriguez-Rosell. Experimental data on how program behavior affects the choice of scheduler parameters. Operating Systems Review, 6(1/2):156–163, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rashid:1981:ACO

[RR81] Richard F. Rashid and George G. Robertson. Accent: A communication oriented network operating system kernel. Operating Systems Review, 15(5):64–75, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Regehr:2004:HSA

[RR04] John Regehr and Alastair Reid. HOIST: a system for automatically deriving static analyzers for embedded systems. Operating Systems Review, 38(5):133–143, December 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rafique:2009:SML

[RRBN09] M. Mustafa Rafique, Benjamin Rose, Ali R. Butt, and Dimitrios S. Nikolopoulos. Supporting MapReduce on large-scale asymmetric multi-core clusters. Operating Systems Review, 43 (2):25–34, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romano:2010:CTH

[RRCC10] Paolo Romano, Luis Rodrigues, Nuno Carvalho, and João Cachopo. Cloud-TM: harnessing the cloud with distributed transactional memories. Operating Systems Review, 44(2):1–6, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reddy:2006:UPB

[RRP06] Vimal K. Reddy, Eric Rotenberg, and Sailashri Parthasarathy. Understanding prediction-based partial redundant threading for low-overhead, high-coverage fault tolerance. Operating Systems Review, 40(5):83–94, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Raghavendra:2008:NPS

[RRT^{+}08] Ramya Raghavendra, Parthasarathy Ranganathan, Vanish Talwar, Zhikui Wang, and Xiaoyun Zhu. No ‘power’ struggles: coordinated multi-level power management for the data center. Operating Systems Review, 42(2):48–59, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rattan:1986:TFV

[RS86] I. Rattan and L. P. S. Singh. On table fragmentation in virtual memory management. Operating Systems Review, 20(3):28–30, July 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rodeheffer:1991:ARA

[RS91] Thomas L. Rodeheffer and Michael D. Schroeder. Automatic reconfiguration in Autonet. Operating Systems Review, 25(5): 183–197, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Regehr:2000:HSP

[RS00] John Regehr and John A. Stankovic. Hierarchical schedulers, performance guarantee, and resource management. Operating Systems Review, 34(2):31, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rauch:2002:CTU

[RS02] Felix Rauch and Thomas M. Stricker. Comments on “Transparent User-Level Process Checkpoint and Restore for Migration” by Bozyigit and Wasiq. Operating Systems Review, 36(3):8–9, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [BW01].

Raj:2008:OEO

[RS08] Himanshu Raj and Karsten Schwan. O2S2: enhanced objectbased virtualized storage. Operating Systems Review, 42(6):24– 29, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rabbah:2004:COP

[RSEW04] Rodric M. Rabbah, Hariharan Sandanagobalane, Mongkol Ekpanyapong, and Weng-Fai Wong. Compiler orchestrated prefetching via speculation and predication. Operating Systems Review, 38(5):189–198, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ronda:2008:IUA

[RSW08] Troy Ronda, Stefan Saroiu, and Alec Wolman. Itrustpage: a user-assisted anti-phishing tool. Operating Systems Review, 42 (4):261–272, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ritchie:1973:UTSb

[RT73] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system. Operating Systems Review, 7(4):27, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X

(electronic). URL http://citeseerx.ist.psu.edu/viewdoc/
summary?doi=10.1.1.100.7314.

Rashid:1987:MIV

[RTY^{+}87] Richard Rashid, Avadis Tevanian, Michael Young, David Golub, and Robert Baron. Machine-independent virtual memory management for paged uniprocessor and multiprocessor architectures. Operating Systems Review, 21(4):31–39, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Russell:1977:PBP

[Rus77] David L. Russell. Process backup in producer-consumer systems. Operating Systems Review, 11(5):151–157, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rushby:1981:DVS

[Rus81] J. M. Rushby. Design and verification of secure systems. Operating Systems Review, 15(5):12–21, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Russell:1988:SUC

[Rus88] Stephen Russell. Single-user capabilities in interprocess communication. Operating Systems Review, 22(2):45–52, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Russell:2008:VTF

[Rus08] Rusty Russell. virtio: towards a de-facto standard for virtual I/O devices. Operating Systems Review, 42(5):95–103, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rangan:1991:DFS

[RV91] P. Venkat Rangan and Harrick M. Vin. Designing file systems for digital video and audio. Operating Systems Review, 25(5):81– 94, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ricci:2015:APR

[RWS^{+}15] Robert Ricci, Gary Wong, Leigh Stoller, Kirk Webb, Jonathon Duerig, Keith Downie, and Mike Hibler. Apt: a platform for repeatable research in computer science. Operating Systems Review, 49(1):100–107, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ryan:1998:SPD

[Rya98] Stein J. Ryan. Synchronization in portable device drivers. Operating Systems Review, 32(4):62–69, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ryan:1999:SPD

[Rya99] Stein J. Ryan. Synchronization in portable device drivers. Operating Systems Review, 33(1):18–25, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Romanovsky:1997:DCA

[RZ97] A. Romanovsky and A. F. Zorzo. On distribution of coordinated atomic actions. Operating Systems Review, 31(4):63–71, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sadeh:1975:APP

[Sad75] E. Sadeh. An analysis of the performance of the page fault frequency (PFF) replacement algorithm. Operating Systems Review, 9(5):6–13, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sivathanu:2002:ERA

[SADAD02] Muthian Sivathanu, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Evolving RPC for active storage. Operating Systems Review, 36(5):264–276, December 2002. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Su:2007:AIC

[SAF07] Ya-Yunn Su, Mona Attariyan, and Jason Flinn. AutoBash: improving configuration management with operating system causality analysis. Operating Systems Review, 41(6):237–250, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soundararajan:2006:DRP

[SAG06] Gokul Soundararajan, Cristiana Amza, and Ashvin Goel. Database replication policies for dynamic content applications. Operating Systems Review, 40(4):89–102, October 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1973:PCI

[Sal73] Jerome H. Saltzer. Protection and control of information sharing in Multics. Operating Systems Review, 7(4):119, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1974:ORD

[Sal74] Jerome H. Saltzer. Ongoing research and development on information protection. Operating Systems Review, 8(3):8–24, July 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1978:DS

[Sal78a] Jerome H. Saltzer. On digital signatures. Operating Systems Review, 12(2):12–14, April 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1978:RPD

[Sal78b] Jerome H. Saltzer. Research problems of decentralized systems with largely autonomous nodes. Operating Systems Review, 12 (1):43–52, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1991:FTV

[Sal91] Jerome H. Saltzer. Fault-tolerance in very large archival systems. Operating Systems Review, 25(1):81–82, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:1993:NSS

[Sal93] Jerome H. Saltzer. Needed: a systematic structuring paradigm for distributed data. Operating Systems Review, 27(2):77–81, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saltzer:2000:CCI

[Sal00] Jerry Saltzer. Copying with complexity (invited talk) (summary only). Operating Systems Review, 34(2):7–8, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Setty:2020:VSM

[SAL20] Srinath Setty, Sebastian Angel, and Jonathan Lee. Verifiable state machines: Proofs that untrusted services operate correctly. Operating Systems Review, 54(1):40–46, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL

https://dl.acm.org/doi/10.1145/3421473.3421479.

Sanguinetti:1981:UMC

[San81] John Sanguinetti. The use of the monitor call instruction to implement domain switching in the IBM 370 architecture. Operating Systems Review, 15(4):55–61, October 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sansom:1986:BRC

[San86] Robert Sansom. Book review: Computer Security: A Global Challenge — Proceedings of the Second IFIP International Conference on Computer Security, IFIP/Sec’84, Toronto, Ontario, Canada, 10–12 September, 1984 : (Elsevier Science Publishing Co. 1984). Operating Systems Review, 20(3):9, July 1986. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1981:SFS

[Sat81] M. Satyanarayanan. A study of file sizes and functional lifetimes. Operating Systems Review, 15(5):96–108, December 1981. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1995:WMC

[Sat95] M. Satyanarayanan. Workshop on mobile computing systems and applications, December 1994: digest of proceedings. Operating Systems Review, 29(2):4–12, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1999:DPS

[Sat99] M. Satyanarayanan. Digest of proceedings, Seventh IEEE Workshop on Hot Topics in Operating Systems, March 29–30, 1999, Rio Rico, AZ. Operating Systems Review, 33(4):4–21, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:2000:CTR

[Sat00] M. Satyanarayanan. Caching trust rather than content. Operating Systems Review, 34(4):32–33, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saha:2007:ESP

[SATG^{+}07] Bratin Saha, Ali-Reza Adl-Tabatabai, Anwar Ghuloum, Mohan Rajagopalan, Richard L. Hudson, Leaf Petersen, Vijay Menon, Brian Murphy, Tatiana Shpeisman, Eric Sprangle, Anwar Rohillah, Doug Carmean, and Jesse Fang. Enabling scalability and performance in a large scale CMP environment. Operating Systems Review, 41(3):73–86, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schneider:1978:SCP

[SB78] F. B. Schneider and A. J. Bernstein. Scheduling in Concurrent Pascal. Operating Systems Review, 12(2):15–20, April 1978. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1989:PFR

[SB89] M. Schroeder and M. Burrows. Performance of Firefly RPC. Operating Systems Review, 23(5):83–90, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stephenson:1991:FCM

[SB91] Pat Stephenson and Kenneth Birman. Fast causal multicast. Operating Systems Review, 25(2):75–79, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Simons:2010:VHP

[SB10a] Joshua E. Simons and Jeffrey Buell. Virtualizing high performance computing. Operating Systems Review, 44(4):136–145, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Slowinska:2010:PTS

[SB10b] Asia Slowinska and Herbert Bos. Pointer tainting still pointless: (but we all see the point of tainting). Operating Systems Review, 44(3):88–92, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Singleton:1986:SMF

[SBB86] P. Singleton, K. H. Bennett, and O. P. Brereton. A single model for files and processes. Operating Systems Review, 20(1):12–18, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Szalay:2010:LPA

[SBH^{+}10] Alexander S. Szalay, Gordon C. Bell, H. Howie Huang, Andreas Terzis, and Alainna White. Low-power Amdahl-balanced blades for data intensive computing. Operating Systems Review, 44 (1):71–75, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saito:1999:MAP

[SBL99] Yasushi Saito, Brian N. Bershad, and Henry M. Levy. Manageability, availability and performance in Porcupine: a highly scalable, cluster-based mail service. Operating Systems Review, 33(5):1–15, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saito:2000:MAP

[SBL00] Yasushi Saito, Brian N. Bershad, and Henry M. Levy. Manageability, availability and performance in Porcupine: a highly scalable, cluster-based mail service (summary only). Operating Systems Review, 34(2):9–11, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1983:EGS

[SBN83] Michael D. Schroeder, Andrew D. Birrell, and Roger M. Needham. Experience with Grapevine (summary): the growth of a distributed system. Operating Systems Review, 17(5):141–142, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Savage:1997:EDD

[SBN^{+}97] Stefan Savage, Michael Burrows, Greg Nelson, Patrick Sobalvarro, and Thomas Anderson. Eraser: a dynamic data race detector for multi-threaded programs. Operating Systems Review, 31(5):27–37, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shahar:2018:ACS

[SBS18] Sagi Shahar, Shai Bergman, and Mark Silberstein. ActivePointers: a case for software address translation on GPUs. Operating Systems Review, 52(1):84–95, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strout:1998:SIS

[SCFS98] Michelle Mills Strout, Larry Carter, Jeanne Ferrante, and Beth Simon. Schedule-independent storage mapping for loops. Operating Systems Review, 32(5):24–33, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Snoeren:2001:MBC

[SCG01] Alex C. Snoeren, Kenneth Conley, and David K. Gifford. Meshbased content routing using XML. Operating Systems Review, 35(5):160–173, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scheffier:1973:OFP

[Sch73a] Lee J. Scheffier. Optimal folding of a paging drum in a three level memory system. Operating Systems Review, 7(4):58–65, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1973:BRS

[Sch73b] M. D. Schroeder. A brief report on the SIGPLAN/SIGOPS interface meeting. Operating Systems Review, 7(3):4–9, July 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1975:ESK

[Sch75] Michael D. Schroeder. Engineering a security kernel for Multics. Operating Systems Review, 9(5):25–32, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schrimpf:1995:MPF

[Sch95] Harald Schrimpf. Migration of processes, files, and virtual devices in the MDX operating system. Operating Systems Review, 29(2):70–81, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:2007:SWM

[Sch07] Michael D. Schroeder. Systems work at Microsoft Research. Operating Systems Review, 41(2):1–2, April 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saghir:1996:EDD

[SCL96] Mazen A. R. Saghir, Paul Chow, and Corinna G. Lee. Exploiting dual data-memory banks in digital signal processors. Operating Systems Review, 30(5):234–243, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steinke:2005:NPF

[SCM05] Robert Steinke, Micah Clark, and Elihu McMahon. A new pattern for flexible worker threads with in-place consumption message queues. Operating Systems Review, 39(2):71–73, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scott:1996:SCT

[Sco96] Steven L. Scott. Synchronization and communication in the T3E multiprocessor. Operating Systems Review, 30(5):26–36, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scott:2004:CIB

[Sco04] Michael Scott. Cryptanalysis of an ID-based password authentication scheme using Smart Cards and fingerprints. Operating Systems Review, 38(2):73–75, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sapuntzakis:2002:OMV

[SCP^{+}02] Constantine P. Sapuntzakis, Ramesh Chandra, Ben Pfaff, Jim Chow, Monica S. Lam, and Mendel Rosenblum. Optimizing the migration of virtual computers. Operating Systems Review, 36 (5S):377–390, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shyam:2006:ULC

[SCP^{+}06] Smitha Shyam, Kypros Constantinides, Sujay Phadke, Valeria Bertacco, and Todd Austin. Ultra low-cost defect protection for microprocessor pipelines. Operating Systems Review, 40(5):73– 82, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1977:MKD

[SCS77] Michael D. Schroeder, David D. Clark, and Jerome H. Saltzer. The Multics kernel design project. Operating Systems Review, 11(5):43–56, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spector:1986:ARD

[SD86] Alfred Z. Spector and Dean Daniels. An algorithm for replicated directories. Operating Systems Review, 20(1):24–43, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spector:1985:DTR

[SDD^{+}85] Alfred Z. Spector, Dean Daniels, Daniel Duchamp, Jeffrey L. Eppinger, and Randy Pausch. Distributed transactions for reliable systems. Operating Systems Review, 19(5):127–146, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stonebraker:1985:PSD

[SDE85] Michael Stonebraker, Deborah DuBourdieux, and William Edwards. Problems in supporting data base transactions in an operating system transaction manager. Operating Systems Review, 19(1):6–14, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stets:1997:CSC

[SDH^{+}97] Robert Stets, Sandhya Dwarkadas, Nikolaos Hardavellas, Galen Hunt, Leonidas Kontothanassis, Srinivasan Parthasarathy, and Michael Scott. Cashmere-2L: software coherent shared memory on a clustered remote-write network. Operating Systems Review, 31(5):170–183, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scott:2000:IOC

[SDP^{+}00] Michael L. Scott, Sandhya Dwarkadas, Srinivasan Parthasarathy, Rajeev Balasubramonian, DeQing Chen, Grigorios Magklis, Athanasios Papathanasiou, Eduardo Pinheiro, Umit Rencuzogullari, and Chunquiang Tang. Interweave: object caching meets software distributed shared memory. Operating Systems Review, 34(2):32, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smith:1987:ZCP

[SDV^{+}87] J. E. Smith, G. E. Dermer, B. D. Vanderwarn, S. D. Klinger, and C. M. Rozewski. The ZS-1 central processor. Operating Systems Review, 21(4):199–204, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schenck:2016:EPM

[SEF^{+}16] Wolfram Schenck, Salem El Sayed, Maciej Foszczynski, Wilhelm Homberg, and Dirk Pleiter. Evaluation and performance modeling of a burst buffer solution. Operating Systems Review, 50 (3):12–26, December 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seigh:1990:DSR

[Sei90] Joseph W. Seigh. A distributed solution to the reader–writer problem. Operating Systems Review, 24(2):66–68, April 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stark:1998:VLP

[SEP98] Jared Stark, Marius Evers, and Yale N. Patt. Variable length path branch prediction. Operating Systems Review, 32(5):170– 179, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Serafini:2021:SGN

[Ser21] Marco Serafini. Scalable graph neural network training: The case for sampling. Operating Systems Review, 55(1):68–76, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3469379.3469387.

Seltzer:1996:DDS

[SESS96] Margo I. Seltzer, Yasuhiro Endo, Christopher Small, and Keith A. Smith. Dealing with disaster: surviving misbehaved kernel extensions. Operating Systems Review, 30(SI):213–227, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sutherland:2008:AVO

[SETB08] Iain Sutherland, Jon Evans, Theodore Tryfonas, and Andrew Blyth. Acquiring volatile operating system data tools and techniques. Operating Systems Review, 42(3):65–73, April 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sincoskie:1980:SDO

[SF80] W. David Sincoskie and David J. Farber. SODS/OS: a distributed operating system for the IBM Series/1. Operating Systems Review, 14(3):46–54, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sohi:1991:HBD

[SF91] Gurindar S. Sohi and Manoj Franklin. High-bandwidth data memory systems for superscalar processors. Operating Systems Review, 25(3S):53–62, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sen:2012:CCS

[SF12] Siddhartha Sen and Michael J. Freedman. Commensal cuckoo: secure group partitioning for large-scale services. Operating Systems Review, 46(1):33–39, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shelepov:2009:HSH

[SFB^{+}09] Daniel Shelepov, Alexandra Fedorova, Sergey Blagodurov, Juan Carlos Saez Alcaide, Nestor Perez, Viren Kumar, Stacey Jeffery, and Zhi Feng Huang. HASS: a scheduler for heterogeneous multicore systems. Operating Systems Review, 43(2):66–75, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Santry:1999:DWF

[SFH^{+}99] Douglas S. Santry, Michael J. Feeley, Norman C. Hutchinson, Alistair C. Veitch, Ross W. Carton, and Jacob Ofir. Deciding when to forget in the Elephant file system. Operating Systems Review, 33(5):110–123, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Santry:2000:DWF

[SFH^{+}00] Douglas S. Santry, Michael J. Feeley, Norman C. Hutchinson, Alistair C. Veitch, Ross W. Carton, and Jacob Ofir. Deciding when to forget in the Elephant file system. Operating Systems Review, 34(2):18–19, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schoinas:1994:FGA

[SFL^{+}94] Ioannis Schoinas, Babak Falsafi, Alvin R. Lebeck, Steven K. Reinhardt, James R. Larus, and David A. Wood. Fine-grain access control for distributed shared memory. Operating Systems Review, 28(5):297–306, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shue:2013:FIM

[SFS13] David Shue, Michael J. Freedman, and Anees Shaikh. Fairness and isolation in multi-tenant storage as optimization decomposition. Operating Systems Review, 47(1):16–21, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saito:2004:FBD

[SFV^{+}04] Yasushi Saito, Svend Frølund, Alistair Veitch, Arif Merchant, and Susan Spence. FAB: building distributed enterprise disk arrays from commodity components. Operating Systems Review, 38(5):48–58, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1999:VPE

[SFW99] M. Satyanarayanan, Jason Flinn, and Kevin R. Walker. Visual proxy: exploiting OS customizations without application source code. Operating Systems Review, 33(3):14–18, July 1999. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scales:1997:TTE

[SG97] Daniel J. Scales and Kourosh Gharachorloo. Towards transparent and efficient software distributed shared memory. Operating Systems Review, 31(5):157–169, December 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Serral:2004:LNS

[SG04] René Serral and Marisa Gil. A Linux networking study. Operating Systems Review, 38(3):1–11, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soules:2005:CUC

[SG05] Craig A. N. Soules and Gregory R. Ganger. Connections: using context to enhance file search. Operating Systems Review, 39(5): 119–132, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmidt:2010:VSB

[SG10a] René W. Schmidt and Steffen Grarup. vApp: a standards-based container for cloud providers. Operating Systems Review, 44(4): 115–123, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soundararajan:2010:CBS

[SG10b] Vijayaraghavan Soundararajan and Kinshuk Govil. Challenges in building scalable virtualized datacenter management. Operating Systems Review, 44(4):95–102, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stewart:2014:WDS

[SG14] Christopher Stewart and Vishakha Gupta. The Workshop on Diversity in Systems Research 2013. Operating Systems Review, 48(1):103–106, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saroiu:2002:AIC

[SGD^{+}02] Stefan Saroiu, Krishna P. Gummadi, Richard J. Dunn, Steven D. Gribble, and Henry M. Levy. An analysis of Internet content delivery systems. Operating Systems Review, 36(5S): 315–327, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sirer:1999:DID

[SGGB99] Emin Gün Sirer, Robert Grimm, Arthur J. Gregory, and Brian N. Bershad. Design and implementation of a distributed virtual machine for networked computers. Operating Systems Review, 33(5):202–216, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sirer:2000:DID

[SGGB00] Emin Gün Sirer, Robert Grimm, Arthur J. Gregory, and Brian N. Bershad. Design and implementation of a distributed virtual machine for networked computers. Operating Systems Review, 34(2):23, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smolens:2004:FBS

[SGK^{+}04] Jared C. Smolens, Brian T. Gold, Jangwoo Kim, Babak Falsafi, James C. Hoe, and Andreas G. Nowatzyk. Fingerprinting: bounding soft-error detection latency and bandwidth. Operating Systems Review, 38(5):224–234, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1985:CFS

[SGN85] Michael D. Schroeder, David K. Gifford, and Roger M. Needham. A caching file system for a programmer’s workstation. Operating Systems Review, 19(5):25–34, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schlosser:2000:DCS

[SGNG00] Steven W. Schlosser, John Linwood Griffin, David F. Nagle, and Gregory R. Ganger. Designing computer systems with MEMSbased storage. Operating Systems Review, 34(5):1–12, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scales:1996:SLO

[SGT96] Daniel J. Scales, Kourosh Gharachorloo, and Chandramohan A. Thekkath. Shasta: a low overhead, software-only approach for supporting fine-grain shared memory. Operating Systems Review, 30(5):174–185, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steenkiste:1987:TTC

[SH87] Peter Steenkiste and John Hennessy. Tags and type checking in LISP: hardware and software approaches. Operating Systems Review, 21(4):50–59, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sarkar:1996:ECC

[SH96] Prasenjit Sarkar and John Hartman. Efficient cooperative caching using hints. Operating Systems Review, 30(SI):35–46, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sonntag:2000:MAS

[SH00] Michael Sonntag and Rudolf Hörmanseder. Mobile agent security based on payment. Operating Systems Review, 34(4):48–55, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shapiro:1995:SSS

[Sha95] Marc Shapiro. Summary of the Sixth SIGOPS European Workshop on “Matching Operating Systems to Application Needs”. Operating Systems Review, 29(1):47–51, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shapiro:2000:LLW

[Sha00] Marc Shapiro. Lessons learned from a wide area sharing platform. Operating Systems Review, 34(2):29, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sasanka:2002:JLG

[SHA02] Ruchira Sasanka, Christopher J. Hughes, and Sarita V. Adve. Joint local and global hardware adaptations for energy. Operating Systems Review, 36(5):144–155, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spier:1973:EIK

[SHC73] Michale J. Spier, Thomas N. Hastings, and David N. Cutler. An experimental implementation of the kernel/domain architecture. Operating Systems Review, 7(4):8–21, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shinjo:2000:DCEb

[Shi00] Yasushi Shinjo. Developing correct and efficient multithreaded programs with thread-specific data and a partial evaluator. Operating Systems Review, 34(2):40, April 2000. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1985:IDF

[SHN^{+}85] M. Satyanarayanan, John H. Howard, David A. Nichols, Robert N. Sidebotham, Alfred Z. Spector, and Michael J. West. The ITC distributed file system: principles and design. Operating Systems Review, 19(5):35–50, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shalev:2016:CCS

[SHP^{+}16] Noam Shalev, Eran Harpaz, Hagar Porat, Idit Keidar, and Yaron Weinsberg. CSR: Core surprise removal in commodity operating systems. Operating Systems Review, 50(2):773–787, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shrivastava:1994:CCL

[Shr94] Santosh K. Shrivastava. To CATOCS or not to CATOCS, that is the . . .. Operating Systems Review, 28(4):11–14, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spier:1975:TME

[SHSB75] Michael J. Spier, Richard L. Hill, Timothy J. Stein, and Daniel Bricklin. The TYPESET-10 Message Exchange Facility: a case study in systemic design. Operating Systems Review, 9(1):10– 18, January 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shi:1997:ICP

[SHT97] Weisong Shi, Weiwu Hu, and Zhimin Tang. An interaction of coherence protocols and memory consistency models in DSM systems. Operating Systems Review, 31(4):41–54, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saunders:2001:RBA

[SHV01] G. Saunders, M. Hitchens, and V. Varadharajan. Role-based access control and the access control matrix. Operating Systems Review, 35(4):6–20, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sard:2015:PPC

[SHW^{+}15] Petter Särd, Benoit Hudzia, Steve Walsh, Johan Tordsson, and Erik Elmroth. Principles and performance characteristics of algorithms for live VM migration. Operating Systems Review, 49 (1):142–155, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sibley:1976:EJO

[Sib76] E. H. Sibley. Economic justification of an OSCL/OSRL. Operating Systems Review, 10(4):7–15, October 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Silverman:1983:RVS

[Sil83] Jonathan M. Silverman. Reflections on the verification of the security of an operating system kernel. Operating Systems Review, 17(5):143–154, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Singh:1985:IPS

[Sin85] Kamaljit Singh. On improvements to password security. Operating Systems Review, 19(1):53–60, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sirer:2006:I

[Sir06] Emin Gün Sirer. Introduction. Operating Systems Review, 40 (3):8, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steensgaard:1995:ONC

[SJ95] B. Steensgaard and E. Jul. Object and native code thread mobility among heterogeneous computers (includes sources). Operating Systems Review, 29(5):68–77, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Srinivas:2005:MCS

[SJ05] A. Vijay Srinivas and D. Janakiram. A model for characterizing the scalability of distributed systems. Operating Systems Review, 39(3):64–71, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seung-Ju:1994:SBS

[SJGY94] Jang Seung-Ju and Kim Gil-Yong. Spin-block synchronization algorithm in the shared memory multiprocessor system. Operating Systems Review, 28(4):15–30, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sauer:1987:RPD

[SJL^{+}87] Charles H. Sauer, Don W. Johnson, Larry K. Loucks, Amal A. Shaheen-Gouda, and Todd A. Smith. RT PC distributed services overview. Operating Systems Review, 21(3):18–29, July 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saxena:2023:NPE

[SJS^{+}23] Divyanshu Saxena, Tao Ji, Arjun Singhvi, Junaid Khalid, and Aditya Akella. Navigating performance–efficiency tradeoffs in serverless computing: Deduplication to the rescue! Operating Systems Review, 57(1):47–53, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3606557.3606564.

Seznec:1996:MBA

[SJSM96] André Seznec, Stéphan Jourdan, Pascal Sainrat, and Pierre Michaud. Multiple-block ahead branch predictors. Operating Systems Review, 30(5):116–127, December 1996. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Samadzadeh:1996:DAT

[SK96] M. H. Samadzadeh and B. S. Koshy. A display and analysis tool for process-resource graphs. Operating Systems Review, 30 (1):39–62, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stabell-Kulo:1997:SLS

[SK97] Tage Stabell-Kulø. Security and log structured file systems. Operating Systems Review, 31(2):9–10, April 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shraer:2013:DSR

[SK13] Alexander Shraer and Rüdiger Kapitza. Dagstuhl seminar report: security and dependability for federated cloud platforms, 2012. Operating Systems Review, 47(2):4–5, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seo:2017:FAS

[SKB^{+}17] Jihye Seo, Wook-Hee Kim, Woongki Baek, Beomseok Nam, and Sam H. Noh. Failure-atomic slotted paging for persistent memory. Operating Systems Review, 51(2):91–104, June 2017. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Srikantaiah:2008:ASP

[SKI08] Shekhar Srikantaiah, Mahmut Kandemir, and Mary Jane Irwin. Adaptive set pinning: managing shared caches in chip multiprocessors. Operating Systems Review, 42(2):135–144, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Song:2017:HBA

[SKJ^{+}17] Wonjun Song, Gwangsun Kim, Hyungjoon Jung, Jongwook Chung, Jung Ho Ahn, Jae W. Lee, and John Kim. History-based arbitration for fairness in processor-interconnect of NUMA servers. Operating Systems Review, 51(2):765–777, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saito:2002:TAR

[SKKM02] Yasushi Saito, Christos Karamanolis, Magnus Karlsson, and Mallik Mahalingam. Taming aggressive replication in the Pangaea wide-area file system. Operating Systems Review, 36(5S): 15–30, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spalink:2001:BRS

[SKPG01] Tammo Spalink, Scott Karlin, Larry Peterson, and Yitzchak Gottlieb. Building a robust software-based router using network processors. Operating Systems Review, 35(5):216–229, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stewart:2007:ENP

[SKZ07] Christopher Stewart, Terence Kelly, and Alex Zhang. Exploiting nonstationarity for performance prediction. Operating Systems Review, 41(3):31–44, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schnarr:1998:FOP

[SL98] Eric Schnarr and James R. Larus. Fast out-of-order processor simulation using memoization. Operating Systems Review, 32 (5):283–294, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shih:1989:STR

[SLCG89] W. Shih, J. S. W. Liu, J. Chung, and D. W. Gillies. Scheduling tasks with ready times and deadlines to minimize average error. Operating Systems Review, 23(3):14–28, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stanisic:2015:EGO

[SLD15] Luka Stanisic, Arnaud Legrand, and Vincent Danjean. An effective git and org-mode based workflow for reproducible research. Operating Systems Review, 49(1):61–70, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sui:2016:PCA

[SLFP16] Xin Sui, Andrew Lenharth, Donald S. Fussell, and Keshav Pingali. Proactive control of approximate programs. Operating Systems Review, 50(2):607–621, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strauss:2010:DTN

[SLLP^{+}10] Jacob Strauss, Chris Lesniewski-Laas, Justin Mazzola Paluska, Bryan Ford, Robert Morris, and Frans Kaashoek. Device transparency: a new model for mobile storage. Operating Systems Review, 44(1):5–9, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saha:2011:FRR

[SLM11] Suman Saha, Julia Lawall, and Gilles Muller. Finding resourcerelease omission faults in Linux. Operating Systems Review, 45 (3):5–9, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmidt:1999:IPS

[SLN99] Brian K. Schmidt, Monica S. Lam, and J. Duane Northcutt. The interactive performance of SLIM: a stateless, thin-client architecture. Operating Systems Review, 33(5):32–47, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schimdt:2000:IPS

[SLN00] Brian K. Schimdt, Monica S. Lam, and J. Duane Northcut. The interactive performance of SLIM: a stateless, thin-client architecture. Operating Systems Review, 34(2):12–13, April 2000.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seshadri:2007:STH

[SLQP07] Arvind Seshadri, Mark Luk, Ning Qu, and Adrian Perrig. SecVisor: a tiny hypervisor to provide lifetime kernel code integrity for commodity OSes. Operating Systems Review, 41(6):335–350, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seshadri:2005:PVC

[SLS^{+}05] Arvind Seshadri, Mark Luk, Elaine Shi, Adrian Perrig, Leendert van Doorn, and Pradeep Khosla. Pioneer: verifying code integrity and enforcing untampered code execution on legacy systems. Operating Systems Review, 39(5):1–16, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Solar-Lezama:2006:CSF

[SLTB^{+}06] Armando Solar-Lezama, Liviu Tancau, Rastislav Bodik, Sanjit Seshia, and Vijay Saraswat. Combinatorial sketching for finite programs. Operating Systems Review, 40(5):404–415, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Suh:2004:SPE

[SLZD04] G. Edward Suh, Jae W. Lee, David Zhang, and Srinivas Devadas. Secure program execution via dynamic information flow tracking. Operating Systems Review, 38(5):85–96, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Srinivasan:1989:SNE

[SM89] V. Srinivasan and J. Mogul. Spritely NFS: experiments with cache-consistency protocols. Operating Systems Review, 23(5): 44–57, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stanton:2010:FAD

[SMBA10] Paul T. Stanton, Benjamin McKeown, Randal Burns, and Giuseppe Ateniese. FastAD: an authenticated directory for billions of objects. Operating Systems Review, 44(1):45–49, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smith:1978:BPR

[Smi78] Alan Jay Smith. Bibliography on paging and related topics. Operating Systems Review, 12(4):39–56, October 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sturgis:1980:IDU

[SMI80] H. Sturgis, J. Mitchell, and J. Israel. Issues in the design and use of a distributed file system. Operating Systems Review, 14(3): 55–69, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smith:1988:SPM

[Smi88] Jonathan M. Smith. A survey of process migration mechanisms. Operating Systems Review, 22(3):28–40, July 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1993:LRV

[SMK^{+}93] M. Satyanarayanan, Henry H. Mashburn, Puneet Kumar, David C. Steere, and James J. Kistler. Lightweight recoverable virtual memory. Operating Systems Review, 27(5):146–160, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strong:2009:FST

[SMM^{+}09] Richard Strong, Jayaram Mudigonda, Jeffrey C. Mogul, Nathan Binkert, and Dean Tullsen. Fast switching of threads between cores. Operating Systems Review, 43(2):35–45, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smolik:1995:OOF

[Smo95] Tomas Smolik. An object-oriented file system—an example of using the class hierarchy framework concept. Operating Systems Review, 29(2):33–53, April 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Singh:2006:UQD

[SMRD06] Atul Singh, Petros Maniatis, Timothy Roscoe, and Peter Druschel. Using queries for distributed monitoring and forensics. Operating Systems Review, 40(4):389–402, October 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schiper:2011:SMK

[SMS11] André Schiper, Zarko Milosevic, and Omid Shahmirzadi. Student mini-kernel project based on an FPGA board. Operating Systems Review, 45(2):54–58, July 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shen:2009:SHP

[SMTZ09] Xipeng Shen, Feng Mao, Kai Tian, and Eddy Zheng Zhang. The study and handling of program inputs in the selection of garbage collectors. Operating Systems Review, 43(3):48–61, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmidt:1994:PHA

[SN94] William J. Schmidt and Kelvin D. Nilsen. Performance of a hardware-assisted real-time garbage collector. Operating Systems Review, 28(5):76–85, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Saur:2013:RFW

[SN13] Karla Saur and Iulian Neamtiu. Report on the Fourth Workshop on Hot Topics in Software Upgrades (HotSWUp 2012). Operating Systems Review, 47(1):55–62, January 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1995:AAA

[SNKP95] M. Satyanarayanan, Brian Noble, Puneet Kumar, and Morgan Price. Application-aware adaptation for mobile computing. Operating Systems Review, 29(1):52–55, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Scales:2010:DPS

[SNV10] Daniel J. Scales, Mike Nelson, and Ganesh Venkitachalam. The design of a practical system for fault-tolerant virtual machines. Operating Systems Review, 44(4):30–39, December 2010. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Snyder:1977:SAP

[Sny77] Lawrence Snyder. On the synthesis and analysis of protection systems. Operating Systems Review, 11(5):141–150, November 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Song:2005:SBP

[Son05] Jia Song. Segment-based proxy caching for distributed cooperative media content servers. Operating Systems Review, 39 (1):22–33, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sopka:1984:NPP

[Sop84] John R. Sopka. National parallel processing research council executive committee charter. Operating Systems Review, 18(3): 25–27, July 1984. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sorenson:1973:ICR

[Sor73] P. G. Sorenson. Interprocess communication in real-time systems. Operating Systems Review, 7(4):1–7, October 1973. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shinjo:2000:DCEa

[SP00] Yasushi Shinjo and Calton Pu. Developing correct and efficient multithreaded programs with thread-specific data and a partial evaluator. Operating Systems Review, 34(2):33, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spring:2006:UPN

[SPBP06] Neil Spring, Larry Peterson, Andy Bavier, and Vivek Pai. Using PlanetLab for network research: myths, realities, and best practices. Operating Systems Review, 40(1):17–24, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spector:1981:PRO

[Spe81] Alfred Z. Spector. Performing remote operations efficiently on a local computer network. Operating Systems Review, 15(5):76– 77, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soltesz:2007:CBO

[SPF^{+}07] Stephen Soltesz, Herbert Pötzl, Marc E. Fiuczynski, Andy Bavier, and Larry Peterson. Container-based operating system virtualization: a scalable, high-performance alternative to hypervisors. Operating Systems Review, 41(3):275–287, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sherwood:2002:ACL

[SPHC02] Timothy Sherwood, Erez Perelman, Greg Hamerly, and Brad Calder. Automatically characterizing large scale program behavior. Operating Systems Review, 36(5):45–57, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Singaravelu:2006:RTC

[SPHH06] Lenin Singaravelu, Calton Pu, Hermann Härtig, and Christian Helmuth. Reducing TCB complexity for security-sensitive applications: three case studies. Operating Systems Review, 40 (4):161–174, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spier:1974:CLS

[Spi74] Michael J. Spier. A critical look at the state of our science. Operating Systems Review, 8(2):9–15, April 1974. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spinellis:1994:TTL

[Spi94] Diomidis Spinellis. Trace: a tool for logging operating system call transactions. Operating Systems Review, 28(4):56–63, October 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spratt:1985:TRJ

[Spr85] Lindsey L. Spratt. The transaction resolution journal: extending the before journal. Operating Systems Review, 19(3):55–62, July 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sundaramoorthy:2000:SPI

[SPR00] Karthik Sundaramoorthy, Zach Purser, and Eric Rotenburg. Slipstream processors: improving both performance and fault tolerance. Operating Systems Review, 34(5):257–268, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Suleman:2008:FDT

[SQP08] M. Aater Suleman, Moinuddin K. Qureshi, and Yale N. Patt. Feedback-driven threading: power-efficient and highperformance execution of multi-threaded workloads on CMPs. Operating Systems Review, 42(2):277–286, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stankovic:1989:SKN

[SR89] J. A. Stankovic and K. Ramamritham. The Spring kernel: a new paradigm for real-time operating systems. Operating Systems Review, 23(3):54–71, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Srinivasan:2004:CFP

[SRA^{+}04] Srikanth T. Srinivasan, Ravi Rajwar, Haitham Akkary, Amit Gandhi, and Mike Upton. Continual flow pipelines. Operating Systems Review, 38(5):107–119, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spear:2006:SSP

[SRH^{+}06] Michael F. Spear, Tom Roeder, Orion Hodson, Galen C. Hunt, and Steven Levi. Solving the starting problem: device drivers as self-describing artifacts. Operating Systems Review, 40(4):45– 57, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shkuro:2022:PPS

[SRS22] Yuri Shkuro, Benjamin Renard, and Atul Singh. Positional paper: Schema-first application telemetry. Operating Systems Review, 56(1):8–17, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3544497.3544500.

Santana:2015:FSS

[SRTH15] Ricardo Santana, Raju Rangaswami, Vasily Tarasov, and Dean Hildebrand. A fast and slippery slope for file systems. Operating Systems Review, 49(2):27–34, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schroeder:1972:HAI

[SS72] Michael D. Schroeder and Jerome H. Saltzer. A hardware architecture for implementing protection rings. Operating Systems Review, 6(1/2):42–54, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmutz:1983:WSN

[SS83a] H. Schmutz and P. Silberbusch. Working sets and nearoptimality. Operating Systems Review, 17(3):24–29, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spector:1983:TCR

[SS83b] Alfred Z. Spector and Peter M. Schwarz. Transactions: a construct for reliable distributed computing. Operating Systems Review, 17(2):18–35, April 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Skeppstedt:1994:SCA

[SS94] Jonas Skeppstedt and Per Stenström. Simple compiler algorithms to reduce ownership overhead in cache coherence protocols. Operating Systems Review, 28(5):286–296, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shyu:1995:VAT

[SS95] Ing-Jye Shyu and Shiuh-Pyng Shieh. Virtual address translation for wide-address architectures. Operating Systems Review, 29 (4):37–46, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Satyanarayanan:1997:AWC

[SS97] M. Satyanarayanan and Mirjana Spasojevic. AFS and the Web: competitors or collaborators? Operating Systems Review, 31 (1):18–23, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sodani:1998:EAI

[SS98] Avinash Sodani and Gurindar S. Sohi. An empirical analysis of instruction repetition. Operating Systems Review, 32(5):35–45, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sierra:2000:NPE

[SS00] J. M. Sierra and S. J. Shepherd. New phase 1 exchange mode for IKE framework. Operating Systems Review, 34(4):34–40, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shalev:2006:PLS

[SS06] Ori Shalev and Nir Shavit. Predictive log-synchronization. Operating Systems Review, 40(4):305–315, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

So:2007:LBM

[SS07] Kelvin C. W. So and Emin Gün Sirer. Latency and bandwidthminimizing failure detectors. Operating Systems Review, 41(3): 89–99, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Soundararajan:2017:SFC

[SS17] Vijayaraghavan Soundararajan and Joshua Schnee. Sustainability as a first-class metric for developers and end-users. Operating Systems Review, 51(1):60–66, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shapiro:1999:EFC

[SSF99] Jonathan S. Shapiro, Jonathan M. Smith, and David J. Farber. EROS: a fast capability system. Operating Systems Review, 33 (5):170–185, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shapiro:2000:EFC

[SSF00] Jonathan S. Shapiro, Jonathan M. Smith, and David J. Farber. EROS: a fast capability system. Operating Systems Review, 34(2):21–22, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sundararajah:2017:LTN

[SSK17] Kirshanthan Sundararajah, Laith Sakka, and Milind Kulkarni. Locality transformations for nested recursive iteration spaces. Operating Systems Review, 51(2):281–295, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sripanidkulchai:2010:CRL

[SSR^{+}10a] Kunwadee Sripanidkulchai, Sambit Sahu, Yaoping Ruan, Anees Shaikh, and Chitra Dorai. Are clouds ready for large distributed applications? Operating Systems Review, 44(2):18–23, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sundararaman:2010:WPI

[SSR^{+}10b] Swaminathan Sundararaman, Sriram Subramanian, Abhishek Rajimwale, Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, and Michael M. Swift. Why panic()?: improving reliability with restartable file systems. Operating Systems Review, 44(1):25–29, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shriram:2001:IMP

[SSS01] Alok Shriram, Anuraag Sarangi, and Avinash S. ICHU model for processor allocation in distributed operating systems. Operating Systems Review, 35(3):16–21, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Spreitzer:1993:PLI

[ST93] Mike Spreitzer and Marvin Theimer. Providing location information in a ubiquitous computing environment (panel session). Operating Systems Review, 27(5):270–283, December 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Snavely:2000:SJS

[ST00] Allan Snavely and Dean M. Tullsen. Symbiotic jobscheduling for a simultaneous multithreaded processor. Operating Systems Review, 34(5):234–244, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Suranauwarat:2001:DII

[ST01] Sukanya Suranauwarat and Hideo Taniguchi. The design, implementation and initial evaluation of an advanced knowledgebased process scheduler. Operating Systems Review, 35(4):61– 81, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Staff:1983:RPS

[Sta83] Staff. Review of “Probability and statistics with reliability, queueing and computer science applications” by Kishor S. Trivedi. Prentice-Hall, Englewood-Cliffs, 1982. Operating Systems Review, 17(1):9, January 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Tri82, Wai83b, Tri02].

Stephenson:1973:SCC

[Ste73] C. J. Stephenson. On the structure and control of commands. Operating Systems Review, 7(4):22–26, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stephenson:1983:NMD

[Ste83] C. J. Stephenson. New methods for dynamic storage allocation (Fast Fits). Operating Systems Review, 17(5):30–32, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steere:1997:END

[Ste97] David C. Steere. Exploiting the non-determinism and asynchrony of set iterators to reduce aggregate file I/O latency. Operating Systems Review, 31(5):252–263, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sheth:2007:SDL

[STM^{+}07] Anmol Sheth, Chandramohan A. Thekkath, Prakshep Mehta, Kalyan Tejaswi, Chandresh Parekh, Trilok N. Singh, and Uday B. Desai. Senslide: a distributed landslide prediction system. Operating Systems Review, 41(2):75–87, April 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stonebraker:1984:VMT

[Sto84] Michael Stonebraker. Virtual memory transaction management. Operating Systems Review, 18(2):8–16, April 1984. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stoess:2007:TEU

[Sto07] Jan Stoess. Towards effective user-controlled scheduling for microkernel-based systems. Operating Systems Review, 41(4): 59–68, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stroustrup:1978:UMI

[Str78] Bjarne Stroustrup. On unifying module interfaces. Operating Systems Review, 12(1):90–98, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stroud:1993:TRD

[Str93] Robert Stroud. Transparency and reflection in distributed systems. Operating Systems Review, 27(2):99–103, April 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strunk:2012:HAC

[Str12] John D. Strunk. Hybrid aggregates: combining SSDs and HDDs in a single storage pool. Operating Systems Review, 46(3):50– 56, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steiner:1995:REE

[STW95] Michael Steiner, Gene Tsudik, and Michael Waidner. Refinement and extension of encrypted key exchange. Operating Systems Review, 29(3):22–30, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shen:2002:IRM

[STYC02] Kai Shen, Hong Tang, Tao Yang, and Lingkun Chu. Integrated resource management for cluster-based Internet services. Operating Systems Review, 36(5S):225–238, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Subramanian:2011:OAS

[Sub11] Lakshminarayanan Subramanian. Overview of the 3rd ACM SOSP Workshop on Networking, Systems and Applications on Mobile Handhelds. Operating Systems Review, 45(3):63–64, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schuchman:2006:PTA

[SV06] Ethan Schuchman and T. N. Vijaykumar. A program transformation and architecture support for quantum uncomputation. Operating Systems Review, 40(5):252–263, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Svigals:1983:PFM

[Svi83] Jerome Svigals. Planning for future market events using data processing support: a five-step growth plan process. Macmillan Publishing Company, New York, NY, USA, 1983. ISBN 0-02-949740-X. xii + 180 pp. LCCN HG1709 .S87 1983.

Svobodova:1973:OSP

[Svo73] Liba Svobodova. Online system performance measurements with software and hybrid monitors. Operating Systems Review, 7(4):45–53, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Svobodova:1981:PMC

[Svo81a] Liba Svobodova. Performance monitoring in computer systems: a structured approach. Operating Systems Review, 15(3):39–50, July 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Svobodova:1981:ROO

[Svo81b] Liba Svobodova. A reliable object-oriented data repository for a distributed computer system. Operating Systems Review, 15 (5):47–58, December 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schmuck:1991:ETQ

[SW91] Frank Schmuck and Jim Wylie. Experience with transactions in QuickSilver. Operating Systems Review, 25(5):239–253, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sun:2000:AFA

[SW00] Yongxing Sun and Xinmei Wang. An approach to finding the attacks on the cryptographic protocols. Operating Systems Review, 34(3):19–28, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Siegenthaler:2010:CSC

[SW10] Michael Siegenthaler and Hakim Weatherspoon. Cloudifying source code repositories: how much does it cost? Operating Systems Review, 44(2):24–28, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Steinder:2008:SVA

[SWC08] Małgorzata Steinder, Ian Whalley, and David Chess. Server virtualization in autonomic management of heterogeneous workloads. Operating Systems Review, 42(1):94–95, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shaw:1977:AVA

[SWL77] Mary Shaw, Wm A. Wulf, and Ralph L. London. Abstraction and verification in Alphard: Defining and specifying iteration and generators. Operating Systems Review, 11(2):139, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sun:1988:DHD

[SXZ^{+}88] Zhongxiu Sun, Xing Xue, Jianqiang Zhou, Peigen Yang, and Xihao Xu. Developing a heterogeneous distributed operating system. Operating Systems Review, 22(2):24–31, April 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shieh:1996:AKD

[SY96] Shiuh-Pyng Shieh and Wen-Her Yang. An authentication and key distribution system for open network systems. Operating Systems Review, 30(2):32–41, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Syverson:1993:KDP

[Syv93] Paul Syverson. On key distribution protocols for repeated authentication. Operating Systems Review, 27(4):24–30, October 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Syverson:1996:NLO

[Syv96] Paul F. Syverson. A new look at an old protocol. Operating Systems Review, 30(3):1–4, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schwan:1992:MRT

[SZ92] Karsten Schwan and Hongyi Zhou. Multiprocessor real-time threads. Operating Systems Review, 26(1):54–65, January 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Seidl:1998:SHO

[SZ98] Matthew L. Seidl and Benjamin G. Zorn. Segregating heap objects by reference behavior and lifetime. Operating Systems Review, 32(5):12–23, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shen:2004:LPP

[SZD04] Xipeng Shen, Yutao Zhong, and Chen Ding. Locality phase prediction. Operating Systems Review, 38(5):165–176, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shen:2008:HCD

[SZD^{+}08] Kai Shen, Ming Zhong, Sandhya Dwarkadas, Chuanpeng Li, Christopher Stewart, and Xiao Zhang. Hardware counter driven on-the-fly request signatures. Operating Systems Review, 42 (2):189–200, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schwan:1991:RTT

[SZG91] Karsten Schwan, Hongyi Zhou, and Ahmed Gheith. Real-time threads. Operating Systems Review, 25(4):35–46, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Srinivasan:2011:EIB

[SZII11] Sadagopan Srinivasan, Li Zhao, Ramesh Illikkal, and Ravishankar Iyer. Efficient interaction between OS and architecture in heterogeneous platforms. Operating Systems Review, 45(1):62– 72, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Schwartz:1987:NSE

[SZN87] M. Schwartz, J. Zahorjan, and D. Notkin. A name service for evolving heterogeneous systems. Operating Systems Review, 21 (5):52–62, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tay:1990:SRP

[TA90] B. H. Tay and A. L. Ananda. A survey of remote procedure calls. Operating Systems Review, 24(3):68–79, July 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tuck:2008:SSE

[TACT08] James Tuck, Wonsun Ahn, Luis Ceze, and Josep Torrellas. Soft-Sig: software-exposed hardware signatures for code analysis and optimization. Operating Systems Review, 42(2):145–156, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Taft:1982:OBV

[Taf82] S. Tucker Taft. An object-based virtual operating system for the Ada programming support environment. Operating Systems Review, 16(1):14–25, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Toslali:2022:VVD

[TAH^{+}22] Mert Toslali, Emre Ates, Darby Huye, Zhaoqi Zhang, Lan Liu, Samantha Puterman, Ayse K. Coskun, and Raja R. Sambasivan. VAIF: Variance-driven automated instrumentation framework. Operating Systems Review, 56(1):42–50, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL

https://dl.acm.org/doi/10.1145/3544497.3544504.

Taifi:2013:BDB

[Tai13] Moussa Taifi. Banking on decoupling: budget-driven sustainability for HPC applications on auction-based clouds. Operating Systems Review, 47(2):41–50, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:1979:MIP

[Tan79] Andrew S. Tanenbaum. A method for implementing paged, segmented virtual memories on microprogrammable computers. Operating Systems Review, 13(2):26–32, April 1979. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:1987:UCS

[Tan87] Andrew S. Tanenbaum. A UNIX clone with source code for operating systems courses. Operating Systems Review, 21(1):20– 29, January 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:1997:RSA

[Tan97] Andrew S. Tanenbaum. Report on the Seventh ACM SIGOPS European Workshop: Systems Support for Worldwide Applications. Operating Systems Review, 31(1):5–17, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tam:2007:TCS

[TAS07] David Tam, Reza Azimi, and Michael Stumm. Thread clustering: sharing-aware scheduling on SMP–CMP–SMT multiprocessors. Operating Systems Review, 41(3):47–58, June 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tournier:2006:TFD

[TBM^{+}06] Jean-Charles Tournier, Patrick G. Bridges, Arthur B. MacCabe, Patrick M. Widener, Zaid Abudayyeh, Ron Brightwell, Rolf Riesen, and Trammel Hudson. Towards a framework for dedicated operating systems development in high-end computing systems. Operating Systems Review, 40(2):16–21, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tai:1996:VNO

[TC96] K. C. Tai and Richard H. Carver. VP: a new operation for semaphores. Operating Systems Review, 30(3):5–11, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tangney:1991:SIS

[TCH^{+}91] Brendan Tangney, Vinny Cahill, Chris Horn, Dominic Herity, Alan Judge, Gradimir Starovic, and Mark Sheppard. Some ideas on support for fault tolerance in COMANDOS, an object oriented distributed system. Operating Systems Review, 25 (2):130–135, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Takouna:2012:EES

[TDM12] Ibrahim Takouna, Wesam Dawoud, and Christoph Meinel. Energy efficient scheduling of HPC-jobs on virtualize clusters using host and VM dynamic configuration. Operating Systems Review, 46(2):19–27, July 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thekkath:1994:EMH

[TE94] Radhika Thekkath and Susan J. Eggers. The effectiveness of multiple hardware contexts. Operating Systems Review, 28(5): 328–337, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Temam:1998:IOL

[Tem98] Olivier Temam. Investigating optimal local memory performance. Operating Systems Review, 32(5):218–227, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tennenhouse:1996:ANA

[Ten96] David Tennenhouse. Active networks (abstract). Operating Systems Review, 30(SI):89, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tennenhouse:2017:RV

[Ten17] David Tennenhouse. Research at VMware. Operating Systems Review, 51(1):1–4, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Terry:2014:RFT

[Ter14] Doug Terry. A report on the First TRIOS Conference. Operating Systems Review, 48(2):26–34, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tetzlaff:2014:SPT

[Tet14] William Tetzlaff. SOSP Professional Travel Scholarship: Reflections by recipient William Tetzlaff. Operating Systems Review, 48(2):23, July 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tsegaye:2004:CLW

[TF04] Melekam Tsegaye and Richard Foss. A comparison of the Linux and Windows device driver architectures. Operating Systems Review, 38(2):8–33, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Toinard:1999:FMP

[TFC99] C. Toinard, G. Florin, and C. Carrez. A formal method to prove ordering properties of multicast systems. Operating Systems Review, 33(4):75, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tucker:1989:PCS

[TG89] A. Tucker and A. Gupta. Process control and scheduling issues for multiprogrammed shared-memory multiprocessors. Operating Systems Review, 23(5):159–166, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thomas:2021:PFL

[TGR^{+}21] Luis Thomas, Sebastien Gougeaud, Stéphane Rubini, Philippe Deniel, and Jalil Boukhobza. Predicting file lifetimes for data placement in multi-tiered storage systems for HPC. Operating Systems Review, 55(1):99–107, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3469379.3469392.

Talluri:1994:STP

[TH94] Madhusudhan Talluri and Mark D. Hill. Surpassing the TLB performance of superpages with less operating system support. Operating Systems Review, 28(5):171–182, December 1994. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:2006:FSD

[THB06] Andrew S. Tanenbaum, Jorrit N. Herder, and Herbert Bos. File size distribution on UNIX systems: then and now. Operating Systems Review, 40(1):100–104, January 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Talluri:1995:NPT

[THK95] M. Talluri, M. D. Hill, and Y. A. Khalidi. A new page table for 64-bit address spaces. Operating Systems Review, 29(5):184–200, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Toll:2008:CSE

[TKP^{+}08] David C. Toll, Paul A. Karger, Elaine R. Palmer, Suzanne K. McIntosh, and Sam Weber. The Caernarvon secure embedded operating system. Operating Systems Review, 42(1):32–39, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thekkath:1994:HSS

[TL94] Chandramohan A. Thekkath and Henry M. Levy. Hardware and software support for efficient exception handling. Operating Systems Review, 28(5):110–119, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tsai:1996:RTS

[TL96] Wen-Jiin Tsai and Suh-Yin Lee. Real-time scheduling of multimedia data retrieval to minimize buffer requirement. Operating Systems Review, 30(3):67–80, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Theimer:1985:PRE

[TLC85] Marvin M. Theimer, Keith A. Lantz, and David R. Cheriton. Preemptable remote execution facilities for the V-system. Operating Systems Review, 19(5):2–12, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tartler:2011:CCA

[TLD^{+}11] Reinhard Tartler, Daniel Lohmann, Christian Dietrich, Christoph Egger, and Julio Sincero. Configuration coverage in the analysis of large-scale system software. Operating Systems Review, 45 (3):10–14, December 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tucek:2007:TDP

[TLH^{+}07] Joseph Tucek, Shan Lu, Chengdu Huang, Spiros Xanthos, and Yuanyuan Zhou. Triage: diagnosing production run failures at the user’s site. Operating Systems Review, 41(6):131–144, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thekkath:1994:SDC

[TLL94] Chandramohan A. Thekkath, Henry M. Levy, and Edward D. Lazowska. Separating data and control transfer in distributed operating systems. Operating Systems Review, 28(5):2–11, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tse:2003:RAO

[TLL03] K. W. Tse, W. K. Lam, and P. K. Lun. Reservation aware operating system for grid economy. Operating Systems Review, 37(3):36–42, July 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tanenbaum:1981:OAD

[TM81] Andrew S. Tanenbaum and Sape J. Mullender. An overview of the Amoeba distributed operating system. Operating Systems Review, 15(3):51–64, July 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tokuda:1989:ADR

[TM89] H. Tokuda and C. W. Mercer. ARTS: a distributed real-time kernel. Operating Systems Review, 23(3):29–53, July 1989. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thekkath:1997:FSD

[TML97] Chandramohan A. Thekkath, Timothy Mann, and Edward K. Lee. Frangipani: a scalable distributed file system. Operating Systems Review, 31(5):224–237, December 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thekkath:2000:ASC

[TML^{+}00] David Lie Chandramohan Thekkath, Mark Mitchell, Patrick Lincoln, Dan Boneh, John Mitchell, and Mark Horowitz. Architectural support for copy and tamper resistant software. Operating Systems Review, 34(5):168–177, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Trippel:2017:TMM

[TML^{+}17] Caroline Trippel, Yatin A. Manerkar, Daniel Lustig, Michael Pellauer, and Margaret Martonosi. TriCheck: Memory model verification at the trisection of software, hardware, and ISA. Operating Systems Review, 51(2):119–133, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Triplett:2010:SCH

[TMW10] Josh Triplett, Paul E. McKenney, and Jonathan Walpole. Scalable concurrent hash tables via relativistic programming. Operating Systems Review, 44(3):102–109, July 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tran:2012:TSA

[TNA12] Viet-Trung Tran, Bogdan Nicolae, and Gabriel Antoniu. Towards scalable array-oriented active storage: the pyramid approach. Operating Systems Review, 46(1):19–25, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tucek:2007:SLE

[TNL^{+}07] Joseph Tucek, James Newsome, Shan Lu, Chengdu Huang, Spiros Xanthos, David Brumley, Yuanyuan Zhou, and Dawn Song. Sweeper: a lightweight end-to-end system for defending against fast worms. Operating Systems Review, 41(3):115–128, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Taki:1987:PAE

[TNNI87] Kazuo Taki, Katzuto Nakajima, Hiroshi Nakashima, and Morihiro Ikeda. Performance and architectural evaluation of the PSI machine. Operating Systems Review, 21(4):128–135, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Toinard:1992:NWD

[Toi92] G. Florin C. Toinard. A new way to design causally and totally ordered multicast protocols. Operating Systems Review, 26(4):77–83, October 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tomlinson:1975:SSN

[Tom75] Raymond S. Tomlinson. Selecting sequence numbers. Operating Systems Review, 9(3):11–23, July 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Teorey:1972:CAD

[TP72] Toby J. Teorey and Tad B. Pinkerton. A comparative analysis of disk scheduling policies. Operating Systems Review, 6(1/2): 114–121, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tatemura:2012:MDA

[TPH12] Junichi Tatemura, Oliver Po, and Hakan Hacgümüs. Microsharding: a declarative approach to support elastic OLTP workloads. Operating Systems Review, 46(1):4–11, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tarditi:2006:AUD

[TPO06] David Tarditi, Sidd Puri, and Jose Oglesby. Accelerator: using data parallelism to program GPUs for general-purpose uses.

Operating Systems Review, 40(5):325–335, December 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Traiger:1982:VMM

[Tra82] Irving L. Traiger. Virtual memory management for database systems. Operating Systems Review, 16(4):26–48, October 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Trivedi:1982:PSR

[Tri82] Kishor Shridharbhai Trivedi. Probability and statistics with reliability, queuing, and computer science applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-711564-4. x + 624 pp. LCCN QA273.19.E4 T74 1982.

Trivedi:2002:PSR

[Tri02] Kishor Shridharbhai Trivedi. Probability and statistics with reliability, queuing, and computer science applications. Wiley, New York, NY, USA, second edition, 2002. ISBN 0-471-33341-7 (cloth). xv + 830 pp. LCCN QA273.19.E4 T74

2002. URL http://www.loc.gov/catdir/bios/wiley046/
2001026951.html;   http://www.loc.gov/catdir/description/
wiley0310/2001026951.html;   http://www.loc.gov/catdir/
toc/onix05/2001026951.html.

Trono:2000:CTS

[Tro00] John A. Trono. Comments on “tagged semaphores”. Operating Systems Review, 34(4):7–11, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Xu00].

Terry:1987:MSV

[TS87a] D. Terry and D. Swinehart. Managing stored voice in the Etherphone system. Operating Systems Review, 21(5):103–104, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Thacker:1987:FMW

[TS87b] Charles P. Thacker and Lawrence C. Stewart. Firefly: a multiprocessor workstation. Operating Systems Review, 21(4):164– 172, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Teller:2006:IPD

[TS06] Patricia J. Teller and Seetharami R. Seelam. Insights into providing dynamic adaptation of operating system policies. Operating Systems Review, 40(2):83–89, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tsafrir:2016:SAW

[Tsa16] Dan Tsafrir. Synopsis of the ASPLOS ’16 Wild and Crazy Ideas (WACI) invited-speakers session. Operating Systems Review, 50 (2):291–294, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Theimer:2000:TMW

[TSE^{+}00] Mark Theimer, M. Satyanarayanan, Maria Ebling, Mary Baker, Frans Kaashoek, Jay Lepreau, Andrew Black, and Carla Ellis. Tribute to Mark Weiser (summary only). Operating Systems Review, 34(2):6–7, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tam:1990:TBC

[TSF90] Ming-Chit Tam, Jonathan M. Smith, and David J. Farber. A taxonomy-based comparison of several distributed shared memory systems. Operating Systems Review, 24(3):40–67, July 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ta-Shma:2008:VMT

[TSLBYF08] Paula Ta-Shma, Guy Laden, Muli Ben-Yehuda, and Michael Factor. Virtual machine time travel using continuous data protection and checkpointing. Operating Systems Review, 42(1): 127–134, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tu:2017:BEO

[TSP17] Cheng-Chun Tu, Joe Stringer, and Justin Pettit. Building an extensible Open vSwitch datapath. Operating Systems Review, 51(1):72–77, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Trono:2000:FCC

[TT00] John A. Trono and William E. Taylor. Further comments on “a correct and unrestrictive implementation of general semaphores”. Operating Systems Review, 34(3):5–10, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Hem88, Kea88, Hem89].

Terry:1995:MUC

[TTP^{+}95] D. B. Terry, M. M. Theimer, Karin Petersen, A. J. Demers, M. J. Spreitzer, and C. H. Hauser. Managing update conflicts in Bayou, a weakly connected replicated storage system. Operating Systems Review, 29(5):172–182, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tuggle:1983:RPF

[Tug83] J. L. Tuggle. Review of “planning for future market events using data processing support by Jerome Svigals.’’ Macmillan Inc. 1983. Operating Systems Review, 17(3):8, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Svi83].

Turton:1980:MOS

[Tur80] Trevor Turton. The management of operating system state data. Operating Systems Review, 14(2):21–24, April 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Turnbull:1987:SHG

[Tur87] Martin Turnbull. Support for heterogeneity in the global distributed operating system. Operating Systems Review, 21(2): 11–22, April 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tang:2005:DIL

[TWL05] Yan Tang, Tao Wang, and Xiaoming Li. The design and implementation of LilyTask in shared memory. Operating Systems Review, 39(3):52–63, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tan:2007:IBB

[TYKZ07] Lin Tan, Ding Yuan, Gopal Krishna, and Yuanyuan Zhou.

/*icomment:  bugs or bad comments?*/. Operating Systems

Review, 41(6):145–158, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tseng:2018:MEP

[TZZ^{+}18] Hung-Wei Tseng, Qianchen Zhao, Yuxiao Zhou, Mark Gahagan, and Steven Swanson. Morpheus: Exploring the potential of near-data processing for creating application objects in heterogeneous computing. Operating Systems Review, 52(1):71–83, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Uhlig:2007:MKS

[Uhl07] Volkmar Uhlig. The mechanics of in-kernel synchronization for a scalable microkernel. Operating Systems Review, 41(4):49–58, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Upton:1994:RAH

[UHMB94] Michael Upton, Thomas Huff, Trevor Mudge, and Richard Brown. Resource allocation in a high clock rate microprocessor. Operating Systems Review, 28(5):98–109, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ugur:2022:OPF

[UJE^{+}22] Muhammed Ugur, Cheng Jiang, Alex Erf, Tanvir Ahmed Khan, and Baris Kasikci. One profile fits all: Profile-guided Linux kernel optimizations for data center applications. Operating Systems Review, 56(1):26–33, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3544497.3544502.

Ullman:1973:PCS

[Ull73] J. D. Ullman. Polynomial complete scheduling problems. Operating Systems Review, 7(4):96–101, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Uhlig:1994:TDS

[UNMS94] Richard Uhlig, David Nagle, Trevor Mudge, and Stuart Sechrest. Trap-driven simulation with Tapeworm II. Operating Systems Review, 28(5):132–144, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Urgaonkar:2002:ROA

[USR02] Bhuvan Urgaonkar, Prashant Shenoy, and Timothy Roscoe. Resource overbooking and application profiling in shared hosting platforms. Operating Systems Review, 36(5S):239–254, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varadharajan:1996:JAB

[VA96] Vijay Varadharajan and Phillip Allen. Joint actions based authorization schemes. Operating Systems Review, 30(3):32–45, July 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vaghani:2010:VMF

[Vag10] Satyam B. Vaghani. Virtual machine file system. Operating Systems Review, 44(4):57–70, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vasudevan:2011:COE

[VAK^{+}11] Vijay Vasudevan, David G. Andersen, Michael Kaminsky, Jason Franklin, Michael A. Kozuch, Iulian Moraru, Padmanabhan Pillai, and Lawrence Tan. Challenges and opportunities for efficient computing with FAWN. Operating Systems Review, 45 (1):34–44, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Valentic:1994:BRH

[Val94] Todd Valentic. Book review: High-Speed Windows Applications: Multitasking Design Methods by Bruce E. Krell: (Bantam Books, New York 1993). Operating Systems Review, 28(2):4–5, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanMeter:1996:BSC

[Van96] Rodney Van Meter. A brief survey of current work on network attached peripherals (extended abstract). Operating Systems Review, 30(1):63–70, January 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanHensbergen:2006:PRP

[Van06] Eric Van Hensbergen. P.R.O.S.E.: partitioned reliable operating system environment. Operating Systems Review, 40(2):12–15, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varney:1972:PSH

[Var72] R. C. Varney. Process selection in a hierarchical operating system. Operating Systems Review, 6(1/2):106–108, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varadharajan:1997:ESP

[Var97] Vijay Varadharajan. Extending the Schematic Protection Model II: revocation. Operating Systems Review, 31(1):64–77, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vigfusson:2010:OIF

[VBHN10] Ymir Vigfusson, Ken Birman, Qi Huang, and Deepak P. Nataraj. Optimizing information flow in the gossip objects platform. Operating Systems Review, 44(2):71–76, April 2010. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vandiver:2007:TBF

[VBLM07] Ben Vandiver, Hari Balakrishnan, Barbara Liskov, and Sam Madden. Tolerating Byzantine faults in transaction processing systems using commit barrier scheduling. Operating Systems Review, 41(6):59–72, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Verghese:1996:OSS

[VDGR96] Ben Verghese, Scott Devine, Anoop Gupta, and Mendel Rosenblum. Operating system support for improving data locality on CC-NUMA compute servers. Operating Systems Review, 30(5): 279–289, December 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanderWijngaart:2011:LWC

[vdWMH11] Rob F. van der Wijngaart, Timothy G. Mattson, and Werner Haas. Light-weight communications on Intel’s single-chip cloud computer processor. Operating Systems Review, 45(1):73–83, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanErtvelde:2008:DPA

[VE08] Luk Van Ertvelde and Lieven Eeckhout. Dispersing proprietary applications as benchmarks through code mutation. Operating Systems Review, 42(2):201–210, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vonEicken:1995:UNU

[vEBBV95] T. von Eicken, A. Basu, V. Buch, and W. Vogels. U-Net: a user-level network interface for parallel and distributed computing (includes URL). Operating Systems Review, 29(5):40–53, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanHensbergen:2010:UEM

[VESM10] Eric Van Hensbergen, Noah Paul Evans, and Phillip Stanley-Marbell. A unified execution model for cloud computing. Operating Systems Review, 44(2):12–17, April 2010. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanMeter:1998:VNV

[VFH98] Rodney Van Meter, Gregory G. Finn, and Steve Hotz. VISA: Netstation’s virtual Internet SCSI adapter. Operating Systems Review, 32(5):71–80, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

VanHensbergen:2008:HAR

[VFMM08] Eric Van Hensbergen, Charles Forsyth, Jim McKie, and Ron Minnich. Holistic aggregate resource environment. Operating Systems Review, 42(1):85–91, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vippagunta:2022:POP

[VFP22] Srinivas Vippagunta, Ken Finnigan, and Kishore Pusukuri. Pharos: The observability platform at workday. Operating Systems Review, 56(1):51–54, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.

acm.org/doi/10.1145/3544497.3544505.

Vogt:2014:TEM

[VGBT14] Dirk Vogt, Cristiano Giuffrida, Herbert Bos, and Andrew S. Tanenbaum. Techniques for efficient in-memory checkpointing.

Operating Systems Review, 48(1):21–25, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Verghese:1998:PIS

[VGR98] Ben Verghese, Anoop Gupta, and Mendel Rosenblum. Performance isolation: sharing and isolation in shared-memory multiprocessors. Operating Systems Review, 32(5):181–192, December 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vora:2017:KFA

[VGX17] Keval Vora, Rajiv Gupta, and Guoqing Xu. KickStarter: Fast and accurate computations on streaming graphs via trimmed approximations. Operating Systems Review, 51(2):237–251, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vonk:2025:GPS

[VHAA25] Robin Vonk, Joost Hoozemans, and Zaid AI-Ars. GSST: Parallel string decompression at 191 GB/s on GPU. Operating Systems Review, 59(2):55–61, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vengerov:2019:MLA

[VJ19] David Vengerov and Sesh Jalagam. A machine learning approach to recommending files in a collaborative work environment. Operating Systems Review, 53(1):46–51, July 2019. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Venkataramani:2002:TNM

[VKD02] Arun Venkataramani, Ravi Kokku, and Mike Dahlin. TCP Nice: a mechanism for background transfers. Operating Systems Review, 36(5S):329–343, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varghese:1987:HHT

[VL87] G. Varghese and T. Lauck. Hashed and hierarchical timing wheels: data structures for the efficient implementation of a timer facility. Operating Systems Review, 21(5):25–38, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vazhkudai:2007:RTD

[VM07] Sudharshan Vazhkudai and Xiaosong Ma. Recovering transient data: automated on-demand data reconstruction and offloading for supercomputers. Operating Systems Review, 41(1):14–18, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vicente:2012:ECS

[VMBM12] Elder Vicente, Rivalino Matias, Lúcio Borges, and Autran Macêdo. Evaluation of compound system calls in the Linux kernel. Operating Systems Review, 46(1):53–63, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vrable:2005:SFC

[VMC^{+}05] Michael Vrable, Justin Ma, Jay Chen, David Moore, Erik Vandekieft, Alex C. Snoeren, Geoffrey M. Voelker, and Stefan Savage. Scalability, fidelity, and containment in the Potemkin Virtual Honeyfarm. Operating Systems Review, 39(5):148–162, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vasudevan:2020:UPP

[VMM20] Amit Vasudevan, Petros Maniatis, and Ruben Martins. überSpark: Practical, provable, end-to-end guarantees on commodity heterogeneous interconnected computing platforms. Operating Systems Review, 54(1):8–22, August 2020. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL

https://dl.acm.org/doi/10.1145/3421473.3421476.

Voelker:1998:RSP

[Voe98] Geoffrey M. Voelker. Report on the SIGMETRICS’98/ PERFORMANCE’98 Joint International Conference on Measurement and Modeling of Computer Systems. Operating Systems Review, 32(4):3–8, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vogt:1997:VUS

[Vog97] Carsten Vogt. Visualizing UNIX synchronization operations. Operating Systems Review, 31(3):52–64, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vogels:1999:FSU

[Vog99] Werner Vogels. File system usage in Windows NT 4.0. Operating Systems Review, 33(5):93–109, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vogels:2000:FSU

[Vog00] Werner Vogels. File system usage in Windows NT 4.0. Operating Systems Review, 34(2):17–18, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varia:2015:AAS

[VPH^{+}15] Mayank Varia, Benjamin Price, Nicholas Hwang, Ariel Hamlin, Jonathan Herzog, Jill Poland, Michael Reschly, Sophia Yakoubov, and Robert K. Cunningham. Automated assessment of secure search systems. Operating Systems Review, 49(1):22–30, January 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanRennesse:1992:DIM

[vR92] Robbert van Rennesse. Design and Implementation of a Multicast Transport Service. Operating Systems Review, 26(2):31, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanRenesse:1993:CCM

[vR93] Robbert van Renesse. Causal controversy at Le Mont St.-Michel. Operating Systems Review, 27(2):44–53, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanRenesse:1994:WBC

[vR94] Robbert van Renesse. Why bother with CATOCS? Operating Systems Review, 28(1):22–27, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanRenesse:2014:SBF

[vR14] Robbert van Renesse. The story behind the first SIGOPS Dennis M. Ritchie Doctoral Dissertation Award. Operating Systems Review, 48(1):98–102, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vanRenesse:1988:PWF

[vRvST88] Robbert van Renesse, Hans van Staveren, and Andrew S. Tanenbaum. Performance of the world’s fastest distributed operating system. Operating Systems Review, 22(4):25–34, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Venkat:2016:HHI

[VSST16] Ashish Venkat, Sriskanda Shamasunder, Hovav Shacham, and Dean M. Tullsen. HIPStR: Heterogeneous-ISA program state relocation. Operating Systems Review, 50(2):727–741, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vakali:2001:MDS

[VT01] Athena Vakali and Evimaria Terzi. Multimedia data storage and representation issues on tertiary storage subsystems: an overview. Operating Systems Review, 35(2):61–77, April 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vora:2017:CCR

[VTGH17] Keval Vora, Chen Tian, Rajiv Gupta, and Ziang Hu. CoRAL: Confined recovery in distributed asynchronous graph processing. Operating Systems Review, 51(2):223–236, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varman:2008:SVP

[VW08] Peter Varman and Jun Wang. Storage and I/O virtualization, performance, energy, evaluation and dependability (SPEED08). Operating Systems Review, 42(6):1–2, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vahdat:2002:SAL

[VYW^{+}02] Amin Vahdat, Ken Yocum, Kevin Walsh, Priya Mahadevan, Dejan Kostić, Jeff Chase, and David Becker. Scalability and accuracy in a large-scale network emulator. Operating Systems Review, 36(5S):271–284, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vaswani:1991:ICA

[VZ91] Raj Vaswani and John Zahorjan. The implications of cache affinity on processor scheduling for multiprogrammed, shared memory multiprocessors. Operating Systems Review, 25(5):26– 40, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vaid:2014:FFC

[VZ14] Kushagra Vaid and Lin Zhong. Fuel, fans, and cores — an introduction to selected papers from HotPower 2013. Operating Systems Review, 48(1):32–33, January 2014. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wentzlaff:2009:FOS

[WA09] David Wentzlaff and Anant Agarwal. Factored operating systems (fos): the case for a scalable operating system for multicores. Operating Systems Review, 43(2):76–85, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wedde:1989:OSS

[WAB^{+}89] H. F. Wedde, G. S. Alijani, W. G. Brown, S. Chen, and G. Kang. Operating system support for adaptive distributed real-time systems in DRAGON SLAYER. Operating Systems Review, 23(3): 126–140, July 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wobber:1993:ATO

[WABL93] Edward Wobber, Martín Abadi, Michael Burrows, and Butler Lampson. Authentication in the Taos operating system. Operating Systems Review, 27(5):256–269, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkinson:1981:PAB

[WAC^{+}81] A. L. Wilkinson, D. H. Anderson, D. P. Chang, Lee Hock Hin, A. J. Mayo, I. T. Viney, R. Williams, and W. Wright. A penetration analysis of a Burroughs Large System. Operating Systems Review, 15(1):14–25, January 1981. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wagner:1998:CCP

[Wag98] Bernhard Wagner. Controlling CGI programs. Operating Systems Review, 32(4):40–46, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1983:RMD

[Wai83a] William M. Waite. Reviews of “Medusa, A Distributed Operating System by John K. Ousterhout”, Harold S. Stone, Series Editor. UMI Research Pres, University Microfilms International, Ann Arbor, Michigan, 1981. Operating Systems Review, 17(2):9–10, April 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Ous81].

Waite:1983:RPS

[Wai83b] William M. Waite. Reviews of “Probability and Statistics with Reliability, Queueing and Computer Science Applications by Kishor S. Trivedi”, Prentice-Hall, Englewood-Cliffs, 1982. Operating Systems Review, 17(2):10, April 1983. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Tri82, Sta83, Tri02].

Waite:1986:BRH

[Wai86] Joanne L. Waite. Book review: A Handbook of Software Development and Operating Procedures for Micro computers by Paul Holliday (Macmillan Publishing Company 1985). Operating Systems Review, 20(2):6–7, April 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1994:BRP

[Wai94] William M. Waite. Book review: Programming with MOTIF by Keith D. Gregory: (Springer-Verlag, New York 1992). Operating Systems Review, 28(1):100, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wainer:1995:IRT

[Wai95a] Gabriel A. Wainer. Implementing real-time services in MINIX. Operating Systems Review, 29(3):75–84, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1995:BRB

[Wai95b] William M. Waite. Book review: Building in Big Brother: The Cryptographic Policy Debate, Lance J. Hoffman. Operating Systems Review, 29(3):2, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1997:BRP

[Wai97a] W. M. Waite. Book review: PostScript & Acrobat/PDF, Thomas Merz. Operating Systems Review, 31(2):1, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1997:BRW

[Wai97b] W. M. Waite. Book reviews: The Web Publisher’s Illustrated Quick Reference, Ralph Grabowski. Operating Systems Review, 31(2):2, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1997:BRR

[Wai97c] W. M. Waite. Book reviews: Robert Slade’s Guide to Computer Viruses. Operating Systems Review, 31(2):1, April 1997. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waite:1998:OSR

[Wai98] W. M. Waite. Is operating systems review obsolete? Operating Systems Review, 32(2):1, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Walther:1973:MSD

[Wal73] W. Walther. Multiprocessor self diagnosis, surgery, and recovery in air terminal traffic control. Operating Systems Review, 7 (4):38–44, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wall:1991:LIL

[Wal91] David W. Wall. Limits of instruction-level parallelism. Operating Systems Review, 25(3S):176–188, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waldspurger:2002:MRM

[Wal02] Carl A. Waldspurger. Memory resource management in VMware ESX server. Operating Systems Review, 36(5S):181–194, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ward:1976:SAO

[War76] Mitchel R. Ward. A simple approach to operating system generation and initialization. Operating Systems Review, 10(1):61– 71, January 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wuu:1986:ESR

[WB86] Gene T. J. Wuu and Arthur J. Bernstein. Efficient solutions to the replicated log and dictionary problems. Operating Systems Review, 20(1):57–66, January 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Walfield:2007:CGH

[WB07] Neal H. Walfield and Marcus Brinkmann. A critique of the GNU Hurd multi-server operating system. Operating Systems Review, 41(4):30–39, July 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weissel:2002:CNS

[WBB02] Andreas Weissel, Björn Beutel, and Frank Bellosa. Cooperative I/O: a novel I/O semantics for energy-aware applications. Operating Systems Review, 36(5S):117–129, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Witten:1983:JDS

[WBC^{+}83] Ian H. Witten, Graham M. Birtwistle, John Cleary, David R. Hill, Danny Levinson, Greg Lomow, Radford Neal, Murray Peterson, Brian W. Unger, and Brian Wyvill. Jade: a distributed software prototyping environment. Operating Systems Review, 17(3):10–23, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wallach:1997:ESA

[WBDF97] Dan S. Wallach, Dirk Balfanz, Drew Dean, and Edward W. Felten. Extensible security architectures for Java. Operating Systems Review, 31(5):116–128, December 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolf:2012:CSO

[WBR^{+}12] Joel Wolf, Andrey Balmin, Deepak Rajan, Kirsten Hildrum, Rohit Khandekar, Sujay Parekh, Kun-Lung Wu, and Rares Vernica. CIRCUMFLEX: a scheduling optimizer for MapReduce workloads with shared scans. Operating Systems Review, 46 (1):26–32, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wada:2002:EDM

[WC02] Yutaka Wada and Zixue Cheng. An efficient distributed method for allocating resources based on an unobstructed squeezing technique. Operating Systems Review, 36(3):33–45, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Witchel:2002:MMP

[WCA02] Emmett Witchel, Josh Cates, and Krste Asanović. Mondrian memory protection. Operating Systems Review, 36(5):304–316, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Welsh:2001:SAW

[WCB01] Matt Welsh, David Culler, and Eric Brewer. SEDA: an architecture for well-conditioned, scalable Internet services. Operating Systems Review, 35(5):230–243, December 2001. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1992:DPS

[WCE^{+}92] John Wilkes, Chia Chao, Robert English, David Jacobson, Bart Sears, Carl Staelin, and Alex Stepanov. DataMesh parallel storage servers. Operating Systems Review, 26(2):11, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:2004:RDS

[WCL^{+}04] Qingsong Wei, Bo Chen, Xianliang Lu, Liyong Ren, and Xu Zhou. The research of the distributed stripped storage spatial model. Operating Systems Review, 38(1):90–96, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wen:2017:REV

[WCL17] Shasha Wen, Milind Chabbi, and Xu Liu. REDSPY: Exploring value locality in software. Operating Systems Review, 51(2): 47–61, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wells:2008:AIF

[WCS08] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Adapting to intermittent faults in multicore systems. Operating Systems Review, 42(2):255–264, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wells:2009:DHN

[WCS09] Philip M. Wells, Koushik Chakraborty, and Gurindar S. Sohi. Dynamic heterogeneity and the need for multicore virtualization. Operating Systems Review, 43(2):5–14, April 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2004:HTV

[WCW^{+}04] Perry H. Wang, Jamison D. Collins, Hong Wang, Dongkeun Kim, Bill Greene, Kai-Ming Chan, Aamir B. Yunus, Terry Sych, Stephen F. Moore, and John P. Shen. Helper threads via virtual multithreading on an experimental Itanium-2 processor-based platform. Operating Systems Review, 38(5):144–155, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2005:SCU

[WCYJ05] S. C. Wang, M. L. Chiang, K. Q. Yan, and K. F. Jea. Streets of consensus under unknown unreliable network. Operating Systems Review, 39(4):80–96, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weinsberg:2008:TFC

[WDA^{+}08] Yaron Weinsberg, Danny Dolev, Tal Anker, Muli Ben-Yehuda, and Pete Wyckoff. Tapping into the fountain of CPUs: on operating system support for programmable devices. Operating Systems Review, 42(2):179–188, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weiser:1989:PCR

[WDH89] M. Weiser, A. Demers, and C. Hauser. The portable common runtime approach to interoperability. Operating Systems Review, 23(5):114–122, December 1989. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wisniewski:2008:KLC

[WdSA^{+}08] Robert W. Wisniewski, Dilma da Silva, Marc Auslander, Orran Krieger, Michal Ostrowski, and Bryan Rosenburg. K42: lessons for the OS community. Operating Systems Review, 42(1):5–12, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weatherspoon:2007:AES

[WECK07] Hakim Weatherspoon, Patrick Eaton, Byung-Gon Chun, and John Kubiatowicz. Antiquity: exploiting a secure log for widearea distributed storage. Operating Systems Review, 41(3):371– 384, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wedekind:1988:UNK

[Wed88] H. Wedekind. Ubiquity and need-to-know: two principles of data distribution. Operating Systems Review, 22(4):39–45, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weihl:1985:DDC

[Wei85] William E. Weihl. Data-dependent concurrency control and recovery. Operating Systems Review, 19(1):19–31, January 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weihl:1992:PTB

[Wei92] William E. Weihl. Prelude: Tools for Building Portable Parallel Programs. Operating Systems Review, 26(2):24, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weissman:1998:PCS

[Wei98] Boris Weissman. Performance counters and state sharing annotations: a unified approach to thread locality. Operating Systems Review, 32(5):127–138, December 1998. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wells:1988:NPI

[Wel88] Codie Wells. A note on “Protection Imperfect”. Operating Systems Review, 22(4):35, October 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [Hog88].

Welch:1995:SPM

[Wel95] Gregory F. Welch. A survey of power management techniques in mobile computing operating systems. Operating Systems Review, 29(4):47–56, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wettstein:1978:PNM

[Wet78] Horst Wettstein. The problem of nested monitor calls revisited. Operating Systems Review, 12(1):19–23, January 1978. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wetherall:1999:ANV

[Wet99] David Wetherall. Active network vision and reality: lessons from a capsule-based system. Operating Systems Review, 33(5): 64–79, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wetherall:2000:ANV

[Wet00] David Wetherall. Active network vision and reality: lessons from a capsule-based system. Operating Systems Review, 34(2): 15–16, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wires:2007:SFS

[WF07] Jake Wires and Michael J. Feeley. Secure file system versioning at the block level. Operating Systems Review, 41(3):203–215, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2007:PCA

[WFHJ07] Helen J. Wang, Xiaofeng Fan, Jon Howell, and Collin Jackson. Protection and communication abstractions for Web browsers in MashupOS. Operating Systems Review, 41(6):1–16, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wampler:2008:NBM

[WG08] Doug Wampler and James H. Graham. A normality based method for detecting kernel rootkits. Operating Systems Review, 42(3):59–64, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2008:HAF

[WGL^{+}08] Xi Wang, Zhenyu Guo, Xuezheng Liu, Zhilei Xu, Haoxiang Lin, Xiaoge Wang, and Zheng Zhang. Hang analysis: fighting responsiveness bugs. Operating Systems Review, 42(4):177–190, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1995:HAH

[WGSS95] J. Wilkes, R. Golding, C. Staelin, and T. Sullivan. The HP AutoRAID hierarchical storage system. Operating Systems Review, 29(5):96–108, December 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolf:1994:SAH

[WH94] Lars C. Wolf and R. G. Herrtwich. The system architecture of the Heidelberg transport system. Operating Systems Review, 28(2):51–64, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waddington:1999:RPG

[WH99] D. G. Waddington and D. Hutchison. Resource partitioning in general purpose operating systems: experimental results in Windows NT. Operating Systems Review, 33(4):52–74, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weinhold:2008:VBV

[WH08] Carsten Weinhold and Hermann Härtig. VPFS: building a virtual private file system with a small trusted computing base. Operating Systems Review, 42(4):81–93, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2017:GSM

[WHZ^{+}17] Kai Wang, Aftab Hussain, Zhiqiang Zuo, Guoqing Xu, and Ardalan Amiri Sani. Graspan: a single-machine disk-based graph system for interprocedural static analyses of large-scale systems code. Operating Systems Review, 51(2):389–404, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wiecek:1992:VM

[Wie92] Cheryl A. Wiecek. VMS on Mach. Operating Systems Review, 26 (2):15, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1980:NHC

[Wil80] M. V. Wilkes. A new hardware capability architecture. Operating Systems Review, 14(2):17–20, April 1980. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1993:DHB

[Wil93] John Wilkes. DataMesh, house-building, and distributed systems technology. Operating Systems Review, 27(2):104–108, April 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1994:OSC

[Wil94] Maurice Wilkes. Operating systems in a changing world. Operating Systems Review, 28(2):9–21, April 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:2009:TRR

[Wil09] John Wilkes. Traveling to Rome: a retrospective on the journey. Operating Systems Review, 43(1):10–15, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Williams:2016:BIC

[Wil16] R. Stanley Williams. Brain inspired computing. Operating Systems Review, 50(2):295, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Winfree:2008:TMP

[Win08] Erik Winfree. Toward molecular programming with DNA. Operating Systems Review, 42(2):1, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wirth:1977:TDR

[Wir77] N. Wirth. Towards a discipline of real-time programming. Operating Systems Review, 11(2):142, April 1977. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wirth:1987:HAP

[Wir87] Niklaus Wirth. Hardware architectures for programming languages and programming languages for hardware architectures. Operating Systems Review, 21(4):2–8, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wiseman:2005:ABS

[Wis05] Yair Wiseman. ARC based superpaging. Operating Systems Review, 39(2):74–78, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Witchel:2016:PPW

[Wit16] Emmett Witchel. Programmer productivity in a world of mushy interfaces: Challenges of the post-ISA reality. Operating Systems Review, 50(2):591, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:1998:SCC

[WJ98] Xiaohui Wei and Jiubin Ju. SFT: a consistent checkpointing algorithm with shorter freezing time. Operating Systems Review, 32(4):70–76, October 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wu:2004:FOM

[WJMC04] Qiang Wu, Philo Juang, Margaret Martonosi, and Douglas W. Clark. Formal online methods for voltage/frequency control in multiple clock domain microprocessors. Operating Systems Review, 38(5):248–259, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2005:ECS

[WK05] Chih-Hung Wang and Yan-Sheng Kuo. An efficient contract signing protocol using the aggregate signature scheme to protect signers’ privacy and promote reliability. Operating Systems Review, 39(4):66–79, October 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wegiel:2008:MCV

[WK08] Michal Wegiel and Chandra Krintz. The Mapping Collector: virtual memory support for generational, parallel, and concurrent compaction. Operating Systems Review, 42(2):91–102, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2007:DCS

[WKL07] Yin Wang, Terence Kelly, and Stéphane Lafortune. Discrete control for safe execution of IT automation workflows. Operating Systems Review, 41(3):305–314, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2013:PTD

[WKT^{+}13] Chengwel Wang, Soila P. Kavulya, Jiaqi Tan, Liting Hu, Mahendra Kutare, Mike Kasick, Karsten Schwan, Priya Narasimhan, and Rajeev Gandhi. Performance troubleshooting in data centers: an annotated bibliography? Operating Systems Review, 47 (3):50–62, December 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weatherly:1982:ESM

[WL82] Richard M. Weatherly and James F. Leathrum. Efficient semaphore management using read/modify/write memory cycles. Operating Systems Review, 16(1):10–13, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Woo:1994:LAP

[WL94] Thomas Y. C. Woo and Simon S. Lam. A lesson on authentication protocol design. Operating Systems Review, 28(3):24–37, July 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wen:2002:DQG

[WL02] Jun Wen and Xianliang Lu. The design of QoS guarantee network subsystem. Operating Systems Review, 36(1):81–87, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Woo:2009:PGO

[WL09] Dong Hyuk Woo and Hsien-Hsin S. Lee. PROPHET: goaloriented provisioning for highly tunable multicore processors in cloud computing. Operating Systems Review, 43(2):102–103, April 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2015:MCC

[WL15] Qiuyun Wang and Benjamin C. Lee. Modeling communication costs in blade servers. Operating Systems Review, 49(2):75–79, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wahbe:1993:ESB

[WLAG93] Robert Wahbe, Steven Lucco, Thomas E. Anderson, and Susan L. Graham. Efficient software-based fault isolation. Operating Systems Review, 27(5):203–216, December 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wen:2001:RNS

[WlL01] Jun Wen and Xiang liang Lu. Realize network subsystem QoS guarantee. Operating Systems Review, 35(3):67–71, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wulf:1975:OHO

[WLP75] W. Wulf, R. Levin, and C. Pierson. Overview of the Hydra Operating System development. Operating Systems Review, 9 (5):122–131, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:2003:NDQ

[WLRZ03] Qingsong Wei, Xianliang Lu, Liyong Ren, and Xu Zhou. A novel disk queue to reduce disk I/O of messaging system. Operating Systems Review, 37(3):55–60, July 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

White:2002:IEE

[WLS^{+}02] Brian White, Jay Lepreau, Leigh Stoller, Robert Ricci, Shashi Guruprasad, Mac Newbold, Mike Hibler, Chad Barb, and Abhijeet Joglekar. An integrated experimental environment for distributed systems and networks. Operating Systems Review, 36 (5S):255–270, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wei:2003:DND

[WLZ03] Qingsong Wei, Xianliang Lu, and Xu Zhou. DFTS: a novel distributed high fault-tolerance storage mechanism. Operating Systems Review, 37(2):19–24, April 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wu:2017:FEF

[WLZJ17] Bo Wu, Xu Liu, Xiaobo Zhou, and Changjun Jiang. FLEP: Enabling flexible and efficient preemption on GPUs. Operating Systems Review, 51(2):483–496, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wettstein:1980:CA

[WM80] H. Wettstein and G. Merbeth. The concept of asynchronization. Operating Systems Review, 14(4):50–70, October 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2016:RTE

[WM16] Xiaodong Wang and José F. Martínez. ReBudget: Trading off efficiency vs. fairness in market-based multicore resource allocation via runtime budget reassignment. Operating Systems Review, 50(2):19–32, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Winograd:1972:SSV

[WMH72] J. Winograd, S. J. Morganstein, and R. Herman. Simulation studies of a virtual memory, time-shared, demand paging operating system. Operating Systems Review, 6(1/2):149–155, June 1972. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weatherspoon:2007:GSS

[WMI^{+}07] Hakim Weatherspoon, Hugo Miranda, Konrad Iwanicki, Ali Ghodsi, and Yann Busnel. Gossiping over storage systems is practical. Operating Systems Review, 41(5):75–81, October 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1980:CMD

[WN80] Maurice V. Wilkes and Roger M. Needham. The Cambridge Model Distributed System. Operating Systems Review, 14(1): 21–29, January 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolthusen:2002:AUC

[Wol02] Stephen D. Wolthusen. Access and use control using externally controlled reference monitors. Operating Systems Review, 36 (1):58–69, January 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wong:1993:DSP

[Won93] K. C. Wong. Determining the shortest process migration paths for program compilation using a dynamic programming approach. Operating Systems Review, 27(2):1–6, April 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wood:1973:ESC

[Woo73] David C. M. Wood. An example in synchronization of cooperating processes: theory and practice. Operating Systems Review, 7(3):10–18, July 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wood:1985:RVV

[Woo85] David C. M. Wood. Review of “VAX/VMS Internals and Data Structures by Lawrence J. Kenah and Simon F. Bate”, Digital Press, Educational Services, Digital Equipment Corporation, Bedford, Massachusetts (1984), ISBN 0-932376-52-5. Operating Systems Review, 19(1):5, January 1985. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic). See [KB84, KGB88, GKD91].

Woo:1990:NLM

[Woo90] Tai-Kuo Woo. A note on Lamport’s mutual exclusion algorithm. Operating Systems Review, 24(4):78–80, October 1990. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wall:1987:MEU

[WP87] David W. Wall and Michael L. Powell. The Mahler experience: using an intermediate language as the machine description. Operating Systems Review, 21(4):100–104, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wagner:1991:BRA

[WP91] David B. Wagner and Carl Ponder. Book review: The Art of Computer Systems Performance Analysis by Raj Jain: (John Wiley and Sons, Inc., New York 1991). Operating Systems Review, 25(3):7–9, July 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wehrmeister:2012:SEV

[WPC12] Marco A. Wehrmeister, Joao G. Packer, and Luis M. Ceron. Support for early verification of embedded real-time systems through UML models simulation. Operating Systems Review, 46(1):73–81, January 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Walker:1983:LDO

[WPE^{+}83] Bruce Walker, Gerald Popek, Robert English, Charles Kline, and Greg Thiel. The LOCUS distributed operating system. Operating Systems Review, 17(5):49–70, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weinstein:1985:TSD

[WPLP85] Matthew J. Weinstein, Thomas W. Page, Jr., Brian K. Livezey, and Gerald J. Popek. Transactions and synchronization in a distributed operating system. Operating Systems Review, 19(5): 115–126, December 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2002:ERR

[WPP02] Limin Wang, Vivek Pai, and Larry Peterson. The effectiveness of request redirection on CDN robustness. Operating Systems Review, 36(5S):345–360, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wu:2024:TTA

[WQA^{+}24] Chenyuan Wu, Haoyun Qin, Mohammad Javad Amiri, Boon Thau Loo, Dahlia Malkhi, and Ryan Marcus. Towards truly adaptive Byzantine fault-tolerant consensus. Operating Systems Review, 58(1):15–22, June 2024. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:/

/dl.acm.org/doi/10.1145/3689051.3689055.

Witchel:2005:MMI

[WRA05] Emmett Witchel, Junghwan Rhee, and Krste Asanović. Mondrix: memory isolation for Linux using Mondrian memory protection. Operating Systems Review, 39(5):31–44, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weiss:1987:SSC

[WS87] Shlomo Weiss and James E. Smith. A study of scalar compilation techniques for pipelined supercomputers. Operating Systems Review, 21(4):105–109, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilkes:1991:SDA

[WS91a] John Wilkes and Raymie Stata. Specifying data availability in multi-device file systems. Operating Systems Review, 25(1):56– 59, January 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolfe:1991:VIS

[WS91b] Andrew Wolfe and John P. Shen. A variable instruction stream extension to the VLIW architecture. Operating Systems Review, 25(3S):2–14, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Walpole:1992:SAS

[WS92] Jonathan Walpole and Richard Staehli. Supporting Access to Stored Multimedia Data in Large Distributed Systems Work in Progress. Operating Systems Review, 26(2):28, April 1992. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wong:2006:CCO

[WS06] Bernard Wong and Emin Gün Sirer. ClosestNode.com: an open access, scalable, shared geocast service for distributed systems. Operating Systems Review, 40(1):62–64, January 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Whitaker:2002:SPD

[WSG02] Andrew Whitaker, Marianne Shaw, and Steven D. Gribble. Scale and performance in the Denali isolation kernel. Operating Systems Review, 36(5S):195–209, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Woo:1994:PAI

[WSH94] Steven Cameron Woo, Jaswinder Pal Singh, and John L. Hennessy. The performance advantages of integrating block data transfer in cache-coherent multiprocessors. Operating Systems Review, 28(5):219–229, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wiseman:2005:EEE

[WSW05] Yair Wiseman, Karsten Schwan, and Patrick Widener. Efficient end to end data exchange using configurable compression. Operating Systems Review, 39(3):4–23, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2010:OCU

[WTB10] Zhikui Wang, Niraj Tolia, and Cullen Bash. Opportunities and challenges to unify workload, power, and cooling management in data centers. Operating Systems Review, 44(3):41–46, July 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weatherspoon:2009:SAS

[WTC09] Hakim Weatherspoon, Doug Terry, and Gregory Chockler. Summary of the 3rd ACM SIGOPS Workshop on Large-Scale Distributed Systems and Middleware (LADIS 2009). Operating Systems Review, 43(4):3–4, December 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wong:2008:TAF

[WTKW08] Chee Siang Wong, Ian Tan, Rosalind Deena Kumari, and Fun Wey. Towards achieving fairness in the Linux scheduler. Operating Systems Review, 42(5):34–43, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wood:2009:MBE

[WTLS^{+}09] Timothy Wood, Gabriel Tarasuk-Levin, Prashant Shenoy, Peter Desnoyers, Emmanuel Cecchet, and Mark D. Corner. Memory buddies: exploiting page sharing for smart colocation in virtualized data centers. Operating Systems Review, 43(3):27–36, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Waddington:2002:IPE

[WV02] Daniel G. Waddington and Ramesh Viswanathan. Interaction points: exploiting operating system mechanisms for intercomponent communications. Operating Systems Review, 36(2):

19–35, April 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolman:1999:SPC

[WVS^{+}99] Alec Wolman, M. Voelker, Nitin Sharma, Neal Cardwell, Anna Karlin, and Henry M. Levy. On the scale and performance of cooperative Web proxy caching. Operating Systems Review, 33 (5):16–31, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wolman:2000:SPC

[WVS^{+}00] Alec Wolman, Geoffrey M. Voelker, Nitin Sharma, Neal Cardwell, Anna Karlin, and Henry M. Levy. On the scale and performance of cooperative Web proxy caching. Operating Systems Review, 34(2):11–12, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wiedenhoft:2008:PME

[WWGF08] Geovani Ricardo Wiedenhoft, Lucas Francisco Wanner, Giovani Gracioli, and Antônio Augusto Fröhlich. Power management in the EPOS system. Operating Systems Review, 42(6):71–80, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wu:2008:DNL

[WXX08] Fengguang Wu, Hongsheng Xi, and Chenfeng Xu. On the design of a new Linux readahead framework. Operating Systems Review, 42(5):75–84, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2004:RFD

[WY04] S. C. Wang and K. Q. Yan. Revisiting fault diagnosis agreement in a new territory. Operating Systems Review, 38(2):41–61, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wobber:2007:AAS

[WYA^{+}07] Ted Wobber, Aydan Yumerefendi, Martín Abadi, Andrew Birrell, and Daniel R. Simon. Authorizing applications in singularity. Operating Systems Review, 41(3):355–368, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2003:ACP

[WYC03a] S. C. Wang, K. Q. Yan, and C. F. Cheng. Asynchronous consensus protocol for the unreliable un-fully connected network. Operating Systems Review, 37(3):43–54, July 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2003:RAU

[WYC03b] S. C. Wang, K. Q. Yan, and C. F. Cheng. Reaching agreement on an unknown network with partial graphic information. Operating Systems Review, 37(4):70–89, October 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2021:GFG

[WYD^{+}21] Yiqiu Wang, Shangdi Yu, Laxman Dhulipala, Yan Gu, and Julian Shun. GeoGraph: a framework for graph processing on geometric data. Operating Systems Review, 55(1):38–46, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3469379.3469384.

Wang:2025:TLL

[WYW^{+}25] Yuan Wang, Zhenyuan Yang, Zhanbo Wang, Mingyu Li, Zhilin Wu, and Haibo Chen. Towards large language model-friendly APls. Operating Systems Review, 59(2):17–23, July 2025. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wu:1994:ENV

[WZ94] Michael Wu and Willy Zwaenepoel. eNVy: a non-volatile, main memory storage system. Operating Systems Review, 28(5):86– 97, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:2008:PIM

[WZWS08] Kun Wang, Yu Zhang, Huayong Wang, and Xiaowei Shen. Parallelization of IBM Mambo system simulator in functional modes. Operating Systems Review, 42(1):71–76, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

West:2010:OCM

[WZWZ10] Richard West, Puneet Zaroo, Carl A. Waldspurger, and Xiao Zhang. Online cache modeling for commodity multicore processors. Operating Systems Review, 44(4):19–29, December 2010.

CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wang:1993:GCC

[WZZ93] Xingwei Wang, Hong Zhao, and Jiakeng Zhu. GRPC: a communication cooperation mechanism in distributed systems. Operating Systems Review, 27(3):75–86, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2017:HAE

[XD17] Xin Xu and Bhavesh Davda. A hypervisor approach to enable live migration with passthrough SR-IOV network devices. Operating Systems Review, 51(1):15–23, August 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xie:1995:IIO

[XDC^{+}95] Li Xie, Xing Du, Jun Chen, Yuhua Zheng, and Zhongxiu Sun. An introduction to intelligent operating system KZ2. Operating Systems Review, 29(1):29–46, January 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xie:2018:EDS

[XDM^{+}18] Shaolin Xie, Scott Davidson, Ikuo Magaki, Moein Khazraee, Luis Vega, Lu Zhang, and Michael B. Taylor. Extreme datacenter specialization for planet-scale computing: ASIC clouds. Operating Systems Review, 52(1):96–108, July 2018. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xia:2008:PBM

[XFO08] Ying Xia, Kevin Fairbanks, and Henry Owen. A program behavior matching architecture for probabilistic file system forensics. Operating Systems Review, 42(3):4–13, April 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2006:RTR

[XHB06] Min Xu, Mark D. Hill, and Rastislav Bodik. A regulated transitive reduction (RTR) for longer memory race recording. Operating Systems Review, 40(5):49–60, December 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xiao-Hui:1999:SAS

[XHJB99] Wei Xiao-Hui and Ju Jiu-Bin. SCR algorithm: saving/restoring states of file systems. Operating Systems Review, 33(1):26–33, January 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xia:2009:IVP

[XLDB09] Lei Xia, Jack Lange, Peter Dinda, and Chang Bae. Investigating virtual passthrough I/O on commodity devices. Operating Systems Review, 43(3):83–94, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2000:TS

[Xu00] Baowen Xu. Tagged semaphores. Operating Systems Review, 34(3):11–15, July 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comments [Tro00].

Ximing:2000:RIC

[XX00] Chen Ximing and Lu Xianliang. Runtime incremental concentrated scheduling on NOW(NRICS). Operating Systems Review, 34(2):84–96, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2004:DDC

[XXM04] Zhou Xu, Lu Xialiang, and Hou Mengshu. DCFS: distributed cooperative fault-tolerance storage mechanism. Operating Systems Review, 38(3):18–25, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:2005:SBA

[XXMC05] Zhou Xu, Lu Xianliang, Hou Mengshu, and Zhan Chuan. A speed-based adaptive dynamic parallel downloading technique. Operating Systems Review, 39(1):63–69, January 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:1997:PCP

[XZZ97] Shouhuai Xu, Gendu Zhang, and Hong Zhu. On the properties of cryptographic protocols and the weaknesses of the BAN-like logics. Operating Systems Review, 31(4):12–23, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xu:1998:STP

[XZZ98] Shouhuai Xu, Gendu Zhang, and Hong Zhu. On the security of three-party cryptographic protocols. Operating Systems Review, 32(3):7–20, July 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comments [Ng99].

Yeo:1993:TIN

[YAK93] A. K. Yeo, A. L. Ananda, and E. K. Koh. A taxonomy of issues in name systems design and implementation. Operating Systems Review, 27(3):4–18, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:1992:DCN

[Yan92] Cui-Qing Yang. Distributed computing in a NUMP (Non-Uniform Message-Passing) environment. Operating Systems Review, 26(2):82–91, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yuan-bo:2004:ITA

[YbJf04] Guo Yuan-bo and Ma Jian-feng. An intrusion-tolerant authorization and authentication scheme in distributed environments. Operating Systems Review, 38(4):45–51, October 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:2025:CIM

[YBQ^{+}25] Zhenning Yang, Archit Bhatnagar, Yiming Qiu, Tongyuan Miao, Patrick Tser Jern Kon, Yunming Xiao, Yibo Huang, Martin Casado, and Ang Chen. Cloud infrastructure management in the age of AI agents. Operating Systems Review, 59(2):1–8, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:1996:CPD

[YD96] Zhonghua Yang and Keith Duddy. CORBA: a platform for distributed object computing. Operating Systems Review, 30(2): 4–31, April 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Youhui:2002:CBH

[YD02] Zhang Youhui and Wang Dongsheng. A checkpoint-based high availability run-time system for Windows NT clusters. Operating Systems Review, 36(2):12–18, April 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:2025:DDR

[YG25] Rui Yang and Rajiv Gupta. DREAM: Distributed regional efficient agent management with LLMs for online multi-agent pathfinding. Operating Systems Review, 59(2):24–33, July 2025. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:2003:TER

[YGG^{+}03] Kun Yang, Xin Guo, Alex Galis, Bo Yang, and Dayou Liu. Towards efficient resource on-demand in Grid Computing. Operating Systems Review, 37(2):37–43, April 2003. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2016:CWM

[YJX^{+}16] Xiao Yu, Pallavi Joshi, Jianwu Xu, Guoliang Jin, Hui Zhang, and Guofei Jiang. CloudSeer: Workflow monitoring of cloud infrastructures via interleaved logs. Operating Systems Review, 50(2):489–502, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2002:IWS

[yKPR02] Hyong youb Kim, Vijay S. Pai, and Scott Rixner. Increasing Web server throughput with network interface data caching. Operating Systems Review, 36(5):239–250, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2006:TOT

[yKR06] Hyong youb Kim and Scott Rixner. TCP offload through connection handoff. Operating Systems Review, 40(4):279–290, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lam:1991:ISD

[yL91] Kwok yan Lam. An implementation for small databases with high availability. Operating Systems Review, 25(4):77, October 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yeo:2002:PAU

[YLE02] C. K. Yeo, B. S. Lee, and M. H. Er. A peering architecture for ubiquitous IP multicast streaming. Operating Systems Review, 36(3):82–95, July 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yuan:2006:AKP

[YLW^{+}06] Chun Yuan, Ni Lao, Ji-Rong Wen, Jiwei Li, Zheng Zhang, Yi-Min Wang, and Wei-Ying Ma. Automated known problem diagnosis with event traces. Operating Systems Review, 40(4): 375–388, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:1993:ABG

[YM93] Zhonghua Yang and T. Anthony Marsland. Annotated bibliography on global states and times in distributed systems. Operating Systems Review, 27(3):55–74, July 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yasa:2012:SSD

[YN12] Giridhar Appaji Nag Yasa and P. C. Nagesh. Space savings and design considerations in variable length deduplication. Operating Systems Review, 46(3):57–64, December 2012. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yanok:2015:TLV

[YN15] Ilya Yanok and Nathaniel Nystrom. Tapir: a language for verified OS kernel probes. Operating Systems Review, 49(2):51–56, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Young:1992:ELT

[You92] Michael Wayne Young. Episode: Lazy Transactions for Filesystem Meta-Data Updates. Operating Systems Review, 26(2):20, April 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2005:RED

[YRC05] Yuan Yu, Tom Rodeheffer, and Wei Chen. RaceTrack: efficient detection of data race conditions via adaptive tracking. Operating Systems Review, 39(5):221–234, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yoon:2004:SUA

[YRY04] Eun-Jun Yoon, Eun-Kyung Ryu, and Kee-Young Yoo. A secure user authentication scheme using hash functions. Operating Systems Review, 38(2):62–68, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Young:1994:IAS

[YS94] Cliff Young and Michael D. Smith. Improving the accuracy of static branch prediction using branch correlation. Operating Systems Review, 28(5):232–241, December 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:1998:CHR

[YS98] Zhonghua Yang and Chengzheng Sun. CORBA for hard real time applications: some critical issues. Operating Systems Review, 32(3):64–71, July 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yeh:2002:SAK

[YS02] Her-Tyan Yeh and Hung-Min Sun. Simple authenticated key agreement protocol resistant to password guessing attacks. Operating Systems Review, 36(4):14–22, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yoon:2016:PPI

[YSCC16] Man-Ki Yoon, Negin Salajegheh, Yin Chen, and Mihai Christodorescu. PIFT: Predictive information-flow tracking. Operating Systems Review, 50(2):713–725, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yokote:1991:MOA

[YTM^{+}91] Yasuhiko Yokote, Fumio Teraoka, Atsushi Mitsuzawa, Nobuhisa Fujinami, and Mario Tokoro. The muse object architecture: a new operating system structuring concept. Operating Systems Review, 25(2):22–46, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Young:1987:DMC

[YTR^{+}87] M. Young, A. Tevanian, R. Rashid, D. Golub, and J. Eppinger. The duality of memory and communication in the implementation of a multiprocessor operating system. Operating Systems Review, 21(5):63–76, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2000:TTAa

[Yu00a] Haifeng Yu. TACT: tunable availability and consistency tradeoffs for replicated Internet services. Operating Systems Review, 34(2):33, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2000:TTAb

[Yu00b] Haifeng Yu. TACT: tunable availability and consistency tradeoffs for replicated Internet services (poster session). Operating Systems Review, 34(2):40, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yuen:1985:PTP

[Yue85] C. K. Yuen. On programs, tasks and processes. Operating Systems Review, 19(3):7–8, July 1985. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yuval:1976:ONS

[Yuv76] G. Yuval. An operating non-system. Operating Systems Review, 10(3):9–10, July 1976. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2001:CLA

[YV01] Haifeng Yu and Amin Vahdat. The costs and limits of availability for replicated services. Operating Systems Review, 35 (5):29–42, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yan:2018:HTC

[YVCB18] Zi Yan, Ján Veselý, Guilherme Cox, and Abhishek Bhattacharjee. Hardware translation coherence for virtualized systems. Operating Systems Review, 52(1):57–70, July 2018. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yazd:2013:BEE

[YVM13] Sara Arbab Yazd, Subbarayan Venkatesan, and Neeraj Mittal. Boosting energy efficiency with mirrored data block replication policy and energy scheduler. Operating Systems Review, 47(2): 33–40, July 2013. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yang:2004:ISE

[YW04] Chou-Chen Yang and Ren-Ching Wang. An improvement of security enhancement for the timestamp-based password authentication scheme using Smart Cards. Operating Systems Review, 38(3):91–96, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yan:2005:BFC

[YW05] Kuo-Qin Yan and Shu-Ching Wang. The bounds of faulty components on consensus with dual failure modes. Operating Systems Review, 39(3):82–89, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yao:2006:RNR

[YW06] Xiaoyu Yao and Jun Wang. RIMAC: a novel redundancy-based hierarchical cache architecture for energy efficient, high performance storage systems. Operating Systems Review, 40(4):249– 262, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yan:2004:NAR

[YWC04] Kuo-Qin Yan, Shu-Ching Wang, and Mao-Lun Chiang. New application of reliable agreement: underlying an unsecured business environment. Operating Systems Review, 38(3):42–57, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yuan:2015:MLF

[YWKYS15] Xinhao Yuan, David Williams-King, Junfeng Yang, and Simha Sethumadhavan. Making lock-free data structures verifiable with artificial transactions. Operating Systems Review, 49(2): 57–63, December 2015. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yan:2011:OSV

[YZG^{+}11] Shoumeng Yan, Xiaocheng Zhou, Ying Gao, Hu Chen, Gansha Wu, Sai Luo, and Bratin Saha. Optimizing a shared virtual memory system for a heterogeneous CPU-accelerator platform. Operating Systems Review, 45(1):92–100, January 2011. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yao:2002:PNA

[YZJ02] Nian-Min Yao, Ming-Yang Zheng, and Jiu-Bin Ju. Pipeline: a new architecture of high performance servers. Operating Systems Review, 36(4):55–64, October 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Yu:2006:UUB

[YZZZ06] Hongliang Yu, Dongdong Zheng, Ben Y. Zhao, and Weimin Zheng. Understanding user behavior in large-scale video-on-demand systems. Operating Systems Review, 40(4):333–344, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zayas:1987:APM

[Zay87] E. Zayas. Attacking the process migration bottleneck. Operating Systems Review, 21(5):13–24, November 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zheng:2007:ACI

[ZBN07] Wei Zheng, Ricardo Bianchini, and Thu D. Nguyen. Automatic configuration of Internet services. Operating Systems Review, 41(3):219–229, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhai:2002:COS

[ZCSM02] Antonia Zhai, Christopher B. Colohan, J. Gregory Steffan, and Todd C. Mowry. Compiler optimization of scalar value communication between speculative threads. Operating Systems Review, 36(5):171–183, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2005:HHD

[ZCT^{+}05] Qingbo Zhu, Zhifeng Chen, Lin Tan, Yuanyuan Zhou, Kimberly Keeton, and John Wilkes. Hibernator: helping disk arrays sleep through the winter. Operating Systems Review, 39(5):177–190, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhongxiu:1983:ZDO

[ZDP83] Sun Zhongxiu, Zhang Du, and Yan Peigen. ZCZOS: a distributed operating system for a LSI-11 microcomputer network.

Operating Systems Review, 17(3):30–34, July 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2016:DEQ

[ZE16] Haishan Zhu and Mattan Erez. Dirigent: Enforcing QoS for latency-critical tasks on shared multicore systems. Operating Systems Review, 50(2):33–47, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zeadally:1997:ERT

[Zea97] Sherali Zeadally. An evaluation of the real-time performances of SVR4.0 and SVR4.2. Operating Systems Review, 31(1):78– 87, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zelkowitz:1974:ITD

[Zel74] Marvin V. Zelkowitz. It is not time to define “structured programming”. Operating Systems Review, 8(2):7–8, April 1974. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zeng:2002:EME

[ZELV02] Heng Zeng, Carla S. Ellis, Alvin R. Lebeck, and Amin Vahdat. ECOSystem: managing energy as a first class operating system resource. Operating Systems Review, 36(5):123–132, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2021:TPF

[ZFP^{+}21] Xiaowei Zhu, Zhisong Fu, Zhenxuan Pan, Jin Jiang, Chuntao Hong, Yongchao Liu, Yang Fang, Wenguang Chen, and Changhua He. Taking the pulse of financial activities with online graph processing. Operating Systems Review, 55(1):84–87, June 2021. CODEN OSRED8. ISSN 0163-5980 (print), 1943-

586X (electronic). URL https://dl.acm.org/doi/10.1145/
3469379.3469389.

Zhu:2010:ILS

[ZFW10] Kenny Q. Zhu, Kathleen Fisher, and David Walker. Incremental learning of system log formats. Operating Systems Review, 44 (1):85–90, January 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2007:HHF

[ZG07] Zhihui Zhang and Kanad Ghose. hFS: a hybrid file system prototype for improving small file and metadata performance. Operating Systems Review, 41(3):175–187, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2016:MPU

[ZH16] Huazhe Zhang and Henry Hoffmann. Maximizing performance under a power cap: a comparison of hardware, software, and hybrid techniques. Operating Systems Review, 50(2):545–559, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2019:LOS

[ZH19] Yiying Zhang and Yutong Huang. “Learned”: Operating systems. Operating Systems Review, 53(1):40–45, July 2019. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2023:MIR

[Zha23] Yiying Zhang. Make it real: an end-to-end implementation of a physically disaggregated data center. Operating Systems Review, 57(1):1–9, June 2023. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https://dl.acm.

org/doi/10.1145/3606557.3606559.

Zheng:2006:PEA

[ZHK06] Gengbin Zheng, Chao Huang, and Laxmikant V. Kalé. Performance evaluation of automatic checkpoint-based fault tolerance for AMPI and Charm++. Operating Systems Review, 40(2): 90–99, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhou:2010:VN

[Zho10] Shudong Zhou. Virtual networking. Operating Systems Review, 44(4):80–85, December 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhou:2016:PUH

[Zho16] Yuanyuan Zhou. Programming uncertain <T> hings. Operating Systems Review, 50(2):1–2, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhou:1996:PET

[ZIL96] Yuanyuan Zhou, Liviu Iftode, and Kai Li. Performance evaluation of two home-based lazy release consistency protocols for shared virtual memory systems. Operating Systems Review, 30 (SI):75–88, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zimmermann:1994:MDM

[Zim94] Chris Zimmermann. Making distributed multimedia systems secure: the switchboard approach. Operating Systems Review, 28(1):88–100, January 1994. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2017:PPD

[ZJL17] Tong Zhang, Changhee Jung, and Dongyoon Lee. ProRace: Practical data race detection for production use. Operating Systems Review, 51(2):149–162, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2011:TPS

[ZJS^{+}11] David (Yu) Zhu, Jaeyeon Jung, Dawn Song, Tadayoshi Kohno, and David Wetherall. TaintEraser: protecting sensitive data leaks using application-level taint tracking. Operating Systems Review, 45(1):142–154, January 2011. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zobel:1988:RTC

[ZK88] Dieter Zöbel and Christoph Koch. Resolution techniques and complexity results with deadlocks: a classifying and annotated bibliography. Operating Systems Review, 22(1):52–72, January 1988. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:1986:PNR

[ZL86] Du Zhang and Meiliu Lu. Process name resolution in faultintolerant CSP programs. Operating Systems Review, 20(4):9– 15, October 1986. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2004:AFV

[ZL04a] Yuqing Zhang and Xiuying Liu. An approach to the formal verification of the three-principal cryptographic protocols. Operating Systems Review, 38(1):35–42, January 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2004:RMA

[ZL04b] Yuqing Zhang and Xiuying Liu. Running-mode analysis of the Security Socket Layer protocol. Operating Systems Review, 38 (2):34–40, April 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2016:TED

[ZLJ16] Tong Zhang, Dongyoon Lee, and Changhee Jung. TxRace: Efficient data race detection using commodity hardware transactional memory. Operating Systems Review, 50(2):159–173, June 2016. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2007:BHR

[ZLL^{+}07] Zheng Zhang, Qiao Lian, Shiding Lin, Wei Chen, Yu Chen, and Chao Jin. BitVault: a highly reliable distributed data retention platform. Operating Systems Review, 41(2):27–36, April 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhongxiu:1980:IDX

[ZLX^{+}80] Sun Zhongxiu, Xie Li, Fei Xianglin, Yi Wenguo, and Tan Yaoming. An introduction to DJS200/XT1. Operating Systems Review, 14(3):70–74, July 1980. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:1999:AFV

[ZLX99] Yuqing Zhang, Jihong Li, and Guozhen Xiao. An approach to the formal verification of the two-party cryptographic protocols. Operating Systems Review, 33(4):48–51, October 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). See comments [JW01].

Zhu:2001:MMO

[ZLX01a] Ming-Yuan Zhu, Lei Luo, and Guang-Ze Xiong. The minimal model of operating systems. Operating Systems Review, 35(3): 22–29, July 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2001:PCO

[ZLX01b] Ming-Yuan Zhu, Lei Luo, and Guang-Zhe Xiong. A provably correct operating system: δ-core. Operating Systems Review, 35 (1):17–33, January 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zadok:2000:FLS

[ZN00] Erez Zadok and Jason Nieh. FIST: a language for stackable file systems. Operating Systems Review, 34(2):38, April 2000. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zobel:1983:DPC

[Zöb83] Dieter Zöbel. The Deadlock problem: a classifying bibliography. Operating Systems Review, 17(4):6–15, October 1983. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zuberi:1999:ESM

[ZPS99] Khawar M. Zuberi, Padmanabhan Pillai, and Kang G. Shin. EMERALDS: a small-memory real-time microkernel. Operating Systems Review, 33(5):277–299, December 1999. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zuberi:2000:ESM

[ZPS00] Khawar M. Zuberi, Padmanabhan Pillai, and Kang G. Shin. EMERALDS: a small-memory real-time microkernel. Operating Systems Review, 34(2):28–29, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhou:2004:DTP

[ZPS^{+}04] Pin Zhou, Vivek Pandey, Jagadeesan Sundaresan, Anand Raghuraman, Yuanyuan Zhou, and Sanjeev Kumar. Dynamic tracking of page miss ratio curve for memory management. Operating Systems Review, 38(5):177–188, December 2004. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zahir:2000:CCD

[ZRMH00] Rumi Zahir, Jonathan Ross, Dale Morris, and Drew Hess. OS and compiler considerations in the design of the IA-64 architecture. Operating Systems Review, 34(5):212–221, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhong:2006:RWB

[ZS06] Ming Zhong and Kai Shen. Random walk based node sampling in self-organizing networks. Operating Systems Review, 40(3): 49–55, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2017:ISC

[ZSG^{+}17] Rui Zhang, Natalie Stanley, Christopher Griggs, Andrew Chi, and Cynthia Sturton. Identifying security critical properties for the dynamic verification of a processor. Operating Systems Review, 51(2):541–554, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:1997:MIL

[ZSK97] Weiping Zhu, Piotr Socko, and Bartek Kiepuszewski. Migration impact on load balancing—an experience on Amoeba. Operating Systems Review, 31(1):43–53, January 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhong:2008:RDC

[ZSS08] Ming Zhong, Kai Shen, and Joel Seiferas. Replication degree customization for high availability. Operating Systems Review, 42(4):55–68, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhu:2009:WDC

[ZUW^{+}09] Xiaoyun Zhu, Mustafa Uysal, Zhikui Wang, Sharad Singhal, Arif Merchant, Pradeep Padala, and Kang Shin. What does control theory bring to systems research? Operating Systems Review, 43(1):62–69, January 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:1997:SSA

[ZWG^{+}97] Xiaolan Zhang, Zheng Wang, Nicholas Gloy, J. Bradley Chen, and Michael D. Smith. System support for automatic profiling and optimization. Operating Systems Review, 31(5):15–26, December 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhao:2009:DMB

[ZWL09] Weiming Zhao, Zhenlin Wang, and Yingwei Luo. Dynamic memory balancing for virtual machines. Operating Systems Review, 43(3):37–47, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2001:USC

[ZWWL01] Yuqing Zhang, Chunling Wang, Jianping Wu, and Xing Li. Using SMV for cryptographic protocol analysis: a case study. Operating Systems Review, 35(2):43–50, April 2001. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2001:TCR

[ZWZ01] Youhui Zhang, Dongsheng Wang, and Weimin Zheng. Transparent checkpointing and rollback recovery mechanism for Windows NT applications. Operating Systems Review, 35(2):78–85, April 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2005:ULC

[ZWZ05] Youhui Zhang, Dongsheng Wong, and Weimin Zheng. Userlevel checkpoint and recovery for LAM/MPI. Operating Systems Review, 39(3):72–81, July 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

ZhouXu:2004:DDR

[ZXMJ04] ZhouXu, Lu Xianliang, Hou Mengshu, and Wu Jin. A dynamic distributed replica management mechanism based on accessing frequency detecting. Operating Systems Review, 38(3):26–34, July 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2000:FVL

[ZYG00] Youtao Zhang, Jun Yang, and Rajiv Gupta. Frequent value locality and value-centric data cache design. Operating Systems Review, 34(5):150–159, December 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhang:2003:ULC

[ZZ03] Youhui Zhang and Weimin Zheng. User-level communication based cooperative caching. Operating Systems Review, 37(1):23– 33, January 2003. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zdancewic:2001:UHC

[ZZNM01] Steve Zdancewic, Lantian Zheng, Nathaniel Nystrom, and Andrew C. Myers. Untrusted hosts and confidentiality: secure program partitioning. Operating Systems Review, 35(5):1–14, December 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Zhuang:2004:HIE

[ZZP04] Xiaotong Zhuang, Tao Zhang, and Santosh Pande. HIDE: an infrastructure for efficiently protecting information leakage on the address bus. Operating Systems Review, 38(5):72–84, December 2004. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).